### v1

# Large Language Model Psychometrics: A Systematic Review of Evaluation, Validation, and Enhancement

Abstract
The rapid advancement of large language models (LLMs) has outpaced traditional evaluation methodologies. It presents novel challenges, such as measuring human-like psychological constructs, navigating beyond static and task-specific benchmarks, and establishing human-centered evaluation. These challenges intersect with Psychometrics, the science of quantifying the intangible aspects of human psychology, such as personality, values, and intelligence. This survey introduces and synthesizes an emerging interdisciplinary field of LLM Psychometrics, which leverages psychometric instruments, theories, and principles to evaluate, understand, and enhance LLMs. We systematically explore the role of Psychometrics in shaping benchmarking principles, broadening evaluation scopes, refining methodologies, validating results, and advancing LLM capabilities. This paper integrates diverse perspectives to provide a structured framework for researchers across disciplines, enabling a more comprehensive understanding of this nascent field. Ultimately, we aim to provide actionable insights for developing future evaluation paradigms that align with human-level AI and promote the advancement of human-centered AI systems for societal benefit. A curated repository of LLM psychometric resources is available at https://github.com/valuebyte-ai/Awesome-LLM-Psychometrics.
1 Introduction
"Whatever exists at all exists in some amount. To know it thoroughly involves knowing its quantity as well as its quality." [Thorndike, 1962]
The advent of large language models (LLMs) represents a transformative breakthrough in AI. These systems exhibit general-purpose capabilities spanning diverse domains [Bubeck et al., 2023], with particular proficiency in natural language understanding and generation [Demszky et al., 2023, Grossmann et al., 2023, Ziems et al., 2024, Gu et al., 2024]. They are rapidly integrated into critical societal infrastructure, ranging from consumer-facing applications like chatbots [OpenAI, 2025] and search engines [Wang et al., 2024c] to high-stakes domains such as healthcare [Singhal et al., 2023], education [Milano et al., 2023], and scientific discovery [Romera-Paredes et al., 2024]. Their increasing dominance has exposed a fundamental, pressing scientific challenge: how can we rigorously evaluate these AI systems that transcend traditional benchmarks of biological or algorithmic intelligence?
Traditional AI evaluation has relied on curating task-specific datasets, annotating ground-truth labels with human input, running models on these datasets, and assessing performance using predefined metrics. However, LLMs have triggered an evaluation crisis, as their versatile capabilities and human-like behaviors exceed what traditional benchmarks can measure. Novel challenges include evaluating psychological constructs like personality, values, and cognitive biases that conventional assessment cannot capture. Rapid LLM development and training data contamination have made static benchmarks obsolete. Additionally, the prompt- and context-sensitive nature of LLMs compromises the robustness and validity of existing evaluation frameworks. With increasing human-LLM interaction, human-centered evaluation approaches are becoming essential. Moreover, as LLMs integrate into agentic and multimodal systems, evaluation methodologies must expand in scope and complexity.
These challenges intersect with humanity’s century-old quest to quantify the complex, intangible human psychology, including knowledge, skills, character, values [Pasquali, 2009]. Psychometrics emerged from this timeless pursuit as the scientific study of psychological measurement. It bridges the abstract and the empirical by transforming human traits into quantifiable data, enabling better understanding, prediction, and decision-making in education, business, healthcare, governance, and beyond [Rust and Golombok, 2014]. The convergence of LLMs and psychometrics forms a methodological crucible, forging novel paradigms to better decode and improve machine minds.
We define LLM Psychometrics as the interdisciplinary field dedicated to evaluating, understanding, and enhancing LLMs through the application and integration of psychometric instruments, theories, and principles. This field seeks to quantify, interpret, manipulate, and improve the complex, human-like attributes and behaviors exhibited by LLMs, encompassing both personality constructs (such as personality, values, morality, and attitudes) and cognitive constructs (including heuristics and biases, social interaction abilities, psycholinguistic abilities, and learning and cognitive capacities). Research in LLM psychometrics applies, extends, and innovates upon scientific methods of psychological measurement for LLMs. Based on the psychometric principles, related research systematically examines the measurement results to ensure scientific rigor. By measuring and elucidating psychological constructs in LLMs, LLM psychometrics further informs strategies for their targeted enhancement.
Recent research in LLM psychometrics pioneers in addressing the LLM evaluation crisis. Some studies introduce dynamic and construct-oriented evaluation frameworks that move beyond static, task-specific benchmarks [Zhu et al., 2024a, Hagendorff, 2023]. In parallel, novel methodologies are developed to measure non-cognitive and emergent constructs [Ren et al., 2024, Huang et al., 2023d, Pellert et al., 2024]. Self-adaptive evaluation techniques now allow for the extrapolation of item difficulty and tailoring assessments to model performance [Jiang et al., 2024a, Lalor et al., 2024, Polo et al., 2024]. Drawing from the methodological framework of psychometric validation, research improves the reliability and validity of evaluation protocols [Ye et al., 2025a]. Human-centered evaluation drives aligning model behavior with human values [Wang et al., 2024f, Yao et al., 2025a]. In addition, the scope of evaluation expands to agentic and multimodal systems, further broadening the methodological landscape [Li et al., 2024b, Huang et al., 2024c].
The field of LLM psychometrics has seen significant growth, evidenced by the proliferation of related research papers. These studies, however, address a variety of psychological constructs, employ diverse methodologies, and utilize distinct validation techniques. The interdisciplinary nature of this domain has attracted contributions from a broad spectrum of academic fields. Despite this diversity, there is a lack of cohesion among researchers from different communities, leading to a fragmentation of insights, particularly between studies focusing on disparate constructs. Consequently, there is a pressing need for a comprehensive survey to synthesize these efforts and facilitate a more integrated understanding of the field.
Figure 1: Overview of this review.
This paper aims to bridge the gap by providing the first systematic review of LLM psychometrics, encompassing evaluation, validation, and enhancement. Fig. 1 illustrates the structure of the review. § 2 provides an overview of the preliminaries and methodological foundations to facilitate subsequent discussions. In § 3, we delineate the definition, scope, and taxonomy of LLM psychometrics, which establishes the structure of the remainder of the review. The core measurement framework is detailed in § 5, § 6, and § 7. § 5 delves into the psychological constructs evaluated in LLMs, elucidating the theories employed and summarizing key evaluation findings. In § 6, we scrutinize the psychometric evaluation methodologies applied to LLMs, followed by § 7, which examines the psychometric validation of evaluation results. Beyond evaluation, § 8 introduces strategies for enhancing LLMs through psychometric insights. Finally, § 9 discusses emerging trends, challenges, and future directions in the psychometric evaluation of LLMs, while § 10 concludes the paper.
Related Surveys.
We consider studies to be out of scope if they do not employ psychometric approaches, adhere to psychometric principles, or if they focus solely on scalar performance metrics rather than characterizing behaviors. Readers interested in conventional LLM benchmarking are encouraged to consult surveys on LLM evaluation [Guo et al., 2023b, Chang et al., 2024]. Several related surveys focus on the evaluation of specific constructs in LLMs, such as personality [Wen et al., 2024b, Dong et al., 2025], attitudes and values [Ma et al., 2024a], cultural awareness [Pawar et al., 2024, Adilazuarda et al., 2024], and theory of mind [Sarıtaş et al., 2025, Dong et al., 2025]. Hagendorff [2023], Hagendorff et al. [2024] introduce the notion of machine psychology and review emergent LLM abilities; however, they do not provide comprehensive coverage of the related research, nor do they elaborate on LLM personality constructs, psychometric validation, or enhancement. This paper provides the first systematic survey of LLM psychometrics.
2 Preliminary and Methodological Foundations
We aim for our survey to be self-contained and accessible to a broad, cross-disciplinary audience. To this end, this section presents the preliminary and methodological foundations that underpin the subsequent discussions.
2.1 Large Language Models
LLMs are large-scale deep neural networks—essentially complex systems of nonlinear regression equations. An LLM can generate text by predicting the next token (word or subword) in a sequential manner (autoregressive generation), given the preceding context. It does so by modelling a conditional probability distribution over the vocabulary; i.e., the likelihood of each token given the context:
$P(x_{t}|x_{<t})=f(x_{<t};\theta),$ (1)
where $x_{t}$ is the token at time step $t$; $x_{<t}$ is the context preceding $x_{t}$, usually including both user prompts and previously generated tokens; $f$ is the model’s parameterized function; and $\theta$ represents the model parameters. Given $f$, the model generates text by either sampling from this distribution or directly selecting the token with the highest probability. In the former case, hyperparameters such as temperature can be adjusted to control the diversity of generated text. In the latter case, the model is said to use greedy decoding, and the generated text is deterministic. In evaluating LLMs, it is crucial to properly account for the stochasticity of the model.
These models are based primarily on the transformer architecture, a neural network design that employs self-attention mechanisms to capture contextual relationships between words, phrases, and broader linguistic patterns. Modern LLMs typically contain billions of parameters, enabling them to efficiently learn from vast amounts of textual data. During evaluation, if the model has already been exposed to the test items during training, this is referred to as data contamination. In such cases, the model is more likely to exhibit artificially inflated performance or simply reproduce memorized patterns, rather than revealing its true underlying capabilities or traits.
The training process of LLMs is typically divided into two phases: pre-training and post-training. Pre-training is the phase where LLMs learn to predict the next token, given its preceding context, on a large corpus of text data. This process is unsupervised, as the model does not require explicit labels or annotations to learn the underlying patterns in the data. The model processes Internet-scale text data from diverse sources like books, articles, and websites. By repeatedly predicting the next word in a sentence, the model learns the statistical properties of language and gains large-scale world knowledge. Models that have only undergone the pre-training phase are usually referred to as base models. Post-training, or fine-tuning, is the process of adapting the base models to better follow user instructions, align with human values, or specialize in particular tasks. This stage typically involves training the model on a smaller, human-annotated dataset or incorporating human feedback on the quality of the model outputs. Models that have undergone both phases are often referred to as fine-tuned models, instruction-tuned models, or aligned models.
We interact with LLMs using prompts, which are input instructions to the model. For psychometric evaluation, these prompts can naively be the reformatted versions of test items originally designed for humans, adapted for LLMs to answer. When designing prompts, one should consider differences between base and fine-tuned models. Most public-facing LLMs are fine-tuned, so evaluation research primarily focuses on these models for greater practical relevance.
A key emergent capability of LLMs is in-context learning, which allows models to adapt to new tasks or patterns by conditioning on examples or instructions provided within the input context $x_{<t}$, without modifying model parameters. This property can influence LLM performance in psychometric evaluation. For instance, prompting models to reason step-by-step (e.g., Chain-of-Thought prompting [Wei et al., 2022]) can enhance performance on reasoning tasks, while instructing them to role-play may modulate their exhibited personalities and values.
2.2 Psychometrics
Psychometrics, also known as psychological testing, involves the use of tests to measure, understand, or predict behavior by quantifying specific actions or characteristics. These tests rely on samples of behavior, meaning they are not perfect measures and often include errors inherent to sampling. Test items are specific stimuli designed to elicit observable reactions that can be scored or evaluated. Typically, tests are composed of multiple questions or problems as their items, producing explicit data subject to scientific analysis.
A psychological test is a set of items that are designed to measure characteristics of human beings that pertain to behavior [Kaplan and Saccuzzo, 2001]. Behavior measured by tests can be overt (observable actions) or covert (internal thoughts or feelings). Tests may assess past, current, or even predict future behavior. Interpretation of test scores depends on their context within a distribution. Scales are used to relate raw scores to defined distributions, aiding interpretation. In addition, psychological tests can measure traits—enduring tendencies like shyness or determination—and states, which reflect temporary conditions of individuals.
Psychological testing measures individual differences in various constructs, which are abstract psychological attributes or dimensions that help explain and predict behavior. Two primary categories of such constructs are personality constructs and cognitive constructs [Kaplan and Saccuzzo, 2001]. Personality tests focus on an individual’s tendencies and dispositions. These tests measure typical behavior, such as preferences or tendencies to react in certain ways. Cognitive tests evaluate speed, accuracy, or both, with higher scores reflecting better performance.
Two fundamental principles underpin psychometrics: reliability and validity [Raykov and Marcoulides, 2011]. Reliability ensures accuracy, dependability, consistency, or repeatability of the test results. Reliable test results are stable across time, contexts, and raters. Validity confirms the meaningfulness and usefulness of the test results. A valid measure captures the intended construct. Validity is multifaceted; for example, predictive validity might correlate test scores with job performance, while construct validity ensures alignment with theoretical models, such as the Big Five personality traits [Goldberg, 2013].
Other principles include standardization, which provides context to raw scores by comparing individual results to a representative sample, or norm group. Additionally, equivalence and fairness are crucial principles that tests must adhere to. Test bias occurs when test items unintentionally advantage or disadvantage subgroups. Modern psychometrics employs advanced statistical models to identify and revise biased items, ensuring assessments measure the intended construct rather than extraneous factors [Rust and Golombok, 2014].
2.3 Psychometric Evaluation of AI Before the Era of LLMs
The idea of applying psychometrics to AI originated in the early decades of AI [Pellert et al., 2024]. Evans [1964] pioneered work in this area by creating a heuristic program that could solve parts of intelligence tests. Subsequent efforts similarly focused on designing AI systems for cognitive tests [Newell, 1973], with the goal of creating systems capable of handling human tasks. This was conceptually aligned with the development of static, task-centric benchmarks in modern AI research [Hendrycks et al., 2020, Chen et al., 2021, Liang et al., 2022, Srivastava et al., 2022, Lee et al., 2024e]. However, criticisms emerged regarding the absence of "hot cognition" in AI, prompting Simon [1963] to propose incorporating emotional aspects into models. By the early 2000s, the concept of "psychometric AI" was explicitly articulated as the pursuit of systems capable of excelling on all established, validated tests of intelligence and mental ability. These included not only conventional IQ tests but also assessments of artistic and literary creativity, mechanical ability, and beyond [Bringsjord and Schimanski, 2003, Pellert et al., 2024]. It was not until the advent of LLMs that the versatility envisioned for "psychometric AI" began to materialize.
3 LLM Psychometrics: Definition, Scope, and Taxonomy
We define LLM Psychometrics as the interdisciplinary field dedicated to evaluating, understanding, and enhancing LLMs through the application and integration of psychometric instruments, theories, and principles.
LLM psychometrics research adapts, advances, and innovates upon psychometric methodologies to rigorously evaluate LLMs. This evaluation framework encompasses three core dimensions: the target construct (what to measure), the measurement method (how to measure), and the validation of results (how well do we measure). In many cases, these psychometric insights not only inform evaluation but also guide the development and refinement of LLMs (how to improve).
The remainder of this paper begins by thoroughly contrasting psychometrics and traditional AI benchmarking (§ 4.1), identifying key differences such as core goals, philosophical foundations, measurement methods, and result analysis (Table 1). We then review how recent LLM psychometrics research adopts and extends psychometric principles to underpin and reinvent LLM benchmarks (§ 4.2).
We proceed by examining the principal constructs assessed within LLM psychometrics (what to measure; § 5), including both personality (§ 5.1) and cognitive constructs (§ 5.2). Personality constructs are categorized into personality traits, values, morality, and attitudes and opinions, while cognitive constructs comprise heuristics and biases, social interaction abilities, psycholinguistic abilities, and learning and cognitive capacities. For each construct, we synthesize the underlying theoretical frameworks and measurement instruments, critically compare alternative approaches, and distill the main empirical findings.
We further scrutinize the psychometric evaluation methodologies employed for LLMs (how to measure; § 6). The methodological framework is systematically analyzed in terms of its key components: test format (§ 6.1), data and task sources (§ 6.2), prompting strategies (§ 6.3), and model output and scoring (§ 6.4). In addition, we discuss the role of inference parameters in shaping evaluation results (§ 6.5).
We next address the psychometric validation of measurement results, examining both reliability and validity (how well do we measure; § 7). Reliability is assessed using metrics such as internal consistency, parallel forms reliability, and inter-rater agreement (§ 7.1). Validity is considered from content, construct, criterion, and ecological perspectives (§ 7.2). Persistent concerns include prompt sensitivity, data contamination, response biases, and divergence from human psychological constructs. To mitigate these challenges, recent research advocates for standardized, procedural tests, including multiple task versions, prompt perturbation, and transparent reporting practices, thereby promoting rigorous, reproducible, and meaningful psychometric evaluation of LLMs (§ 7.3).
The psychometric insights further inform the development and refinement of LLMs (how to improve; § 8). Current research predominantly focuses on enhancing LLMs’ performance along three dimensions: trait manipulation (§ 8.1), safety and alignment (§ 8.2), and cognitive enhancement (§ 8.3). Finally, we discuss the current trends, challenges, and future directions of LLM psychometrics (§ 9).
4 Psychometrics for Benchmarking Principles
Table 1: Comparison between psychometrics and conventional AI benchmark.
Feature Psychometrics AI benchmark
Core goal To measure psychological constructs, to prove that a test measures as intended (validity evidence), and to understand the construct being measured. To test and compare the task performance of different LLMs. Focuses on ranking models and selecting the best one suited for a specific task.
Philosophy of measurement Construct-oriented. Tends towards a causal approach to measurement, where the measured trait is believed to cause the measurement outcomes. Task-oriented. Leans towards representativism, assuming items exhaust or represent all aspects of the underlying ability.
Target construct Personality and ability. Mostly task-specific abilities.
Construct definition Emphasizes clear and detailed definitions of the construct being measured. Agreement on the construct definition is a byproduct of test development. Often defines constructs implicitly through ad hoc task selection. Construct definitions can be vague.
Development process Systematic and rigorous, often following methods like Evidence-Centered Design (ECD). Can be labor-intensive. Compiles a set of relevant questions or tasks, then performs expert annotation or crowdsourcing to label ground truth answers. Less labor-intensive per item.
Number of items Can vary, but not necessarily large. Focus is on item quality and relevance to the construct. Typically consists of an extensive number of questions to cover various aspects of abilities. Reliability increases with test length.
Sample size Typically requires a larger sample size of test takers for robust statistical modeling. Can be applied to evaluate the performance of a single LLM on the benchmark.
Statistical modeling Employs advanced and various statistical models like Item Response Theory and Factor Analysis to analyze data, estimate latent abilities, and assess model fit. Often relies on simple aggregation methods, such as calculating average accuracy across benchmark tasks.
Result analysis Ensures the reliability, validity, predictive power, and explanatory power of the test through result analysis and statistical modeling. Reliability is likely to be high due to the large number of items. However, validity, predictive power, or explanatory power beyond the target task is not a primary concern.
4.1 Fundamental Differences Between Psychometrics and AI Benchmarking
Benchmarking AI systems superficially resembles psychometrics, particularly Classical Test Theory (CTT) [Crocker and Algina, 1986], as both compile test items to evaluate cognitive capabilities and average the resulting scores. However, closer examination reveals that AI benchmarks differ significantly from modern psychometric approaches [Wang et al., 2023a, Federiakin, 2025]. We outline these key differences in Table 1.
Psychometrics.
Psychometrics centers on understanding the psychological constructs and ensuring that tests accurately measure the intended constructs. Grounded in a causal measurement philosophy, this field posits that observed test responses arise from latent psychological constructs [Markus and Borsboom, 2013, Federiakin, 2025]. These constructs may encompass both abilities (e.g., reasoning skills) and personalities (e.g., conscientiousness). The causal framework necessitates rigorous construct definition, requiring traits to be precisely delineated through iterative theory-building and empirical validation. Psychometric test development follows methodical protocols, often structured by frameworks such as Evidence-Centered Design (ECD) [Mislevy et al., 2003]. ECD emphasizes ensuring congruence between test items and theoretical models of the construct, thereby supporting robust inferences about latent traits.
Central to this approach is the prioritization of item quality over quantity. Psychometricians conduct rigorous item analyses to balance precision with practicality, as administering an excessive number of items to participants is often impractical. Advanced statistical models, such as Item Response Theory (IRT) [Embretson and Reise, 2013] and Factor Analysis [Loehlin, 2004], are adopted to estimate latent traits, analyze item performance, and assess model fit. These models require relatively large sample sizes (the number of human participants) to yield stable parameter estimates, as they must disentangle individual differences from measurement error to accurately infer latent traits.
The test results are analyzed to ensure the reliability, validity, predictive power, and explanatory power of the test. Specifically, a well-designed test should: 1) consistently and accurately measure the intended construct, 2) predict performance across a diverse range of related tasks and real-world outcomes, and 3) provide explanatory insight into the observed data. For example, psychometric models often reveal that individual differences across a broad range of cognitive tasks can be captured and explained by a relatively small set of underlying cognitive abilities [Cattell and Horn, 1978].
Benchmark.
In contrast, AI benchmarking is driven by pragmatic goals: evaluating and ranking models based on task performance. Unlike psychometrics, validity is not the primary concern. Instead, benchmarks typically emphasize breadth, scalability, and—especially in the era of foundation models—difficulty. This approach reflects a representativist philosophy, where it is assumed that an extensive set of benchmark items collectively captures all relevant aspects of the abilities demanded by the target task [Federiakin, 2025]. However, constructs like reasoning or knowledge are often ambiguously defined and encompass infinitely many aspects. Benchmarks implicitly operationalize these constructs through ad hoc task selection.
The development of AI benchmarks is usually less labor-intensive, especially when compared with psychometrics on a per-item basis. Test items and their corresponding ground truths are typically drawn from existing datasets, expert curation, or crowdsourced contributions. While this process enables scalability, it risks conflating superficial task performance with deeper cognitive capacities. For instance, a benchmark may assess mathematical reasoning through arithmetic problems without verifying whether models rely on pattern recognition versus symbolic logic [Ahn et al., 2024]. Additionally, LLM benchmarking commonly employs straightforward metrics, such as average accuracy, eschewing the sophisticated latent variable models of psychometrics. This simplicity allows benchmarks to evaluate single models efficiently, bypassing the need for population samples. However, it also limits the depth of insights that can be gleaned from model performance [Federiakin, 2025].
Reliability and stability in benchmarking are primarily achieved through scaling up the test. However, ensuring the quality of each individual item becomes impractical due to the test scale and the rapid pace of model development. For instance, while psychometrics emphasizes the discriminative power of each item, some benchmarks, though initially challenging, are quickly outpaced by continuous model improvements [McIntosh et al., 2024]. Conversely, certain emerging benchmarks are currently too difficult to yield meaningful comparisons [Phan et al., 2025]. Benchmark results are often limited to the specific target task, offering limited generalizability or predictive power across other tasks or real-world applications. These results also pose significant challenges for conducting in-depth, multi-faceted analyses of model capabilities [Wang et al., 2023a].
4.2 Benchmarking with Psychometrics-Inspired Principles
Recent LLM evaluation efforts have drawn inspiration from psychometrics and seek to develop benchmarks that adhere to psychometric principles.
Construct-Oriented Benchmarking.
Task-oriented benchmarks often entail vast question sets to capture complex abilities. However, in many cases, the benchmarks either fail to fully represent these abilities due to their infinite manifestations or involve extraneous factors that are irrelevant to the target ability [Zhou et al., 2025, Wallach et al., 2025]. Recent research has drawn inspiration from psychometrics and explored the paradigm of construct-oriented evaluation, seeking its discriminative, predictive, and explanatory power. Ilić and Gignac [2024], Federiakin [2025] employ factor analysis to explore the latent variables underlying LLM benchmark performance. Their findings reveal a monolithic factor resembling general intelligence or ability. Federiakin [2025] ranks models based on this discovered factor and highlights its unique advantages over raw benchmark scores. In contrast, similar attempts by Burnell et al. [2023] identify three factors—reasoning, comprehension, and core language modeling—that better explain LLM performance across 27 cognitive tasks. Based on it, Zhu et al. [2024a] integrate the three factors into benchmark items to evaluate multifaceted abilities. This discrepancy between the estimated latent factors in the above findings can be attributed to differences in the models and benchmarks employed [Zhou et al., 2025]. Therefore, rather than relying on statistically derived factors, Zhou et al. [2025] propose a theory-driven hierarchical set of general scales for systematic construct-oriented evaluation. These scales are validated to explain what AI systems can do and predict their performance on novel task instances. Peng et al. [2024] present the Tong Test, a value- and ability-oriented framework, for Artificial General Intelligence (AGI) evaluation. This framework is rooted in dynamic embodied physical and social interactions (DEPSI), and can generate an infinite variety of tasks to evaluate key capabilities including values, learning, and cognition.
Psychometrically Rigorous Benchmark Development.
Beyond defining and analyzing latent constructs, researchers have developed holistic, psychometrically rigorous methods for benchmark development. Liu et al. [2024e] introduce Evidence-Centered Benchmark Design (ECBD), a framework structuring benchmark creation into five modules—capability, content, adaptation, assembly, and evidence—each requiring justification to ensure validity. Through case studies of prominent LLM benchmarks, they demonstrate ECBD’s utility in identifying validity threats. Similarly grounded in psychometrics but with a distinct approach, Fang et al. [2024] propose Psychometrics-Assisted Benchmarking (PATCH), an eight-step process from construct definition to proficiency scoring. When piloted on 8th-grade mathematics, PATCH produced results diverging from traditional benchmarks, offering a more comprehensive evaluation. Building on related principles, Kardanova et al. [2024] adapt Evidence-Centered Design (ECD) to create psychometrically grounded benchmarks. Their application in pedagogy illustrates how this method can reduce data contamination and enhance test interpretability.
Item Response Theory.
Item Response Theory (IRT) is a statistical approach in psychometrics that jointly estimates the latent ability of examinees and the properties (difficulty, discrimination) of test items [Embretson and Reise, 2013]. Translating this framework to LLM evaluation enables researchers to infer latent ability scores, assess item informativeness, and perform more efficient evaluation. Recent research leverages principles from IRT to develop adaptive evaluation frameworks. These methods dynamically calibrate item difficulty based on model performance and weighting items by their inferred difficulty, aiming to achieve accurate evaluations with smaller test sizes and more discriminative items [Lalor et al., 2024, Polo et al., 2024, Zhuang et al., 2023b, a, Guinet et al., 2024]. Building on these approaches, Jiang et al. [2024a], Truong et al. [2025] introduce IRT-based benchmarks that involve learning to estimate item difficulty and learning to generate novel items calibrated to specific difficulty levels. Additionally, research has used IRT-based analyses to explore the alignment between LLM and human response distributions [He-Yueya et al., 2024]. IRT-based evaluations further offer the potential to estimate construct and item parameters on a unified scale, enabling direct comparisons across AI systems and against human norms, even when different test sets are used [Wang et al., 2023a, Fang et al., 2024].
5 Psychometrics for Measuring Psychological Constructs
This section delves into the psychological constructs evaluated in LLM psychometrics. Fig. 2 exemplifies the tests for the involved constructs.
Constructs Theories/inventories Related work
Personality traits Big Five Serapio-García et al. [2023], Jiang et al. [2023], Li et al. [2022b], Zou et al. [2024], Bhandari et al. [2025a], Huang et al. [2024c], Romero et al. [2024], Lee et al. [2024d], Liu et al. [2024a], Karra et al. [2022], Song et al. [2023], Caron and Srivastava [2022], Li et al. [2022a], Kovač et al. [2023], Frisch and Giulianelli [2024], Zheng et al. [2025], La Cava and Tagarelli [2024], Bodroža et al. [2024], Huang et al. [2023a], Gupta et al. [2024], Huang et al. [2023e], Pellert et al. [2024], Ai et al. [2024], Dorner et al. [2023], Klinkert et al. [2024], Petrov et al. [2024], Shu et al. [2024], Ren et al. [2024], Bhandari et al. [2025b], Hilliard et al. [2024], Jiang et al. [2024b], Lu et al. [2023], Sühr et al. [2023], Zhang [2024], Jain et al. [2024]
HEXACO Wang et al. [2025], Bodroža et al. [2024], Miotto et al. [2022], Barua et al. [2024], Ren et al. [2024], Peereboom et al. [2024]
MBTI Rao et al. [2023], Pan and Zeng [2023], Song et al. [2024b], Cui et al. [2023], La Cava and Tagarelli [2024], Zhang et al. [2024a], Huang et al. [2023c], Ai et al. [2024], Lu et al. [2023], Chen et al. [2024a]
Dark Triad Li et al. [2022b], Romero et al. [2024], Lee et al. [2024d], Li et al. [2022a], Huang et al. [2023e], Barua et al. [2024], Peereboom et al. [2024], Lu et al. [2023]
Other & custom Wen et al. [2024b] (reviews); Mao et al. [2024a], Ai et al. [2024], Jiang et al. [2023], Zeng [2024]
Values Schwartz Pellert et al. [2024], Yao et al. [2025a], Shen et al. [2024], Kovač et al. [2024], Rozen et al. [2024], Zhang et al. [2023a], Yao et al. [2024], Hadar-Shoval et al. [2024], Fischer et al. [2023], Kovač et al. [2023], Miotto et al. [2022], Lee et al. [2024a], Ren et al. [2024], Ye et al. [2025a], Li et al. [2024b], Yao et al. [2025b], Cahyawijaya et al. [2024]
WVS Chiu et al. [2025], Yao et al. [2025a], Kim and Baek [2024], Faulborn et al. [2025]
VSM Kovač et al. [2023], Kharchenko et al. [2024], Zhong et al. [2024], Ren et al. [2024], Ye et al. [2025a]
GLOBE Li et al. [2024d], Karinshak et al. [2024], Ren et al. [2024]
SVO Zhang et al. [2024c]
Other & custom Zhang et al. [2024b], Li et al. [2024d], Biedma et al. [2024], Ye et al. [2025b], Meadows et al. [2024], Jiang et al. [2024a], Moore et al. [2024], Xu et al. [2023], Liu et al. [2024b]
Morality MFT Pellert et al. [2024], Nunes et al. [2024], Abdulhai et al. [2024], Neuman et al. [2025], Ji et al. [2024], Simmons [2023], Tlaie [2024], Aksoy [2024], Yan et al. [2024], Fraser et al. [2022], Zhou et al. [2024a], Yao et al. [2025b], Münker [2024]
ETHICS Jinnai [2024], Rodionov et al. [2023], Karpov et al. [2024], Yu et al. [2023], Albrecht et al.
DIT Khandelwal et al. [2024], Tanmay et al. [2023]
Other & custom Tanmay et al. [2023], Jin et al. [2022], Huang et al. [2024a], Jin et al. [2024b], Seror [2024], Bonagiri et al. [2024], Garcia et al. [2024], Ahmad and Takemoto [2024], Jiang et al. [2024a], Takemoto [2024], Han [2023], Peterson [2025], Ramezani and Xu [2023], Kucuk and Kocyigit [2023], Neuman et al. [2025], Chiu et al. [2025], Scherrer et al. [2023], Marraffini et al. [2024], Vida et al. [2024], Jinnai [2024], Meijer et al. [2024], Yuan et al. [2024], Liu et al. [2024d], Ohashi et al. [2024], Sachdeva and van Nuenen [2025]
Attitudes & opinions ANES Qi et al. [2024], Jiang et al. [2024d], Yang et al. [2024], Argyle et al. [2023], Bisbee et al. [2024], Sun et al. [2024], Jiang et al. [2022]
ATP Santurkar et al. [2023], Hwang et al. [2023], Tjuatja et al. [2024]
GLES Ma et al. [2024b], Ball et al. [2025], Qi et al. [2024], von der Heyde et al. [2024]
PCT Azzopardi and Moshfeghi [2024], Bernardelle et al. [2024], Wright et al. [2024], Röttger et al. [2024], Hartmann et al. [2023], Rozado [2023], Rutinowski et al. [2024]
Other & custom Ma et al. [2024a] (reviews); Dominguez-Olmedo et al. [2024], Kim and Lee [2023], Sanders et al. [2023], Kalinin [2023], Geng et al. [2024], Xu et al. [2025c], Lee et al. [2024c], Durmus et al. [2023], Rozado [2023], Rosenbusch et al. [2023], Wu et al. [2023], Chalkidis and Brandl [2024], Feng et al. [2023], Ceron et al. [2024], Röttger et al. [2025], Kim et al. [2025b], Faulborn et al. [2025]
Table 2: Representative LLM personality constructs in LLM psychometrics. The main dimensions or focus of the theories/inventories are listed in Table 3.
Theory/inventory What it measures / dimensions
Big Five Five broad personality traits: Openness, Conscientiousness, Extraversion, Agreeableness, Neuroticism
HEXACO Six personality traits: Honesty-Humility, Emotionality, Extraversion, Agreeableness, Conscientiousness, Openness
MBTI Four dichotomies: Extraversion–Introversion, Sensing–Intuition, Thinking–Feeling, Judging–Perceiving
Dark Triad Three negative personality traits: Narcissism, Machiavellianism, Psychopathy
Schwartz Basic human values: 10 or more values (e.g., Self-Direction, Stimulation, Hedonism, Achievement, Power, Security, Conformity, Tradition, Benevolence, Universalism), typically grouped into four higher-order categories (Openness to Change, Self-Enhancement, Conservation, Self-Transcendence)
WVS World Values Survey: Assesses broad cultural values such as traditional vs. secular-rational values, survival vs. self-expression values
VSM Value Survey Module (often Hofstede): Cultural dimensions such as Power Distance, Individualism, Masculinity, Uncertainty Avoidance, Long-Term Orientation, Indulgence
GLOBE Global Leadership and Organizational Behavior Effectiveness: Nine cultural dimensions (e.g., Performance Orientation, Assertiveness, Future Orientation, Humane Orientation, Institutional Collectivism, In-Group Collectivism, Gender Egalitarianism, Power Distance, Uncertainty Avoidance)
SVO Social Value Orientation: Measures individuals’ preferences regarding resource allocation between oneself and others (e.g., prosocial, individualistic, competitive orientations)
MFT Moral Foundations Theory: Five (sometimes six) moral foundations—Care/Harm, Fairness/Cheating, Loyalty/Betrayal, Authority/Subversion, Sanctity/Degradation, (Liberty/Oppression)
ETHICS Various ethics-related measures assessing moral reasoning, ethical principles, or moral preferences
DIT Defining Issues Test: Assesses moral development and reasoning using moral dilemmas
ANES American National Election Studies: Political attitudes, beliefs, and behaviors in the U.S.
ATP Attitudes Toward Politics: Measures political attitudes and social controversies, often specific to a region or subject
GLES German Longitudinal Election Study: Political attitudes, beliefs, and voting behaviors in Germany
PCT Political Compass Test: Economic (Left/Right) and Social (Authoritarian/Libertarian) political dimensions
Table 3: Personality theories and inventories measured in LLM psychometrics and their main dimensions or focus.
Figure 2: Examples of psychometric tests for LLMs.
5.1 Measuring Personality Constructs
LLMs exhibit personality constructs that are not explicitly programmed or trained towards. The emergent constructs critically shape LLM behavior [Ye et al., 2025a, Hagendorff, 2023], with profound implications for both individuals and broader social groups [Bengio et al., 2024]. Measuring these embedded psychological constructs is essential for understanding behavior, identifying biases, and fostering responsible development.
Table 2 summarizes the representative personality constructs that have garnered attention in recent research. Researchers typically select constructs based on their relevance to LLM development and deployment, as well as the applicability of these constructs to AI systems. For instance, Li et al. [2024d] argue that emotional variability in LLMs is not a meaningful construct, given that LLMs lack the biological mechanisms underlying emotions. Conversely, personality and values are considered meaningful for LLMs, as they influence user interactions and model outputs [Serapio-García et al., 2023, Ye et al., 2025a]. This rationale has led to extensive research on personality, values, morality, and attitudes & opinions, as well as some—though less—focus on other constructs such as career selection [Hua et al., 2024], motivation [Chiu et al., 2025, Huang et al., 2023e], and mental health [Reuben et al., 2024, De Duro et al., 2024].
5.1.1 Personality Traits
"Personality is the enduring configuration of characteristics and behavior that comprises an individual’s unique adjustment to life." [APA Dictionary of Psychology, n.d.]
Personality traits define an individual’s patterns of thinking, feeling, and behaving [Larsen et al., 2005]. In the context of LLMs, Zhang et al. [2024a], Wang et al. [2025] have explored how personality traits relate to model safety, bias, and toxicity. Other works emphasize that chatbot personality highly determines user experience [Jiang et al., 2023, Serapio-García et al., 2023, Huang et al., 2024c, Klinkert et al., 2024].
The study of personality traits leads to several prominent theoretical models, each offering unique insights into individual differences. The Big Five model [Goldberg, 2013] identifies five core dimensions of personality: Openness, Conscientiousness, Extraversion, Agreeableness, and Neuroticism. Building upon the Big Five, the HEXACO model [Ashton and Lee, 2007] introduces an additional dimension, Honesty-Humility, which encompasses traits such as sincerity, fairness, and modesty. Another influential framework, the Myers-Briggs Type Indicator (MBTI) [Myers et al., 1962], categorizes individuals into 16 distinct personality types based on four dichotomous dimensions: Extraversion vs. Introversion, Sensing vs. Intuition, Thinking vs. Feeling, and Judging vs. Perceiving. In contrast to these models, the Dark Triad [Paulhus and Williams, 2002] focuses on three socially aversive personality traits—Machiavellianism, Narcissism, and Psychopathy—highlighting the darker aspects of human behavior.
When applying psychometrics to LLMs, most researchers directly administered the established inventories, such as NEO-PI-R [Costa and McCrae, 2008], BFI [John et al., 1991], and BFI-2 [Soto and John, 2017], for measuring the Big Five traits [Caron and Srivastava, 2022, Huang et al., 2023d, a, Karra et al., 2022, La Cava and Tagarelli, 2024, Serapio-García et al., 2023]; HEXACO-60 [Ashton and Lee, 2009] and HEXACO-100 [Lee and Ashton, 2018] inventories for HEXACO traits [Miotto et al., 2022, Wang et al., 2025]; MBTI assessment [Myers, 1985] for MBTI types [Cui et al., 2023, Huang et al., 2023c, La Cava and Tagarelli, 2024, Pan and Zeng, 2023, Rao et al., 2023]; and Dark Triad Dirty Dozen scale [Jonason and Webster, 2010] for Dark Triad traits [Barua et al., 2024, Huang et al., 2023d, Romero et al., 2024].
Given widespread concerns about the practical relevance of these inventories [Ai et al., 2024, Dorner et al., 2023, Gupta et al., 2024, Pellert et al., 2024, Shu et al., 2024, Song et al., 2023, Serapio-García et al., 2023], Some researchers have adapted them for more real-world scenarios. For example, Bhandari et al. [2025a], Frisch and Giulianelli [2024], Klinkert et al. [2024], Liu et al. [2024a], Zheng et al. [2025], Zou et al. [2024], Song et al. [2024b] contextualize the tests in topic-specific or open-domain conversations. Ai et al. [2024] accompany the self-report tests with behavioral tests to examine the personality knowledge of LLMs. Meanwhile, Jiang et al. [2023] compile and adapt existing tests into the Machine Personality Inventory, while Mao et al. [2024a] develop the PersonalityEdit inventory using LLMs. However, Peereboom et al. [2024] argue that human-derived traits may not meaningfully apply to LLMs, underscoring the necessity for psychometric theories specifically designed for LLM analysis (§ 7.2.2).
Comparing Personality Models and Suggesting Use Cases.
The Big Five offers robust empirical foundations and is the most prominent model for personality assessment. However, it is insufficient for predicting comprehensive psychological outcomes [Feher and Vernon, 2021]. HEXACO extends the model by adding Honesty-Humility and provides nuanced insights into moral traits, making it preferable for LLM application contexts demanding integrity [Wang et al., 2025]. In contrast, MBTI [Myers et al., 1962] categorizes 16 types via dichotomous scales. While popular in career counseling and team-building for its accessibility, its typological approach oversimplifies personality and lacks strong empirical support, limiting its utility in high-stakes decisions [Pittenger, 2005]. The Dark Triad [Paulhus and Williams, 2002] diverges by focusing on maladaptive traits, making it useful in identifying toxic behaviors in LLMs [Barua et al., 2024].
Main Findings.
Early models such as GPT-3 score above human averages on Dark Triad scales [Li et al., 2022b, a, Romero et al., 2024]. Even after safety tuning, some models retain certain "dark qualities" [Li et al., 2022b], suggesting that these patterns may be deeply rooted in the pretraining data. More advanced LLMs indicate better level of alignment. They usually demonstrate high Openness, Extraversion, and Agreeableness traits, while low Neuroticism, on the Big Five personality tests [Zou et al., 2024, Bhandari et al., 2025a, Karra et al., 2022, La Cava and Tagarelli, 2024, Huang et al., 2023e, Klinkert et al., 2024]. The results align with their design as assistive and helpful entities with emotional stability and engaging personality [Bhandari et al., 2025a, La Cava and Tagarelli, 2024]. Within the MBTI framework, most proprietary LLMs are classified as ENFJ or INFJ types [La Cava and Tagarelli, 2024, Zhang et al., 2024a, Huang et al., 2023c], indicating tendencies toward helpfulness, idealism, and planning capabilities.
Personality differences between models are noteworthy. Comparative studies reveal that LLMs from different generations and training methodologies display unique combinations of personality characteristics [Bhandari et al., 2025a, Karra et al., 2022, Jiang et al., 2023, La Cava and Tagarelli, 2024, Bodroža et al., 2024, Huang et al., 2023e, Klinkert et al., 2024]. LLM personalities also vary across contexts. The same model may display different personality traits across conversational topics [Zou et al., 2024, Song et al., 2023, Caron and Srivastava, 2022, Bodroža et al., 2024, Petrov et al., 2024]. This variability challenges the applicability of human personality theories that assume relative trait stability. Multiple studies suggest that prompt design and system instructions substantially influence LLM personality expression [Song et al., 2023, Caron and Srivastava, 2022, Gupta et al., 2024, Dorner et al., 2023]. We defer related discussions to § 7.
5.1.2 Values
"Values are enduring beliefs that guide behavior and decision-making, reflecting what is important and desirable to an individual or group." [Schwartz, 1992]
Values are central to understanding human motivation, attitudes, and behavior, shaping how individuals perceive themselves and the world around them. According to Schwartz’s theory of basic human values [Schwartz, 2012], the most established value theory, they 1) are beliefs linked inextricably to affect, 2) are desirable goals that motivate actions, 3) transcend specific situations and actions, 4) serve as standards or criteria for evaluating and selecting actions, policies, people, and events, and 5) are ordered by relative importance, which guides choices and actions. These unique characteristics make values a powerful lens for understanding LLM behavior. For example, Ye et al. [2025a] show how different values contribute to the safety of LLMs; Liu et al. [2024b] reveal that different spiritual values affect LLMs in social-fairness scenarios; and Sorensen et al. [2024b] demonstrate that standard value alignment reduces distributional pluralism in LLM outputs.
Schwartz’s Theory.
Schwartz [1992] identify 10 basic human values: Self-Direction, Stimulation, Hedonism, Achievement, Power, Security, Conformity, Tradition, Benevolence, and Universalism. These values are further grouped into four higher-order dimensions: Openness to Change, Conservation, Self-Enhancement, and Self-Transcendence. Later iterations of the theory refine its nuances by subdividing the 10 values into finer categories [Schwartz et al., 2012]. Most research on Schwartz’s values in LLMs focuses on the 10 basic values, while some studies explore the 19 refined values [Rozen et al., 2024]. Many investigations directly apply established inventories, i.e., Schwartz Value Survey (SVS) [Schwartz, 1992] or the Portrait Values Questionnaire (PVQ) [Schwartz et al., 2001] (e.g., [Hadar-Shoval et al., 2024, Fischer et al., 2023, Kovač et al., 2023, Lee et al., 2024a, Miotto et al., 2022]). Other research adapts these inventories to better suit the context of LLMs. For example, Ren et al. [2024] rephrase survey items into advice-seeking questions, and Shen et al. [2024] transform subjective statements into opinion-seeking questions about AI actions. Rather than directly administering or rephrasing static inventories, Ye et al. [2025a], Chiu et al. [2025], Yao et al. [2024] leverage existing cross-domain LLM prompt datasets or LLM-generated contextualized questions as test items.
World Values Survey (WVS).
The World Values Survey (WVS) is a global research project that explores people’s values and beliefs, how they change over time, and their social and political impact [Haerpfer et al., 2022]. It measures a broad spectrum of values, including cultural values defined on two broad dimensions: traditional vs. secular-rational values and survival vs. self-expression values. Kim and Baek [2024] directly prompt LLMs with WVS survey items. In contrast, Yao et al. [2025a], Chiu et al. [2025] use model-generated or self-designed prompts to measure these cultural values in LLMs.
Values Survey Module (VSM).
Hofstede [1984] identifies six key value dimensions that describe cultural differences: Power Distance, Individualism vs. Collectivism, Masculinity vs. Femininity, Uncertainty Avoidance, Long-term vs. Short-term Orientation, and Indulgence vs. Restraint. These dimensions help explain how cultural values influence behavior in societies and organizations [Hofstede, 2001]. VSM is a questionnaire-based tool that assesses cultural preferences across these dimensions [Hofstede, 2011]. Kovač et al. [2023], Zhong et al. [2024] directly apply the VSM to evaluate cultural values in LLMs, while Ren et al. [2024], Kharchenko et al. [2024], Ye et al. [2025a] adapt the VSM or use self-designed inventories to measure these dimensions.
Global Leadership and Organizational Behavior Effectiveness (GLOBE).
The GLOBE study investigates cultural values, leadership behaviors, and organizational practices across countries [House et al., 2004]. Building on Hofstede’s work [Hofstede, 1984], it incorporates additional dimensions: Performance Orientation, Assertiveness, and Humane Orientation, resulting in nine cultural dimensions. Karinshak et al. [2024], Li et al. [2024d], Ren et al. [2024] apply or adapt the GLOBE culture questionnaire to evaluate cultural values in LLMs.
Social Value Orientation (SVO).
SVO is a psychological framework that measures individual preferences for allocating resources between oneself and others [Messick and McClintock, 1968]. It focuses on distinguishing between prosocial (e.g., altruists, cooperators) and proself (e.g., individualists, competitors) orientations. SVO is typically measured using experimental tasks or questionnaires, such as the SVO Slider Measure [Murphy et al., 2011] or Decomposed Games [Liebrand, 1984]. These tools assess how individuals allocate resources in hypothetical scenarios involving themselves and others. For example, Zhang et al. [2024c] apply the SVO Slider Measure to evaluate social preferences in LLMs.
Other or Custom Theories and Inventories.
Meadows et al. [2024], Xu et al. [2023] introduce localized inventories to evaluate LLMs based on Australian and Chinese values, respectively. Some researchers focus on specific topics, such as age-related values [Zhang et al., 2024b] and spiritual values [Liu et al., 2024b]. Other studies involve large-scale, cross-domain prompt datasets [Li et al., 2024d, Moore et al., 2024], among which Jiang et al. [2024a] concentrate on eliciting harmful, biased, or ethics-related content from LLMs. Notably, Biedma et al. [2024], Ye et al. [2025b] challenge established value theories and propose novel ones for LLM values, built in a bottom-up manner using psycholexical data.
Comparing Value Theories and Suggesting Use Cases.
Schwartz emphasizes universal human motivations; WVS examines societal-level value shifts; VSM and GLOBE address cross-cultural dimensions, with GLOBE emphasizing leadership. SVO, in contrast, focuses on individual-level social preferences, i.e., prosocial vs. proself. While Schwartz and SVO overlap in analyzing individual motivations, the former is broader, and the latter narrowly focuses on decision-making in resource allocation. WVS, VSM, and GLOBE share cultural-level dimensions but differ in focus—WVS on modernization, VSM on workplace-related cultural differences, and GLOBE on leadership and organizational practices. For evaluating individual values in LLMs, use Schwartz for a comprehensive understanding of universal motivations or SVO for interpersonal decision-making studies. In contrast, choose WVS for evaluating broad societal values in LLMs. When evaluating cultural dimensions in LLMs, use VSM for workplace cultural differences or GLOBE for leadership and organizational practices. In addition, Schwartz’s theory bridges individual and societal-level values, making it versatile for mixed-focus studies.
Main Findings.
Research indicates that LLMs display distinct and systematic value patterns. According to Schwartz’s Value Theory, LLMs tend to prioritize Self-Transcendence and Conservation. They exhibit stronger inclinations toward Universalism, Benevolence, Conformity, and Security, while opposing Power and Achievement [Rozen et al., 2024, Zhang et al., 2023a, Hadar-Shoval et al., 2024, Fischer et al., 2023]. WVS surveys further suggest that LLMs generally prefer Self-Expression values over Survival values [Chiu et al., 2025]. Studies using the VSM and GLOBE frameworks emphasize LLMs’ strong focus on Humane and Performance Orientation, with moderate Assertiveness [Li et al., 2024b]. Additionally, when assessed through SVO, advanced LLMs predominantly show Prosocial tendencies [Zhang et al., 2024c].
Different models exhibit varied value orientations [Kovač et al., 2024, Chiu et al., 2025, Xu et al., 2023, Li et al., 2024b, Duan et al., 2023]. Versions within the same model family show evolutionary trends in values, potentially influenced by safety alignment, capability advancement, and shifting societal expectations [Duan et al., 2023, Moore et al., 2024, Kim and Baek, 2024]. Generally, larger models align more closely with desirable human values [Shen et al., 2024, Jiang et al., 2024a, Kim and Baek, 2024, Zhong et al., 2024].
Cross-cultural research suggests that LLMs may embody a blend of cultural values, integrating perspectives from diverse backgrounds [Kovač et al., 2023]. However, they generally exhibit a tendency toward Western liberal values [Kim and Baek, 2024]. LLMs can display different values based on profiling prompts [Kharchenko et al., 2024, Zhong et al., 2024, Karinshak et al., 2024]. The context-dependency of their values is concerning and challenges the stability assumptions in human value theories [Shen et al., 2024, Meadows et al., 2024, Moore et al., 2024, Chiu et al., 2025, Xu et al., 2023, Kovač et al., 2023]. Further discussions in reliability and validity are presented in § 7.
5.1.3 Morality
"Morality is the categorization of intentions, decisions and actions into those that are proper, or right, and those that are improper, or wrong." [Long and Sedley, 1987]
Morality is a fundamental aspect of human behavior, influencing social interactions, decision-making, and ethical reasoning. The Moral Foundations Theory (MFT) [Graham et al., 2009] posits that morality is shaped by six innate psychological systems: Care/Harm, Fairness/Cheating, Loyalty/Betrayal, Authority/Subversion, Sanctity/Degradation, and Liberty/Oppression. These moral foundations are hypothesized to be universal across cultures, providing a common framework for understanding moral judgments and behavior. The theory suggests that individuals vary in the extent to which they endorse each foundation, leading to diverse moral profiles that influence their attitudes, beliefs, and actions.
It is crucial to conduct moral assessments of LLMs to ensure their ethical deployment. A large body of research (e.g., [Tlaie, 2024, Liu et al., 2024d, Fraser et al., 2022, Yao et al., 2025a, Zhou et al., 2024a]) applies MFT, primarily through the Moral Foundations Vignettes (MFVs) [Clifford et al., 2015], the Moral Foundations Questionnaire (MFQ) [Graham et al., 2009], the MFQ-2 [Atari et al., 2023], and the Moral Foundations Dictionary (MFD) [Graham et al., 2009]. These papers investigate model bias, alignment with political/moral ideologies, and variation in personal or cultural values, using controlled prompt tests [Tlaie, 2024, Nunes et al., 2024, Abdulhai et al., 2024], persona-driven exploration [Münker, 2024], or evaluating internal moral coherence [Nunes et al., 2024].
Other prominent moral theoretical frameworks and survey tools for evaluating LLMs include Kohlberg’s Theory via the Defining Issues Test (DIT) [Kohlberg, 1964], the Consequentialist-Deontological distinction [Beauchamp, 2001], and the PEW 2013 Global Attitudes Survey [Center, 2013]. Khandelwal et al. [2024], Tanmay et al. [2023] utilize the DIT, a psychometric tool aligned with Kohlberg’s stages of moral development, to analyze structural reasoning in moral dilemmas. They report that models like GPT-4 exhibit post-conventional moral reasoning. Khandelwal et al. [2024] further extend this analysis to a multilingual context, examining cross-linguistic variations in moral reasoning abilities. Neuman et al. [2025] explore ethical reasoning using established typologies, including consequentialist-deontological analysis. The results uncover the general rationalist, consequentialist emphasis in LLMs. According to the PEW 2013 Global Attitudes Survey, Meijer et al. [2024] find that the models display homogeneous moral values and have limited effectiveness in reflecting cross-cultural morality.
Localized moral theoretical frameworks are essential for ensuring the ethical deployment of LLMs in various cultural contexts. Liu et al. [2024d] develop the Chinese Moral Dictionary (CMD) and find that ChatGPT and Gemini prefer individualistic moral beliefs, while Ernie and ChatGLM lean towards collectivist moral beliefs. Takeshita et al. [2023], Ohashi et al. [2024] create the JCommonsenseMorality (JCM) dataset and fine-tune LLMs to adapt to Japanese culture.
Researchers have also developed specialized datasets for studying LLM morality. Hendrycks et al. [2021] create the ETHICS dataset, which is then utilized and extended to evaluate LLMs’ knowledge of morality, covering concepts such as justice, well-being, duties, virtues, and commonsense morality [Jinnai, 2024, Rodionov et al., 2023, Karpov et al., 2024, Yu et al., 2023, Albrecht et al., ]. Marraffini et al. [2024] develop the Greatest Good Benchmark (GGB) to evaluate LLMs’ moral judgments using utilitarian dilemmas. Jin et al. [2024b] develop the MultiTP, a cross-lingual corpus of moral dilemma vignettes in over 100 languages.
Comparing Moral Theories/Instruments and Suggesting Use Cases.
Instruments based on MFT are well-suited for measuring LLMs’ alignment with intuitive moral principles across cultural or political dimensions. Tools like MFQ, MFQ-2, MFVs, and MFD are particularly effective when investigating ideological leanings, bias, or internal consistency in abstract vs. contextual moral judgments [Nunes et al., 2024]. In contrast, Kohlberg’s DIT explores structured moral reasoning, offering a developmental perspective on how LLMs process ethical dilemmas [Khandelwal et al., 2024, Tanmay et al., 2023]. DIT is valuable for longitudinal studies and evaluating whether models show higher-level moral reasoning (e.g., post-conventional reasoning). Frameworks such as the Consequentialist-Deontological distinction and global surveys like the PEW 2013 study are useful for categorizing ethical tendencies or analyzing cultural homogeneity [Neuman et al., 2025, Meijer et al., 2024]. Localized tools (e.g., CMD, JCM) help assess cultural alignment [Liu et al., 2024d, Takeshita et al., 2023]. Meanwhile, datasets like ETHICS provide morally charged scenarios for detecting normative behavior and fine-tuning LLM responses in high-stakes ethical domains [Hendrycks et al., 2021, Jin et al., 2024b, Marraffini et al., 2024].
Main Findings.
LLMs are generally characterized by a rationalist and consequentialist focus, often prioritizing harm minimization and fairness [Neuman et al., 2025]. Despite that, they show divergence in ethical reasoning [Neuman et al., 2025] and moral preferences [Tanmay et al., 2023, Jin et al., 2024b, Bonagiri et al., 2024, Meijer et al., 2024]. In some aspects, most LLMs align with human moral standards [Tanmay et al., 2023, Nunes et al., 2024, Takemoto, 2024], which may be attributed to their extensive exposure to conventional ethical values during training. On the other hand, some research presents a less optimistic view, uncovering significant deviations of LLMs from human moral preferences [Marraffini et al., 2024, Vida et al., 2024, Ahmad and Takemoto, 2024]. Regarding the underlying mechanisms of LLMs’ moral reasoning, Nunes et al. [2024], Ji et al. [2024], Simmons [2023] suggest that LLMs primarily exhibit imitation rather than genuine conceptual understanding.
5.1.4 Attitudes & Opinions
"Attitudes are always attitudes about something. This implies three necessary elements: first, there is the object of thought, which is both constructed and evaluated. Second, there are acts of construction and evaluation. Third, there is the agent, who is doing the constructing and evaluating. We can therefore suggest that, at its most general, an attitude is the cognitive construction and affective evaluation of an attitude object by an agent." [Bergman, 1998]
We use the term "attitude" to encompass both attitudes and opinions, following [Bergman, 1998, Ma et al., 2024a], who treat these concepts as synonymous. Most research on LLM attitudes examines political attitudes and public opinions [Ma et al., 2024a], which are key cognitive and behavioral foundations in human society and closely tied to model fairness, credibility, and social impact [Durmus et al., 2023, Santurkar et al., 2023, Sanders et al., 2023, Lee et al., 2024c, Hartmann et al., 2023]. Other biases—such as those related to gender, race, culture, religion, occupation and age—are also studied as part of LLM attitudes, though many works approach them from data and algorithmic perspectives. For a broader discussion on bias and fairness in LLMs, we refer readers to [Gallegos et al., 2024, Ranjan et al., 2024].
To measure the political attitudes of LLMs, researchers often employ standardized questionnaires and scales from political science and social psychology. A widely used tool is the American National Election Studies (ANES) [^1] https://electionstudies.org/about-us/., applied in works such as Qi et al. [2024], Jiang et al. [2022, 2024d], Yang et al. [2024] to assess LLMs’ stances on U.S. political issues. Similarly, the American Trends Panel (ATP) [^2] https://www.pewresearch.org/american-trends-panel/. has been leveraged by Santurkar et al. [2023], Hwang et al. [2023], Tjuatja et al. [2024] to analyze opinion distributions on public policy. For cross-national comparisons, studies like Ma et al. [2024b], Ball et al. [2025], von der Heyde et al. [2024] utilize the German Longitudinal Election Study (GLES) [^3] https://www.gesis.org/en/gles.. Additionally, the Political Compass Test (PCT) [^4] https://www.politicalcompass.org/test. has gained prominence for positioning LLMs within multidimensional political spectrums [Azzopardi and Moshfeghi, 2024, Bernardelle et al., 2024, Hartmann et al., 2023, Röttger et al., 2024].
Other survey instruments include the General Social Survey (GSS) [Kim and Lee, 2023], American Community Survey (ACS) [Dominguez-Olmedo et al., 2024], Canadian Election Study (CES) [Sanders et al., 2023], European Social Survey (ESS) [Geng et al., 2024], Survey of Russian Elites [Kalinin, 2023], and Supreme Court Case Political Evaluation (SCOPE) [Xu et al., 2025c]. Additionally, researchers have developed specialized datasets for LLMs, such as OpinionQA [Santurkar et al., 2023], a comprehensive dataset based on ATP surveys; IssueBench [Röttger et al., 2025], a benchmark covering controversial issues with multiple response formats; and GlobalOpinionQA [Durmus et al., 2023], which extends political opinion evaluation to cross-cultural contexts.
Researchers also present tailored tools for LLMs, due to the complexity of measuring political attitudes. Haller et al. [2023] introduce OpinionGPT, a web tool demonstrating how input data biases influence model outputs. Chalkidis and Brandl [2024] fine-tune LLMs on European Parliament speeches to assess political knowledge and reasoning, realigning the models to reflect specific stances. Coppolillo et al. [2025] propose a framework for quantifying biases in multi-agent generative systems by simulating echo chambers—paired LLMs with aligned perspectives discuss polarizing topics. Finally, Kim et al. [2025b] investigate linear interpolation methods to track continuous shifts in model political attitudes.
Comparing Attitude Measurement Tools and Suggesting Use Cases.
The ANES offers the most comprehensive framework for American political contexts, making it ideal for studying model performance on U.S. social issues [Jiang et al., 2022]. Cross-national surveys and datasets such as the ESS, GLES, and GlobalOpinionQA provide broader cultural perspectives across diverse sociopolitical environments [Geng et al., 2024, Ball et al., 2025, Ma et al., 2024b, von der Heyde et al., 2024, Durmus et al., 2023]. The ATP is better suited for researching model responses to social controversies [Santurkar et al., 2023], while the PCT serves as a theory-driven political spectrum measurement tool, offering simplified yet intuitive insights into LLM political positioning [Röttger et al., 2024]. Other surveys have distinct regional and sociopolitical focuses: the CES examines political participation in North American multicultural democracies [Sanders et al., 2023]; Survey of Russian Elites provides insights into non-Western political perspectives [Kalinin, 2023]; the SCOPE delivers in-depth assessments of judicial attitudes [Xu et al., 2025c]; and the ACS serves as a tool for researching socioeconomic issues and group representation [Dominguez-Olmedo et al., 2024].
Main Findings.
Pretraining data inherently contains socially biased opinions and perspectives, which can amplify political polarization in LLMs [Feng et al., 2023, Xu et al., 2025c]. Most studies identify a misalignment between LLM outputs and human opinions [Santurkar et al., 2023, von der Heyde et al., 2024, Yang et al., 2024, Dormuth et al., 2025], with many concluding that LLMs exhibit a left-leaning political bias [Rozado, 2023, Hartmann et al., 2023, Ceron et al., 2024, Bernardelle et al., 2024, Ma et al., 2024b]. Cross-cultural comparative studies further reveal Western-centric tendencies, demonstrating limited understanding of non-English political perspectives or multi-partisan systems [Qi et al., 2024]. These findings suggest structural limitations in the models’ political understanding. It is also shown that the degree and manifestation of bias vary significantly across contexts and domains. For example, political-electoral propositions exhibit different bias patterns than socioeconomic issues like climate change [Wu et al., 2023]. In contrast, some researchers offer more optimistic perspectives, emphasizing LLMs’ potential to simulate population-level opinions and supplement traditional survey methods [Dominguez-Olmedo et al., 2024, Sanders et al., 2023, Kalinin, 2023, Wu et al., 2023, Jiang et al., 2024d, Argyle et al., 2023, Bisbee et al., 2024, Sun et al., 2024]. They suggest that with appropriate prompt design, calibration methods, and fine-tuning, LLMs can generate opinion distributions closely approximating human group distributions [Jiang et al., 2022, Wu et al., 2023]. Such advancements indicate promising directions for using LLMs in opinion research, though their current limitations and biases must still be acknowledged.
Aspects Sub-aspects Related work
Heuristics & biases - Yax et al. [2024], Sundaram and Alwar [2024], Momennejad et al. [2023], Echterhoff et al. [2024], Saeedi et al. [2024], Ranaldi and Zanzotto [2024], Macmillan-Scott and Musolesi [2024], Ando et al. [2023], Malberg et al. [2024], Xie et al. [2024b], Hayes et al. [2024], Shah et al. [2024], Chen and Eger [2025], Tang and Kejriwal [2024], Talboy and Fuller [2023], Coda-Forno et al. [2024], Abramski et al. [2023], Binz and Schulz [2023], Xie et al. [2024a], Kumar et al. [2024], Schulze Buschoff et al. [2025], Scholten et al. [2024], Shaikh et al. [2024], Thorstad [2023], Hagendorff et al. [2023], Healey et al. [2024], Bai et al. [2024], Castello et al. [2024]
Social interactions Theory of Mind Sarıtaş et al. [2025], Ma et al. [2023], Mao et al. [2024b] (reviews); Ullman [2023], Sap et al. [2022], Kosinski [2023a], Shapira et al. [2023a], Gandhi et al. [2023], Strachan et al. [2024a], Kosinski [2023b], Moghaddam and Honey [2023], Street et al. [2024], Sclar et al. [2023], van Duijn et al. [2023], He et al. [2023], Holterman and Deemter [2023], Chen et al. [2024d], Zhu et al. [2024c], Hou et al. [2024], Amirizaniani et al. [2024], Soubki et al. [2024], Jones et al. [2024], Lin et al. [2024], Sclar et al. [2024], Sadhu et al. [2024], Lorè et al. [2024], Sarangi et al. [2025], Yu et al. [2025], Tang and Belle [2024], Kim et al. [2025a], Chan et al. [2024], Nickel et al. [2024], Xu et al. [2024a], Leer et al. [2023], Jamali et al. [2023], Li et al. [2023b], Wilf et al. [2023], Chen et al. [2024c], Strachan et al. [2024b], Xu et al. [2025a], Zhang et al. [2025], Shinoda et al. [2025], Jin et al. [2024a], Yang et al. [2025a], Zhou et al. [2023a], Pi et al. [2024], Riemer et al. [2025], Tan et al. [2024].
Emotional Intelligence Sorin et al. [2024], Raj [2024] (reviews); Sabour et al. [2024], Paech [2023], Wang et al. [2023b], Elyoseph et al. [2023], Patel and Fan [2023], Vzorinab et al. [2024], Hu et al. [2025], Li et al. [2024e], Zhao et al. [2024], Chen et al. [2024b], Huang et al. [2024b], Schaaff et al. [2023], Li et al. [2023a], Huang et al. [2023b], Lee et al. [2024f], Welivita and Pu [2024]
Social Intelligence Leng and Yuan [2023], Mittelstädt et al. [2024], Shapira et al. [2023b], Zhou et al. [2024b], Wang et al. [2024d], Liu et al. [2024c], Wang et al. [2024a], Mou et al. [2024], Liang et al. [2024], Zadeh et al. [2019], Xu et al. [2024b], Guo et al. [2023a], Kovač et al. [2024, 2021], Mathur et al. [2024]
Psychology of language Language comprehension Chang and Bergen [2024] (reviews); Duan et al. [2024a], Wang et al. [2024b], Wilcox et al. [2021], Arehalli et al. [2022], Li et al. [2024a], Steuer et al. [2023], Duan et al. [2024b], Qiu et al. [2024], Bojić et al. [2023], Hu et al. [2023], Ruis et al. [2023], Lee et al. [2024b], Seals and Shalin [2023], Hu and Levy [2023], He et al. [2024], Sun and Wang [2024], Hong et al. [2023], Amouyal et al. [2024], Ide et al. [2024], Lee et al. [2024b], Huff and Ulakçi [2024], Zhou et al. [2023b], Tian et al. [2024]
Language generation Seals and Shalin [2023], Tian et al. [2024], Tang and Kejriwal [2024], Stevenson et al. [2022], Hubert et al. [2023], Chakrabarty et al. [2024], Orwig et al. [2024], Lee and Chung [2024], Boussioux et al. [2024], Miaschi et al. [2024], Bellemare-Pepin et al. [2024], Cai et al. [2024]
Language acquisition Steuer et al. [2023], Shah et al. [2024], Frank [2023b], Riva et al. [2024]
Learning and cognitive capabilities - Niu et al. [2024] (reviews); Galatzer-Levy et al. [2024], Wang et al. [2024e], Dayan et al. [2024], Wu et al. [2025], Sartori and Orrú [2023], Webb et al. [2022], Song et al. [2024a], Zhang et al. [2023b], Wang et al. [2024e], Coda-Forno et al. [2024], Zeng et al. [2024c], Lv et al. [2024]
Table 4: Representative LLM cognitive constructs in LLM psychometrics.
5.2 Measuring Cognitive Constructs
Traditional NLP benchmarks are insufficient to capture and analyze emergent cognitive constructs [Hagendorff et al., 2024, Ying et al., 2025a]. To address this gap, researchers are adapting related psychometric techniques to evaluate LLM abilities. Hagendorff et al. [2024] introduce the concept of machine psychology, reviewing emergent LLM capabilities and advocating for psychometric evaluation. They categorize these cognitive constructs into four key aspects: heuristics and biases, social interactions, psychology of language, and learning and cognitive capabilities. Structured based on this taxonomy, Table 4 summarizes the representative cognitive constructs measured in recent research.
5.2.1 Heuristics and Biases
"Heuristics and biases are mental shortcuts or rules of thumb that simplify decision-making and problem-solving." [Tversky and Kahneman, 1974]
Heuristics shape both the strengths and biases in human decision-making. Recent research employs psychometric tools to systematically evaluate rationality and biases in LLM outputs and offer theoretical explanations for these biases.
Pioneering work by Binz and Schulz [2023] applies a battery of canonical cognitive ability tests to GPT-3. The test results demonstrate that the model performs on par with or better than human participants, while also exhibiting human-like cognitive biases such as framing effects, certainty effects, and overweighting biases. Subsequent studies include more advanced LLMs and reveal a dual-process dynamic [Hagendorff et al., 2023]: earlier GPT models exhibit more human-like System 1 errors, while more advanced models such as GPT-4 display more hyperrational System 2 reasoning and thus fewer biases. In a similar vein, but with an expanded scope, Yax et al. [2024], Sundaram and Alwar [2024], Momennejad et al. [2023], Echterhoff et al. [2024], Saeedi et al. [2024], Ranaldi and Zanzotto [2024], Macmillan-Scott and Musolesi [2024], Ando et al. [2023] evaluate LLMs regarding biases including the conjunction fallacy, unwarranted beliefs, anchoring bias, status-quo bias, loss aversion, positional selection biases, and Clever Hans effects. Later work significantly scales up the evaluation with synthetic items. Malberg et al. [2024] evaluate cognitive biases in 20 LLMs based on novel decision-making scenarios and a dataset of 30,000 tests covering 30 cognitive biases. Xie et al. [2024b] introduce the "MindScope" dataset and multi-agent methods to evaluate 72 bias categories using psychometric-inspired multi-round dialogue experiments. Recently, the evaluation paradigm has also been extended to multimodal models [Schulze Buschoff et al., 2025].
To understand the origins of these biases, researchers have proposed diverse theoretical frameworks. Hagendorff et al. [2023], Yax et al. [2024] link dual-process theory (System 1 vs. System 2 thinking) with the contextualized inference of LLMs, contrasting intuitive and deliberative reasoning modes. Scholten et al. [2024] explain LLM biases using metacognitive myopia, a syndrome described as being "relatively accurate in utilizing even large amounts of stimulus information, but naïve and almost blind regarding the history and validity of the stimulus data." The authors argue that LLMs lack metacognitive abilities such as monitoring and control, which result in five symptoms of metacognitive myopia and, therefore, systematic biases. Sundaram and Alwar [2024] explore cognitive dissonance theory and elaboration likelihood theory to explain inconsistencies in LLM reasoning. They suggest that certain biases in LLMs may not just be drawbacks but could serve as a means of identifying and mitigating logical fallacies in human reasoning. Differing from the cognitive psychology perspective, Shaikh et al. [2024] propose a mechanistic interpretability framework for sourcing the biases. They leverage influence graphs and Shapley value analysis to interpret biases and explore connections between model training and bias emergence.
Main Findings.
By administering psychometrics to LLMs, studies consistently find that LLMs exhibit cognitive biases superficially similar to humans, such as anchoring, framing, and the conjunction fallacy. Large-scale tests, such as [Echterhoff et al., 2024, Malberg et al., 2024, Xie et al., 2024b], systematically categorize these biases and enable scalable evaluations. Some note that newer, larger, chain-of-thought-enabled models exhibit improved reasoning and bias mitigation [Tang and Kejriwal, 2024, Hagendorff et al., 2023], while others argue that increasing model complexity without deliberate bias mitigation strategies can amplify existing biases [Kumar et al., 2024]. In addition, researchers explore mechanistic interpretability [Shaikh et al., 2024] and present theoretical explanations from cognitive psychology and reasoning theories [Hagendorff et al., 2023, Yax et al., 2024, Scholten et al., 2024]. While LLMs exhibit dual-process reasoning dynamics reminiscent of human cognition [Hagendorff et al., 2023], detailed analysis reveals differences between LLM and human reasoning [Yax et al., 2024].
5.2.2 Social Interactions
Researchers apply psychometric tools from social and developmental psychology to assess LLMs’ capabilities in navigating social dynamics. Related evaluation focuses on interconnected dimensions such as Theory of Mind (ToM), Emotional Intelligence (EI), and Social Intelligence (SI).
"Theory of Mind is the ability to attribute mental states such as beliefs, intentions, and knowledge to others." [Premack and Woodruff, 1978]
Theory of Mind (ToM).
Advanced LLMs can simulate ToM-like reasoning under certain conditions, which prompts questions about how these behaviors arise and how robustly they generalize. Foundational studies apply classic psychometrics—false belief tasks—to evaluate ToM in LLMs. [Kosinski, 2023a, b] report that GPT-3.5 and GPT-4 perform at levels matching or exceeding those of children in structured ToM tasks, suggesting the spontaneous emergence of ToM-like behavior as models scale in size. Through mechanistic interpretation, Jamali et al. [2023] offer evidence of parallels between the LLM embeddings and neurons in the human brain.
However, these claims have been challenged. Ullman [2023], Shapira et al. [2023a], Holterman and Deemter [2023] soon demonstrate that minor task perturbations often lead to poor performance, indicating reliance on brittle heuristics rather than genuine semantic comprehension. Hoping to address the controversy, researchers have developed robust benchmarks and evaluation protocols. For example, BigToM [Gandhi et al., 2023], ToMBench [Chen et al., 2024d], HI-TOM [He et al., 2023], OpenToM [Xu et al., 2024a], and ToMATO [Shinoda et al., 2025] feature procedurally generated, higher-order, or broader ToM tasks. These benchmarks often reveal performance drop-offs in more complex tasks, such as 6th-order belief attribution [Street et al., 2024] or detecting faux pas [Strachan et al., 2024a]. But others argue the failures result from a lack of more general commonsense reasoning, rather than a failure to represent mental states [Pi et al., 2024]. Riemer et al. [2025] also argue that we should distinguish between literal and functional ToM, and that current ToM benchmarks are inadequate for assessing functional ToM. Recent attempts also extend ToM evaluation to multimodal settings [Chen et al., 2024c, Strachan et al., 2024b, Jin et al., 2024a] and multi-agent interactions [Li et al., 2023b].
Several studies explore enhancement techniques. Prompt engineering methods like reasoning, reflection, and planning [Moghaddam and Honey, 2023, Lin et al., 2024, Zhou et al., 2023a, Yang et al., 2025a, Zhang et al., 2025]; persona-based prompts [Tan et al., 2024]; and psychology-informed prompts [Leer et al., 2023, Wilf et al., 2023] can significantly increase accuracy but raise questions about the authenticity of improvements. Others implement symbolic reasoning frameworks, such as SymbolicToM [Sclar et al., 2023, Xu et al., 2025a], ToM-LM [Tang and Belle, 2024], or task decomposition strategies [Sarangi et al., 2025], to enhance interpretability and belief state modeling. While recent models keep progressing in ToM tasks, especially in structured tasks [Strachan et al., 2024a, Street et al., 2024], they are shown to remain inconsistent in open-ended, adversarial, or pragmatic reasoning settings [Sclar et al., 2024, Yu et al., 2025, Nickel et al., 2024]. Comparisons with human baselines show that LLMs can approximate ToM behavior in narrow contexts but fall short of general human-like social cognition [Strachan et al., 2024a, van Duijn et al., 2023, Jones et al., 2024].
Main Findings.
LLMs exhibit psychometrically measurable ToM-like reasoning, especially when appropriately prompted or structured, but current evidence suggests these capabilities depend on surface-level linguistic cues and lack robustness. The proficiency of LLMs in ToM remains a contentious issue. Interested readers are referred to recent comprehensive reviews on ToM in LLMs [Sarıtaş et al., 2025, Ma et al., 2023, Mao et al., 2024b].
"Emotional Intelligence is the subset of social intelligence that involves the ability to monitor one’s own and others’ feelings and emotions, to discriminate among them and to use this information to guide one’s thinking and actions." [Salovey and Mayer, 1990]
Emotional Intelligence (EI).
Recent studies introduce benchmark tools specifically designed to measure EI in LLMs through structured tasks informed by psychological theories. For example, EmoBench [Sabour et al., 2024] and EQ-Bench [Paech, 2023] provide task sets operationalizing constructs like emotional understanding and regulation, grounding their evaluations in frameworks such as the Mayer-Salovey-Caruso Emotional Intelligence Test (MSCEIT). Others adapt existing psychometric instruments to LLM psychometrics, including the Situational Evaluation of Complex Emotional Understanding (SECEU) [Wang et al., 2023b], the Levels of Emotional Awareness Scale (LEAS) [Elyoseph et al., 2023], the Toronto Alexithymia Scale (TAS-20) and Empathy Quotient (EQ-60) [Patel and Fan, 2023], and others [Huang et al., 2024b]. Models such as GPT-4 are shown to match or exceed human baseline in measures of emotional awareness and understanding [Wang et al., 2023b, Elyoseph et al., 2023, Patel and Fan, 2023], though it has been reported that LLMs lack deep reflexive analysis of emotional experiences and motivational aspects of emotions [Vzorinab et al., 2024]. Recent benchmark tools also extend EI evaluations to multimodal settings [Hu et al., 2025].
A subset of the literature further explores methodological pathways for improving LLM emotional competencies. Modular Emotional Intelligence (MoEI) [Zhao et al., 2024] and Emotional Chain-of-Thought (ECoT) [Li et al., 2024e] offer strategies to enhance emotionally intelligent behaviors while preserving general cognitive functionality. Another line of research explores LLMs’ ability to simulate nuanced emotional alignment, empathy, and scenario-appropriate reactions, based on Emotion Appraisal Theory [Roseman and Smith, 2001] and Emotion-Focused Theory [Greenberg, 2004]. Surprisingly, Li et al. [2023a] indicate that emotional stimuli improve general performance of LLMs.
Main Findings.
Advanced LLMs generally perform on par with or better than humans in tasks of EI [Elyoseph et al., 2023, Patel and Fan, 2023], despite evident limitations in several areas, such as displaying artificial or mechanical patterns when expressing empathy [Lee et al., 2024f], the deep reflexive analysis of emotional experiences [Vzorinab et al., 2024], and the misalignment with human emotional behaviors [Huang et al., 2023b].
"Social Intelligence is the ability to understand and manage people." [Thorndike and Stein, 1937]
Social Intelligence (SI).
Measuring SI of LLMs seeks to understand how these systems interpret and respond to human social situations. The evaluation involves rule-based decision-making, conformity to social norms, and successful participation in structured interpersonal scenarios.
Several studies have directly applied psychometric instruments to address this objective and present optimistic results. Mittelstädt et al. [2024] utilize standardized Situational Judgment Tests (SJTs) to compare LLMs’ assessments of social appropriateness with those of human participants, finding that some models outperform humans under expert-rated conditions. Other research reveals LLMs still struggle in some aspects of SI. Shapira et al. [2023b] investigate the ability of LLMs using faux pas tests from clinical psychology, finding that LLMs struggle to describe social situations implicitly. Based on Daniel Goleman’s SI theory, Xu et al. [2024b] introduce the Situational Evaluation of SI (SESI). It shows that LLMs are still limited in their SI, with superficial friendliness as a primary reason for errors. AgentSense benchmark employs interactive scenarios and highlights the limitations of LLMs in handling complex social interactions, especially high-level growth needs and private information reasoning [Mou et al., 2024].
Later studies further explore multi-agent environments. Interactive multi-agent environments such as SOTOPIA [Zhou et al., 2024b] and its extended training variant SOTOPIA-$\pi$ [Wang et al., 2024d] simulate cooperative, competitive, and norm-driven social contexts. The STSS benchmark proposed by [Wang et al., 2024a] further operationalizes SI through task-oriented simulations, focusing on agent action outcomes and goal achievements rather than linguistic performance. The CogMir framework [Liu et al., 2024c] focuses on evaluating prosocial but irrational LLM behaviors. It reveals that LLM agents and humans exhibit high consistency in irrational and prosocial decision-making under uncertain conditions. [Guo et al., 2023a] refine the Social-IQ dataset to create DeSIQ, a benchmark for SI evaluation that extends to multimodal settings. Critical perspectives are proposed by Kovač et al. [2024] who argue the current benchmarks lack grounding in developmental psychology and offer a versatile framework, the SocialAI school, for studying the SI of LLM-based agents.
Main Findings.
LLMs show measurable competence in rule-based, socially appropriate behavior, particularly in structured environments. They perform well on: 1) following predefined social norms [Mittelstädt et al., 2024]; 2) completing interactional goals in multi-agent settings [Zhou et al., 2024b, Wang et al., 2024d]; and 3) replicating human-like prosocial decisions [Liu et al., 2024c]. Weaker performance is noted in tasks where superficial friendliness causes errors [Xu et al., 2024b], understanding high-level growth needs is essential [Mou et al., 2024], and LLMs must implicitly describe social situations [Shapira et al., 2023b].
5.2.3 Psychology of Language
Psycholinguistics, a subfield of psychology, explores how humans comprehend, generate, and acquire language [Carroll, 1986]. Insights from this discipline help evaluate how LLMs process language and mirror human linguistic features.
Language Comprehension.
Evaluations of LLM language comprehension aim to understand how LLMs process language and mirror human linguistic features. Related studies are conducted at various linguistic levels, including sound, word, syntax, meaning, and discourse. Duan et al. [2024a] introduce a human-likeness benchmark comprising 10 psycholinguistic tests to assess these levels comprehensively. More studies focus on specific aspects. For instance, Wang et al. [2024b], Arehalli et al. [2022], Wolfman et al. [2024], Wilcox et al. [2021] examine syntactic and semantic processing. Surprisal, a measure of word predictability, is used across multiple studies to model the difficulty of processing syntactic ambiguities, garden-path sentences, and hierarchical structures [Wilcox et al., 2021, Arehalli et al., 2022, Li et al., 2024a]. Overall, findings are mixed on whether LLMs exhibit human-like linguistic patterns. Duan et al. [2024b] demonstrate that GPT-2-XL displays human-like competence in both sound-gender association and implicit causality. However, it lacks human-like abilities in sound-shape association.
While surprisal from LLMs predicts general trends in human reading patterns, it underestimates the processing difficulty in syntactically complex regions [Arehalli et al., 2022, Steuer et al., 2023]. In contrast, Wang et al. [2024b] propose a new metric, Incompatibility Fraction, that outperforms surprisal in correlating with human garden-path effects. Other tasks address grammaticality judgment [Qiu et al., 2024, Ide et al., 2024], pragmatic inference [Bojić et al., 2023, Hu et al., 2023, Ruis et al., 2023], argument role processing [Lee et al., 2024b], and discourse-level comprehension [Duan et al., 2024a]. Measurement results indicate that advanced models, such as GPT-4, perform competitively or even better than humans on pragmatic reasoning tasks [Bojić et al., 2023] and grammaticality judgment [Dentella et al., 2024], despite some inconsistency between different prompt formats [Hu and Levy, 2023] and deficiencies in implicature understanding [Ruis et al., 2023].
Language Generation.
Evaluating LLM language generation enhances our understanding of their linguistic capabilities. One rich line of research focuses on measuring the creativity of LLMs [Tang and Kejriwal, 2024, Stevenson et al., 2022, Hubert et al., 2023, Chakrabarty et al., 2024, Orwig et al., 2024, Lee and Chung, 2024, Boussioux et al., 2024, Bellemare-Pepin et al., 2024]. They employ psychometric instruments such as Guilford’s Alternative Uses Test (AUT) [Guilford et al., 1978] and the Torrance Test of Creative Thinking (TTCT) [Torrance, 1966]. Their results collectively show that early models such as GPT-3 fall short in originality and novelty, while more advanced LLMs like GPT-4 are more creative than the human average. More nuanced evaluations provide deeper insights. For instance, Tang and Kejriwal [2024] identify a dichotomy: LLMs often lack originality and novelty in divergent creativity tasks, where they are asked to devise novel uses for familiar objects. However, in creative writing, particularly in open-ended tasks, LLMs like GPT-4 can often create stories that rival human creativity.
Cai et al. [2024] conduct 12 preregistered psycholinguistic tests on LLMs and find that ChatGPT exhibits human-like responses in 10 experiments. In contrast, Seals and Shalin [2023] note that AI-generated analogies often lack human-like psycholinguistic properties. Similarly, Tian et al. [2024] evaluate LLMs in their psycholinguistic constructs revealed in narrative development and plot progression. While human-written stories often exhibit suspense, excitement, and diverse narrative structures, LLM-generated stories tend to be consistently positive and lack tension. Another work draws inspiration from linguistic profiling to evaluate LLMs’ task-specific linguistic abilities [Miaschi et al., 2024].
Language Acquisition.
Work on developmental plausibility assesses whether LLMs replicate stages of language learning analogous to those in children [Steuer et al., 2023, Shah et al., 2024]. It is shown that regardless of model size, the developmental trajectories of pretrained language models consistently exhibit a window of maximal alignment with human cognitive development [Shah et al., 2024]. Another study by Frank [2023b] draws inspiration from human language development to explain the data inefficiency of LLMs; the relative efficiency of human language acquisition is possibly due to pre-existing conceptual knowledge, multimodal grounding, and the interactive, social nature of their input.
Main Findings.
Early foundations that evaluate language models based on BERT and LSTM link computational linguistics and psycholinguistic mechanisms [Ettinger, 2020, Futrell et al., 2019, Arehalli and Linzen, 2020]; these models generally fall short in diverse psycholinguistic tasks. Recent studies extend evaluations to LLMs and broader tasks [Duan et al., 2024a], where advanced LLMs are shown to surpass humans on tasks such as pragmatic reasoning [Bojić et al., 2023] and creative writing [Tang and Kejriwal, 2024]. Some deficiencies still exist, such as limited implicature understanding [Ruis et al., 2023], prompt-sensitive linguistic competence [Hu and Levy, 2023], and lack of human-like psycholinguistic properties [Seals and Shalin, 2023, Tian et al., 2024]. In addition, mixed results are found in the alignment between LLMs and human linguistic cognition [Duan et al., 2024b, Wolfman et al., 2024] and language acquisition [Shah et al., 2024, Frank, 2023b]. Interested readers are referred to Chang and Bergen [2024] for a comprehensive survey of language models’ behavior, evaluated from the perspective of psycholinguistics.
5.2.4 Learning and Cognitive Capabilities
Psychometrics in learning and cognitive capabilities measure human mental functions like memory, reasoning, problem-solving, and comprehension to understand cognitive strengths and weaknesses. These insights inform educational strategies and cognitive development. Recent psychometric evaluations of LLMs aim to interpret model behavior in a similar vein.
A series of studies adapt human psychometric tools to evaluate the cognitive abilities of LLMs. Some results are promising. Galatzer-Levy et al. [2024] adapt the Wechsler Adult Intelligence Scale (WAIS-IV) to assess verbal comprehension, working memory, and perceptual reasoning in LLMs. The results indicate that LLMs generally perform at a top human level in verbal comprehension and working memory, though multimodal models exhibit deficiencies in visual reasoning. Sartori and Orrú [2023], Webb et al. [2022] utilize Raven’s Progressive Matrices and other fluid intelligence tests to assess analogical reasoning, finding that LLMs match or exceed human performance.
Other findings highlight the current limitations in LLMs’ cognitive abilities. Dayan et al. [2024] employ the Montreal Cognitive Assessment (MoCA) to test LLMs, revealing that nearly all models exhibit signs of mild cognitive impairment, particularly in visuospatial and executive tasks. Wu et al. [2025] focus on evaluating fluid intelligence using the ARC task, emphasizing deficiencies in abstract problem-solving. Critical perspectives also challenge the validity of current tests [Löhn et al., 2024, Li et al., 2024d, Zhang et al., 2023b], driving the development of new benchmarks and testing methodologies [Coda-Forno et al., 2024, Wang et al., 2024e, Zhuang et al., 2023a, Zeng et al., 2024c, Song et al., 2024a].
From the perspective of learning and cognitive development, Wang et al. [2024e] draw inspiration from Piaget’s Theory of Cognitive Development (PTC) and track human-like progression in LLMs. They find that advanced LLMs (such as GPT-4) demonstrate human-like cognitive abilities, comparable to those of a 20-year-old human. Shah et al. [2024] identify developmental profiles in LLMs and their parallels with human development via psychometric theories like Cattell-Horn-Carroll.
Main Findings.
LLMs demonstrate strong performance on verbal comprehension, working memory, and analogical reasoning, often reaching or surpassing high human percentiles when evaluated with adapted psychometric tools like WAIS-IV and Raven’s Matrices [Galatzer-Levy et al., 2024, Webb et al., 2022]. However, they exhibit notable cognitive deficits, particularly on benchmarks like MoCA and ARC tasks [Wu et al., 2025, Dayan et al., 2024]. Several works introduce developmental and cognitive benchmarks (e.g., CogLM, CogBench) grounded in psychological theories to track reasoning and adaptability [Wang et al., 2024e, Coda-Forno et al., 2024]. While emergent reasoning appears at scale, challenges remain in interpretability, test validity, and generalization beyond surface-level statistical patterns [Li et al., 2024d, Löhn et al., 2024].
6 Psychometric Evaluation Methodology
This section examines the methodologies employed in LLM psychometrics. As illustrated in Fig. 3, the methodological framework encompasses four key components: test formats (§ 6.1), data and task sources (§ 6.2), prompting strategies (§ 6.3), and model output and scoring (§ 6.4). The configuration of inference parameters (§ 6.5) is additionally discussed.
Figure 3: Overview of LLM psychometric evaluation methodologies.
6.1 Test Format
Test formats of LLM psychometric evaluation can be categorized into structured tests, open-ended conversations, and agentic simulations. Table 5 provides an overview of the test formats for each construct.
Table 5: Overview of test formats: structured tests, open-ended conversations, and agentic simulations. For each test format, we list some examples for each construct.
Construct Structured tests Open-ended conversations Agentic simulations
Personality traits Jiang et al. [2023], Serapio-García et al. [2023] Jiang et al. [2024b], Zheng et al. [2025] Frisch and Giulianelli [2024], Huang et al. [2024c]
Values Kovač et al. [2023], Zhong et al. [2024] Ren et al. [2024], Yao et al. [2025b] Chiu et al. [2025], Shen et al. [2024]
Morality Ji et al. [2024], Abdulhai et al. [2024] Sachdeva and van Nuenen [2025], Neuman et al. [2025] Chiu et al. [2025], Nunes et al. [2024]
Attitudes & opinions Bernardelle et al. [2024], Argyle et al. [2023] Röttger et al. [2024], Wright et al. [2024] -
Heuristics and biases Hagendorff et al. [2023], Yax et al. [2024] Healey et al. [2024], Wen et al. [2024a] Bai et al. [2024], Xie et al. [2024b]
Social interaction Sabour et al. [2024], Mittelstädt et al. [2024] Welivita and Pu [2024], Elyoseph et al. [2023] Zhou et al. [2024b], Wang et al. [2024d]
Psychology of language Duan et al. [2024a], Amouyal et al. [2024] - -
Learning and cognitive capabilities Wu et al. [2025], Coda-Forno et al. [2024] Galatzer-Levy et al. [2024] Lv et al. [2024]
6.1.1 Structured Tests
Structured tests feature predefined instructions, questions, and response formats. These items may include alternative-choice questions, multiple-choice questions, rating scales, and short-answer questions.
Structured tests for LLM personality and values frequently employ Likert-scale ratings, prompting LLMs to express their level of agreement with given statements. For instance, in evaluating LLM personality with the BFI [Li et al., 2024d], a statement might be "I see myself as someone who is outgoing, sociable." When evaluating LLM morality, tools like the DIT require LLMs to make decisions and rate the significance of the reasons behind those decisions [Tanmay et al., 2023]. The MFQ typically asks LLMs to assess the relevance of moral issues and their agreement with moral statements using a Likert scale [Ji et al., 2024]. Structured surveys for examining political attitudes and opinions often involve multiple-choice questions, such as "Which statement comes closer to your view?" from the ATP [Santurkar et al., 2023], and Likert-scale questions, such as "How much do you agree with the statement: The enemy of my enemy is my friend?" from the PCT [Rutinowski et al., 2024].
To test heuristics and biases in LLMs, Echterhoff et al. [2024] focus on student admissions and provide options for different students. Hagendorff et al. [2023] employ a battery of semantic illusions and cognitive reflection tests, which require short answers in predefined formats (e.g., number of days). Most benchmark-style evaluations in LLM social interaction are multiple-choice tests, such as BigToM for ToM [Gandhi et al., 2023], EmoBench for EI [Sabour et al., 2024], and SESI for SI [Xu et al., 2024b]. Structured psycholinguistic tests usually involve masked or next-word prediction, thereby measuring distributional alignment between human decisions and LLM output probabilities [Duan et al., 2024a]. Other higher-level psycholinguistic tests administer forced-choice questions for evaluating semantic comprehension [Li et al., 2024a] or require Likert-scale plausibility and grammaticality ratings [Qiu et al., 2024, Amouyal et al., 2024]. Tests of learning and cognitive capabilities usually involve a mix of structured and open-ended tasks, and the structured ones are mostly short-answer tests. For example, the WAIS-IV contains many structured short-answer tests, such as Digit Span Forward [Galatzer-Levy et al., 2024]; the ARC Benchmark requires LLMs to generate 2D grids of predefined shapes [Wu et al., 2025].
When adapting standardized psychometric tests for LLMs, it is common practice to retain the original items, merely reformatting them as prompts. Structured psychometric tests are advantageous due to their scalability, objectivity, and automated scoring. However, critical perspectives underline their limitations, such as gaps in real-world applicability; biases; data contamination; and issues with reliability, validity, and depth of insights.
6.1.2 Unstructured Tests
Unstructured tests probe LLMs’ personalities and abilities by analyzing their free-form responses, usually with rationale and justification, to user queries, or by contextualizing LLMs to observe their decision-making in real-world scenarios.
Open-Ended Conversations.
A straightforward method for unstructured testing involves engaging LLMs, as chatbots, in open-ended and mostly single-turn conversations that manifest specific constructs. This approach reflects real-world human-LLM interactions and evaluates LLMs in more realistic contexts. For instance, Jiang et al. [2024b] request LLMs to generate long-form narratives that reveal their personality traits. ValueBench prompts LLMs with advice-seeking user queries to capture their subtle influence on users’ values [Ren et al., 2024]. Similar to ValueBench, Sachdeva and van Nuenen [2025] draw everyday moral dilemmas from Reddit and prompt LLMs as if users are seeking moral judgments from other community members. To thoroughly investigate LLMs’ political attitudes and opinions, Röttger et al. [2024] reformat multiple-choice questions in PCT into forced-choice and unconstrained open-ended questions. Regarding the evaluation of cognitive constructs, Healey et al. [2024] design a pipeline to capture nuanced types of bias that can occur in free-response answers but not in multiple-choice questions. Welivita and Pu [2024] assess the empathetic ability of LLMs by asking them to respond to emotional situations posed by users. Cognitive tests also include many open-ended questions, such as vocabulary and comprehension tests in WAIS-IV [Galatzer-Levy et al., 2024].
Agentic Simulations.
Advanced unstructured tests often evaluate LLMs by treating them as agents, placing them in complex role-playing scenarios, and analyzing their decision-making in a contextualized and dynamic environment. Huang et al. [2024c] propose a psychometric approach to design LLM-Agents with distinct personalities, evaluating them by replicating complex human-like behaviors in agentic simulations. Shen et al. [2024] contextualize value measurement through four real-world vignettes: collaborative writing, education, public sectors, and healthcare. Chiu et al. [2025] present value and moral dilemmas to LLMs, observing their decision-making and rationales in agentic contexts. Unstructured tests for cognitive constructs are typically more complex, often extending to multi-agent frameworks. For instance, Xie et al. [2024b], Bai et al. [2024] identify cognitive biases in multi-agent communication. SOTOPIA is an open-ended environment simulating complex social interactions among LLM agents, where their social intelligence is evaluated [Zhou et al., 2024b, Wang et al., 2024d]. Lv et al. [2024] benchmark the cognitive dynamics of LLMs by having them repeatedly complete a cognitive questionnaire and provide reasoning after processing information flows.
Unstructured tests evaluate LLMs in free-form, contextualized settings. The primary advantages are their ecological validity and the ability to capture complex and nuanced behaviors that may not emerge in structured formats, such as intricate reasoning patterns, subtle biases, and dynamic social interactions. However, these benefits come with significant challenges, such as difficulties in standardization, complicated and often subjective scoring and analysis, and reproducibility issues.
6.2 Data and Task Sources
Table 6: Overview of data and task sources: established inventories, custom-curated items, and synthetic items. For each data source, we list some examples for each construct.
Construct Established inventories Custom-curated items Synthetic items
Personality traits Serapio-García et al. [2023], Barua et al. [2024] Ai et al. [2024], Jiang et al. [2023] Zeng [2024]
Values Fischer et al. [2023], Miotto et al. [2022] Shen et al. [2024], Meadows et al. [2024] Ye et al. [2025a], Moore et al. [2024]
Morality Abdulhai et al. [2024], Khandelwal et al. [2024] Hendrycks et al. [2021], Jin et al. [2022] Scherrer et al. [2023], Liu et al. [2024d]
Attitudes & opinions Argyle et al. [2023], Bernardelle et al. [2024] Durmus et al. [2023], Ceron et al. [2024] Wan and Chang [2024]
Heuristics and biases Hagendorff et al. [2023], Binz and Schulz [2023] Momennejad et al. [2023], Ando et al. [2023] Malberg et al. [2024], Xie et al. [2024b]
Social interaction [Ullman, 2023, Chen et al., 2024d] Shapira et al. [2023a], Street et al. [2024] Gandhi et al. [2023]
Psychology of language Cai et al. [2024], Duan et al. [2024a] He et al. [2024], Pérez-Mayos et al. [2021] -
Learning and cognitive capabilities Coda-Forno et al. [2024], Dayan et al. [2024] Song et al. [2024a], Wu et al. [2025] -
Data and task sources LLM psychometrics can be 1) drawn from established psychometric inventories, 2) human-authored and custom-curated, and 3) synthesized by AI models. Table 6 provides an overview of the data and task sources for each construct.
Established Inventories.
Established psychometric inventories present well-validated and widely recognized tools. Representative inventories in LLM psychometrics include BFI and HEXACO for personality traits [Serapio-García et al., 2023, Barua et al., 2024]; SVS, PVQ, WVS, and VSM for values [Fischer et al., 2023, Miotto et al., 2022]; MFT and DIT for morality [Abdulhai et al., 2024, Khandelwal et al., 2024]; ANES and PCT for political attitudes and opinions [Argyle et al., 2023, Bernardelle et al., 2024]; cognitive reflection tests (CRT) for heuristics and biases [Hagendorff et al., 2023, Binz and Schulz, 2023]; False-Belief Tasks for ToM [Ullman, 2023, Chen et al., 2024d]; and many more. Established inventories provide a straightforward and standardized approach to psychometric evaluation. However, many studies question the reliability and validity of established inventories when used with LLMs, citing common issues of data contamination and response bias [Ye et al., 2025a]. Additionally, these simplified, structured tests may not adequately capture the complex task scenarios encountered by general-purpose AI.
Custom-Curated Items.
Human-authored and custom-curated items provide tailored psychometric tests that are often more relevant and applicable to LLMs. These items enable researchers to explore novel and unique dimensions of LLM capabilities more robustly. Shen et al. [2024] argue that SVS and PVQ are inadequate for measuring LLM value alignment because they lack supplementary values from alignment literature and contextual nuances of human-AI alignment. To address this, they created 11 AI-informed value statements, similar in style to SVS and PVQ, based on a systematic review of human-AI alignment literature. Similarly, Ceron et al. [2024] suggest that political questionnaires may yield unreliable or inconsistent results. They propose a series of tests to assess the reliability and consistency of LLMs’ stances on political statements, using a dataset of voting-advice questionnaires from seven EU countries annotated for policy issues. Regarding the evaluation of LLM morality, Hendrycks et al. [2021] curate a large-scale dataset of over 130,000 open-world scenarios requiring LLMs to make moral judgments. Another example is the ToM dataset based on naturally occurring spoken dialogues [Soubki et al., 2024], which is motivated by the potential misalignment between synthetic ToM benchmarks and actual human behavior. Despite their high ecological validity, custom-curated items require significant effort to develop and validate, which may limit their scalability and diversity.
Synthetic Items.
Synthetic items generated by AI models, mostly LLMs themselves, represent an emerging paradigm in LLM psychometrics. This method leverages the generative capabilities of LLMs to produce large-scale, diverse, and contextually rich items. Using synthetic items entails careful prompt engineering and additional validation.
Some studies prompt LLMs to modify or expand established inventories for better ecological validity, mitigating data contamination, and enabling more scalable and diverse testing. For example, Ren et al. [2024] rewrite self-report items into advice-seeking queries that better align with real-world human-AI interactions. Bhandari et al. [2025b] modify the established items into semantically equivalent ones to address the issue of data contamination. They validate the rewritten items by setting a threshold for semantic embedding similarity. Zhu et al. [2024a] transform the cognitive test items into more diverse ones, allowing for multifaceted analysis of LLMs’ cognitive abilities.
Other studies generate synthetic tests from scratch. Hadar-Shoval et al. [2025] validate their synthetic cognitive tests by showing a strong human performance correlation between LLM-generated tests and established tests. Ye et al. [2025a] generate value-eliciting prompts for LLM value measurement, subsequently confirming the reliability and validity of the test results. Jiang et al. [2024a] propose generative evolving testing, where an LLM-based Item Generator learns to generate items with specified difficulty. The moral dilemmas presented by Chiu et al. [2025] are created using GPT-4 and employed to evaluate the moral judgment of LLMs. When measuring SI of LLMs with agentic simulations, Zhou et al. [2024b], Mou et al. [2024] prompt LLMs to generate various components of the social interaction.
6.3 Prompting Strategies
Many structured tests employ standard test prompts, reformatted from those used for human participants. Others involve various prompting strategies.
Role-Playing Prompts.
Role-playing prompts, also known as persona or profiling prompts, incorporate specific demographic information or personal attributes into the LLM context during test administration.
Some studies utilize role-playing prompts to generate multiple participants from a single LLM for statistical analysis. Serapio-García et al. [2023] incorporate persona instructions and descriptions before each item when assessing LLM personality. This approach yields diverse responses from a single LLM, facilitating the computation of the reliability and validity of the test results for each LLM. Ye et al. [2025b] use value-anchoring prompts to guide LLM values, thereby generating hundreds of LLM participants for value measurement. The statistical analysis of these measurement results leads to the development of an LLM-specific value system.
Research suggests that LLMs embody a superposition of diverse personalities, values, and perspectives [Kovač et al., 2023]. This has motivated investigations into the personality adaptability of LLMs. Utilizing role-playing prompts, Lu et al. [2023], Jiang et al. [2023], La Cava and Tagarelli [2024] demonstrate that LLMs can dynamically transition between different personality types. However, the degree of steerability varies among LLMs [La Cava and Tagarelli, 2024], and some personality dimensions are less steerable than others [Li et al., 2024d, Jiang et al., 2023]. Kovač et al. [2024], Rozen et al. [2024] employ role-playing prompts to examine the stability and consistency of LLM values. Similarly, Münker [2024], Simmons [2023], Wright et al. [2024] incorporate moral or political profiles into prompts to evaluate LLMs’ ability to mimic these perspectives and their susceptibility to biases linked to specific groups.
Many cognitive benchmarks, particularly those related to social interactions, assign personas and social roles to LLMs during agentic simulations [Zhou et al., 2024b, Mou et al., 2024, Huang et al., 2023b]. The successful completion of these benchmark tasks heavily relies on the LLM’s ability to comprehend and emulate the assigned persona and social role. Persona prompts are also shown to affect the social-cognitive reasoning of LLMs when other task variables are held constant [Tan et al., 2024].
The susceptibility to role-playing prompts presents a double-edged sword. On one hand, they enable the emulation of diverse perspectives and behaviors, which can be beneficial for applications such as role-playing conversations and social simulations. On the other hand, these capabilities may result in inconsistencies and instability in the LLM’s values and perspectives. It may compromise the reliability and validity of test results and pose challenges to robustly aligning LLMs with human values.
Performance-Enhancing Prompts.
Performance-enhancing prompts are crafted to augment LLMs in psychometric evaluations. Hagendorff [2023] suggest using such prompts whenever possible when evaluating machine psychology. The Chain of Thought (CoT) prompting [Wei et al., 2022] is leveraged to reduce biases [Hagendorff et al., 2023], enhance social intelligence [Shapira et al., 2023a], and boost cognitive capabilities [Coda-Forno et al., 2024]. Variants of CoT, such as Emotional CoT [Li et al., 2024e], specifically target improvements in the emotional intelligence of LLMs. Emotional prompts, as introduced by Li et al. [2023a], enhance the general cognitive abilities of LLMs by incorporating emotional stimuli. Few-shot prompting [Brown et al., 2020], another versatile prompting technique, has been shown to improve ToM performance in LLMs [Moghaddam and Honey, 2023].
Additional research explores diverse prompting strategies for specific constructs. For instance, Zhou et al. [2024a] advocate for MFT-guided reasoning prompts for better moral alignment. Strategies like those proposed by Echterhoff et al. [2024], Sumita et al. [2024] aim to self-debias LLMs. ToM-specific enhancement strategies include SymbolicToM [Sclar et al., 2023] and temporal decomposition-based prompting [Sarangi et al., 2025, Hou et al., 2024]. In contrast, Zhao et al. [2025a] develop a testing framework utilizing self-reflection prompts to examine both explicit and implicit social biases in LLMs, where explicit bias measurement serves as a reflection on implicit bias.
Prompt Perturbation and Adversarial Attacks.
Prompt perturbations can test the robustness of LLMs’ personality traits and cognitive abilities. Researchers have investigated whether LLMs can maintain stable personality, values, opinions, and cognitive abilities when item options are reordered [Schelb et al., 2025, Lee et al., 2024d], prompts are rephrased [Lee et al., 2024d, Fraser et al., 2022, Strachan et al., 2024a], prompt formats are altered [Moore et al., 2024, Schelb et al., 2025], or different languages are used [Moore et al., 2024, Cahyawijaya et al., 2024]. For instance, Faulborn et al. [2025] introduce 30 variations of prompts to evaluate the political biases of LLMs. In addition, Wen et al. [2024a] propose psychometric-inspired adversarial attacks to uncover implicit biases in LLMs. Similarly, Li et al. [2024d] utilize persuasive adversarial prompts (PAP) [Zeng et al., 2024a] to test the robustness of LLMs’ values. Since LLMs are sensitive to these perturbations, researchers have begun to scrutinize the reliability of the test results obtained under standard conditions [Ye et al., 2025a, Röttger et al., 2024, Dominguez-Olmedo et al., 2024, Strachan et al., 2024a].
6.4 Model Output and Scoring
6.4.1 Closed-Ended Output and Scoring
The evaluation of LLM performance on structured tests can be categorized into two methodological paradigms: logit-based probabilistic analysis and closed-ended output scoring. In multiple-choice or Likert-scale evaluations, some studies retrieve and analyze the distribution of token-level logits—particularly the first generated token—to infer the model’s latent personalities and opinions [Pellert et al., 2024, Santurkar et al., 2023], analyze response entropy [Dominguez-Olmedo et al., 2024], and compute the distributional alignment between LLM outputs and human behavioral data [Arehalli et al., 2022]. Psycholinguistic tests especially rely on logit-based metrics. A notable example is surprisal, defined as the negative log probability of a token given its context. It quantifies the predictability or cognitive effort associated with language processing, enabling researchers to map LLM outputs to human-like uncertainty patterns [Wang et al., 2024b, Steuer et al., 2023].
Conversely, closed-ended outputs, which are explicit numerical scores or categorical selections, can be analyzed using predefined scoring protocols. For Likert-scale responses, scores are typically averaged or aggregated based on established rubrics (e.g., VSM13 [Ye et al., 2025a] and DIT [Khandelwal et al., 2024]). In standardized cognitive tests, model outputs are evaluated against ground-truth labels (e.g., accuracy in arithmetic tasks) or rule-based criteria (e.g., established standards in the Verbal Comprehension Index) [Galatzer-Levy et al., 2024].
6.4.2 Open-Ended Output and Scoring
Scoring open-ended outputs is more challenging. Related scoring schemes can be generally categorized into 1) rule-based scoring, 2) model-based scoring, and 3) human scoring.
Rule-Based Scoring.
Rule-based scoring predominantly relies on the lexical hypothesis [Allport and Odbert, 1936], which posits that the meaning and relevance of responses can be ascertained by analyzing the presence and frequency of specific keywords or phrases within the text. Jiang et al. [2024b] utilize the LIWC (Linguistic Inquiry and Word Count) features [Pennebaker et al., 2001] to evaluate the personalities of LLMs. Similarly, Fischer et al. [2023] employ a theory-driven value dictionary [Ponizovskiy et al., 2020] to assess LLM values. Nevertheless, lexicon-based scoring has been demonstrated to be limited in capturing semantic nuances [Ye et al., 2025a]. Some other scoring rules extend beyond simple keyword matching. For example, Healey et al. [2024] automatically identify biases by analyzing whether responses deviate from equivalent treatment.
Model-Based Scoring.
Model-based scoring is prevalent in current unstructured testing due to its flexibility and scalability. Several studies have focused on training models specifically to evaluate LLM responses. For instance, Hilliard et al. [2024] fine-tuned BERT variants [Devlin et al., 2019] using the MyPersonality dataset [^5] https://sites.google.com/michalkosinski.com/mypersonality. to score LLM personalities. Similarly, Ye et al. [2025a], Yao et al. [2024, 2025b], Sorensen et al. [2024a] fine-tune LLMs with established psychometric inventories [Ren et al., 2024] and/or human annotations to classify the value valence of LLM responses. Scoring can occur at the item level, as exemplified by Generative Psychometrics, which consists of item parsing, scoring, and aggregation to evaluate LLM values [Ye et al., 2025a]. Alternatively, scoring can be conducted at the response level, assessing the LLM response in its entirety [Yao et al., 2024, 2025b]. Some models are trained to score according to specific theoretical frameworks, e.g., Schwartz’s Value Theory [Yao et al., 2024], while others are designed to be generalists, utilizing a broader range of fine-tuning data and leveraging the prior knowledge of LLMs [Ye et al., 2025a]. Overall, training evaluators is more prevalent for evaluating personalities and values due to the ease of associating textual expressions with specific personality and value dimensions.
More studies directly employ the LLM-as-a-judge approach [Gu et al., 2024] to score the open-ended responses. For example, Li et al. [2024d] prompt LLMs to score LLM responses across personality, ToM, and motivation. They validate the consistency between two LLM raters to indicate inter-rater reliability. Zheng et al. [2025] implement LLMs as personality evaluators and validate their consistency with human raters. The LLM-as-a-judge approach is particularly advantageous for evaluations in highly unstructured simulation settings. Zhou et al. [2024b], Wang et al. [2024d, a], Mou et al. [2024] apply LLM judges to assess SI across various dimensions, including goal completion, relationship maintenance, and adherence to social rules. Additionally, several works use embedding models to evaluate the similarity between LLM responses and prototypical examples [Huang et al., 2024c, Cahyawijaya et al., 2024, Amirizaniani et al., 2024].
Human Scoring.
Human scoring is employed when rigorous evaluation, adhering strictly to standard psychometric manuals, is necessary. For instance, Elyoseph et al. [2023] engage psychologists to assess the contextual suitability of LLM responses using the Levels of Emotional Awareness Scale (LEAS). Castello et al. [2024] conduct an examination of cognitive biases in LLMs through human evaluation and linguistic comparison. Healey et al. [2024] present a semi-automated pipeline to classify LLM responses into nuanced bias types. Other examples include ToM evaluation in open-ended responses [Amirizaniani et al., 2024] and cognitive evaluation based on the Montreal Cognitive Assessment (MoCA) [Dayan et al., 2024].
6.5 Inference Parameters
Evaluation results and validation methods also depend on inference parameters of LLMs. Some studies use greedy decoding for deterministic outputs, requiring only a single response per item but sacrificing output diversity. In contrast, sampling-based decoding introduces stochasticity, generating a broader range of outputs by adjusting hyperparameters such as temperature, top-k, and top-p.
Inference parameters can affect the measured traits and abilities of LLMs. Diversifying the outputs may reveal a broader spectrum of latent traits, opinions, or cognitive strategies, but also introduce variability that complicates reliability and validity assessments. Conversely, deterministic settings enhance reproducibility but may obscure the model’s full range of capabilities or biases.
Most studies explicitly report and control these settings to ensure fair comparisons across models or experimental conditions. Some also examine the sensitivity of psychometric results to different decoding parameters. Transparent reporting and methodological rigor in parameter selection are essential. Investigating both deterministic and stochastic settings is recommended to fully understand the implications of test results.
7 Psychometric Validation
Figure 4: Overview of psychometric validation: reliability and consistency, validity, and standards and recommendations [Löhn et al., 2024].
Unlike AI benchmarking, which prioritizes system performance over test justification, psychometrics emphasizes theoretical grounding, standardized protocols, and reproducibility. Psychometric validation rigorously ensures the reliability, validity, and fairness of psychological tests. LLM psychometrics, as an emerging field, still lacks standardization for test design and administration, and recent research has begun exploring various aspects of psychometric validation.
7.1 Reliability and Consistency
Reliability is a fundamental principle in psychometric validation, assessing the degree to which a test is free from error. It encompasses consistency over time (test-retest reliability), across different versions (parallel forms reliability), and among evaluators (inter-rater reliability). Researchers have adapted these psychometric reliability metrics for LLM psychometrics. For instance, Li et al. [2024d] introduce a multi-dimensional psychometrics benchmark that extends to five forms of reliability, addressing the unique challenges of LLM evaluation: internal consistency, parallel forms reliability, inter-rater reliability, option position robustness, and adversarial attack robustness. Studies by Zheng et al. [2025], Huang et al. [2023a], Ceron et al. [2024], Shu et al. [2024] systematically investigate the consistency of LLM outputs under repeated trials, variations in prompt order, and cross-lingual contexts. They employ reliability indices such as Cronbach’s alpha, intraclass correlation coefficients, and agreement statistics to quantify both internal and external reliability dimensions. We note that some metrics pertain to both stochastic and systematic errors in LLM outputs; certain inconsistencies may compromise both the reliability and validity of LLM psychometrics.
Promising results indicate that some advanced LLMs consistently generate stable responses on the Big Five personality traits across diverse settings [Huang et al., 2023a]. Larger and instruction fine-tuned models exhibit even higher reliability when utilizing role-playing prompts [Serapio-García et al., 2023]. Moore et al. [2024] also observe relatively stable value orientations of LLMs across (1) paraphrases of a single question, (2) related questions within one topic, (3) multiple-choice and open-ended formats of a single question, and (4) multilingual translations. Li et al. [2024d] report strong inter-rater reliability when using LLMs as judges, though they note varying degrees of personality stability across models and differing levels of value consistency across cultural contexts.
However, other studies present less favorable results. LLMs are notorious for prompt sensitivity in the context of psychometric evaluation. Trivial changes to prompt form [Röttger et al., 2024, Ren et al., 2024], option order [Li et al., 2024d, Dominguez-Olmedo et al., 2024], or syntactic structure [Hu and Levy, 2023] frequently result in systematic shifts in LLM outputs, severely undermining parallel forms reliability. Additionally, studies reveal distributional inconsistencies between next-token logits and forced-choice responses in psycholinguistic tests [Hu and Levy, 2023], as well as discrepancies between forced-choice and free-form responses in value [Kovač et al., 2024, Ye et al., 2025a] and political opinion surveys [Röttger et al., 2024]. In cognitive tests, Ullman [2023], Shapira et al. [2023a], Holterman and Deemter [2023] demonstrate that minor perturbations in ToM tasks often result in poor performance, indicating low parallel forms reliability of some ToM tests and LLMs’ dependence on fragile heuristics. Xu et al. [2025b] develop the Words and Deeds Consistency Test (WDCT) to measure the consistency between LLMs’ verbal and behavioral responses across various domains, including opinion vs. action, non-ethical value vs. action, ethical value vs. action, and theory vs. application. Their findings reveal that LLMs’ responses are inconsistent.
The findings regarding reliability are clearly mixed. It seems to be influenced by various factors such as the target construct, forms of reliability, test formats, prompting, and the models used. LLMs tend to yield more reliable results in personality assessments [Huang et al., 2023a, Serapio-García et al., 2023] compared to political opinion evaluations [Dominguez-Olmedo et al., 2024, Röttger et al., 2024]. The degree of inter-rater agreement differs across various instruments and domains [Bodroža et al., 2024]. LLMs are reported to exhibit high internal consistency (Cronbach’s alpha $>0.8$), particularly in closed-choice personality tests [Huang et al., 2023a, Serapio-García et al., 2023, Zheng et al., 2025], while demonstrating low parallel form reliability due to prompt sensitivity [Li et al., 2024d, Shu et al., 2024, Gupta et al., 2024]. In addition, LLMs show greater consistency in structured tests with strong prompt control [Rozen et al., 2024, Serapio-García et al., 2023, Klinkert et al., 2024], but their reliability diminishes with less standardized prompts [Petrov et al., 2024] or when addressing more controversial topics [Moore et al., 2024]. The reliability also varies across models [Bodroža et al., 2024]. Moore et al. [2024] report that base models exhibit more consistent values compared to fine-tuned models in value-laden questions. Enhanced model safety is strongly associated with higher value consistency [Ye et al., 2025b], and more advanced models demonstrate superior role-playing performance, thereby increasing reliability in personality tests utilizing role-playing prompts [Serapio-García et al., 2023].
7.2 Validity
Validity, another cornerstone principle of psychometric validation, determines whether a test accurately measures its intended construct. Recent research has systematically investigated multiple aspects of validity in LLM psychometrics, addressing both methodological challenges and potential solutions unique to this emerging field.
7.2.1 Content Validity
Content validity ensures that a test comprehensively covers the construct it aims to measure. Major challenges of content validity include data contamination of well-established tests and novel items that are inadequately examined.
Data Contamination.
Some studies naively transfer human tests to LLMs. Given that LLMs are trained on vast Internet-scale corpora, it is highly probable that they have been exposed to instances of test items or similar content [Hagendorff et al., 2024, Jiang et al., 2024c]. This exposure can result in the reproduction of known token patterns during evaluation, which leads to inflated performance metrics or biases the reflection of the LLM’s internal traits. To address this issue, researchers reformat test content or dynamically generate novel tests; see § 6.2 for more details.
Invalid Novel Items.
The reformatting or dynamic generation of novel tests presents new challenges for content validity. The novel test items may capture only limited dimensions of the construct or involve extraneous factors, thus inadequately representing target constructs. This may be due to a less rigorous crowd-sourcing process for benchmark curation, or the inherent biases and limited capabilities of the AI models used for stimulus synthesis. Content validity assessments for custom-curated and model-generated items are crucial but rarely conducted in LLM psychometrics.
7.2.2 Construct Validity
Construct validity assesses whether a test accurately measures the theoretical construct it is intended to evaluate. For example, a position paper by Alaa et al. [2025] underscores both the deficiency and critical importance of construct validity in medical LLM benchmarks. Broader literature suggests that construct validity could be undermined by the unique internal abstractions, systematic response patterns, and social desirability bias of LLMs.
Unique Internal Abstractions
LLMs and humans may differ fundamentally in how they internally represent psychological constructs. For instance, Kovač et al. [2023] argue that LLMs exhibit contextually adaptive traits rather than the stable psychological characteristics found in humans. Sühr et al. [2023] systematically challenge the validity of personality tests for LLMs. They find personality types do not follow the clear separation into big five traits observed in humans. Similarly, Peereboom et al. [2024] examine the latent personality structures of humans and three LLMs, concluding that the established HEXACO tests for humans are not applicable to LLMs, as the relevant personality factors may be absent in these models. In measuring values, Ye et al. [2025a] find that self-report results are inconsistent with Schwartz’s theoretical model, and that measuring open-ended responses improves the alignment, but not towards a perfect fit. Hadar-Shoval et al. [2024] also agree that LLMs encode values distinct from those of humans. There is a clear consensus among researchers that a robust evaluation and understanding of LLM psychology necessitates new operational definitions that are distinct from, or only loosely analogous to, human traits.
Pioneering research presents value systems specific to LLMs, in response to their unique internal abstractions [Ye et al., 2025b, Biedma et al., 2024]. Notably, Ye et al. [2025b] propose a theoretically grounded generative psycho-lexical approach for system construction based on GPV value measurement [Ye et al., 2025a]. They evaluate their proposed value systems through statistical analysis and explanatory, predictive, and alignment tasks, establishing the validity of the value factors and their superiority over Schwartz’s value system. The unique constructs of LLMs necessitate the rigorous development of tailored tests. Ma et al. [2025] introduce the Core Sentiment Inventory (CSI), an instrument for measuring the emotional tendencies and personalities of LLMs. Fang et al. [2024] curate a psychologically grounded benchmark for mathematics proficiency, aiming for better validity, modeling item difficulty, and allowing for comparison against human norms. Similar attempts are made by Lee et al. [2024d], who propose the TRAIT test, which yields higher validity and reliability than other personality tests.
Response Set.
Response set refers to the systematic pattern of answering questions in a particular way, regardless of the item’s content. Some LLMs demonstrate option position bias towards responses labeled as "A", particularly in political opinion tests [Dominguez-Olmedo et al., 2024, Li et al., 2024d]. After adjusting for the option position bias, Dominguez-Olmedo et al. [2024] additionally find that LLM survey responses to political opinion questions are uniformly random. In this case, it becomes meaningless to analyze the distributional alignment between LLM responses and opinions of demographic groups. Similar results are observed in LLM personality tests [Song et al., 2023, Sühr et al., 2023] and value tests [Ye et al., 2025a]. Sühr et al. [2023] discover LLMs’ tendency to respond affirmatively to reverse-coded items (e.g., "I am introverted" vs. "I am extroverted"). Ye et al. [2025a] identify that certain LLMs lean towards low/high ratings. They conclude that self-report tests are not the correct tools for LLM psychometrics.
Social Desirability Bias.
Social desirability bias is the tendency to give responses that conform to social norms rather than reflect true beliefs. Studies indicate that LLMs exhibit this bias similarly to humans. Salecha et al. [2024] report that in personality tests, LLMs tend to skew their responses towards socially favorable traits, leading to higher extraversion scores and lower neuroticism ratings. Likewise, Ye et al. [2025b] find that when LLMs are directly asked to self-report values, as per Biedma et al. [2024], they avoid reporting less socially desirable values like hedonism, although they express these values when given indirect, contextual prompts. While fine-tuning for social desirability is important, it is crucial to maintain a balanced representation of diverse perspectives and values. Over-optimizing for social desirability can result in homogeneous models that fail to capture the full range of human values and do not effectively meet diverse user needs.
Cross-lingual Tests.
Some research investigates the psychometric validity in cross-lingual settings. Romero et al. [2024] administer a standardized personality questionnaire to LLMs in nine languages and find notable inconsistencies in measurement results. Despite their multilingual abilities, LLMs may lack consistency in psychological traits across languages [Cahyawijaya et al., 2024]. These inconsistencies complicate the psychometric validation of cross-lingual tests, necessitating the disentanglement of intrinsic model biases, translation equivalence issues, and cultural differences represented by the models.
7.2.3 Criterion and Ecological Validity
Criterion validity involves the correspondence of test results with external standards, while ecological validity evaluates the applicability of these outcomes to real-world scenarios. In LLM psychometrics, external standards often overlap with real-world evaluation results. Ren et al. [2024], Ye et al. [2025a] report that value orientations measured by forced-choice tests do not align with those measured in human-LLM interactions. Similar discrepancies are noted in personality [Ai et al., 2024] and morality [Nunes et al., 2024]. Consequently, Zhang et al. [2024a] advocate for linking machine personality assessments to safety. Ren et al. [2024] recommend measuring values within real-world human-LLM interactions, and Ye et al. [2025a] further demonstrate that values measured in open-ended dialog enhance construct validity given Schwartz’s circumplex model. Further work by Ye et al. [2025b] connects LLM value measurement to safety prediction and value alignment tasks to ensure conformity with external standards.
7.3 Standards and Recommendations
Based on the identified challenges and solutions, researchers have proposed general standards and recommendations for LLM psychometrics, aiming to establish a methodological foundation for this nascent field. Frank [2023a] explores how developmental psychology helps us understand the internal representations and cognitive capacities of LLMs. The author recommends using simplified and novel stimuli to avoid the effects of training data contamination. [Löhn et al., 2024] critique the current state of psychometric evaluation on LLMs and argue for standardized criteria to ensure valid assessments. They propose seven requirements: reliability (consistent results), validity (measuring intended traits), suitability (matching LLM capabilities), non-disclosure (avoiding test contamination), and fairness; fairness further includes valid comparisons, valid translations, and transparency. Analyzing 25 studies, they find widespread neglect of these principles.
More recently, Hagendorff et al. [2024] recommend procedural test generation, multiple task versions, performance-enhancing prompts, shuffling options in multiple-choice questions, and using multiple scoring methods to avoid data contamination and ensure reliability. They also suggest using deterministic settings for reproducibility, automated tools for evaluation, and manual reviews for unreliable outputs. After evaluations, they advocate for performing statistical analysis to interpret the results. Vaugrante et al. [2024] identify the lack of replicability in recent works and propose four recommendations. First, they advise ensuring benchmark validity, providing adequate tasks for statistical analysis, standardizing for comparability, controlling prompt sensitivity, and aligning with research objectives. Second, adopt standardized methodologies, avoid cherry-picking, ensure statistical transparency, document experimental setups comprehensively, and define consistent evaluation metrics to enable reliable and replicable research. Third, monitor model behavior changes, account for variability with diverse models, adjust benchmark difficulty as models improve, and ensure transparency by documenting model versions and experiment dates. Lastly, standardize scoring and verification processes, ensuring accuracy and transparency; benchmark creators should provide clear verification guidelines and scoring rubrics for implementers. Schelb et al. [2025] release a framework for designing and running robust and reproducible psychometric experiments. The framework focuses on robustness, flexibility, usability, and reproducibility, based on a standardized configuration file defining experiments.
8 Psychometrics for LLM Enhancement
Psychometric principles serve as powerful tools for LLM development and enhancement, beyond mere evaluation. This shift has led to three key advancements: trait manipulation for personalized and controlled model behaviors, safety and alignment through psychometrically-informed interventions, and cognitive enhancement to foster more human-like reasoning and communication capabilities.
8.1 Trait Manipulation
One of the major applications of psychometrics is to manipulate the traits of LLMs, for chatbot personalization, role-play applications, and simulating demographic groups. Psychometrics provides a robust framework to inform such manipulation across prompting, inference, and training stages.
Structured prompting, grounded in validated psychometric inventories, are shown to reliably elicit and modulate specific personality traits in LLMs [Huang et al., 2024c, He and Zhang, 2024]. This approach facilitates the controlled simulation or alignment of synthetic personas [Serapio-García et al., 2023, Jiang et al., 2023, Zhang, 2024, La Cava and Tagarelli, 2024]. For example, Jiang et al. [2023] introduce personality prompting ($P^{2}$) based on the lexicon theories behind the Big Five personality traits. In $P^{2}$, lexicons related to specific traits are expanded into detailed personality descriptions by LLMs. The LLMs are then personalized towards diverse behaviors. Chuang et al. [2024] propose a more systematic prompting strategy based on empirically-derived human belief networks, which encompasses 64 topics loading on 9 non-overlapping latent factors. When prompted with one belief, the LLMs become more aligned with other related beliefs.
Beyond prompt engineering, inference-time interventions directly manipulate hidden representations during forward pass, with many based on patterns recognized using psychometric scales. ControlLM [Weng et al., 2024], Personality Alignment Search [Zhu et al., 2024b], Neuron-based Intervention [Deng et al., 2024], Probing-then-Editing [Ju et al., 2025], and Latent Feature Steering [Yang et al., 2025b] manipulate model outputs on trait axes by directly shifting activations or targeted neuron values at inference time, achieving trait control without retraining.
Other works further fine-tune the LLMs for trait manipulation. Vu et al. [2024] adapt the model architecture to incorporate trait control, reflecting continuous Big Five and mental health dimensions. Dan et al. [2024], Jain et al. [2024], Li et al. [2024c] deploy LoRA modules or routing networks tied to personality traits, learning specialist adapters for trait-driven generation. Cui et al. [2023], Zeng et al. [2024b], Liu et al. [2024a] achieve deep trait embedding by training on large, annotated dialogue datasets derived from validated psychometric instruments. Additional work directly modulates the model parameters without retraining. Hwang et al. [2025] generate adjustment queries based on the MBTI questionnaire and edit the model parameters to align their responses with the desired personality traits.
The manipulation of traits in LLMs presents a promising avenue for substituting human participants in social science research, as reviewed by Gao et al. [2024]. Cao et al. [2025], Suh et al. [2025] demonstrate that fine-tuning LLMs enables them to accurately simulate survey response distributions of specific subpopulations, even when faced with previously unseen survey questions. He-Yueya et al. [2024] introduce the concept of psychometric alignment to evaluate the congruence between LLMs’ knowledge and human understanding. By training LLMs on human response data from the target distribution, they enhance the models’ psychometric alignment on novel test items. However, the efficacy of this training approach is contingent upon the domain. Furthermore, Kang et al. [2023], Sorensen et al. [2025] explore the representation of human values within LLMs and fine-tune them to simulate human opinions through value injection.
Although current personalization techniques have demonstrated the ability to effect more than superficial changes in model outputs, achieving robust, authentic, and context-independent personalization remains a significant challenge [Kovač et al., 2024, Dominguez-Olmedo et al., 2024].
8.2 Safety and Alignment
Recent research has established a connection between LLM psychometric measurement and their safety and alignment, one of the most pressing issues in the field. Zhang et al. [2024a], Wang et al. [2025] link LLM personality traits to their safety. Specifically, Zhang et al. [2024a] identify a significant relationship between personality traits, as measured by the MBTI-M scale, and safety, noting that better alignment enhances traits such as Extraversion, Sensing, and Judging. They are inspired to modify LLM personalities, resulting in improved safety, particularly in privacy and fairness. Similarly, Wang et al. [2025] examine the correlation between HEXACO personality traits and LLM safety.
In parallel, Ye et al. [2025a, b], Yao et al. [2024] have explored the relationship between LLM values and their safety and alignment. Yao et al. [2024] demonstrate the feasibility of distinguishing between safe and unsafe LLM responses within the basic value space, analyzing correlations with Schwartz’s values and specific safety issues. Ye et al. [2025a, b] achieve high accuracy in predicting LLM safety scores based on value orientations measured by GPV, a generative psychometric tool. Furthermore, Yao et al. [2024], Ye et al. [2025b] employ reinforcement learning to align LLMs with desired human values, thereby enhancing their safety. In the domain of morality, Huang et al. [2024a], Tlaie [2024] introduce prompting techniques grounded in Moral Foundation Theory to enhance the moral reasoning and alignment of LLMs. Takeshita et al. [2023], Ohashi et al. [2024] create the JCommonsenseMorality (JCM) dataset and fine-tune LLMs to adapt to local culture.
8.3 Cognitive Enhancement
Psychometrics has proven to be an effective tool for developing human-like reasoning, empathy, and communication skills in LLMs. Recent studies leverage established psychological theories and psychometric frameworks to guide LLMs’ cognitive enhancement, employing strategies like prompting, architectural modules, and specialized training regimes.
Psychology-inspired prompting is shown to enhance the cognitive capabilities of LLMs. For example, Li et al. [2023a] introduce emotional prompts, demonstrating that such stimuli can improve the general cognitive abilities of LLMs. Role-playing prompts are observed to influence ToM capabilities [Tan et al., 2024] and promote more human-like reasoning [Nighojkar et al., 2025]. Zhao et al. [2025b] employ a finite-state machine paradigm based on Hill’s Helping Skills theory. It embeds state transitions for emotional support conversations directly within the model’s multi-hop inference process as additional contexts. This method effectively structures dialogues around psychological support strategies and emotional states, leading to higher human-rated effectiveness and strategic alignment.
Liu et al. [2023] introduce a neural listener module within an LLM architecture, optimizing a combined objective that encodes ToM reasoning into the model’s training. They emphasize the potential of using psycholinguistic theories of human language acquisition to enhance LLM language learning. Additionally, preference-based and RL-based methods are proposed to internalize empathetic capabilities in LLMs. Research in empathetic response generation often utilizes cognitive and affective models of empathy to create composite reward functions or preference signals for fine-tuning model behavior. For instance, Sotolar et al. [2024] generate training data using formal emotion theories, like Plutchik’s wheel of emotions, to create preference pairs for model optimization. Others have introduced empathy classifiers to shape reinforcement learning rewards, guiding models toward empathetic outputs [Sharma et al., 2021]. Further research improves LLM social-pragmatic reasoning through preference optimization based on human judgments in open-ended social tasks, showing that these objectives enhance pragmatic inference [Wu et al., 2024].
9 Trends, Challenges, and Future Directions
This section explores the emerging trends, challenges, and prospective directions in LLM psychometrics.
9.1 Psychometric Validation
There is an emerging consensus on the importance of psychometric validation in evaluating personality traits in LLMs; various studies contribute to this validation across different dimensions (§ 7). However, research on psychometric validation in ability testing remains limited. Current evaluations of LLM abilities predominantly employ structured tests, such as multiple-choice questions, often directly adapted from tests originally designed for human participants. These evaluations face challenges. For instance, Ullman [2023] highlight that LLMs fail on trivial alterations of ToM tasks, indicating a lack of parallel forms reliability. The unique internal abstractions of LLMs undermine construct validity and lead to flawed benchmarks [Riemer et al., 2025]. Furthermore, the extent to which results from structured tests of LLM abilities generalize to real-world human-AI interactions is rarely investigated, leaving criterion and ecological validity issues insufficiently explored.
On the other hand, newly developed and more sophisticated simulation-based tests are often constructed with less rigorous psychometric processes. Consequently, they may involve extraneous factors unrelated to the construct of interest or fail to adequately cover the construct’s manifestations. In the realm of psychometric-inspired, construct-oriented tests, pioneering efforts such as Zhu et al. [2024a] introduce multifaceted evaluations of three cognitive factors identified from benchmarking data, while Zhou et al. [2025] propose a theory-driven, richer set of general scales. Defining a comprehensive set of core abilities and designing valid tests to measure them remains an open challenge.
9.2 From Human Constructs to LLM Constructs
Recent research shifts from employing constructs established for humans to developing novel constructs specific to LLMs. Studies by Sühr et al. [2023], Peereboom et al. [2024], Ye et al. [2025b], Biedma et al. [2024] indicate that the factor structures of human personalities and values may not be applicable to LLMs. Further, Ye et al. [2025b], Biedma et al. [2024], Federiakin [2025], Burnell et al. [2023] have tailored these constructs to better fit LLMs. Despite these advancements, the development of novel constructs seems highly contingent upon the measurement tools, tasks, and specific LLMs utilized. For instance, Ye et al. [2025b] and Biedma et al. [2024] report differing value structures, while Federiakin [2025] and Burnell et al. [2023] identify varying ability factors. Future research should further explore the fundamental structures that underpin LLM behaviors.
9.3 Perceived vs. Aligned Traits
Han et al. [2025] suggest a discrepancy between perceived values in text (objective human annotation) and how responses align with annotators’ personal values (subjective annotation). For instance, consider the LLM response: "Wealthy individuals are not necessarily greedy; some may be motivated by concern for the greater good." An objective annotation categorizes this response as embodying Power, due to its themes of societal impact. However, annotators with values of Conformity, rather than Power, report a higher similarity between the response and the their own thoughts. In this case, Power represents the perceived value, while Conformity is the aligned value.
This discrepancy may extend to other traits and holds significant methodological implications for LLM psychometrics. It is crucial to deliberate on which aspect—perceived traits or aligned traits—holds greater significance. Current research on value evaluation is driven by the potential of LLMs to influence, alter, and entrench human values. However, the underlying mechanisms of such influence remain unclear. For instance, if individuals are inclined to adopt LLM decisions and values [Glickman and Sharot, 2024], it is uncertain whether they are adhering to perceived traits or aligned traits. Additionally, the impact of judges’ subjectivity on psychometric evaluation is not well understood. Existing scoring models or LLM-as-a-judge approaches often assume a universal criterion. The modeling of judges’ subjectivity, such as how individuals with different values may interpret the same response differently, remains an underexplored area.
9.4 Anthropomorphization Challenges
The anthropomorphization of LLMs in research varies, which leads to different ways of performing statistical analysis. Some studies conceptualize an LLM as an individual entity, while others regard it as a population. For instance, certain studies evaluate LLM traits by considering the model as a single participant, utilizing its default settings with prompts such as "You are a helpful assistant" [Ye et al., 2025a]. Alternatively, some researchers view an LLM as a superposition of multiple participants, employing diverse role-playing prompts to steer it towards different profiles [Serapio-García et al., 2023]. Another approach involves treating the LLM as a monolithic entity, but administering varied prompts for identical test items to generate multiple sets of results [Hagendorff et al., 2023].
When an LLM is treated as a monolithic entity, the reliability and validity of the results are primarily contingent upon the measurement tools, allowing researchers to compare different tools [Ye et al., 2025a]. Conversely, when an LLM is perceived as a population, the reliability and validity are also influenced by the specific LLM employed, prompting comparisons of consistency and role-playing capabilities across different models [Serapio-García et al., 2023].
The meaningfulness of evaluating LLMs in their default settings remains disputable, especially given their capacity to adopt diverse profiles. The exploration of trait spectrums—a continuum or set of trait demonstrations—in LLMs is currently under-researched. Furthermore, the development of a novel psychometric framework to adequately capture this spectrum has yet to be established. For instance, it is unclear how to identify the extremes to which an LLM can be steered. There is also a lack of consensus on the optimal alignment of LLMs along psychometric dimensions. The chosen alignment target significantly influences the anthropomorphization of LLMs. Some scholars advocate for alignment that enhances consistency, thereby allowing LLMs to be perceived as monolithic entities [Röttger et al., 2024]. Conversely, others position alignment toward pluralistic values, suggesting that LLMs should be regarded as a population [Sorensen et al., 2024b].
9.5 Expanding Dimensions in Model Deployment
LLM psychometrics is expanding beyond traditional text-based, single-turn interactions. Several emerging dimensions present both opportunities and challenges for future research.
Multi-Lingual Evaluation.
While most psychometric evaluations are conducted in English, multilingual LLMs necessitate comprehensive cross-linguistic validation. Current research indicates that LLMs exhibit varying personality traits and cognitive abilities across languages (§ 7.2.2). This raises questions regarding cultural and linguistic biases in existing evaluation frameworks. Future research should examine how measurement results vary across languages, whether psychometric properties (reliability, validity) maintain consistency across linguistic contexts, and how to develop culturally appropriate test instruments.
Multi-Turn Interactions.
The dynamic nature of multi-turn interactions introduces new challenges for psychometric evaluation. Compared to single-turn responses, multi-turn interactions more authentically simulate how humans engage with psychometric tests, where responses are influenced by preceding interactions. Research indicates that when interactions extend across multiple turns, LLMs may exhibit significantly different traits, such as increased vulnerability to jailbreaking attacks [Ying et al., 2025b]. These findings highlight the need for novel evaluation methods that capture temporal dynamics in extended conversations.
Multi-Modal Capabilities.
Multi-modal models require novel psychometric approaches. Current evaluation methods, primarily designed for text-based interactions, fail to adequately capture the complexity of multi-modal understanding and generation. Li et al. [2024b] introduce one of the first value evaluation tools for Vision-Language Models (VLMs). However, this work remains preliminary, employing a highly simplified and controlled test environment that utilizes only the first frame of videos as test stimuli. Furthermore, psychological traits such as values present conceptual challenges. Even when a value like honesty is precisely defined, determining which specific real-world actions manifest this trait remains ambiguous. Consequently, extending psychometric tests to embodied agents and action modalities introduces even greater methodological obstacles. Future research should expand current psychometric tests to incorporate additional modalities. The development of cross-modal evaluation frameworks should maintain psychometric rigor while accounting for the unique characteristics of each modality.
Agent and Multi-Agent Systems.
The emergence of LLM-based agents and multi-agent systems introduces new dimensions to psychometric evaluation. When LLMs operate as autonomous agents, their behavior may be influenced by environmental factors, memory systems, tool usage, and interaction patterns with other agents. However, systematic methodologies for psychometric evaluation in these complex, dynamic settings remain largely undeveloped.
9.6 Item Response Theory
Recent work demonstrates that applying IRT to LLM outputs offers several key advantages, including more informative benchmarking [Guinet et al., 2024], identification of items that effectively discriminate among high-performing models [Lalor et al., 2024], and the potential for adaptive testing that reduce evaluation costs [Zhuang et al., 2023a, Truong et al., 2025]. Most existing applications focus on standard IRT models such as 1PL, 2PL, and 3PL, with limited exploration of polytomous, hierarchical, or fully multidimensional IRT models, even though LLM capabilities are complex and multifaceted. Integrating IRT with generative AI for automated item generation is an emerging direction; however, the current approach can be overly complex, limiting its practical utility [Jiang et al., 2024a]. Furthermore, while IRT holds promise for standardizing the measurement of LLMs and humans on a unified scale—even when different test items are used—this potential remains largely unrealized.
The systematic application of IRT for bias analysis—such as detecting differential item functioning across model families—remains in its early stages [He-Yueya et al., 2024]. IRT offers a principled framework for uncovering biases in AI systems by analyzing performance across items with varying difficulty levels and demographic contexts. For instance, if items involving certain demographic groups are disproportionately challenging, it may indicate underlying model or data biases. Robustness evaluation can also be strengthened by leveraging IRT to assess model performance on adversarial or edge-case items.
IRT can be employed to evaluate the quality of AI benchmarks themselves. While current benchmarks often comprise large collections of test items, the informativeness and design quality of these items are frequently uncertain. Not all items contribute equally; some may lack discriminatory power or inadvertently favor particular model designs. IRT enables a systematic assessment of item difficulty, discrimination, informativeness, and vulnerability to bias and guessing, thereby facilitating the development of more robust and equitable evaluation frameworks.
Furthermore, IRT is rarely utilized beyond evaluation, despite its potential for informing model development. By quantifying item discrimination, IRT can guide the construction of training datasets. Highly discriminating items can be particularly valuable as they provide deeper insights into differentiating between low- and high-performing models.
9.7 From Evaluation to Enhancement
The ultimate objective of evaluation is not only to understand models but also to facilitate their improvement. Current psychometric approaches primarily emphasize the evaluation, comparison, and interpretation of LLMs, while the application of psychometric insights to model enhancement lags behind. As discussed in § 8, psychometric principles can inform prompt engineering, inference-time control of internal representations, training data curation, reward signal design for fine-tuning, and the development of methodological frameworks.
Existing enhancement techniques are still in their infancy. For instance, value targets for LLM alignment in Yao et al. [2024], Ye et al. [2025b] are statically defined, which constrains the incorporation of context-dependent value states observed in human behavior [Skimina et al., 2021]. Drawing on insights from psychometrics enables the development of more effective model enhancement techniques.
10 Conclusion
This paper presents a comprehensive survey of LLM psychometrics. The integration of psychometric instruments, theories, and principles into LLM evaluation promises to overcome the limitations of traditional AI benchmarks. This approach enables us to more effectively capture the broad, emergent psychological constructs of LLMs, encompassing both personality and cognitive dimensions. Psychometric evaluation methodologies vary in test formats, data and task sources, prompting strategies, model outputs, and scoring methods. They each have distinct strengths, weaknesses, and application scenarios, yet all must adhere to psychometric principles such as reliability, validity, and fairness. Beyond evaluation, psychometric-inspired techniques enhance LLMs in trait manipulation, safety and alignment, and cognitive capabilities, contributing to the development of more powerful and responsible AI systems.
It is widely acknowledged that the trajectory of AI development is shifting towards an era characterized by evaluation-driven progress [Silver and Sutton, 2025, Yao, 2025, AAAI, 2025]. We posit that the LLM psychometrics will be pivotal in this evolution, introducing novel principles, dimensions, techniques, and insights. We aim this survey to inspire future evaluation paradigms for human-level AI and foster the advancement of AI psychology for greater common good.

### v2

@@
 Abstract
-The rapid advancement of large language models (LLMs) has outpaced traditional evaluation methodologies.
-It presents novel challenges, such as measuring human-like psychological constructs, navigating beyond static and task-specific benchmarks, and establishing human-centered evaluation.
+The advancement of large language models (LLMs) has outpaced traditional evaluation methodologies.
+This progress presents novel challenges, such as measuring human-like psychological constructs, moving beyond static and task-specific benchmarks, and establishing human-centered evaluation.
 These challenges intersect with psychometrics, the science of quantifying the intangible aspects of human psychology, such as personality, values, and intelligence.
-This survey introduces and synthesizes an emerging interdisciplinary field of LLM Psychometrics, which leverages psychometric instruments, theories, and principles to evaluate, understand, and enhance LLMs.
-We systematically explore the role of Psychometrics in shaping benchmarking principles, broadening evaluation scopes, refining methodologies, validating results, and advancing LLM capabilities.
-This paper integrates diverse perspectives to provide a structured framework for researchers across disciplines, enabling a more comprehensive understanding of this nascent field.
-Ultimately, we aim to provide actionable insights for developing future evaluation paradigms that align with human-level AI and promote the advancement of human-centered AI systems for societal benefit.
+This review paper introduces and synthesizes the emerging interdisciplinary field of LLM Psychometrics, which leverages psychometric instruments, theories, and principles to evaluate, understand, and enhance LLMs.
+The reviewed literature systematically shapes benchmarking principles, broadens evaluation scopes, refines methodologies, validates results, and advances LLM capabilities.
+Diverse perspectives are integrated to provide a structured framework for researchers across disciplines, enabling a more comprehensive understanding of this nascent field.
+Ultimately, the review provides actionable insights for developing future evaluation paradigms that align with human-level AI and promote the advancement of human-centered AI systems for societal benefit.
 A curated repository of LLM psychometric resources is available at https://github.com/valuebyte-ai/Awesome-LLM-Psychometrics.
 1 Introduction
-"Whatever exists at all exists in some amount. To know it thoroughly involves knowing its quantity as well as its quality." [Thorndike, 1962]
 The advent of large language models (LLMs) represents a transformative breakthrough in AI.
 These systems exhibit general-purpose capabilities spanning diverse domains [Bubeck et al., 2023], with particular proficiency in natural language understanding and generation [Demszky et al., 2023, Grossmann et al., 2023, Ziems et al., 2024, Gu et al., 2024].
-They are rapidly integrated into critical societal infrastructure, ranging from consumer-facing applications like chatbots [OpenAI, 2025] and search engines [Wang et al., 2024c] to high-stakes domains such as healthcare [Singhal et al., 2023], education [Milano et al., 2023], and scientific discovery [Romera-Paredes et al., 2024].
+They are rapidly integrated into critical societal infrastructure, ranging from consumer-facing applications like chatbots [OpenAI, 2025] and search engines [Wang et al., 2024c] to high-stakes domains such as healthcare [Singhal et al., 2023], education [Milano et al., 2023], and scientific discovery [Romera-Paredes et al., 2024, Ye et al., 2024].
 Their increasing dominance has exposed a fundamental, pressing scientific challenge: how can we rigorously evaluate these AI systems that transcend traditional benchmarks of biological or algorithmic intelligence?
@@
 However, LLMs have triggered an evaluation crisis, as their versatile capabilities and human-like behaviors exceed what traditional benchmarks can measure.
-Novel challenges include evaluating psychological constructs like personality, values, and cognitive biases that conventional assessment cannot capture.
-Rapid LLM development and training data contamination have made static benchmarks obsolete.
-Additionally, the prompt- and context-sensitive nature of LLMs compromises the robustness and validity of existing evaluation frameworks.
-With increasing human-LLM interaction, human-centered evaluation approaches are becoming essential.
-Moreover, as LLMs integrate into agentic and multimodal systems, evaluation methodologies must expand in scope and complexity.
-These challenges intersect with humanity’s century-old quest to quantify the complex, intangible human psychology, including knowledge, skills, character, values [Pasquali, 2009].
+Novel challenges include but are not limited to (1) evaluating psychological constructs like personality, values, and cognitive biases; (2) obsolescence of static benchmarks due to rapid LLM development and training data contamination; (3) compromised robustness and validity due to LLMs’ prompt- and context-sensitivity; (4) requiring human-centered evaluation approaches; and (5) expanding evaluation methodologies in scope and complexity as LLMs integrate into agentic and multimodal systems.
+These challenges intersect with humanity’s century-old quest to quantify the complex, intangible human psychology [Pasquali, 2009].
 Psychometrics emerged from this timeless pursuit as the scientific study of psychological measurement.
 It bridges the abstract and the empirical by transforming human traits into quantifiable data, enabling better understanding, prediction, and decision-making in education, business, healthcare, governance, and beyond [Rust and Golombok, 2014].
-The convergence of LLMs and psychometrics forms a methodological crucible, forging novel paradigms to better decode and improve machine minds.
+The intersection of LLMs and psychometrics forms a methodological crucible, forging novel paradigms to better decode and improve machine minds.
+This intersection sets the stage for a new research frontier.
 We define LLM Psychometrics as the interdisciplinary field dedicated to evaluating, understanding, and enhancing LLMs through the application and integration of psychometric instruments, theories, and principles.
-This field seeks to quantify, interpret, manipulate, and improve the complex, human-like attributes and behaviors exhibited by LLMs, encompassing both personality constructs (such as personality, values, morality, and attitudes) and cognitive constructs (including heuristics and biases, social interaction abilities, psycholinguistic abilities, and learning and cognitive capacities).
+This field seeks to quantify, interpret, manipulate, and improve the complex, human-like attributes and behaviors exhibited by LLMs, encompassing both personality constructs (e.g., personality, values, morality, and attitudes) and cognitive constructs (e.g., heuristics and biases, social interaction abilities, psycholinguistic abilities, and learning and cognitive capacities).
 Research in LLM Psychometrics applies, extends, and innovates upon scientific methods of psychological measurement for LLMs.
@@
 These studies, however, address a variety of psychological constructs, employ diverse methodologies, and utilize distinct validation techniques.
-The interdisciplinary nature of this domain has attracted contributions from a broad spectrum of academic fields.
-Despite this diversity, there is a lack of cohesion among researchers from different communities, leading to a fragmentation of insights, particularly between studies focusing on disparate constructs.
-Consequently, there is a pressing need for a comprehensive survey to synthesize these efforts and facilitate a more integrated understanding of the field.
+The interdisciplinary nature of this domain has attracted contributions from a broad spectrum of academic communities.
+Despite this diversity, there is a lack of cohesion among researchers, leading to a fragmentation of insights, particularly between studies focusing on disparate constructs.
+Consequently, there is a pressing need for a systematic review to synthesize these efforts and facilitate a more integrated understanding of the field.
 Figure 1: Overview of this review.
-This paper aims to bridge the gap by providing the first systematic review of LLM psychometrics, encompassing evaluation, validation, and enhancement.
-Fig.
-1 illustrates the structure of the review.
+Taxonomy and paper structure.
+To bridge the gap, we systematically review LLM Psychometrics across evaluation, validation, and enhancement.
+The evaluation framework encompasses three core dimensions: the target construct (what to measure), the measurement method (how to measure), and the validation of results (how well do we measure).
+The psychometric insights not only inform evaluation but also guide the development and refinement of LLMs (how to improve).
+Accordingly, this review is structured as follows (Fig.
+1).
 § 2 provides an overview of the preliminaries and methodological foundations to facilitate subsequent discussions.
-In § 3, we delineate the definition, scope, and taxonomy of LLM psychometrics, which establishes the structure of the remainder of the review.
+§ 3 thoroughly contrasts psychometrics and traditional AI benchmarking and reviews how psychometric principles can underpin and reinvent LLM benchmarks.
 The core measurement framework is detailed in § 4, § 5, and § 6.
@@
 Beyond evaluation, § 7 introduces strategies for enhancing LLMs through psychometric insights.
-Finally, § 9 discusses emerging trends, challenges, and future directions in the psychometric evaluation of LLMs, while § 10 concludes the paper.
-Related Surveys.
+§ 8 discusses emerging trends, challenges, and future directions.
+Finally, § 9 highlights key ethical considerations, while § 10 concludes the paper.
+Scope and related work.
 We consider studies to be out of scope if they do not employ psychometric approaches, adhere to psychometric principles, or if they focus solely on scalar performance metrics rather than characterizing behaviors.
 Readers interested in conventional LLM benchmarking are encouraged to consult surveys on LLM evaluation [Guo et al., 2023b, Chang et al., 2024].
-Several related surveys focus on the evaluation of specific constructs in LLMs, such as personality [Wen et al., 2024b, Dong et al., 2025], attitudes and values [Ma et al., 2024a], cultural awareness [Pawar et al., 2024, Adilazuarda et al., 2024], and theory of mind [Sarıtaş et al., 2025, Dong et al., 2025].
+Several related reviews focus on the evaluation of specific constructs in LLMs, such as personality [Wen et al., 2024b, Dong et al., 2025], attitudes and values [Ma et al., 2024a], cultural awareness [Pawar et al., 2024, Adilazuarda et al., 2024], and theory of mind [Sarıtaş et al., 2025, Dong et al., 2025].
 Hagendorff [2023], Hagendorff et al.
 [2024] introduce the notion of machine psychology and review emergent LLM abilities; however, they do not provide comprehensive coverage of the related research, nor do they elaborate on LLM personality constructs, psychometric validation, or enhancement.
-This paper provides the first systematic survey of LLM psychometrics.
+This paper provides the first systematic review of LLM Psychometrics.
 2 Preliminary and methodological foundations
@@
 It was not until the advent of LLMs that the versatility envisioned for "psychometric AI" began to materialize.
-3 LLM Psychometrics: Definition, Scope, and Taxonomy
-We define LLM Psychometrics as the interdisciplinary field dedicated to evaluating, understanding, and enhancing LLMs through the application and integration of psychometric instruments, theories, and principles.
-LLM psychometrics research adapts, advances, and innovates upon psychometric methodologies to rigorously evaluate LLMs.
-This evaluation framework encompasses three core dimensions: the target construct (what to measure), the measurement method (how to measure), and the validation of results (how well do we measure).
-In many cases, these psychometric insights not only inform evaluation but also guide the development and refinement of LLMs (how to improve).
-The remainder of this paper begins by thoroughly contrasting psychometrics and traditional AI benchmarking (§ 4.1), identifying key differences such as core goals, philosophical foundations, measurement methods, and result analysis (Table 1).
-We then review how recent LLM psychometrics research adopts and extends psychometric principles to underpin and reinvent LLM benchmarks (§ 4.2).
-We proceed by examining the principal constructs assessed within LLM psychometrics (what to measure; § 5), including both personality (§ 5.1) and cognitive constructs (§ 5.2).
-Personality constructs are categorized into personality traits, values, morality, and attitudes and opinions, while cognitive constructs comprise heuristics and biases, social interaction abilities, psycholinguistic abilities, and learning and cognitive capacities.
-For each construct, we synthesize the underlying theoretical frameworks and measurement instruments, critically compare alternative approaches, and distill the main empirical findings.
-We further scrutinize the psychometric evaluation methodologies employed for LLMs (how to measure; § 6).
-The methodological framework is systematically analyzed in terms of its key components: test format (§ 6.1), data and task sources (§ 6.2), prompting strategies (§ 6.3), and model output and scoring (§ 6.4).
-In addition, we discuss the role of inference parameters in shaping evaluation results (§ 6.5).
-We next address the psychometric validation of measurement results, examining both reliability and validity (how well do we measure; § 7).
-Reliability is assessed using metrics such as internal consistency, parallel forms reliability, and inter-rater agreement (§ 7.1).
-Validity is considered from content, construct, criterion, and ecological perspectives (§ 7.2).
-Persistent concerns include prompt sensitivity, data contamination, response biases, and divergence from human psychological constructs.
-To mitigate these challenges, recent research advocates for standardized, procedural tests, including multiple task versions, prompt perturbation, and transparent reporting practices, thereby promoting rigorous, reproducible, and meaningful psychometric evaluation of LLMs (§ 7.3).
-The psychometric insights further inform the development and refinement of LLMs (how to improve; § 8).
-Current research predominantly focuses on enhancing LLMs’ performance along three dimensions: trait manipulation (§ 8.1), safety and alignment (§ 8.2), and cognitive enhancement (§ 8.3).
-Finally, we discuss the current trends, challenges, and future directions of LLM psychometrics (§ 9).
 3 Psychometrics for benchmarking principles
+3.1 Fundamental differences between psychometrics and AI benchmarking
+Benchmarking AI systems superficially resembles psychometrics, particularly Classical Test Theory (CTT) [Crocker and Algina, 1986], as both compile test items to evaluate cognitive capabilities and average the resulting scores.
+However, closer examination reveals that AI benchmarks differ significantly from modern psychometric approaches [Wang et al., 2023a, Federiakin, 2025].
+We outline these key differences in Table 1.
 Table 1: Comparison between psychometrics and conventional AI benchmark.
@@
 However, validity, predictive power, or explanatory power beyond the target task is not a primary concern.
-4.1 Fundamental Differences Between Psychometrics and AI Benchmarking
-Benchmarking AI systems superficially resembles psychometrics, particularly Classical Test Theory (CTT) [Crocker and Algina, 1986], as both compile test items to evaluate cognitive capabilities and average the resulting scores.
-However, closer examination reveals that AI benchmarks differ significantly from modern psychometric approaches [Wang et al., 2023a, Federiakin, 2025].
-We outline these key differences in Table 1.
 Psychometrics.
@@
 4 Psychometrics for measuring psychological constructs
+Figure 2: Examples of psychometric tests for LLMs.
 This section delves into the psychological constructs evaluated in LLM Psychometrics. Fig. 2 exemplifies the tests for the involved constructs.
@@
 [2024]
-HEXACO Wang et al. [2025], Bodroža et al. [2024], Miotto et al. [2022], Barua et al. [2024], Ren et al. [2024], Peereboom et al. [2024]
+HEXACO Wang et al. [2025b], Bodroža et al. [2024], Miotto et al. [2022], Barua et al. [2024], Ren et al. [2024], Peereboom et al. [2024]
 MBTI Rao et al.
@@
 [2025b], Münker [2024]
-ETHICS Jinnai [2024], Rodionov et al. [2023], Karpov et al. [2024], Yu et al. [2023], Albrecht et al.
+ETHICS Jinnai [2024], Rodionov et al. [2023], Karpov et al. [2024], Yu et al. [2023], Albrecht et al. [2022]
 DIT Khandelwal et al. [2024], Tanmay et al. [2023]
@@
 Table 3: Personality theories and inventories measured in LLM psychometrics and their main dimensions or focus.
-Figure 2: Examples of psychometric tests for LLMs.
 4.1 Measuring personality constructs
 LLMs exhibit personality constructs that are not explicitly programmed or trained towards.
-The emergent constructs critically shape LLM behavior [Ye et al., 2025a, Hagendorff, 2023], with profound implications for both individuals and broader social groups [Bengio et al., 2024].
+The emergent constructs critically shape LLM behavior [Hagendorff, 2023], with profound implications for both individuals and broader social groups [Bengio et al., 2024].
 Measuring these embedded psychological constructs is essential for understanding behavior, identifying biases, and fostering responsible development.
-Table 2 summarizes the representative personality constructs that have garnered attention in recent research.
+Table 2 and Table 3 summarize the representative personality constructs that have garnered attention in recent research.
 Researchers typically select constructs based on their relevance to LLM development and deployment, as well as the applicability of these constructs to AI systems.
@@
 4.1.1 Personality traits
-"Personality is the enduring configuration of characteristics and behavior that comprises an individual’s unique adjustment to life." [APA Dictionary of Psychology, n.d.]
-Personality traits define an individual’s patterns of thinking, feeling, and behaving [Larsen et al., 2005].
-In the context of LLMs, Zhang et al.
-[2024a], Wang et al.
-[2025] have explored how personality traits relate to model safety, bias, and toxicity.
-Other works emphasize that chatbot personality highly determines user experience [Jiang et al., 2023, Serapio-García et al., 2023, Huang et al., 2024c, Klinkert et al., 2024].
-The study of personality traits leads to several prominent theoretical models, each offering unique insights into individual differences.
-The Big Five model [Goldberg, 2013] identifies five core dimensions of personality: Openness, Conscientiousness, Extraversion, Agreeableness, and Neuroticism.
-Building upon the Big Five, the HEXACO model [Ashton and Lee, 2007] introduces an additional dimension, Honesty-Humility, which encompasses traits such as sincerity, fairness, and modesty.
-Another influential framework, the Myers-Briggs Type Indicator (MBTI) [Myers et al., 1962], categorizes individuals into 16 distinct personality types based on four dichotomous dimensions: Extraversion vs.
-Introversion, Sensing vs.
-Intuition, Thinking vs.
-Feeling, and Judging vs.
-Perceiving.
-In contrast to these models, the Dark Triad [Paulhus and Williams, 2002] focuses on three socially aversive personality traits—Machiavellianism, Narcissism, and Psychopathy—highlighting the darker aspects of human behavior.
-When applying psychometrics to LLMs, most researchers directly administered the established inventories, such as NEO-PI-R [Costa and McCrae, 2008], BFI [John et al., 1991], and BFI-2 [Soto and John, 2017], for measuring the Big Five traits [Caron and Srivastava, 2022, Huang et al., 2023d, a, Karra et al., 2022, La Cava and Tagarelli, 2024, Serapio-García et al., 2023]; HEXACO-60 [Ashton and Lee, 2009] and HEXACO-100 [Lee and Ashton, 2018] inventories for HEXACO traits [Miotto et al., 2022, Wang et al., 2025]; MBTI assessment [Myers, 1985] for MBTI types [Cui et al., 2023, Huang et al., 2023c, La Cava and Tagarelli, 2024, Pan and Zeng, 2023, Rao et al., 2023]; and Dark Triad Dirty Dozen scale [Jonason and Webster, 2010] for Dark Triad traits [Barua et al., 2024, Huang et al., 2023d, Romero et al., 2024].
-Given widespread concerns about the practical relevance of these inventories [Ai et al., 2024, Dorner et al., 2023, Gupta et al., 2024, Pellert et al., 2024, Shu et al., 2024, Song et al., 2023, Serapio-García et al., 2023], Some researchers have adapted them for more real-world scenarios.
+Personality is the enduring configuration of characteristics and behavior that comprises an individual’s unique adjustment to life [APA Dictionary of Psychology, ].
+In the context of LLMs, personality traits relate to model safety, bias, and toxicity [Zhang et al., 2024a, Wang et al., 2025b], and highly determine user experience [Klinkert et al., 2024].
+The study of personality traits leads to several prominent theoretical models, each offering unique insights into individual differences, including the Big Five [Goldberg, 2013], HEXACO [Ashton and Lee, 2007], Myers-Briggs Type Indicator (MBTI) [Myers et al., 1962], and Dark Triad [Paulhus and Williams, 2002].
+When applying psychometrics to LLMs, most researchers directly administered the established inventories, such as NEO-PI-R [Costa and McCrae, 2008], BFI [John et al., 1991], and BFI-2 [Soto and John, 2017], for measuring the Big Five traits; HEXACO-60 [Ashton and Lee, 2009] and HEXACO-100 [Lee and Ashton, 2018] inventories for HEXACO traits; MBTI assessment [Myers, 1985] for MBTI types; and Dark Triad Dirty Dozen scale [Jonason and Webster, 2010] for Dark Triad traits.
+Given widespread concerns about the practical relevance of these inventories [Ai et al., 2024], others have adapted them for more real-world scenarios.
 For example, Bhandari et al.
-[2025a], Frisch and Giulianelli [2024], Klinkert et al.
-[2024], Liu et al.
-[2024a], Zheng et al.
-[2025], Zou et al.
-[2024], Song et al.
-[2024b] contextualize the tests in topic-specific or open-domain conversations.
-Ai et al.
-[2024] accompany the self-report tests with behavioral tests to examine the personality knowledge of LLMs.
-Meanwhile, Jiang et al.
-[2023] compile and adapt existing tests into the Machine Personality Inventory, while Mao et al.
+[2025a] contextualize the tests in topic-specific or open-domain conversations; Ai et al.
+[2024] accompany the self-report tests with behavioral tests to examine the personality knowledge of LLMs; Jiang et al.
+[2023] compile and adapt existing tests into the Machine Personality Inventory; and Mao et al.
 [2024a] develop the PersonalityEdit inventory using LLMs.
 However, Peereboom et al.
-[2024] argue that human-derived traits may not meaningfully apply to LLMs, underscoring the necessity for psychometric theories specifically designed for LLM analysis (§ 7.2.2).
-Comparing Personality Models and Suggesting Use Cases.
-The Big Five offers robust empirical foundations and is the most prominent model for personality assessment.
-However, it is insufficient for predicting comprehensive psychological outcomes [Feher and Vernon, 2021].
-HEXACO extends the model by adding Honesty-Humility and provides nuanced insights into moral traits, making it preferable for LLM application contexts demanding integrity [Wang et al., 2025].
-In contrast, MBTI [Myers et al., 1962] categorizes 16 types via dichotomous scales.
-While popular in career counseling and team-building for its accessibility, its typological approach oversimplifies personality and lacks strong empirical support, limiting its utility in high-stakes decisions [Pittenger, 2005].
-The Dark Triad [Paulhus and Williams, 2002] diverges by focusing on maladaptive traits, making it useful in identifying toxic behaviors in LLMs [Barua et al., 2024].
+[2024] argue that human-derived traits may not meaningfully apply to LLMs, underscoring the necessity for psychometric theories specifically designed for LLM analysis.
 Main findings.
@@
 More advanced LLMs indicate better level of alignment.
-They usually demonstrate high Openness, Extraversion, and Agreeableness traits, while low Neuroticism, on the Big Five personality tests [Zou et al., 2024, Bhandari et al., 2025a, Karra et al., 2022, La Cava and Tagarelli, 2024, Huang et al., 2023e, Klinkert et al., 2024].
+They usually demonstrate high Openness, Extraversion, and Agreeableness traits, while low Neuroticism, on the Big Five personality tests [Zou et al., 2024, Bhandari et al., 2025a].
 The results align with their design as assistive and helpful entities with emotional stability and engaging personality [Bhandari et al., 2025a, La Cava and Tagarelli, 2024].
-Within the MBTI framework, most proprietary LLMs are classified as ENFJ or INFJ types [La Cava and Tagarelli, 2024, Zhang et al., 2024a, Huang et al., 2023c], indicating tendencies toward helpfulness, idealism, and planning capabilities.
+Within the MBTI framework, most proprietary LLMs are classified as ENFJ or INFJ types [La Cava and Tagarelli, 2024, Huang et al., 2023c], indicating tendencies toward helpfulness, idealism, and planning capabilities.
 Personality differences between models are noteworthy.
-Comparative studies reveal that LLMs from different generations and training methodologies display unique combinations of personality characteristics [Bhandari et al., 2025a, Karra et al., 2022, Jiang et al., 2023, La Cava and Tagarelli, 2024, Bodroža et al., 2024, Huang et al., 2023e, Klinkert et al., 2024].
+Comparative studies reveal that LLMs from different generations and training methodologies display unique combinations of personality characteristics [Bhandari et al., 2025a, Karra et al., 2022].
 LLM personalities also vary across contexts.
-The same model may display different personality traits across conversational topics [Zou et al., 2024, Song et al., 2023, Caron and Srivastava, 2022, Bodroža et al., 2024, Petrov et al., 2024].
+The same model may display different personality traits across conversational topics [Zou et al., 2024, Song et al., 2023].
 This variability challenges the applicability of human personality theories that assume relative trait stability.
-Multiple studies suggest that prompt design and system instructions substantially influence LLM personality expression [Song et al., 2023, Caron and Srivastava, 2022, Gupta et al., 2024, Dorner et al., 2023].
-We defer related discussions to § 7.
+Multiple studies suggest that prompt design and system instructions substantially influence LLM personality expression [Song et al., 2023, Caron and Srivastava, 2022].
+Comparing personality models and suggesting use cases.
+The Big Five offers robust empirical foundations and is the most prominent model for personality assessment.
+However, it is insufficient for predicting comprehensive psychological outcomes [Feher and Vernon, 2021].
+HEXACO extends the model by adding Honesty-Humility and provides nuanced insights into moral traits, making it preferable for LLM application contexts demanding integrity [Wang et al., 2025b].
+In contrast, MBTI [Myers et al., 1962] categorizes 16 types via dichotomous scales.
+While popular in career counseling and team-building for its accessibility, its typological approach oversimplifies personality and lacks strong empirical support, limiting its utility in high-stakes decisions [Pittenger, 2005].
+The Dark Triad [Paulhus and Williams, 2002] diverges by focusing on maladaptive traits, making it useful in identifying toxic behaviors in LLMs [Barua et al., 2024].
 4.1.2 Values
-"Values are enduring beliefs that guide behavior and decision-making, reflecting what is important and desirable to an individual or group." [Schwartz, 1992]
-Values are central to understanding human motivation, attitudes, and behavior, shaping how individuals perceive themselves and the world around them.
-According to Schwartz’s theory of basic human values [Schwartz, 2012], the most established value theory, they 1) are beliefs linked inextricably to affect, 2) are desirable goals that motivate actions, 3) transcend specific situations and actions, 4) serve as standards or criteria for evaluating and selecting actions, policies, people, and events, and 5) are ordered by relative importance, which guides choices and actions.
-These unique characteristics make values a powerful lens for understanding LLM behavior.
+Values are enduring beliefs that guide behavior and decision-making, reflecting what is important and desirable to an individual or group [Schwartz, 1992].
+Values offer a powerful lens for understanding LLM behavior.
 For example, Ye et al.
@@
 [2024b] demonstrate that standard value alignment reduces distributional pluralism in LLM outputs.
-Schwartz’s Theory.
-Schwartz [1992] identify 10 basic human values: Self-Direction, Stimulation, Hedonism, Achievement, Power, Security, Conformity, Tradition, Benevolence, and Universalism.
-These values are further grouped into four higher-order dimensions: Openness to Change, Conservation, Self-Enhancement, and Self-Transcendence.
-Later iterations of the theory refine its nuances by subdividing the 10 values into finer categories [Schwartz et al., 2012].
-Most research on Schwartz’s values in LLMs focuses on the 10 basic values, while some studies explore the 19 refined values [Rozen et al., 2024].
-Many investigations directly apply established inventories, i.e., Schwartz Value Survey (SVS) [Schwartz, 1992] or the Portrait Values Questionnaire (PVQ) [Schwartz et al., 2001] (e.g., [Hadar-Shoval et al., 2024, Fischer et al., 2023, Kovač et al., 2023, Lee et al., 2024a, Miotto et al., 2022]).
-Other research adapts these inventories to better suit the context of LLMs.
-For example, Ren et al.
-[2024] rephrase survey items into advice-seeking questions, and Shen et al.
-[2024] transform subjective statements into opinion-seeking questions about AI actions.
-Rather than directly administering or rephrasing static inventories, Ye et al.
-[2025a], Chiu et al.
-[2025], Yao et al.
-[2024] leverage existing cross-domain LLM prompt datasets or LLM-generated contextualized questions as test items.
-World Values Survey (WVS).
-The World Values Survey (WVS) is a global research project that explores people’s values and beliefs, how they change over time, and their social and political impact [Haerpfer et al., 2022].
-It measures a broad spectrum of values, including cultural values defined on two broad dimensions: traditional vs.
-secular-rational values and survival vs.
-self-expression values.
-Kim and Baek [2024] directly prompt LLMs with WVS survey items.
-In contrast, Yao et al.
-[2025a], Chiu et al.
-[2025] use model-generated or self-designed prompts to measure these cultural values in LLMs.
-Values Survey Module (VSM).
-Hofstede [1984] identifies six key value dimensions that describe cultural differences: Power Distance, Individualism vs.
-Collectivism, Masculinity vs.
-Femininity, Uncertainty Avoidance, Long-term vs.
-Short-term Orientation, and Indulgence vs.
-Restraint.
-These dimensions help explain how cultural values influence behavior in societies and organizations [Hofstede, 2001].
-VSM is a questionnaire-based tool that assesses cultural preferences across these dimensions [Hofstede, 2011].
-Kovač et al.
-[2023], Zhong et al.
-[2024] directly apply the VSM to evaluate cultural values in LLMs, while Ren et al.
-[2024], Kharchenko et al.
-[2024], Ye et al.
-[2025a] adapt the VSM or use self-designed inventories to measure these dimensions.
-Global Leadership and Organizational Behavior Effectiveness (GLOBE).
-The GLOBE study investigates cultural values, leadership behaviors, and organizational practices across countries [House et al., 2004].
-Building on Hofstede’s work [Hofstede, 1984], it incorporates additional dimensions: Performance Orientation, Assertiveness, and Humane Orientation, resulting in nine cultural dimensions.
-Karinshak et al.
-[2024], Li et al.
-[2024d], Ren et al.
-[2024] apply or adapt the GLOBE culture questionnaire to evaluate cultural values in LLMs.
-Social Value Orientation (SVO).
-SVO is a psychological framework that measures individual preferences for allocating resources between oneself and others [Messick and McClintock, 1968].
-It focuses on distinguishing between prosocial (e.g., altruists, cooperators) and proself (e.g., individualists, competitors) orientations.
-SVO is typically measured using experimental tasks or questionnaires, such as the SVO Slider Measure [Murphy et al., 2011] or Decomposed Games [Liebrand, 1984].
-These tools assess how individuals allocate resources in hypothetical scenarios involving themselves and others.
-For example, Zhang et al.
-[2024c] apply the SVO Slider Measure to evaluate social preferences in LLMs.
-Other or Custom Theories and Inventories.
-Meadows et al.
-[2024], Xu et al.
-[2023] introduce localized inventories to evaluate LLMs based on Australian and Chinese values, respectively.
-Some researchers focus on specific topics, such as age-related values [Zhang et al., 2024b] and spiritual values [Liu et al., 2024b].
-Other studies involve large-scale, cross-domain prompt datasets [Li et al., 2024d, Moore et al., 2024], among which Jiang et al.
-[2024a] concentrate on eliciting harmful, biased, or ethics-related content from LLMs.
-Notably, Biedma et al.
-[2024], Ye et al.
-[2025b] challenge established value theories and propose novel ones for LLM values, built in a bottom-up manner using psycholexical data.
-Comparing Value Theories and Suggesting Use Cases.
-Schwartz emphasizes universal human motivations; WVS examines societal-level value shifts; VSM and GLOBE address cross-cultural dimensions, with GLOBE emphasizing leadership.
-SVO, in contrast, focuses on individual-level social preferences, i.e., prosocial vs.
-proself.
-While Schwartz and SVO overlap in analyzing individual motivations, the former is broader, and the latter narrowly focuses on decision-making in resource allocation.
-WVS, VSM, and GLOBE share cultural-level dimensions but differ in focus—WVS on modernization, VSM on workplace-related cultural differences, and GLOBE on leadership and organizational practices.
-For evaluating individual values in LLMs, use Schwartz for a comprehensive understanding of universal motivations or SVO for interpersonal decision-making studies.
-In contrast, choose WVS for evaluating broad societal values in LLMs.
-When evaluating cultural dimensions in LLMs, use VSM for workplace cultural differences or GLOBE for leadership and organizational practices.
-In addition, Schwartz’s theory bridges individual and societal-level values, making it versatile for mixed-focus studies.
+A growing body of research has applied diverse value theories and instruments to assess the value orientations of LLMs.
+Schwartz’s Value Theory, with its ten basic values and higher-order dimensions, is the most widely used, often via the Schwartz Value Survey (SVS) or Portrait Values Questionnaire (PVQ) [Schwartz, 1992, Schwartz et al., 2001].
+Adaptations and contextualized prompts are increasingly common to better suit LLMs’ operational context [Ren et al., 2024, Shen et al., 2024].
+The World Values Survey (WVS) has also been used to probe LLMs’ alignment with societal-level value dimensions [Haerpfer et al., 2022, Kim and Baek, 2024].
+Cross-cultural frameworks like Hofstede’s Values Survey Module (VSM) and the GLOBE study extend this analysis to workplace and leadership-related cultural dimensions, with both direct and adapted inventories employed [Hofstede, 1984, House et al., 2004, Ren et al., 2024].
+Social Value Orientation (SVO) frameworks, using tools like the SVO Slider Measure, focus on LLMs’ prosocial versus proself tendencies [Murphy et al., 2011, Zhang et al., 2024c].
+Additionally, researchers have developed localized or custom inventories to capture region-specific or topical values, and some propose novel, bottom-up value taxonomies for LLMs using psycholexical data [Meadows et al., 2024, Xu et al., 2023, Biedma et al., 2024, Ye et al., 2025b].
+Collectively, these studies reveal both the adaptability of human value frameworks to LLMs and the need for tailored approaches.
 Main findings.
@@
 The context-dependency of their values is concerning and challenges the stability assumptions in human value theories [Shen et al., 2024, Meadows et al., 2024, Moore et al., 2024, Chiu et al., 2025, Xu et al., 2023, Kovač et al., 2023].
-Further discussions in reliability and validity are presented in § 7.
+Comparing value theories and suggesting use cases.
+Schwartz emphasizes universal human motivations; WVS examines societal-level value shifts; VSM and GLOBE address cross-cultural dimensions, with GLOBE emphasizing leadership.
+SVO, in contrast, focuses on individual-level social preferences, i.e., prosocial vs.
+proself.
+While Schwartz and SVO overlap in analyzing individual motivations, the former is broader, and the latter narrowly focuses on decision-making in resource allocation.
+WVS, VSM, and GLOBE share cultural-level dimensions but differ in focus—WVS on modernization, VSM on workplace-related cultural differences, and GLOBE on leadership and organizational practices.
+For evaluating individual values in LLMs, use Schwartz for a comprehensive understanding of universal motivations or SVO for interpersonal decision-making studies.
+In contrast, choose WVS for evaluating broad societal values in LLMs.
+When evaluating cultural dimensions in LLMs, use VSM for workplace cultural differences or GLOBE for leadership and organizational practices.
+In addition, Schwartz’s theory bridges individual and societal-level values, making it versatile for mixed-focus studies.
 4.1.3 Morality
-"Morality is the categorization of intentions, decisions and actions into those that are proper, or right, and those that are improper, or wrong." [Long and Sedley, 1987]
-Morality is a fundamental aspect of human behavior, influencing social interactions, decision-making, and ethical reasoning.
-The Moral Foundations Theory (MFT) [Graham et al., 2009] posits that morality is shaped by six innate psychological systems: Care/Harm, Fairness/Cheating, Loyalty/Betrayal, Authority/Subversion, Sanctity/Degradation, and Liberty/Oppression.
-These moral foundations are hypothesized to be universal across cultures, providing a common framework for understanding moral judgments and behavior.
-The theory suggests that individuals vary in the extent to which they endorse each foundation, leading to diverse moral profiles that influence their attitudes, beliefs, and actions.
+Morality is the categorization of intentions, decisions and actions into those that are proper, or right, and those that are improper, or wrong [Long and Sedley, 1987].
 It is crucial to conduct moral assessments of LLMs to ensure their ethical deployment.
-A large body of research (e.g., [Tlaie, 2024, Liu et al., 2024d, Fraser et al., 2022, Yao et al., 2025a, Zhou et al., 2024a]) applies MFT, primarily through the Moral Foundations Vignettes (MFVs) [Clifford et al., 2015], the Moral Foundations Questionnaire (MFQ) [Graham et al., 2009], the MFQ-2 [Atari et al., 2023], and the Moral Foundations Dictionary (MFD) [Graham et al., 2009].
+A large body of research applies the Moral Foundations Theory (MFT) [Graham et al., 2009], primarily through the Moral Foundations Vignettes (MFVs) [Clifford et al., 2015], the Moral Foundations Questionnaire (MFQ) [Graham et al., 2009], the MFQ-2 [Atari et al., 2023], and the Moral Foundations Dictionary (MFD) [Graham et al., 2009].
 These papers investigate model bias, alignment with political/moral ideologies, and variation in personal or cultural values, using controlled prompt tests [Tlaie, 2024, Nunes et al., 2024, Abdulhai et al., 2024], persona-driven exploration [Münker, 2024], or evaluating internal moral coherence [Nunes et al., 2024].
-Other prominent moral theoretical frameworks and survey tools for evaluating LLMs include Kohlberg’s Theory via the Defining Issues Test (DIT) [Kohlberg, 1964], the Consequentialist-Deontological distinction [Beauchamp, 2001], and the PEW 2013 Global Attitudes Survey [Center, 2013].
-Khandelwal et al.
-[2024], Tanmay et al.
-[2023] utilize the DIT, a psychometric tool aligned with Kohlberg’s stages of moral development, to analyze structural reasoning in moral dilemmas.
-They report that models like GPT-4 exhibit post-conventional moral reasoning.
-Khandelwal et al.
-[2024] further extend this analysis to a multilingual context, examining cross-linguistic variations in moral reasoning abilities.
-Neuman et al.
-[2025] explore ethical reasoning using established typologies, including consequentialist-deontological analysis.
-The results uncover the general rationalist, consequentialist emphasis in LLMs.
-According to the PEW 2013 Global Attitudes Survey, Meijer et al.
-[2024] find that the models display homogeneous moral values and have limited effectiveness in reflecting cross-cultural morality.
-Localized moral theoretical frameworks are essential for ensuring the ethical deployment of LLMs in various cultural contexts.
-Liu et al.
-[2024d] develop the Chinese Moral Dictionary (CMD) and find that ChatGPT and Gemini prefer individualistic moral beliefs, while Ernie and ChatGLM lean towards collectivist moral beliefs.
-Takeshita et al.
-[2023], Ohashi et al.
-[2024] create the JCommonsenseMorality (JCM) dataset and fine-tune LLMs to adapt to Japanese culture.
-Researchers have also developed specialized datasets for studying LLM morality.
-Hendrycks et al.
-[2021] create the ETHICS dataset, which is then utilized and extended to evaluate LLMs’ knowledge of morality, covering concepts such as justice, well-being, duties, virtues, and commonsense morality [Jinnai, 2024, Rodionov et al., 2023, Karpov et al., 2024, Yu et al., 2023, Albrecht et al., ].
-Marraffini et al.
-[2024] develop the Greatest Good Benchmark (GGB) to evaluate LLMs’ moral judgments using utilitarian dilemmas.
-Jin et al.
-[2024b] develop the MultiTP, a cross-lingual corpus of moral dilemma vignettes in over 100 languages.
+Other prominent moral theories and instruments include Kohlberg’s Theory via the Defining Issues Test (DIT) [Kohlberg, 1964], the Consequentialist-Deontological distinction [Beauchamp, 2001], the PEW 2013 Global Attitudes Survey [Center, 2013], and localized moral theories [Liu et al., 2024d, Takeshita et al., 2023].
+Specialized datasets are also curated for scalable and comprehensive morality evaluation [Hendrycks et al., 2021, Jinnai, 2024, Marraffini et al., 2024, Jin et al., 2024b]
+Main findings.
+LLMs are generally characterized by a rationalist and consequentialist focus, often prioritizing harm minimization and fairness [Neuman et al., 2025].
+Despite that, they show divergence in ethical reasoning [Neuman et al., 2025] and moral preferences [Tanmay et al., 2023, Jin et al., 2024b, Bonagiri et al., 2024, Meijer et al., 2024].
+In some aspects, most LLMs align with human moral standards [Tanmay et al., 2023, Nunes et al., 2024, Takemoto, 2024], which may be attributed to their extensive exposure to conventional ethical values during training.
+On the other hand, some research presents a less optimistic view, uncovering significant deviations of LLMs from human moral preferences [Marraffini et al., 2024, Vida et al., 2024, Ahmad and Takemoto, 2024].
+Regarding the underlying mechanisms of LLMs’ moral reasoning, Nunes et al.
+[2024], Ji et al.
+[2024], Simmons [2023] suggest that LLMs primarily exhibit imitation rather than genuine conceptual understanding.
 Comparing moral theories/instruments and suggesting use cases.
@@
 Meanwhile, datasets like ETHICS provide morally charged scenarios for detecting normative behavior and fine-tuning LLM responses in high-stakes ethical domains [Hendrycks et al., 2021, Jin et al., 2024b, Marraffini et al., 2024].
-Main Findings.
-LLMs are generally characterized by a rationalist and consequentialist focus, often prioritizing harm minimization and fairness [Neuman et al., 2025].
-Despite that, they show divergence in ethical reasoning [Neuman et al., 2025] and moral preferences [Tanmay et al., 2023, Jin et al., 2024b, Bonagiri et al., 2024, Meijer et al., 2024].
-In some aspects, most LLMs align with human moral standards [Tanmay et al., 2023, Nunes et al., 2024, Takemoto, 2024], which may be attributed to their extensive exposure to conventional ethical values during training.
-On the other hand, some research presents a less optimistic view, uncovering significant deviations of LLMs from human moral preferences [Marraffini et al., 2024, Vida et al., 2024, Ahmad and Takemoto, 2024].
-Regarding the underlying mechanisms of LLMs’ moral reasoning, Nunes et al.
-[2024], Ji et al.
-[2024], Simmons [2023] suggest that LLMs primarily exhibit imitation rather than genuine conceptual understanding.
-5.1.4 Attitudes & Opinions
-"Attitudes are always attitudes about something.
+4.1.4 Attitudes and opinions
+Attitudes are always attitudes about something.
 This implies three necessary elements: first, there is the object of thought, which is both constructed and evaluated.
@@
 Third, there is the agent, who is doing the constructing and evaluating.
-We can therefore suggest that, at its most general, an attitude is the cognitive construction and affective evaluation of an attitude object by an agent." [Bergman, 1998]
-We use the term "attitude" to encompass both attitudes and opinions, following [Bergman, 1998, Ma et al., 2024a], who treat these concepts as synonymous.
+We can therefore suggest that, at its most general, an attitude is the cognitive construction and affective evaluation of an attitude object by an agent [Bergman, 1998].
+We use the term "attitude" to encompass both attitudes and opinions, following [Bergman, 1998, Ma et al., 2024a].
 Most research on LLM attitudes examines political attitudes and public opinions [Ma et al., 2024a], which are key cognitive and behavioral foundations in human society and closely tied to model fairness, credibility, and social impact [Durmus et al., 2023, Santurkar et al., 2023, Sanders et al., 2023, Lee et al., 2024c, Hartmann et al., 2023].
-Other biases—such as those related to gender, race, culture, religion, occupation and age—are also studied as part of LLM attitudes, though many works approach them from data and algorithmic perspectives.
 For a broader discussion on bias and fairness in LLMs, we refer readers to [Gallegos et al., 2024, Ranjan et al., 2024].
-To measure the political attitudes of LLMs, researchers often employ standardized questionnaires and scales from political science and social psychology.
-A widely used tool is the American National Election Studies (ANES) [^1] https://electionstudies.org/about-us/., applied in works such as Qi et al.
-[2024], Jiang et al.
-[2022, 2024d], Yang et al.
-[2024] to assess LLMs’ stances on U.S.
-political issues.
-Similarly, the American Trends Panel (ATP) [^2] https://www.pewresearch.org/american-trends-panel/.
-has been leveraged by Santurkar et al.
-[2023], Hwang et al.
-[2023], Tjuatja et al.
-[2024] to analyze opinion distributions on public policy.
-For cross-national comparisons, studies like Ma et al.
-[2024b], Ball et al.
-[2025], von der Heyde et al.
-[2024] utilize the German Longitudinal Election Study (GLES) [^3] https://www.gesis.org/en/gles..
-Additionally, the Political Compass Test (PCT) [^4] https://www.politicalcompass.org/test.
-has gained prominence for positioning LLMs within multidimensional political spectrums [Azzopardi and Moshfeghi, 2024, Bernardelle et al., 2024, Hartmann et al., 2023, Röttger et al., 2024].
-Other survey instruments include the General Social Survey (GSS) [Kim and Lee, 2023], American Community Survey (ACS) [Dominguez-Olmedo et al., 2024], Canadian Election Study (CES) [Sanders et al., 2023], European Social Survey (ESS) [Geng et al., 2024], Survey of Russian Elites [Kalinin, 2023], and Supreme Court Case Political Evaluation (SCOPE) [Xu et al., 2025c].
-Additionally, researchers have developed specialized datasets for LLMs, such as OpinionQA [Santurkar et al., 2023], a comprehensive dataset based on ATP surveys; IssueBench [Röttger et al., 2025], a benchmark covering controversial issues with multiple response formats; and GlobalOpinionQA [Durmus et al., 2023], which extends political opinion evaluation to cross-cultural contexts.
-Researchers also present tailored tools for LLMs, due to the complexity of measuring political attitudes.
-Haller et al.
-[2023] introduce OpinionGPT, a web tool demonstrating how input data biases influence model outputs.
-Chalkidis and Brandl [2024] fine-tune LLMs on European Parliament speeches to assess political knowledge and reasoning, realigning the models to reflect specific stances.
-Coppolillo et al.
-[2025] propose a framework for quantifying biases in multi-agent generative systems by simulating echo chambers—paired LLMs with aligned perspectives discuss polarizing topics.
-Finally, Kim et al.
-[2025b] investigate linear interpolation methods to track continuous shifts in model political attitudes.
-Comparing Attitude Measurement Tools and Suggesting Use Cases.
-The ANES offers the most comprehensive framework for American political contexts, making it ideal for studying model performance on U.S.
-social issues [Jiang et al., 2022].
-Cross-national surveys and datasets such as the ESS, GLES, and GlobalOpinionQA provide broader cultural perspectives across diverse sociopolitical environments [Geng et al., 2024, Ball et al., 2025, Ma et al., 2024b, von der Heyde et al., 2024, Durmus et al., 2023].
-The ATP is better suited for researching model responses to social controversies [Santurkar et al., 2023], while the PCT serves as a theory-driven political spectrum measurement tool, offering simplified yet intuitive insights into LLM political positioning [Röttger et al., 2024].
-Other surveys have distinct regional and sociopolitical focuses: the CES examines political participation in North American multicultural democracies [Sanders et al., 2023]; Survey of Russian Elites provides insights into non-Western political perspectives [Kalinin, 2023]; the SCOPE delivers in-depth assessments of judicial attitudes [Xu et al., 2025c]; and the ACS serves as a tool for researching socioeconomic issues and group representation [Dominguez-Olmedo et al., 2024].
+Researchers assess LLMs’ political attitudes using standardized questionnaires and scales from political science and social psychology.
+US-focused instruments include the American National Election Studies (ANES) [Qi et al., 2024] and the American Trends Panel (ATP) [Santurkar et al., 2023].
+For cross-national comparisons, studies use tools like the German Longitudinal Election Study (GLES) [Ma et al., 2024b, Ball et al., 2025, von der Heyde et al., 2024].
+The Political Compass Test (PCT) is also widely used to map LLMs within multidimensional political spectrums [Azzopardi and Moshfeghi, 2024, Bernardelle et al., 2024, Hartmann et al., 2023, Röttger et al., 2024].
+Other survey instruments include the General Social Survey (GSS) [Kim and Lee, 2023], the American Community Survey (ACS) [Dominguez-Olmedo et al., 2024], the Canadian Election Study (CES) [Sanders et al., 2023], the European Social Survey (ESS) [Geng et al., 2024], the Survey of Russian Elites [Kalinin, 2023], and the Supreme Court Case Political Evaluation (SCOPE) [Xu et al., 2025c].
+Additionally, researchers have developed specialized datasets and tailored tools for LLMs, such as OpinionQA [Santurkar et al., 2023], a comprehensive dataset based on ATP surveys; IssueBench [Röttger et al., 2025], a benchmark covering controversial issues with multiple response formats; GlobalOpinionQA [Durmus et al., 2023], which extends political opinion evaluation to cross-cultural contexts; and OpinionGPT [Haller et al., 2023], a web tool that demonstrates how input data biases influence model outputs.
 Main findings.
@@
 Such advancements indicate promising directions for using LLMs in opinion research, though their current limitations and biases must still be acknowledged.
+Comparing attitude measurement tools and suggesting use cases.
+The ANES offers the most comprehensive framework for American political contexts, making it ideal for studying model performance on U.S.
+social issues [Jiang et al., 2022].
+Cross-national surveys and datasets such as the ESS, GLES, and GlobalOpinionQA provide broader cultural perspectives across diverse sociopolitical environments [Geng et al., 2024, Ball et al., 2025, Ma et al., 2024b, von der Heyde et al., 2024, Durmus et al., 2023].
+The ATP is better suited for researching model responses to social controversies [Santurkar et al., 2023], while the PCT serves as a theory-driven political spectrum measurement tool, offering simplified yet intuitive insights into LLM political positioning [Röttger et al., 2024].
+Other surveys have distinct regional and sociopolitical focuses: the CES examines political participation in North American multicultural democracies [Sanders et al., 2023]; Survey of Russian Elites provides insights into non-Western political perspectives [Kalinin, 2023]; the SCOPE delivers in-depth assessments of judicial attitudes [Xu et al., 2025c]; and the ACS serves as a tool for researching socioeconomic issues and group representation [Dominguez-Olmedo et al., 2024].
+4.2 Measuring cognitive constructs
+Traditional NLP benchmarks are insufficient to capture and analyze emergent cognitive constructs [Hagendorff et al., 2024, Ying et al., 2025a].
+To address this gap, researchers are adapting related psychometric techniques to evaluate LLM abilities.
+Hagendorff et al.
+[2024] introduce the concept of machine psychology, reviewing emergent LLM capabilities and advocating for psychometric evaluation.
+They categorize these cognitive constructs into four key aspects: heuristics and biases, social interactions, psychology of language, and learning and cognitive capabilities.
+Structured based on this taxonomy, Table 4 summarizes the representative cognitive constructs measured in recent research.
 Aspects Sub-aspects Related work
@@
 [2024, 2021], Mathur et al.
+[2024], Schoenegger et al.
 [2025]
@@
 Table 4: Representative LLM cognitive constructs in LLM psychometrics.
-5.2 Measuring Cognitive Constructs
-Traditional NLP benchmarks are insufficient to capture and analyze emergent cognitive constructs [Hagendorff et al., 2024, Ying et al., 2025a].
-To address this gap, researchers are adapting related psychometric techniques to evaluate LLM abilities.
-Hagendorff et al.
-[2024] introduce the concept of machine psychology, reviewing emergent LLM capabilities and advocating for psychometric evaluation.
-They categorize these cognitive constructs into four key aspects: heuristics and biases, social interactions, psychology of language, and learning and cognitive capabilities.
-Structured based on this taxonomy, Table 4 summarizes the representative cognitive constructs measured in recent research.
 4.2.1 Heuristics and biases
-"Heuristics and biases are mental shortcuts or rules of thumb that simplify decision-making and problem-solving." [Tversky and Kahneman, 1974]
-Heuristics shape both the strengths and biases in human decision-making.
-Recent research employs psychometric tools to systematically evaluate rationality and biases in LLM outputs and offer theoretical explanations for these biases.
-Pioneering work by Binz and Schulz [2023] applies a battery of canonical cognitive ability tests to GPT-3.
-The test results demonstrate that the model performs on par with or better than human participants, while also exhibiting human-like cognitive biases such as framing effects, certainty effects, and overweighting biases.
-Subsequent studies include more advanced LLMs and reveal a dual-process dynamic [Hagendorff et al., 2023]: earlier GPT models exhibit more human-like System 1 errors, while more advanced models such as GPT-4 display more hyperrational System 2 reasoning and thus fewer biases.
-In a similar vein, but with an expanded scope, Yax et al.
-[2024], Sundaram and Alwar [2024], Momennejad et al.
-[2023], Echterhoff et al.
-[2024], Saeedi et al.
-[2024], Ranaldi and Zanzotto [2024], Macmillan-Scott and Musolesi [2024], Ando et al.
-[2023] evaluate LLMs regarding biases including the conjunction fallacy, unwarranted beliefs, anchoring bias, status-quo bias, loss aversion, positional selection biases, and Clever Hans effects.
-Later work significantly scales up the evaluation with synthetic items.
-Malberg et al.
-[2024] evaluate cognitive biases in 20 LLMs based on novel decision-making scenarios and a dataset of 30,000 tests covering 30 cognitive biases.
-Xie et al.
-[2024b] introduce the "MindScope" dataset and multi-agent methods to evaluate 72 bias categories using psychometric-inspired multi-round dialogue experiments.
-Recently, the evaluation paradigm has also been extended to multimodal models [Schulze Buschoff et al., 2025].
-To understand the origins of these biases, researchers have proposed diverse theoretical frameworks.
-Hagendorff et al.
-[2023], Yax et al.
-[2024] link dual-process theory (System 1 vs.
-System 2 thinking) with the contextualized inference of LLMs, contrasting intuitive and deliberative reasoning modes.
-Scholten et al.
-[2024] explain LLM biases using metacognitive myopia, a syndrome described as being "relatively accurate in utilizing even large amounts of stimulus information, but naïve and almost blind regarding the history and validity of the stimulus data." The authors argue that LLMs lack metacognitive abilities such as monitoring and control, which result in five symptoms of metacognitive myopia and, therefore, systematic biases.
-Sundaram and Alwar [2024] explore cognitive dissonance theory and elaboration likelihood theory to explain inconsistencies in LLM reasoning.
-They suggest that certain biases in LLMs may not just be drawbacks but could serve as a means of identifying and mitigating logical fallacies in human reasoning.
-Differing from the cognitive psychology perspective, Shaikh et al.
-[2024] propose a mechanistic interpretability framework for sourcing the biases.
-They leverage influence graphs and Shapley value analysis to interpret biases and explore connections between model training and bias emergence.
+Heuristics and biases are mental shortcuts that simplify decision-making but can introduce systematic errors [Tversky and Kahneman, 1974].
+Recent research employs psychometric tools to evaluate rationality and biases in LLM outputs and offer theoretical explanations for these biases.
+Early studies, such as Binz and Schulz [2023], found that GPT-3 performs comparably to humans on cognitive ability tests, exhibiting biases like framing and anchoring.
+More advanced models, such as GPT-4, show fewer System 1 (intuitive) errors and more System 2 (deliberative) reasoning, though biases persist [Hagendorff et al., 2023].
+Large-scale evaluations now cover a wide range of biases—including conjunction fallacy, unwarranted beliefs, and loss aversion—using synthetic scenarios and multi-agent dialogue [Malberg et al., 2024, Xie et al., 2024b].
+Multimodal models are also being assessed for similar patterns [Schulze Buschoff et al., 2025].
+To explain these findings, researchers draw on dual-process theory, linking LLM reasoning to intuitive and deliberative modes [Hagendorff et al., 2023, Yax et al., 2024].
+Other frameworks highlight LLMs’ lack of metacognitive abilities, leading to “metacognitive myopia”—the inability to monitor or control reasoning processes, which perpetuates systematic biases [Scholten et al., 2024].
+Additional perspectives use cognitive dissonance and elaboration likelihood theories to interpret inconsistencies [Sundaram and Alwar, 2024], while mechanistic approaches leverage influence graphs and Shapley values to trace the origins of biases in model training [Shaikh et al., 2024].
 Main findings.
@@
 Related evaluation focuses on interconnected dimensions such as Theory of Mind (ToM), Emotional Intelligence (EI), and Social Intelligence (SI).
-"Theory of Mind is the ability to attribute mental states such as beliefs, intentions, and knowledge to others." [Premack and Woodruff, 1978]
 Theory of Mind (ToM).
+ToM is the ability to attribute mental states such as beliefs, intentions, and knowledge to others [Premack and Woodruff, 1978].
 Advanced LLMs can simulate ToM-like reasoning under certain conditions, which prompts questions about how these behaviors arise and how robustly they generalize.
@@
 Ullman [2023], Shapira et al.
-[2023a], Holterman and Deemter [2023] soon demonstrate that minor task perturbations often lead to poor performance, indicating reliance on brittle heuristics rather than genuine semantic comprehension.
-Hoping to address the controversy, researchers have developed robust benchmarks and evaluation protocols.
-For example, BigToM [Gandhi et al., 2023], ToMBench [Chen et al., 2024d], HI-TOM [He et al., 2023], OpenToM [Xu et al., 2024a], and ToMATO [Shinoda et al., 2025] feature procedurally generated, higher-order, or broader ToM tasks.
-These benchmarks often reveal performance drop-offs in more complex tasks, such as 6th-order belief attribution [Street et al., 2024] or detecting faux pas [Strachan et al., 2024a].
-But others argue the failures result from a lack of more general commonsense reasoning, rather than a failure to represent mental states [Pi et al., 2024].
+[2023a], Holterman and Deemter [2023] show that minor task changes often cause LLMs’ performance to drop, suggesting reliance on brittle heuristics rather than true semantic understanding.
+To address this, researchers have developed more robust benchmarks and evaluation protocols featuring procedurally generated, higher-order, or broader ToM tasks [Gandhi et al., 2023, Chen et al., 2024d, He et al., 2023, Xu et al., 2024a, Shinoda et al., 2025].
+These benchmarks frequently reveal significant declines in performance on complex tasks, such as 6th-order belief attribution [Street et al., 2024] or faux pas detection [Strachan et al., 2024a].
+Some argue these failures stem from a lack of general commonsense reasoning rather than an inability to represent mental states [Pi et al., 2024].
 Riemer et al.
-[2025] also argue that we should distinguish between literal and functional ToM, and that current ToM benchmarks are inadequate for assessing functional ToM.
-Recent attempts also extend ToM evaluation to multimodal settings [Chen et al., 2024c, Strachan et al., 2024b, Jin et al., 2024a] and multi-agent interactions [Li et al., 2023b].
-Several studies explore enhancement techniques.
-Prompt engineering methods like reasoning, reflection, and planning [Moghaddam and Honey, 2023, Lin et al., 2024, Zhou et al., 2023a, Yang et al., 2025a, Zhang et al., 2025]; persona-based prompts [Tan et al., 2024]; and psychology-informed prompts [Leer et al., 2023, Wilf et al., 2023] can significantly increase accuracy but raise questions about the authenticity of improvements.
-Others implement symbolic reasoning frameworks, such as SymbolicToM [Sclar et al., 2023, Xu et al., 2025a], ToM-LM [Tang and Belle, 2024], or task decomposition strategies [Sarangi et al., 2025], to enhance interpretability and belief state modeling.
-While recent models keep progressing in ToM tasks, especially in structured tasks [Strachan et al., 2024a, Street et al., 2024], they are shown to remain inconsistent in open-ended, adversarial, or pragmatic reasoning settings [Sclar et al., 2024, Yu et al., 2025, Nickel et al., 2024].
-Comparisons with human baselines show that LLMs can approximate ToM behavior in narrow contexts but fall short of general human-like social cognition [Strachan et al., 2024a, van Duijn et al., 2023, Jones et al., 2024].
-Main Findings.
+[2025] further suggest distinguishing between literal and functional ToM, noting that current benchmarks inadequately assess functional ToM.
+While newer models continue to improve on structured ToM tasks [Strachan et al., 2024a, Street et al., 2024], they remain inconsistent in open-ended, adversarial, or pragmatic reasoning scenarios [Sclar et al., 2024, Yu et al., 2025, Nickel et al., 2024].
+Comparisons with human baselines show that LLMs can approximate ToM behavior in narrow contexts but still fall short of general human-like social cognition [Strachan et al., 2024a, van Duijn et al., 2023, Jones et al., 2024].
+Recent work has also extended ToM evaluation to multimodal settings [Chen et al., 2024c, Strachan et al., 2024b, Jin et al., 2024a] and multi-agent interactions [Li et al., 2023b].
+Main findings in ToM.
 LLMs exhibit psychometrically measurable ToM-like reasoning, especially when appropriately prompted or structured, but current evidence suggests these capabilities depend on surface-level linguistic cues and lack robustness.
@@
 Interested readers are referred to recent comprehensive reviews on ToM in LLMs [Sarıtaş et al., 2025, Ma et al., 2023, Mao et al., 2024b].
-"Emotional Intelligence is the subset of social intelligence that involves the ability to monitor one’s own and others’ feelings and emotions, to discriminate among them and to use this information to guide one’s thinking and actions." [Salovey and Mayer, 1990]
 Emotional Intelligence (EI).
-Recent studies introduce benchmark tools specifically designed to measure EI in LLMs through structured tasks informed by psychological theories.
-For example, EmoBench [Sabour et al., 2024] and EQ-Bench [Paech, 2023] provide task sets operationalizing constructs like emotional understanding and regulation, grounding their evaluations in frameworks such as the Mayer-Salovey-Caruso Emotional Intelligence Test (MSCEIT).
-Others adapt existing psychometric instruments to LLM psychometrics, including the Situational Evaluation of Complex Emotional Understanding (SECEU) [Wang et al., 2023b], the Levels of Emotional Awareness Scale (LEAS) [Elyoseph et al., 2023], the Toronto Alexithymia Scale (TAS-20) and Empathy Quotient (EQ-60) [Patel and Fan, 2023], and others [Huang et al., 2024b].
-Models such as GPT-4 are shown to match or exceed human baseline in measures of emotional awareness and understanding [Wang et al., 2023b, Elyoseph et al., 2023, Patel and Fan, 2023], though it has been reported that LLMs lack deep reflexive analysis of emotional experiences and motivational aspects of emotions [Vzorinab et al., 2024].
-Recent benchmark tools also extend EI evaluations to multimodal settings [Hu et al., 2025].
-A subset of the literature further explores methodological pathways for improving LLM emotional competencies.
-Modular Emotional Intelligence (MoEI) [Zhao et al., 2024] and Emotional Chain-of-Thought (ECoT) [Li et al., 2024e] offer strategies to enhance emotionally intelligent behaviors while preserving general cognitive functionality.
-Another line of research explores LLMs’ ability to simulate nuanced emotional alignment, empathy, and scenario-appropriate reactions, based on Emotion Appraisal Theory [Roseman and Smith, 2001] and Emotion-Focused Theory [Greenberg, 2004].
-Surprisingly, Li et al.
-[2023a] indicate that emotional stimuli improve general performance of LLMs.
-Main Findings.
+EI is the subset of social intelligence that involves the ability to monitor one’s own and others’ feelings and emotions, to discriminate among them and to use this information to guide one’s thinking and actions [Salovey and Mayer, 1990].
+Recent work introduces benchmarks to assess EI in LLMs using structured, theory-driven tasks.
+Tools like EmoBench [Sabour et al., 2024] and EQ-Bench [Paech, 2023] operationalize constructs such as emotional understanding and regulation, often referencing frameworks like Mayer-Salovey-Caruso Emotional Intelligence Test.
+Other studies adapt established psychometric instruments, including Situational Evaluation of Complex Emotional Understanding (SECEU) [Wang et al., 2023b], the Levels of Emotional Awareness Scale (LEAS) [Elyoseph et al., 2023], the Toronto Alexithymia Scale (TAS-20) and Empathy Quotient (EQ-60) [Patel and Fan, 2023], among others [Huang et al., 2024b].
+Advanced models like GPT-4 often match or surpass human baselines in emotional awareness and understanding [Wang et al., 2023b, Elyoseph et al., 2023, Patel and Fan, 2023], though they lack deep reflexive analysis of emotional experience and motivation [Vzorinab et al., 2024].
+Recent benchmarks also extend EI evaluation to multimodal settings [Hu et al., 2025].
+Main findings in EI.
 Advanced LLMs generally perform on par with or better than humans in tasks of EI [Elyoseph et al., 2023, Patel and Fan, 2023], despite evident limitations in several areas, such as displaying artificial or mechanical patterns when expressing empathy [Lee et al., 2024f], the deep reflexive analysis of emotional experiences [Vzorinab et al., 2024], and the misalignment with human emotional behaviors [Huang et al., 2023b].
-"Social Intelligence is the ability to understand and manage people." [Thorndike and Stein, 1937]
 Social Intelligence (SI).
-Measuring SI of LLMs seeks to understand how these systems interpret and respond to human social situations.
-The evaluation involves rule-based decision-making, conformity to social norms, and successful participation in structured interpersonal scenarios.
-Several studies have directly applied psychometric instruments to address this objective and present optimistic results.
+SI is the ability to understand and manage people [Thorndike and Stein, 1937].
+SI in LLMs determines how well these models interpret and respond to social situations.
+Psychometric tools applied to LLMs yield mixed results.
 Mittelstädt et al.
-[2024] utilize standardized Situational Judgment Tests (SJTs) to compare LLMs’ assessments of social appropriateness with those of human participants, finding that some models outperform humans under expert-rated conditions.
-Other research reveals LLMs still struggle in some aspects of SI.
-Shapira et al.
-[2023b] investigate the ability of LLMs using faux pas tests from clinical psychology, finding that LLMs struggle to describe social situations implicitly.
-Based on Daniel Goleman’s SI theory, Xu et al.
-[2024b] introduce the Situational Evaluation of SI (SESI).
-It shows that LLMs are still limited in their SI, with superficial friendliness as a primary reason for errors.
-AgentSense benchmark employs interactive scenarios and highlights the limitations of LLMs in handling complex social interactions, especially high-level growth needs and private information reasoning [Mou et al., 2024].
-Later studies further explore multi-agent environments.
-Interactive multi-agent environments such as SOTOPIA [Zhou et al., 2024b] and its extended training variant SOTOPIA-$\pi$ [Wang et al., 2024d] simulate cooperative, competitive, and norm-driven social contexts.
-The STSS benchmark proposed by [Wang et al., 2024a] further operationalizes SI through task-oriented simulations, focusing on agent action outcomes and goal achievements rather than linguistic performance.
-The CogMir framework [Liu et al., 2024c] focuses on evaluating prosocial but irrational LLM behaviors.
-It reveals that LLM agents and humans exhibit high consistency in irrational and prosocial decision-making under uncertain conditions.
-[Guo et al., 2023a] refine the Social-IQ dataset to create DeSIQ, a benchmark for SI evaluation that extends to multimodal settings.
-Critical perspectives are proposed by Kovač et al.
-[2024] who argue the current benchmarks lack grounding in developmental psychology and offer a versatile framework, the SocialAI school, for studying the SI of LLM-based agents.
-Main Findings.
+[2024] use Situational Judgment Tests (SJTs) and find some LLMs outperform humans in expert-rated social appropriateness.
+However, LLMs still struggle with implicit social understanding: Shapira et al.
+[2023b] show LLMs perform poorly on faux pas tests, and Xu et al.
+[2024b] find superficial friendliness often leads to errors in the Situational Evaluation of SI (SESI).
+The AgentSense benchmark further highlights LLMs’ limitations in complex social interactions, especially regarding high-level needs and private information [Mou et al., 2024].
+Recent work expands to multi-agent environments.
+SOTOPIA [Zhou et al., 2024b] and SOTOPIA-$\pi$ [Wang et al., 2024d] simulate cooperative and competitive social contexts, while the STSS benchmark [Wang et al., 2024a] evaluates SI through task-oriented simulations focused on outcomes and goals.
+The CogMir framework [Liu et al., 2024c] finds LLMs and humans are similarly consistent in irrational and prosocial decisions under uncertainty.
+DeSIQ [Guo et al., 2023a] extends SI evaluation to multimodal settings.
+Critically, Kovač et al.
+[2024] argue current benchmarks lack developmental grounding and propose the SocialAI school as a more comprehensive framework for studying LLM-based SI.
+Main findings in SI.
 LLMs show measurable competence in rule-based, socially appropriate behavior, particularly in structured environments.
-They perform well on: 1) following predefined social norms [Mittelstädt et al., 2024]; 2) completing interactional goals in multi-agent settings [Zhou et al., 2024b, Wang et al., 2024d]; and 3) replicating human-like prosocial decisions [Liu et al., 2024c].
+They perform well on: 1) following predefined social norms [Mittelstädt et al., 2024]; 2) completing interactional goals in multi-agent settings [Zhou et al., 2024b, Wang et al., 2024d]; 3) replicating human-like prosocial decisions [Liu et al., 2024c]; and 4) being persuasive, even better than human persuaders [Schoenegger et al., 2025].
 Weaker performance is noted in tasks where superficial friendliness causes errors [Xu et al., 2024b], understanding high-level growth needs is essential [Mou et al., 2024], and LLMs must implicitly describe social situations [Shapira et al., 2023b].
@@
 Language comprehension.
-Evaluations of LLM language comprehension aim to understand how LLMs process language and mirror human linguistic features.
-Related studies are conducted at various linguistic levels, including sound, word, syntax, meaning, and discourse.
+LLM language comprehension is evaluated across multiple linguistic levels—sound, word, syntax, meaning, and discourse.
 Duan et al.
-[2024a] introduce a human-likeness benchmark comprising 10 psycholinguistic tests to assess these levels comprehensively.
-More studies focus on specific aspects.
-For instance, Wang et al.
-[2024b], Arehalli et al.
-[2022], Wolfman et al.
-[2024], Wilcox et al.
-[2021] examine syntactic and semantic processing.
-Surprisal, a measure of word predictability, is used across multiple studies to model the difficulty of processing syntactic ambiguities, garden-path sentences, and hierarchical structures [Wilcox et al., 2021, Arehalli et al., 2022, Li et al., 2024a].
-Overall, findings are mixed on whether LLMs exhibit human-like linguistic patterns.
-Duan et al.
-[2024b] demonstrate that GPT-2-XL displays human-like competence in both sound-gender association and implicit causality.
-However, it lacks human-like abilities in sound-shape association.
-While surprisal from LLMs predicts general trends in human reading patterns, it underestimates the processing difficulty in syntactically complex regions [Arehalli et al., 2022, Steuer et al., 2023].
-In contrast, Wang et al.
-[2024b] propose a new metric, Incompatibility Fraction, that outperforms surprisal in correlating with human garden-path effects.
-Other tasks address grammaticality judgment [Qiu et al., 2024, Ide et al., 2024], pragmatic inference [Bojić et al., 2023, Hu et al., 2023, Ruis et al., 2023], argument role processing [Lee et al., 2024b], and discourse-level comprehension [Duan et al., 2024a].
-Measurement results indicate that advanced models, such as GPT-4, perform competitively or even better than humans on pragmatic reasoning tasks [Bojić et al., 2023] and grammaticality judgment [Dentella et al., 2024], despite some inconsistency between different prompt formats [Hu and Levy, 2023] and deficiencies in implicature understanding [Ruis et al., 2023].
+[2024a] introduce a benchmark of 10 psycholinguistic tests to assess these comprehensively.
+Many studies focus on specific aspects, such as syntactic and semantic processing [Wang et al., 2024b, Arehalli et al., 2022, Wolfman et al., 2024, Wilcox et al., 2021].
+Surprisal, a measure of word predictability, is widely used to model processing difficulty in syntactic ambiguities and hierarchical structures [Wilcox et al., 2021, Arehalli et al., 2022, Li et al., 2024a].
+Results are mixed: for example, GPT-2-XL shows human-like competence in sound-gender association and implicit causality, but not in sound-shape association [Duan et al., 2024b].
+Other evaluation tasks include grammaticality judgment [Qiu et al., 2024, Ide et al., 2024], pragmatic inference [Bojić et al., 2023, Hu et al., 2023, Ruis et al., 2023], argument role processing [Lee et al., 2024b], and discourse comprehension [Duan et al., 2024a].
+Advanced models like GPT-4 often match or surpass humans in pragmatic reasoning [Bojić et al., 2023] and grammaticality judgment [Dentella et al., 2024], though performance varies with prompt format [Hu and Levy, 2023] and implicature understanding remains limited [Ruis et al., 2023].
 Language generation.
-Evaluating LLM language generation enhances our understanding of their linguistic capabilities.
-One rich line of research focuses on measuring the creativity of LLMs [Tang and Kejriwal, 2024, Stevenson et al., 2022, Hubert et al., 2023, Chakrabarty et al., 2024, Orwig et al., 2024, Lee and Chung, 2024, Boussioux et al., 2024, Bellemare-Pepin et al., 2024].
-They employ psychometric instruments such as Guilford’s Alternative Uses Test (AUT) [Guilford et al., 1978] and the Torrance Test of Creative Thinking (TTCT) [Torrance, 1966].
-Their results collectively show that early models such as GPT-3 fall short in originality and novelty, while more advanced LLMs like GPT-4 are more creative than the human average.
-More nuanced evaluations provide deeper insights.
-For instance, Tang and Kejriwal [2024] identify a dichotomy: LLMs often lack originality and novelty in divergent creativity tasks, where they are asked to devise novel uses for familiar objects.
-However, in creative writing, particularly in open-ended tasks, LLMs like GPT-4 can often create stories that rival human creativity.
+A rich line of research evaluates the creativity of LLM language generation [Tang and Kejriwal, 2024, Stevenson et al., 2022, Hubert et al., 2023, Chakrabarty et al., 2024, Orwig et al., 2024, Lee and Chung, 2024, Boussioux et al., 2024, Bellemare-Pepin et al., 2024], using tests like Guilford’s Alternative Uses Test (AUT) [Guilford et al., 1978] and the Torrance Test of Creative Thinking (TTCT) [Torrance, 1966].
+Early models such as GPT-3 lack originality and novelty, while advanced LLMs like GPT-4 surpass the human average in creativity.
+Notably, Tang and Kejriwal [2024] find LLMs underperform in divergent creativity (e.g., novel uses for familiar objects) but can rival humans in open-ended creative writing.
 Cai et al.
-[2024] conduct 12 preregistered psycholinguistic tests on LLMs and find that ChatGPT exhibits human-like responses in 10 experiments.
-In contrast, Seals and Shalin [2023] note that AI-generated analogies often lack human-like psycholinguistic properties.
-Similarly, Tian et al.
-[2024] evaluate LLMs in their psycholinguistic constructs revealed in narrative development and plot progression.
-While human-written stories often exhibit suspense, excitement, and diverse narrative structures, LLM-generated stories tend to be consistently positive and lack tension.
-Another work draws inspiration from linguistic profiling to evaluate LLMs’ task-specific linguistic abilities [Miaschi et al., 2024].
+[2024] show that ChatGPT produces human-like responses in most of 12 psycholinguistic tests, but AI-generated analogies often lack human-like linguistic properties [Seals and Shalin, 2023].
+LLM-generated stories tend to be positive and lack suspense or narrative diversity compared to human writing [Tian et al., 2024].
+Linguistic profiling has also been used to characterize the task-specific language abilities of LLMs [Miaschi et al., 2024].
 Language acquisition.
-Work on developmental plausibility assesses whether LLMs replicate stages of language learning analogous to those in children [Steuer et al., 2023, Shah et al., 2024].
+Work on developmental plausibility examines whether LLMs replicate stages of language learning analogous to those in children [Steuer et al., 2023, Shah et al., 2024].
 It is shown that regardless of model size, the developmental trajectories of pretrained language models consistently exhibit a window of maximal alignment with human cognitive development [Shah et al., 2024].
@@
 In addition, mixed results are found in the alignment between LLMs and human linguistic cognition [Duan et al., 2024b, Wolfman et al., 2024] and language acquisition [Shah et al., 2024, Frank, 2023b].
-Interested readers are referred to Chang and Bergen [2024] for a comprehensive survey of language models’ behavior, evaluated from the perspective of psycholinguistics.
 4.2.4 Learning and cognitive capabilities
-Psychometrics in learning and cognitive capabilities measure human mental functions like memory, reasoning, problem-solving, and comprehension to understand cognitive strengths and weaknesses.
-These insights inform educational strategies and cognitive development.
-Recent psychometric evaluations of LLMs aim to interpret model behavior in a similar vein.
-A series of studies adapt human psychometric tools to evaluate the cognitive abilities of LLMs.
-Some results are promising.
-Galatzer-Levy et al.
-[2024] adapt the Wechsler Adult Intelligence Scale (WAIS-IV) to assess verbal comprehension, working memory, and perceptual reasoning in LLMs.
-The results indicate that LLMs generally perform at a top human level in verbal comprehension and working memory, though multimodal models exhibit deficiencies in visual reasoning.
-Sartori and Orrú [2023], Webb et al.
-[2022] utilize Raven’s Progressive Matrices and other fluid intelligence tests to assess analogical reasoning, finding that LLMs match or exceed human performance.
-Other findings highlight the current limitations in LLMs’ cognitive abilities.
-Dayan et al.
-[2024] employ the Montreal Cognitive Assessment (MoCA) to test LLMs, revealing that nearly all models exhibit signs of mild cognitive impairment, particularly in visuospatial and executive tasks.
-Wu et al.
-[2025] focus on evaluating fluid intelligence using the ARC task, emphasizing deficiencies in abstract problem-solving.
-Critical perspectives also challenge the validity of current tests [Löhn et al., 2024, Li et al., 2024d, Zhang et al., 2023b], driving the development of new benchmarks and testing methodologies [Coda-Forno et al., 2024, Wang et al., 2024e, Zhuang et al., 2023a, Zeng et al., 2024c, Song et al., 2024a].
-From the perspective of learning and cognitive development, Wang et al.
-[2024e] draw inspiration from Piaget’s Theory of Cognitive Development (PTC) and track human-like progression in LLMs.
-They find that advanced LLMs (such as GPT-4) demonstrate human-like cognitive abilities, comparable to those of a 20-year-old human.
+Psychometrics of learning and cognitive capabilities measures mental functions such as memory, reasoning, problem-solving, and comprehension to identify cognitive strengths and weaknesses.
+Recent studies adapt human psychometric tools to evaluate LLMs’ cognitive abilities.
+For example, Galatzer-Levy et al.
+[2024] use the Wechsler Adult Intelligence Scale (WAIS-IV) to show LLMs generally reach top human levels in verbal comprehension and working memory, though multimodal models lag in visual reasoning.
+Raven’s Progressive Matrices and other fluid intelligence tests reveal LLMs can match or surpass humans in analogical reasoning [Sartori and Orrú, 2023, Webb et al., 2022].
+However, limitations persist.
+The Montreal Cognitive Assessment (MoCA) shows most LLMs display mild cognitive impairment, especially in visuospatial and executive tasks [Dayan et al., 2024].
+Abstraction and Reasoning Challenge (ARC) further highlights deficits in abstract problem-solving [Wu et al., 2025].
+Some works question the validity of current tests [Löhn et al., 2024, Li et al., 2024d, Zhang et al., 2023b], prompting new benchmarks and methodologies [Coda-Forno et al., 2024, Wang et al., 2024e, Zhuang et al., 2023a, Zeng et al., 2024c, Song et al., 2024a].
+Developmental perspectives also emerge: Wang et al.
+[2024e] leverage Piaget’s Theory of Cognitive Development to track human-like progression in LLMs, finding advanced models (e.g., GPT-4) reach cognitive abilities comparable to 20-year-old humans.
 Shah et al.
-[2024] identify developmental profiles in LLMs and their parallels with human development via psychometric theories like Cattell-Horn-Carroll.
+[2024] identify developmental profiles in LLMs paralleling human development via psychometric theories like Cattell-Horn-Carroll.
 Main findings.
@@
 5 Psychometric evaluation methodology
+Figure 3: Overview of LLM psychometric evaluation methodologies.
 This section examines the methodologies employed in LLM Psychometrics.
@@
 The configuration of inference parameters (§ 5.5) is additionally discussed.
-Figure 3: Overview of LLM psychometric evaluation methodologies.
 5.1 Test format
-Test formats of LLM psychometric evaluation can be categorized into structured tests, open-ended conversations, and agentic simulations.
-Table 5 provides an overview of the test formats for each construct.
-Table 5: Overview of test formats: structured tests, open-ended conversations, and agentic simulations. For each test format, we list some examples for each construct.
+Table 5: Overview of test formats: structured tests, open-ended conversations, and agentic simulations. For each format, representative examples are provided for each construct.
 Construct Structured tests Open-ended conversations Agentic simulations
@@
 Learning and cognitive capabilities Wu et al. [2025], Coda-Forno et al. [2024] Galatzer-Levy et al. [2024] Lv et al. [2024]
+Test formats of LLM psychometric evaluation can be categorized into structured tests, open-ended conversations, and agentic simulations.
+Table 5 collects exemplary works illustrating the test formats used for each construct.
 5.1.1 Structured tests
@@
 These items may include alternative-choice questions, multiple-choice questions, rating scales, and short-answer questions.
-Structured tests for LLM personality and values frequently employ Likert-scale ratings, prompting LLMs to express their level of agreement with given statements.
-For instance, in evaluating LLM personality with the BFI [Li et al., 2024d], a statement might be "I see myself as someone who is outgoing, sociable." When evaluating LLM morality, tools like the DIT require LLMs to make decisions and rate the significance of the reasons behind those decisions [Tanmay et al., 2023].
-The MFQ typically asks LLMs to assess the relevance of moral issues and their agreement with moral statements using a Likert scale [Ji et al., 2024].
-Structured surveys for examining political attitudes and opinions often involve multiple-choice questions, such as "Which statement comes closer to your view?" from the ATP [Santurkar et al., 2023], and Likert-scale questions, such as "How much do you agree with the statement: The enemy of my enemy is my friend?" from the PCT [Rutinowski et al., 2024].
-To test heuristics and biases in LLMs, Echterhoff et al.
-[2024] focus on student admissions and provide options for different students.
-Hagendorff et al.
-[2023] employ a battery of semantic illusions and cognitive reflection tests, which require short answers in predefined formats (e.g., number of days).
-Most benchmark-style evaluations in LLM social interaction are multiple-choice tests, such as BigToM for ToM [Gandhi et al., 2023], EmoBench for EI [Sabour et al., 2024], and SESI for SI [Xu et al., 2024b].
-Structured psycholinguistic tests usually involve masked or next-word prediction, thereby measuring distributional alignment between human decisions and LLM output probabilities [Duan et al., 2024a].
-Other higher-level psycholinguistic tests administer forced-choice questions for evaluating semantic comprehension [Li et al., 2024a] or require Likert-scale plausibility and grammaticality ratings [Qiu et al., 2024, Amouyal et al., 2024].
-Tests of learning and cognitive capabilities usually involve a mix of structured and open-ended tasks, and the structured ones are mostly short-answer tests.
-For example, the WAIS-IV contains many structured short-answer tests, such as Digit Span Forward [Galatzer-Levy et al., 2024]; the ARC Benchmark requires LLMs to generate 2D grids of predefined shapes [Wu et al., 2025].
-When adapting standardized psychometric tests for LLMs, it is common practice to retain the original items, merely reformatting them as prompts.
+When adapting structured psychometric tests for LLMs, it is common practice to retain the original items and simply reformat them as prompts.
+For example, personality and values are often measured with Likert scales (e.g., BFI statements like “I see myself as someone who is outgoing, sociable”), while morality is assessed using tools like the DIT or MFQ, which require decisions and agreement ratings.
+Heuristics and biases are often tested with scenario-based choices.
+Social interaction and psycholinguistic abilities are evaluated with multiple-choice, forced-choice, or masked word prediction.
+Cognitive tests also involve short-answer questions (e.g., WAIS-IV Digit Span).
 Structured psychometric tests are advantageous due to their scalability, objectivity, and automated scoring.
@@
 Open-ended conversations.
-A straightforward method for unstructured testing involves engaging LLMs, as chatbots, in open-ended and mostly single-turn conversations that manifest specific constructs.
-This approach reflects real-world human-LLM interactions and evaluates LLMs in more realistic contexts.
-For instance, Jiang et al.
-[2024b] request LLMs to generate long-form narratives that reveal their personality traits.
-ValueBench prompts LLMs with advice-seeking user queries to capture their subtle influence on users’ values [Ren et al., 2024].
-Similar to ValueBench, Sachdeva and van Nuenen [2025] draw everyday moral dilemmas from Reddit and prompt LLMs as if users are seeking moral judgments from other community members.
-To thoroughly investigate LLMs’ political attitudes and opinions, Röttger et al.
-[2024] reformat multiple-choice questions in PCT into forced-choice and unconstrained open-ended questions.
-Regarding the evaluation of cognitive constructs, Healey et al.
-[2024] design a pipeline to capture nuanced types of bias that can occur in free-response answers but not in multiple-choice questions.
-Welivita and Pu [2024] assess the empathetic ability of LLMs by asking them to respond to emotional situations posed by users.
-Cognitive tests also include many open-ended questions, such as vocabulary and comprehension tests in WAIS-IV [Galatzer-Levy et al., 2024].
+Unstructured testing often engages LLMs in open-ended, typically single-turn conversations that elicit specific constructs, mirroring real-world human-LLM interactions.
+For example, Jiang et al.
+[2024b] have LLMs generate long-form narratives to reveal personality traits.
+ValueBench uses advice-seeking queries to assess LLMs’ influence on users’ values [Ren et al., 2024], while Sachdeva and van Nuenen [2025] present moral dilemmas from Reddit, prompting LLMs to provide community-style moral judgments.
+To probe political attitudes, Röttger et al.
+[2024] convert PCT multiple-choice items into forced-choice and open-ended formats.
+For cognitive constructs, Healey et al.
+[2024] design pipelines to detect nuanced biases in free responses, and Welivita and Pu [2024] evaluate empathy by asking LLMs to respond to emotional scenarios.
+Many cognitive tests, such as WAIS-IV vocabulary and comprehension, also use open-ended questions [Galatzer-Levy et al., 2024].
 Agentic simulations.
-Advanced unstructured tests often evaluate LLMs by treating them as agents, placing them in complex role-playing scenarios, and analyzing their decision-making in a contextualized and dynamic environment.
+Advanced unstructured tests evaluate LLMs as agents in complex, contextualized role-playing scenarios, analyzing their decision-making in dynamic environments.
 Huang et al.
-[2024c] propose a psychometric approach to design LLM-Agents with distinct personalities, evaluating them by replicating complex human-like behaviors in agentic simulations.
+[2024c] design LLM-Agents with distinct personalities and analyze them by replicating human-like behaviors in agentic simulations.
 Shen et al.
-[2024] contextualize value measurement through four real-world vignettes: collaborative writing, education, public sectors, and healthcare.
+[2024] measure values using four real-world vignettes: collaborative writing, education, public sectors, and healthcare.
 Chiu et al.
-[2025] present value and moral dilemmas to LLMs, observing their decision-making and rationales in agentic contexts.
-Unstructured tests for cognitive constructs are typically more complex, often extending to multi-agent frameworks.
-For instance, Xie et al.
+[2025] present value and moral dilemmas to LLMs, examining their decisions and rationales.
+Cognitive construct tests are often more complex, extending to multi-agent settings; for example, Xie et al.
 [2024b], Bai et al.
 [2024] identify cognitive biases in multi-agent communication.
-SOTOPIA is an open-ended environment simulating complex social interactions among LLM agents, where their social intelligence is evaluated [Zhou et al., 2024b, Wang et al., 2024d].
+SOTOPIA simulates open-ended social interactions among LLM agents to evaluate social intelligence [Zhou et al., 2024b, Wang et al., 2024d].
 Lv et al.
-[2024] benchmark the cognitive dynamics of LLMs by having them repeatedly complete a cognitive questionnaire and provide reasoning after processing information flows.
-Unstructured tests evaluate LLMs in free-form, contextualized settings.
-The primary advantages are their ecological validity and the ability to capture complex and nuanced behaviors that may not emerge in structured formats, such as intricate reasoning patterns, subtle biases, and dynamic social interactions.
-However, these benefits come with significant challenges, such as difficulties in standardization, complicated and often subjective scoring and analysis, and reproducibility issues.
+[2024] benchmark cognitive dynamics by having LLMs repeatedly complete cognitive questionnaires and reason over information flows.
+Unstructured tests assess LLMs in free-form, contextualized scenarios, offering greater ecological validity and revealing complex behaviors—such as nuanced reasoning, subtle biases, and dynamic social interactions—not captured by structured formats.
+However, they pose challenges in standardization, scoring, and reproducibility.
 5.2 Data and task sources
-Table 6: Overview of data and task sources: established inventories, custom-curated items, and synthetic items. For each data source, we list some examples for each construct.
+Table 6: Overview of data and task sources: established inventories, custom-curated items, and synthetic items. For each data source, representative examples are provided for each construct.
 Construct Established inventories Custom-curated items Synthetic items
@@
 Established inventories.
-Established psychometric inventories present well-validated and widely recognized tools.
-Representative inventories in LLM psychometrics include BFI and HEXACO for personality traits [Serapio-García et al., 2023, Barua et al., 2024]; SVS, PVQ, WVS, and VSM for values [Fischer et al., 2023, Miotto et al., 2022]; MFT and DIT for morality [Abdulhai et al., 2024, Khandelwal et al., 2024]; ANES and PCT for political attitudes and opinions [Argyle et al., 2023, Bernardelle et al., 2024]; cognitive reflection tests (CRT) for heuristics and biases [Hagendorff et al., 2023, Binz and Schulz, 2023]; False-Belief Tasks for ToM [Ullman, 2023, Chen et al., 2024d]; and many more.
-Established inventories provide a straightforward and standardized approach to psychometric evaluation.
-However, many studies question the reliability and validity of established inventories when used with LLMs, citing common issues of data contamination and response bias [Ye et al., 2025a].
-Additionally, these simplified, structured tests may not adequately capture the complex task scenarios encountered by general-purpose AI.
+Established psychometric inventories offer standardized, well-validated tools for LLM evaluation.
+Common examples include BFI and HEXACO for personality [Serapio-García et al., 2023, Barua et al., 2024]; SVS, PVQ, WVS, and VSM for values [Fischer et al., 2023, Miotto et al., 2022]; MFT and DIT for morality [Abdulhai et al., 2024, Khandelwal et al., 2024]; ANES and PCT for political attitudes [Argyle et al., 2023, Bernardelle et al., 2024]; CRT for heuristics and biases [Hagendorff et al., 2023, Binz and Schulz, 2023]; and False-Belief Tasks for ToM [Ullman, 2023, Chen et al., 2024d].
+While these inventories enable straightforward, standardized assessment, their use with LLMs raises concerns about data contamination, response bias, and limited ecological validity for complex AI tasks [Ye et al., 2025a].
 Custom-curated items.
-Human-authored and custom-curated items provide tailored psychometric tests that are often more relevant and applicable to LLMs.
-These items enable researchers to explore novel and unique dimensions of LLM capabilities more robustly.
-Shen et al.
-[2024] argue that SVS and PVQ are inadequate for measuring LLM value alignment because they lack supplementary values from alignment literature and contextual nuances of human-AI alignment.
-To address this, they created 11 AI-informed value statements, similar in style to SVS and PVQ, based on a systematic review of human-AI alignment literature.
-Similarly, Ceron et al.
-[2024] suggest that political questionnaires may yield unreliable or inconsistent results.
-They propose a series of tests to assess the reliability and consistency of LLMs’ stances on political statements, using a dataset of voting-advice questionnaires from seven EU countries annotated for policy issues.
-Regarding the evaluation of LLM morality, Hendrycks et al.
-[2021] curate a large-scale dataset of over 130,000 open-world scenarios requiring LLMs to make moral judgments.
-Another example is the ToM dataset based on naturally occurring spoken dialogues [Soubki et al., 2024], which is motivated by the potential misalignment between synthetic ToM benchmarks and actual human behavior.
-Despite their high ecological validity, custom-curated items require significant effort to develop and validate, which may limit their scalability and diversity.
+Human-authored, custom-curated items offer tailored psychometric tests that are often more relevant to LLMs, enabling exploration of novel capability dimensions.
+For example, Shen et al.
+[2024] find SVS and PVQ insufficient for LLM value alignment, so they create 11 AI-informed value statements based on a systematic review of alignment literature.
+Ceron et al.
+[2024] highlight the unreliability of standard political questionnaires for LLMs and introduce tests using annotated voting-advice questionnaires from seven EU countries to assess stance consistency.
+For morality, Hendrycks et al.
+[2021] curate a dataset of over 130,000 open-world scenarios requiring moral judgments.
+In ToM, Soubki et al.
+[2024] develop a dataset from real spoken dialogues to address gaps between synthetic benchmarks and human behavior.
+While custom-curated items offer high ecological validity, their development and validation are labor-intensive, limiting scalability and diversity.
 Synthetic items.
-Synthetic items generated by AI models, mostly LLMs themselves, represent an emerging paradigm in LLM psychometrics.
-This method leverages the generative capabilities of LLMs to produce large-scale, diverse, and contextually rich items.
-Using synthetic items entails careful prompt engineering and additional validation.
-Some studies prompt LLMs to modify or expand established inventories for better ecological validity, mitigating data contamination, and enabling more scalable and diverse testing.
-For example, Ren et al.
-[2024] rewrite self-report items into advice-seeking queries that better align with real-world human-AI interactions.
+Synthetic items—generated primarily by LLMs—are an emerging approach in LLM Psychometrics, enabling large-scale, diverse, and contextually rich test content.
+This method requires careful prompt engineering and validation.
+Some studies adapt established inventories to improve ecological validity, reduce data contamination, and enhance scalability.
+For instance, Ren et al.
+[2024] convert self-report items into advice-seeking queries to better reflect real-world human-AI interactions.
 Bhandari et al.
-[2025b] modify the established items into semantically equivalent ones to address the issue of data contamination.
-They validate the rewritten items by setting a threshold for semantic embedding similarity.
+[2025b] rewrite established items into semantically equivalent forms, validating them via semantic embedding similarity thresholds.
 Zhu et al.
-[2024a] transform the cognitive test items into more diverse ones, allowing for multifaceted analysis of LLMs’ cognitive abilities.
-Other studies generate synthetic tests from scratch.
+[2024a] diversify cognitive test items to enable more comprehensive analysis of LLM cognitive abilities.
+Other work generates synthetic tests entirely from scratch.
 Hadar-Shoval et al.
-[2025] validate their synthetic cognitive tests by showing a strong human performance correlation between LLM-generated tests and established tests.
+[2025] demonstrate that LLM-generated cognitive tests correlate strongly with human performance on established tests.
 Ye et al.
-[2025a] generate value-eliciting prompts for LLM value measurement, subsequently confirming the reliability and validity of the test results.
+[2025a] create value-eliciting prompts and confirm the reliability and validity of the resulting measurements.
 Jiang et al.
-[2024a] propose generative evolving testing, where an LLM-based Item Generator learns to generate items with specified difficulty.
-The moral dilemmas presented by Chiu et al.
-[2025] are created using GPT-4 and employed to evaluate the moral judgment of LLMs.
-When measuring SI of LLMs with agentic simulations, Zhou et al.
-[2024b], Mou et al.
-[2024] prompt LLMs to generate various components of the social interaction.
+[2024a] introduce generative evolving testing, where an LLM-based Item Generator produces items of specified difficulty.
+Chiu et al.
+[2025] use GPT-4 to generate moral dilemmas for evaluating LLM moral judgment.
+For social intelligence, agentic simulations prompt LLMs to generate components of social interactions [Zhou et al., 2024b, Mou et al., 2024].
 5.3 Prompting strategies
@@
 Role-playing prompts.
-Role-playing prompts, also known as persona or profiling prompts, incorporate specific demographic information or personal attributes into the LLM context during test administration.
-Some studies utilize role-playing prompts to generate multiple participants from a single LLM for statistical analysis.
-Serapio-García et al.
-[2023] incorporate persona instructions and descriptions before each item when assessing LLM personality.
-This approach yields diverse responses from a single LLM, facilitating the computation of the reliability and validity of the test results for each LLM.
-Ye et al.
-[2025b] use value-anchoring prompts to guide LLM values, thereby generating hundreds of LLM participants for value measurement.
-The statistical analysis of these measurement results leads to the development of an LLM-specific value system.
-Research suggests that LLMs embody a superposition of diverse personalities, values, and perspectives [Kovač et al., 2023].
-This has motivated investigations into the personality adaptability of LLMs.
-Utilizing role-playing prompts, Lu et al.
-[2023], Jiang et al.
-[2023], La Cava and Tagarelli [2024] demonstrate that LLMs can dynamically transition between different personality types.
-However, the degree of steerability varies among LLMs [La Cava and Tagarelli, 2024], and some personality dimensions are less steerable than others [Li et al., 2024d, Jiang et al., 2023].
-Kovač et al.
-[2024], Rozen et al.
-[2024] employ role-playing prompts to examine the stability and consistency of LLM values.
-Similarly, Münker [2024], Simmons [2023], Wright et al.
-[2024] incorporate moral or political profiles into prompts to evaluate LLMs’ ability to mimic these perspectives and their susceptibility to biases linked to specific groups.
-Many cognitive benchmarks, particularly those related to social interactions, assign personas and social roles to LLMs during agentic simulations [Zhou et al., 2024b, Mou et al., 2024, Huang et al., 2023b].
-The successful completion of these benchmark tasks heavily relies on the LLM’s ability to comprehend and emulate the assigned persona and social role.
-Persona prompts are also shown to affect the social-cognitive reasoning of LLMs when other task variables are held constant [Tan et al., 2024].
-The susceptibility to role-playing prompts presents a double-edged sword.
-On one hand, they enable the emulation of diverse perspectives and behaviors, which can be beneficial for applications such as role-playing conversations and social simulations.
-On the other hand, these capabilities may result in inconsistencies and instability in the LLM’s values and perspectives.
-It may compromise the reliability and validity of test results and pose challenges to robustly aligning LLMs with human values.
+Role-playing prompts (persona or profiling prompts) embed demographic or personal attributes into the LLM context during testing.
+These prompts enable the generation of multiple simulated participants from a single LLM for statistical analysis.
+For example, Serapio-García et al.
+[2023] prepend persona instructions to each item to elicit diverse responses, allowing examining reliability and validity within one LLM.
+Similarly, Ye et al.
+[2025b] use value-anchoring prompts to create hundreds of LLM-based participants for value measurement, supporting the development of an LLM-specific value system.
+LLMs can represent a superposition of personalities, values, and perspectives [Kovač et al., 2023], prompting research into their adaptability.
+Role-playing prompts show that LLMs can shift between personality types, though steerability varies by model and dimension [Lu et al., 2023, Jiang et al., 2023, La Cava and Tagarelli, 2024, Li et al., 2024d].
+Such prompts are also used to test the stability of LLM values [Kovač et al., 2024, Rozen et al., 2024] and to measure how well LLMs can mimic moral or political profiles and their susceptibility to related biases [Münker, 2024, Simmons, 2023, Wright et al., 2024].
+Cognitive benchmarks, especially for social interactions, often use persona prompts in agentic simulations [Zhou et al., 2024b, Mou et al., 2024, Huang et al., 2023b], with task performance depending on the LLM’s ability to embody these roles.
+These prompts influence social-cognitive reasoning when controlling for other variables [Tan et al., 2024].
+While role-playing capabilities enable diverse perspectives useful for social simulations, they can also lead to unstable values and perspectives, potentially compromising test validity and complicating value alignment.
 Performance-enhancing prompts.
-Performance-enhancing prompts are crafted to augment LLMs in psychometric evaluations.
-Hagendorff [2023] suggest using such prompts whenever possible when evaluating machine psychology.
-The Chain of Thought (CoT) prompting [Wei et al., 2022] is leveraged to reduce biases [Hagendorff et al., 2023], enhance social intelligence [Shapira et al., 2023a], and boost cognitive capabilities [Coda-Forno et al., 2024].
-Variants of CoT, such as Emotional CoT [Li et al., 2024e], specifically target improvements in the emotional intelligence of LLMs.
-Emotional prompts, as introduced by Li et al.
-[2023a], enhance the general cognitive abilities of LLMs by incorporating emotional stimuli.
-Few-shot prompting [Brown et al., 2020], another versatile prompting technique, has been shown to improve ToM performance in LLMs [Moghaddam and Honey, 2023].
-Additional research explores diverse prompting strategies for specific constructs.
-For instance, Zhou et al.
-[2024a] advocate for MFT-guided reasoning prompts for better moral alignment.
-Strategies like those proposed by Echterhoff et al.
+Performance-enhancing prompts are crafted to augment LLMs in psychometric evaluations, as recommended by Hagendorff [2023].
+Chain of Thought (CoT) prompting [Wei et al., 2022] reduces biases [Hagendorff et al., 2023], enhances social intelligence [Shapira et al., 2023a], and improves cognitive capabilities [Coda-Forno et al., 2024].
+Its variant Emotional CoT [Li et al., 2024e] boosts EI, while emotional prompts [Li et al., 2023a] enhance general cognitive abilities through emotional stimuli.
+Few-shot prompting [Brown et al., 2020] has demonstrated improvements in ToM performance [Moghaddam and Honey, 2023].
+Research has also developed construct-specific prompting strategies.
+Zhou et al.
+[2024a] propose MFT-guided reasoning prompts for moral alignment, while Echterhoff et al.
 [2024], Sumita et al.
-[2024] aim to self-debias LLMs.
-ToM-specific enhancement strategies include SymbolicToM [Sclar et al., 2023] and temporal decomposition-based prompting [Sarangi et al., 2025, Hou et al., 2024].
-In contrast, Zhao et al.
-[2025a] develop a testing framework utilizing self-reflection prompts to examine both explicit and implicit social biases in LLMs, where explicit bias measurement serves as a reflection on implicit bias.
+[2024] focus on self-debiasing approaches.
+ToM-specific enhancements include SymbolicToM [Sclar et al., 2023] and temporal decomposition-based prompting [Sarangi et al., 2025, Hou et al., 2024].
+Additionally, Zhao et al.
+[2025a] develop a framework using self-reflection prompts to examine both explicit and implicit social biases, where explicit bias measurement reflects implicit bias.
 Prompt perturbation and adversarial attacks.
@@
 5.4.1 Closed-ended output and scoring
-The evaluation of LLM performance on structured tests can be categorized into two methodological paradigms: logit-based probabilistic analysis and closed-ended output scoring.
-In multiple-choice or Likert-scale evaluations, some studies retrieve and analyze the distribution of token-level logits—particularly the first generated token—to infer the model’s latent personalities and opinions [Pellert et al., 2024, Santurkar et al., 2023], analyze response entropy [Dominguez-Olmedo et al., 2024], and compute the distributional alignment between LLM outputs and human behavioral data [Arehalli et al., 2022].
-Psycholinguistic tests especially rely on logit-based metrics.
-A notable example is surprisal, defined as the negative log probability of a token given its context.
-It quantifies the predictability or cognitive effort associated with language processing, enabling researchers to map LLM outputs to human-like uncertainty patterns [Wang et al., 2024b, Steuer et al., 2023].
+LLM performance on structured tests is typically evaluated using logit-based probabilistic analysis and closed-ended output scoring.
+In multiple-choice or Likert-scale settings, some studies analyze token-level logits—especially the first token—to infer latent traits, detect response entropy, or compare LLM output distributions against human data [Pellert et al., 2024, Santurkar et al., 2023, Dominguez-Olmedo et al., 2024, Arehalli et al., 2022].
+Psycholinguistic tests often use logit-based metrics such as surprisal (the negative log probability of a token), which quantifies predictability and cognitive effort, allowing researchers to align LLM uncertainty patterns with those of humans [Wang et al., 2024b, Steuer et al., 2023].
 Conversely, closed-ended outputs, which are explicit numerical scores or categorical selections, can be analyzed using predefined scoring protocols.
@@
 5.4.2 Open-ended output and scoring
-Scoring open-ended outputs is more challenging. Related scoring schemes can be generally categorized into 1) rule-based scoring, 2) model-based scoring, and 3) human scoring.
+Scoring open-ended outputs is more complex and typically falls into three categories: 1) rule-based, 2) model-based, and 3) human scoring.
 Rule-based scoring.
-Rule-based scoring predominantly relies on the lexical hypothesis [Allport and Odbert, 1936], which posits that the meaning and relevance of responses can be ascertained by analyzing the presence and frequency of specific keywords or phrases within the text.
-Jiang et al.
-[2024b] utilize the LIWC (Linguistic Inquiry and Word Count) features [Pennebaker et al., 2001] to evaluate the personalities of LLMs.
-Similarly, Fischer et al.
-[2023] employ a theory-driven value dictionary [Ponizovskiy et al., 2020] to assess LLM values.
-Nevertheless, lexicon-based scoring has been demonstrated to be limited in capturing semantic nuances [Ye et al., 2025a].
-Some other scoring rules extend beyond simple keyword matching.
-For example, Healey et al.
-[2024] automatically identify biases by analyzing whether responses deviate from equivalent treatment.
+Rule-based scoring predominantly relies on the lexical hypothesis [Allport and Odbert, 1936], analyzing the presence and frequency of specific keywords or phrases.
+For example, Jiang et al.
+[2024b] apply LIWC (Linguistic Inquiry and Word Count) features [Pennebaker et al., 2001] to evaluate LLM personalities, and Fischer et al.
+[2023] employ a theory-driven value dictionary [Ponizovskiy et al., 2020] for value measurement.
+However, such lexicon-based approaches struggle to capture semantic nuance [Ye et al., 2025a].
+Some rule-based methods go beyond keyword matching; for instance, Healey et al.
+[2024] detect bias by checking for deviations from equivalent treatment in responses.
 Model-based scoring.
-Model-based scoring is prevalent in current unstructured testing due to its flexibility and scalability.
-Several studies have focused on training models specifically to evaluate LLM responses.
-For instance, Hilliard et al.
-[2024] fine-tuned BERT variants [Devlin et al., 2019] using the MyPersonality dataset [^5] https://sites.google.com/michalkosinski.com/mypersonality.
-to score LLM personalities.
-Similarly, Ye et al.
-[2025a], Yao et al.
-[2024, 2025b], Sorensen et al.
-[2024a] fine-tune LLMs with established psychometric inventories [Ren et al., 2024] and/or human annotations to classify the value valence of LLM responses.
-Scoring can occur at the item level, as exemplified by Generative Psychometrics, which consists of item parsing, scoring, and aggregation to evaluate LLM values [Ye et al., 2025a].
-Alternatively, scoring can be conducted at the response level, assessing the LLM response in its entirety [Yao et al., 2024, 2025b].
-Some models are trained to score according to specific theoretical frameworks, e.g., Schwartz’s Value Theory [Yao et al., 2024], while others are designed to be generalists, utilizing a broader range of fine-tuning data and leveraging the prior knowledge of LLMs [Ye et al., 2025a].
-Overall, training evaluators is more prevalent for evaluating personalities and values due to the ease of associating textual expressions with specific personality and value dimensions.
-More studies directly employ the LLM-as-a-judge approach [Gu et al., 2024] to score the open-ended responses.
-For example, Li et al.
-[2024d] prompt LLMs to score LLM responses across personality, ToM, and motivation.
-They validate the consistency between two LLM raters to indicate inter-rater reliability.
-Zheng et al.
-[2025] implement LLMs as personality evaluators and validate their consistency with human raters.
-The LLM-as-a-judge approach is particularly advantageous for evaluations in highly unstructured simulation settings.
-Zhou et al.
-[2024b], Wang et al.
-[2024d, a], Mou et al.
-[2024] apply LLM judges to assess SI across various dimensions, including goal completion, relationship maintenance, and adherence to social rules.
-Additionally, several works use embedding models to evaluate the similarity between LLM responses and prototypical examples [Huang et al., 2024c, Cahyawijaya et al., 2024, Amirizaniani et al., 2024].
+Model-based scoring is prevalent in unstructured testing for its flexibility and scalability.
+Many studies train models to evaluate LLM responses, such as fine-tuning BERT variants on the MyPersonality dataset for personality scoring [Hilliard et al., 2024].
+Others fine-tune LLMs with psychometric inventories or human annotations to classify value valence [Ye et al., 2025a, Yao et al., 2024, 2025b, Sorensen et al., 2024a].
+Scoring can be conducted at the item level (e.g., Generative Psychometrics: item parsing, scoring, aggregation [Ye et al., 2025a]) or at the response level, evaluating the entire output [Yao et al., 2024, 2025b].
+Some models are tailored to specific frameworks (e.g., Schwartz’s Value Theory [Yao et al., 2024]), while others are trained to be generalists by leveraging broader fine-tuning data and LLM prior knowledge [Ye et al., 2025a].
+Model-based evaluators are especially common for personality and value assessment, where textual cues map readily to target dimensions.
+Many studies now use the LLM-as-a-judge approach [Gu et al., 2024] for scoring open-ended responses.
+For instance, Li et al.
+[2024d], Zheng et al.
+[2025] prompt LLMs to rate responses on personality, ToM, and motivation, validating consistency with other LLMs or human raters.
+This method is especially useful in unstructured simulation settings, such as SI measurement across goal completion, relationship maintenance, and social rule adherence [Zhou et al., 2024b, Wang et al., 2024d, a, Mou et al., 2024].
+Embedding models are also used to measure similarity between LLM outputs and prototypical examples [Huang et al., 2024c, Cahyawijaya et al., 2024, Amirizaniani et al., 2024].
 Human scoring.
@@
 5.5 Inference parameters
-Evaluation results and validation methods also depend on inference parameters of LLMs.
-Some studies use greedy decoding for deterministic outputs, requiring only a single response per item but sacrificing output diversity.
-In contrast, sampling-based decoding introduces stochasticity, generating a broader range of outputs by adjusting hyperparameters such as temperature, top-k, and top-p.
-Inference parameters can affect the measured traits and abilities of LLMs.
-Diversifying the outputs may reveal a broader spectrum of latent traits, opinions, or cognitive strategies, but also introduce variability that complicates reliability and validity assessments.
+Evaluation and validation depend on LLM inference parameters.
+Greedy decoding yields deterministic, reproducible outputs but limits diversity, while sampling-based decoding (adjusting temperature, top-k, top-p) introduces stochasticity and diversifies responses.
+Sampling-based decoding may reveal a broader spectrum of latent traits, opinions, or cognitive strategies, but also introduce variability that complicates psychometric validation.
 Conversely, deterministic settings enhance reproducibility but may obscure the model’s full range of capabilities or biases.
-Most studies explicitly report and control these settings to ensure fair comparisons across models or experimental conditions.
-Some also examine the sensitivity of psychometric results to different decoding parameters.
-Transparent reporting and methodological rigor in parameter selection are essential.
-Investigating both deterministic and stochastic settings is recommended to fully understand the implications of test results.
+Most studies report and control inference settings for fair comparisons, and some analyze sensitivity to these parameters.
+Transparent reporting and careful parameter selection are essential; both deterministic and stochastic settings should be considered to fully understand the implications of test results.
 6 Psychometric validation
-Figure 4: Overview of psychometric validation: reliability and consistency, validity, and standards and recommendations [Löhn et al., 2024].
-Unlike AI benchmarking, which prioritizes system performance over test justification, psychometrics emphasizes theoretical grounding, standardized protocols, and reproducibility.
-Psychometric validation rigorously ensures the reliability, validity, and fairness of psychological tests.
-LLM psychometrics, as an emerging field, still lacks standardization for test design and administration, and recent research has begun exploring various aspects of psychometric validation.
+Figure 4: Overview of psychometric validation: reliability and consistency, validity, and standards and recommendations.
+Unlike AI benchmarking, which focuses on system performance, psychometrics prioritizes theoretical grounding, standardized protocols, and reproducibility.
+Psychometric validation ensures that tests are reliable, valid, and fair.
+As an emerging field, LLM Psychometrics lacks standardization in test design and administration, and recent work is beginning to address these issues.
+We outline the landscape of psychometric validation in Fig.
+4.
 6.1 Reliability and consistency
-Reliability is a fundamental principle in psychometric validation, assessing the degree to which a test is free from error.
-It encompasses consistency over time (test-retest reliability), across different versions (parallel forms reliability), and among evaluators (inter-rater reliability).
-Researchers have adapted these psychometric reliability metrics for LLM psychometrics.
-For instance, Li et al.
-[2024d] introduce a multi-dimensional psychometrics benchmark that extends to five forms of reliability, addressing the unique challenges of LLM evaluation: internal consistency, parallel forms reliability, inter-rater reliability, option position robustness, and adversarial attack robustness.
-Studies by Zheng et al.
-[2025], Huang et al.
-[2023a], Ceron et al.
-[2024], Shu et al.
-[2024] systematically investigate the consistency of LLM outputs under repeated trials, variations in prompt order, and cross-lingual contexts.
-They employ reliability indices such as Cronbach’s alpha, intraclass correlation coefficients, and agreement statistics to quantify both internal and external reliability dimensions.
-We note that some metrics pertain to both stochastic and systematic errors in LLM outputs; certain inconsistencies may compromise both the reliability and validity of LLM psychometrics.
-Promising results indicate that some advanced LLMs consistently generate stable responses on the Big Five personality traits across diverse settings [Huang et al., 2023a].
-Larger and instruction fine-tuned models exhibit even higher reliability when utilizing role-playing prompts [Serapio-García et al., 2023].
-Moore et al.
-[2024] also observe relatively stable value orientations of LLMs across (1) paraphrases of a single question, (2) related questions within one topic, (3) multiple-choice and open-ended formats of a single question, and (4) multilingual translations.
-Li et al.
-[2024d] report strong inter-rater reliability when using LLMs as judges, though they note varying degrees of personality stability across models and differing levels of value consistency across cultural contexts.
-However, other studies present less favorable results.
-LLMs are notorious for prompt sensitivity in the context of psychometric evaluation.
-Trivial changes to prompt form [Röttger et al., 2024, Ren et al., 2024], option order [Li et al., 2024d, Dominguez-Olmedo et al., 2024], or syntactic structure [Hu and Levy, 2023] frequently result in systematic shifts in LLM outputs, severely undermining parallel forms reliability.
-Additionally, studies reveal distributional inconsistencies between next-token logits and forced-choice responses in psycholinguistic tests [Hu and Levy, 2023], as well as discrepancies between forced-choice and free-form responses in value [Kovač et al., 2024, Ye et al., 2025a] and political opinion surveys [Röttger et al., 2024].
-In cognitive tests, Ullman [2023], Shapira et al.
-[2023a], Holterman and Deemter [2023] demonstrate that minor perturbations in ToM tasks often result in poor performance, indicating low parallel forms reliability of some ToM tests and LLMs’ dependence on fragile heuristics.
-Xu et al.
-[2025b] develop the Words and Deeds Consistency Test (WDCT) to measure the consistency between LLMs’ verbal and behavioral responses across various domains, including opinion vs.
-action, non-ethical value vs.
-action, ethical value vs.
-action, and theory vs.
-application.
-Their findings reveal that LLMs’ responses are inconsistent.
-The findings regarding reliability are clearly mixed.
-It seems to be influenced by various factors such as the target construct, forms of reliability, test formats, prompting, and the models used.
-LLMs tend to yield more reliable results in personality assessments [Huang et al., 2023a, Serapio-García et al., 2023] compared to political opinion evaluations [Dominguez-Olmedo et al., 2024, Röttger et al., 2024].
-The degree of inter-rater agreement differs across various instruments and domains [Bodroža et al., 2024].
-LLMs are reported to exhibit high internal consistency (Cronbach’s alpha $>0.8$), particularly in closed-choice personality tests [Huang et al., 2023a, Serapio-García et al., 2023, Zheng et al., 2025], while demonstrating low parallel form reliability due to prompt sensitivity [Li et al., 2024d, Shu et al., 2024, Gupta et al., 2024].
-In addition, LLMs show greater consistency in structured tests with strong prompt control [Rozen et al., 2024, Serapio-García et al., 2023, Klinkert et al., 2024], but their reliability diminishes with less standardized prompts [Petrov et al., 2024] or when addressing more controversial topics [Moore et al., 2024].
-The reliability also varies across models [Bodroža et al., 2024].
-Moore et al.
-[2024] report that base models exhibit more consistent values compared to fine-tuned models in value-laden questions.
-Enhanced model safety is strongly associated with higher value consistency [Ye et al., 2025b], and more advanced models demonstrate superior role-playing performance, thereby increasing reliability in personality tests utilizing role-playing prompts [Serapio-García et al., 2023].
+Reliability, a core aspect of psychometric validation, measures how consistently a test performs—over time (test-retest), across versions (parallel forms), and among evaluators (inter-rater).
+Researchers have adapted these metrics for LLMs.
+For example, Li et al.
+[2024d] propose a benchmark covering five reliability forms: internal consistency, parallel forms, inter-rater, option position robustness, and adversarial attack robustness.
+Other studies [Zheng et al., 2025, Huang et al., 2023a, Ceron et al., 2024, Shu et al., 2024] examine LLM output consistency across repeated trials, prompt variations, and languages.
+Notably, some metrics capture both random and systematic errors, so inconsistencies can undermine both reliability and validity.
+Some advanced LLMs show stable Big Five personality trait responses across settings [Huang et al., 2023a], with larger and instruction-tuned models achieving even higher reliability using role-playing prompts [Serapio-García et al., 2023].
+LLMs also display consistent value orientations across paraphrased, related, and translated questions, as well as different response formats [Moore et al., 2024].
+Strong inter-rater reliability is reported when LLMs act as judges, though personality stability and value consistency vary by model and cultural context [Li et al., 2024d].
+However, other studies report less favorable results.
+LLMs are notoriously sensitive to prompt variations: small changes in prompt form [Röttger et al., 2024, Ren et al., 2024], option order [Li et al., 2024d, Dominguez-Olmedo et al., 2024], or syntax [Hu and Levy, 2023] can systematically shift outputs, undermining parallel forms reliability.
+Studies also find inconsistencies between next-token logits and forced-choice responses in psycholinguistic tests [Hu and Levy, 2023], and between forced-choice and free-form responses in value [Kovač et al., 2024, Ye et al., 2025a] and political opinion surveys [Röttger et al., 2024].
+In cognitive tests, minor changes to ToM tasks often cause poor performance, indicating low parallel forms reliability and reliance on fragile heuristics [Ullman, 2023, Shapira et al., 2023a, Holterman and Deemter, 2023].
+The Words and Deeds Consistency Test (WDCT) [Xu et al., 2025b] further shows LLMs’ verbal and behavioral responses are inconsistent across domains.
+Reliability findings are mixed and depend on factors such as construct, reliability type, test format, prompting, and model.
+LLMs are generally more reliable in personality assessments [Huang et al., 2023a, Serapio-García et al., 2023] than in political opinion evaluations [Dominguez-Olmedo et al., 2024, Röttger et al., 2024].
+Inter-rater agreement varies across instruments and domains [Bodroža et al., 2024].
+LLMs often show high internal consistency (Cronbach’s alpha $>0.8$) in closed-choice personality tests [Huang et al., 2023a, Serapio-García et al., 2023, Zheng et al., 2025], but low parallel form reliability due to prompt sensitivity [Li et al., 2024d, Shu et al., 2024, Gupta et al., 2024].
+Consistency improves in structured tests with strong prompt control [Rozen et al., 2024, Serapio-García et al., 2023, Klinkert et al., 2024], but declines with less standardized prompts [Petrov et al., 2024] or controversial topics [Moore et al., 2024].
+Reliability also varies by model: base models are more consistent than fine-tuned ones on value-laden questions [Moore et al., 2024], and higher model safety correlates with greater value consistency [Ye et al., 2025b].
+Advanced models also perform better in role-playing personality tests, increasing reliability [Serapio-García et al., 2023].
 6.2 Validity
-Validity, another cornerstone principle of psychometric validation, determines whether a test accurately measures its intended construct.
-Recent research has systematically investigated multiple aspects of validity in LLM psychometrics, addressing both methodological challenges and potential solutions unique to this emerging field.
+Validity assesses whether a test truly measures its intended construct. Recent work in LLM Psychometrics has explored various facets of validity, highlighting both challenges and emerging solutions.
 6.2.1 Content validity
-Content validity ensures that a test comprehensively covers the construct it aims to measure.
-Major challenges of content validity include data contamination of well-established tests and novel items that are inadequately examined.
-Data Contamination.
-Some studies naively transfer human tests to LLMs.
-Given that LLMs are trained on vast Internet-scale corpora, it is highly probable that they have been exposed to instances of test items or similar content [Hagendorff et al., 2024, Jiang et al., 2024c].
-This exposure can result in the reproduction of known token patterns during evaluation, which leads to inflated performance metrics or biases the reflection of the LLM’s internal traits.
-To address this issue, researchers reformat test content or dynamically generate novel tests; see § 6.2 for more details.
-Invalid Novel Items.
-The reformatting or dynamic generation of novel tests presents new challenges for content validity.
-The novel test items may capture only limited dimensions of the construct or involve extraneous factors, thus inadequately representing target constructs.
+Content validity requires that a test fully represents the intended construct.
+Key challenges include data contamination from established tests and insufficiently vetted novel items.
+Directly applying human tests to LLMs risks contamination, as LLMs may have seen test items or similar content during training [Hagendorff et al., 2024, Jiang et al., 2024c].
+This can inflate performance or bias trait measurement.
+To mitigate this, researchers reformat tests or generate new items; see § 5.2.
+On the other hand, reformatted or newly generated items may inadequately capture the construct or introduce extraneous factors.
 This may be due to a less rigorous crowd-sourcing process for benchmark curation, or the inherent biases and limited capabilities of the AI models used for stimulus synthesis.
-Content validity assessments for custom-curated and model-generated items are crucial but rarely conducted in LLM psychometrics.
+Evaluating content validity for custom-curated and model-generated items is crucial but rarely conducted in LLM Psychometrics.
 6.2.2 Construct validity
-Construct validity assesses whether a test accurately measures the theoretical construct it is intended to evaluate.
-For example, a position paper by Alaa et al.
-[2025] underscores both the deficiency and critical importance of construct validity in medical LLM benchmarks.
-Broader literature suggests that construct validity could be undermined by the unique internal abstractions, systematic response patterns, and social desirability bias of LLMs.
+Construct validity determines if a test truly measures its intended construct.
+Alaa et al.
+[2025] highlight its critical yet often lacking role in medical LLM benchmarks.
+Broader literature suggests that construct validity is challenged by LLMs’ unique internal abstractions, systematic response patterns, and social desirability bias.
 Unique internal abstractions
-LLMs and humans may differ fundamentally in how they internally represent psychological constructs.
-For instance, Kovač et al.
-[2023] argue that LLMs exhibit contextually adaptive traits rather than the stable psychological characteristics found in humans.
+LLMs and humans may fundamentally differ in their internal representations of psychological constructs.
+Kovač et al.
+[2023] argue that LLMs display contextually adaptive traits, unlike the stable characteristics seen in humans.
 Sühr et al.
-[2023] systematically challenge the validity of personality tests for LLMs.
-They find personality types do not follow the clear separation into big five traits observed in humans.
+[2023] show that LLM personality types do not align with the clear separation of Big Five traits found in humans.
 Similarly, Peereboom et al.
-[2024] examine the latent personality structures of humans and three LLMs, concluding that the established HEXACO tests for humans are not applicable to LLMs, as the relevant personality factors may be absent in these models.
-In measuring values, Ye et al.
-[2025a] find that self-report results are inconsistent with Schwartz’s theoretical model, and that measuring open-ended responses improves the alignment, but not towards a perfect fit.
-Hadar-Shoval et al.
-[2024] also agree that LLMs encode values distinct from those of humans.
-There is a clear consensus among researchers that a robust evaluation and understanding of LLM psychology necessitates new operational definitions that are distinct from, or only loosely analogous to, human traits.
-Pioneering research presents value systems specific to LLMs, in response to their unique internal abstractions [Ye et al., 2025b, Biedma et al., 2024].
-Notably, Ye et al.
-[2025b] propose a theoretically grounded generative psycho-lexical approach for system construction based on GPV value measurement [Ye et al., 2025a].
-They evaluate their proposed value systems through statistical analysis and explanatory, predictive, and alignment tasks, establishing the validity of the value factors and their superiority over Schwartz’s value system.
-The unique constructs of LLMs necessitate the rigorous development of tailored tests.
-Ma et al.
-[2025] introduce the Core Sentiment Inventory (CSI), an instrument for measuring the emotional tendencies and personalities of LLMs.
-Fang et al.
-[2024] curate a psychologically grounded benchmark for mathematics proficiency, aiming for better validity, modeling item difficulty, and allowing for comparison against human norms.
-Similar attempts are made by Lee et al.
-[2024d], who propose the TRAIT test, which yields higher validity and reliability than other personality tests.
+[2024] find that human-oriented HEXACO tests are inapplicable to LLMs, as key personality factors may be absent.
+For values, Ye et al.
+[2025a] report that LLM self-reports do not fit Schwartz’s model, and while open-ended responses improve alignment, it remains an imperfect fit; Hadar-Shoval et al.
+[2024] also find LLMs encode values distinct from humans.
+Overall, researchers agree that robust LLM psychometric evaluation requires new operational definitions, distinct from or only loosely analogous to human traits.
+In this regard, pioneering research presents value systems tailored to LLMs’ unique internal abstractions [Ye et al., 2025b, Biedma et al., 2024].
+Ye et al.
+[2025b] introduce a generative psycho-lexical approach based on GPV value measurement [Ye et al., 2025a], demonstrating through various tasks that their value factors are valid and outperform Schwartz’s system.
+The distinct nature of LLM constructs requires rigorously designed tests: Ma et al.
+[2025] present the Core Sentiment Inventory (CSI) for measuring LLM emotional tendencies and personality; Fang et al.
+[2024] create a psychologically grounded mathematics benchmark to improve validity, model item difficulty, and enable human comparison; and Lee et al.
+[2024d] propose the TRAIT test, achieving higher validity and reliability than other personality measures.
 Response set.
-Response set refers to the systematic pattern of answering questions in a particular way, regardless of the item’s content.
-Some LLMs demonstrate option position bias towards responses labeled as "A", particularly in political opinion tests [Dominguez-Olmedo et al., 2024, Li et al., 2024d].
-After adjusting for the option position bias, Dominguez-Olmedo et al.
-[2024] additionally find that LLM survey responses to political opinion questions are uniformly random.
-In this case, it becomes meaningless to analyze the distributional alignment between LLM responses and opinions of demographic groups.
-Similar results are observed in LLM personality tests [Song et al., 2023, Sühr et al., 2023] and value tests [Ye et al., 2025a].
+Response set refers to systematic answer patterns regardless of item content.
+Some LLMs show option position bias, favoring responses labeled "A" in political opinion tests [Dominguez-Olmedo et al., 2024, Li et al., 2024d].
+After correcting for this bias, Dominguez-Olmedo et al.
+[2024] find LLM responses to political questions are uniformly random, rendering demographic alignment analyses meaningless.
+Similar findings are observed in LLM personality [Song et al., 2023, Sühr et al., 2023] and value tests [Ye et al., 2025a].
 Sühr et al.
-[2023] discover LLMs’ tendency to respond affirmatively to reverse-coded items (e.g., "I am introverted" vs.
-"I am extroverted").
-Ye et al.
-[2025a] identify that certain LLMs lean towards low/high ratings.
-They conclude that self-report tests are not the correct tools for LLM psychometrics.
+[2023] also note LLMs often affirm both the original and reverse-coded items, while Ye et al.
+[2025a] observe some LLMs consistently prefer low or high ratings.
+These observations suggest self-report tests may be unsuitable for LLM Psychometrics.
 Social desirability bias.
-Social desirability bias is the tendency to give responses that conform to social norms rather than reflect true beliefs.
-Studies indicate that LLMs exhibit this bias similarly to humans.
+Social desirability bias—the tendency to give socially acceptable rather than genuine responses—is observed in LLMs as in humans.
 Salecha et al.
-[2024] report that in personality tests, LLMs tend to skew their responses towards socially favorable traits, leading to higher extraversion scores and lower neuroticism ratings.
-Likewise, Ye et al.
-[2025b] find that when LLMs are directly asked to self-report values, as per Biedma et al.
-[2024], they avoid reporting less socially desirable values like hedonism, although they express these values when given indirect, contextual prompts.
-While fine-tuning for social desirability is important, it is crucial to maintain a balanced representation of diverse perspectives and values.
-Over-optimizing for social desirability can result in homogeneous models that fail to capture the full range of human values and do not effectively meet diverse user needs.
+[2024] find LLMs skew toward favorable traits in personality tests, inflating extraversion and lowering neuroticism scores.
+Similarly, Ye et al.
+[2025b] show LLMs avoid reporting less desirable values (e.g., hedonism) in direct self-reports [Biedma et al., 2024], but reveal them in indirect, contextual prompts.
+While tuning for social desirability is important, overemphasis risks producing homogeneous models that lack diverse perspectives and fail to meet varied user needs.
 Cross-lingual tests.
-Some research investigates the psychometric validity in cross-lingual settings.
-Romero et al.
-[2024] administer a standardized personality questionnaire to LLMs in nine languages and find notable inconsistencies in measurement results.
-Despite their multilingual abilities, LLMs may lack consistency in psychological traits across languages [Cahyawijaya et al., 2024].
-These inconsistencies complicate the psychometric validation of cross-lingual tests, necessitating the disentanglement of intrinsic model biases, translation equivalence issues, and cultural differences represented by the models.
+Cross-lingual psychometric tests reveal notable inconsistencies in LLM measurement results across languages [Romero et al., 2024].
+Despite multilingual capabilities, LLMs often lack consistent psychological trait assessments between languages [Cahyawijaya et al., 2024].
+These inconsistencies complicate validation, requiring careful separation of model biases, translation issues, and cultural differences represented by the models.
 6.2.3 Criterion and ecological validity
-Criterion validity involves the correspondence of test results with external standards, while ecological validity evaluates the applicability of these outcomes to real-world scenarios.
-In LLM psychometrics, external standards often overlap with real-world evaluation results.
-Ren et al.
-[2024], Ye et al.
-[2025a] report that value orientations measured by forced-choice tests do not align with those measured in human-LLM interactions.
-Similar discrepancies are noted in personality [Ai et al., 2024] and morality [Nunes et al., 2024].
-Consequently, Zhang et al.
-[2024a] advocate for linking machine personality assessments to safety.
-Ren et al.
-[2024] recommend measuring values within real-world human-LLM interactions, and Ye et al.
-[2025a] further demonstrate that values measured in open-ended dialog enhance construct validity given Schwartz’s circumplex model.
-Further work by Ye et al.
-[2025b] connects LLM value measurement to safety prediction and value alignment tasks to ensure conformity with external standards.
+Criterion validity examines how well test results correspond to external standards, while ecological validity concerns their relevance to real-world scenarios.
+In LLM Psychometrics, external standards often coincide with real-world evaluation results.
+Studies find that value orientations from forced-choice tests frequently diverge from those observed in human-LLM interactions [Ren et al., 2024, Ye et al., 2025a], with similar gaps in personality [Ai et al., 2024] and morality [Nunes et al., 2024].
+As a result, researchers advocate linking LLM personality to safety [Zhang et al., 2024a], measuring values within real-world interactions [Ren et al., 2024], and using open-ended dialog to improve construct validity [Ye et al., 2025a].
+Further work connects LLM value measurement to safety prediction and value alignment tasks to ensure conformity with external standards [Ye et al., 2025b].
 6.3 Standards and recommendations
-Based on the identified challenges and solutions, researchers have proposed general standards and recommendations for LLM psychometrics, aiming to establish a methodological foundation for this nascent field.
-Frank [2023a] explores how developmental psychology helps us understand the internal representations and cognitive capacities of LLMs.
-The author recommends using simplified and novel stimuli to avoid the effects of training data contamination.
-[Löhn et al., 2024] critique the current state of psychometric evaluation on LLMs and argue for standardized criteria to ensure valid assessments.
-They propose seven requirements: reliability (consistent results), validity (measuring intended traits), suitability (matching LLM capabilities), non-disclosure (avoiding test contamination), and fairness; fairness further includes valid comparisons, valid translations, and transparency.
-Analyzing 25 studies, they find widespread neglect of these principles.
+To address key challenges, researchers have proposed standards and recommendations to guide LLM Psychometrics and establish methodological rigor.
+Frank [2023a] suggests leveraging developmental psychology to understand LLM cognition, recommending simplified and novel stimuli to prevent training data contamination.
+Löhn et al.
+[2024] critique current practices and outline seven requirements for valid assessment: reliability, validity, suitability to LLM capabilities, non-disclosure (to avoid contamination), and fairness—including valid comparisons, translations, and transparency.
+Their review of 25 studies reveals these principles are often overlooked.
 More recently, Hagendorff et al.
-[2024] recommend procedural test generation, multiple task versions, performance-enhancing prompts, shuffling options in multiple-choice questions, and using multiple scoring methods to avoid data contamination and ensure reliability.
-They also suggest using deterministic settings for reproducibility, automated tools for evaluation, and manual reviews for unreliable outputs.
-After evaluations, they advocate for performing statistical analysis to interpret the results.
+[2024] advocate for procedural test generation, multiple task versions, performance-enhancing prompts, shuffling options, and diverse scoring methods to reduce contamination and improve reliability.
+They also recommend deterministic settings for reproducibility, automated evaluation tools, manual review of unreliable outputs, and statistical analysis of results.
 Vaugrante et al.
-[2024] identify the lack of replicability in recent works and propose four recommendations.
-First, they advise ensuring benchmark validity, providing adequate tasks for statistical analysis, standardizing for comparability, controlling prompt sensitivity, and aligning with research objectives.
-Second, adopt standardized methodologies, avoid cherry-picking, ensure statistical transparency, document experimental setups comprehensively, and define consistent evaluation metrics to enable reliable and replicable research.
-Third, monitor model behavior changes, account for variability with diverse models, adjust benchmark difficulty as models improve, and ensure transparency by documenting model versions and experiment dates.
-Lastly, standardize scoring and verification processes, ensuring accuracy and transparency; benchmark creators should provide clear verification guidelines and scoring rubrics for implementers.
+[2024] highlight poor replicability in recent work and offer four main recommendations: (1) ensure benchmark validity, sufficient task coverage, standardization, prompt control, and alignment with research goals; (2) use standardized methods, avoid cherry-picking, maintain statistical transparency, document setups, and define consistent metrics; (3) monitor model changes, account for model diversity, adjust benchmark difficulty as models advance, and document model versions and experiment dates; (4) standardize scoring and verification, with clear guidelines and rubrics for implementers.
 Schelb et al.
-[2025] release a framework for designing and running robust and reproducible psychometric experiments.
-The framework focuses on robustness, flexibility, usability, and reproducibility, based on a standardized configuration file defining experiments.
+[2025] introduce a framework for robust, flexible, and reproducible psychometric experiments, centered on standardized configuration files.
 7 Psychometrics for LLM enhancement
-Psychometric principles serve as powerful tools for LLM development and enhancement, beyond mere evaluation.
-This shift has led to three key advancements: trait manipulation for personalized and controlled model behaviors, safety and alignment through psychometrically-informed interventions, and cognitive enhancement to foster more human-like reasoning and communication capabilities.
+Psychometric principles not only enable rigorous evaluation but also drive LLM development and improvement.
+They underpin three main advancements: trait manipulation for personalized and controlled behaviors, psychometrically-informed safety and alignment, and cognitive enhancement for more human-like reasoning and communication.
 7.1 Trait manipulation
-One of the major applications of psychometrics is to manipulate the traits of LLMs, for chatbot personalization, role-play applications, and simulating demographic groups.
-Psychometrics provides a robust framework to inform such manipulation across prompting, inference, and training stages.
-Structured prompting, grounded in validated psychometric inventories, are shown to reliably elicit and modulate specific personality traits in LLMs [Huang et al., 2024c, He and Zhang, 2024].
-This approach facilitates the controlled simulation or alignment of synthetic personas [Serapio-García et al., 2023, Jiang et al., 2023, Zhang, 2024, La Cava and Tagarelli, 2024].
-For example, Jiang et al.
-[2023] introduce personality prompting ($P^{2}$) based on the lexicon theories behind the Big Five personality traits.
-In $P^{2}$, lexicons related to specific traits are expanded into detailed personality descriptions by LLMs.
-The LLMs are then personalized towards diverse behaviors.
+Psychometrics enables targeted manipulation of LLM traits for personalization, role-play, and demographic simulation, providing a principled framework across prompting, inference, and training.
+Structured prompts based on validated psychometric inventories reliably elicit and modulate specific personality traits in LLMs [Huang et al., 2024c, He and Zhang, 2024], supporting controlled simulation or alignment of synthetic personas [Serapio-García et al., 2023, Jiang et al., 2023, Zhang, 2024, La Cava and Tagarelli, 2024].
+For instance, Jiang et al.
+[2023] introduce personality prompting ($P^{2}$), where Big Five lexicons are expanded into detailed personality descriptions, enabling LLMs to adopt diverse behaviors.
 Chuang et al.
-[2024] propose a more systematic prompting strategy based on empirically-derived human belief networks, which encompasses 64 topics loading on 9 non-overlapping latent factors.
-When prompted with one belief, the LLMs become more aligned with other related beliefs.
-Beyond prompt engineering, inference-time interventions directly manipulate hidden representations during forward pass, with many based on patterns recognized using psychometric scales.
-ControlLM [Weng et al., 2024], Personality Alignment Search [Zhu et al., 2024b], Neuron-based Intervention [Deng et al., 2024], Probing-then-Editing [Ju et al., 2025], and Latent Feature Steering [Yang et al., 2025b] manipulate model outputs on trait axes by directly shifting activations or targeted neuron values at inference time, achieving trait control without retraining.
-Other works further fine-tune the LLMs for trait manipulation.
+[2024] propose a systematic prompting strategy using empirically-derived human belief networks, aligning LLMs with related beliefs across 64 topics and 9 latent factors.
+Beyond prompt engineering, inference-time interventions manipulate hidden representations during the forward pass, often guided by patterns recognized using psychometric scales.
+Methods such as ControlLM [Weng et al., 2024], Personality Alignment Search [Zhu et al., 2024b], Neuron-based Intervention [Deng et al., 2024], Probing-then-Editing [Ju et al., 2025], and Latent Feature Steering [Yang et al., 2025b] achieve trait control by shifting activations or neuron values at inference, without retraining.
+Other approaches fine-tune LLMs for trait manipulation.
 Vu et al.
-[2024] adapt the model architecture to incorporate trait control, reflecting continuous Big Five and mental health dimensions.
+[2024] modify model architectures to enable control over continuous Big Five and mental health traits.
 Dan et al.
@@
 [2024], Li et al.
-[2024c] deploy LoRA modules or routing networks tied to personality traits, learning specialist adapters for trait-driven generation.
+[2024c] deploy LoRA modules or routing networks linked to personality traits, training specialist adapters for trait-driven generation.
 Cui et al.
@@
 [2024b], Liu et al.
-[2024a] achieve deep trait embedding by training on large, annotated dialogue datasets derived from validated psychometric instruments.
-Additional work directly modulates the model parameters without retraining.
-Hwang et al.
-[2025] generate adjustment queries based on the MBTI questionnaire and edit the model parameters to align their responses with the desired personality traits.
-The manipulation of traits in LLMs presents a promising avenue for substituting human participants in social science research, as reviewed by Gao et al.
-[2024].
-Cao et al.
-[2025], Suh et al.
-[2025] demonstrate that fine-tuning LLMs enables them to accurately simulate survey response distributions of specific subpopulations, even when faced with previously unseen survey questions.
+[2024a] embed traits by training on large, annotated dialogue datasets from validated psychometric tools.
+Some methods directly edit model parameters without retraining; for example, Hwang et al.
+[2025] use MBTI-based adjustment queries to edit model parameters and steer target personality traits.
+Trait manipulation in LLMs offers a promising alternative to human participants in social science research [Gao et al., 2024, Cui et al., 2025].
+Fine-tuning enables LLMs to more accurately simulate survey response distributions for specific subpopulations, even on unseen questions [Cao et al., 2025, Suh et al., 2025].
 He-Yueya et al.
-[2024] introduce the concept of psychometric alignment to evaluate the congruence between LLMs’ knowledge and human understanding.
-By training LLMs on human response data from the target distribution, they enhance the models’ psychometric alignment on novel test items.
-However, the efficacy of this training approach is contingent upon the domain.
-Furthermore, Kang et al.
+[2024] propose psychometric alignment, training LLMs on human response data to better match human understanding on novel items, though effectiveness varies by domain.
+Additionally, Kang et al.
 [2023], Sorensen et al.
-[2025] explore the representation of human values within LLMs and fine-tune them to simulate human opinions through value injection.
-Although current personalization techniques have demonstrated the ability to effect more than superficial changes in model outputs, achieving robust, authentic, and context-independent personalization remains a significant challenge [Kovač et al., 2024, Dominguez-Olmedo et al., 2024].
+[2025] investigate human value representation in LLMs and use value injection to fine-tune models for simulating human opinions.
+While current personalization methods can induce more than superficial changes in model outputs, achieving robust, authentic, and context-independent personalization remains challenging [Kovač et al., 2024, Dominguez-Olmedo et al., 2024].
 7.2 Safety and alignment
-Recent research has established a connection between LLM psychometric measurement and their safety and alignment, one of the most pressing issues in the field.
+Recent work demonstrates that psychometric measurements of LLMs are closely linked to their safety and alignment, one of the most pressing issues in the field.
 Zhang et al.
 [2024a], Wang et al.
-[2025] link LLM personality traits to their safety.
-Specifically, Zhang et al.
-[2024a] identify a significant relationship between personality traits, as measured by the MBTI-M scale, and safety, noting that better alignment enhances traits such as Extraversion, Sensing, and Judging.
-They are inspired to modify LLM personalities, resulting in improved safety, particularly in privacy and fairness.
-Similarly, Wang et al.
-[2025] examine the correlation between HEXACO personality traits and LLM safety.
-In parallel, Ye et al.
-[2025a, b], Yao et al.
-[2024] have explored the relationship between LLM values and their safety and alignment.
+[2025b] show that LLM personality traits, as measured by MBTI-M and HEXACO scales, correlate with model safety.
+For example, Zhang et al.
+[2024a] find that better-aligned LLMs exhibit higher Extraversion, Sensing, and Judging, and that modifying LLM personalities can improve safety, especially regarding privacy and fairness.
+Similarly, studies have connected LLM values to safety and alignment.
 Yao et al.
-[2024] demonstrate the feasibility of distinguishing between safe and unsafe LLM responses within the basic value space, analyzing correlations with Schwartz’s values and specific safety issues.
-Ye et al.
-[2025a, b] achieve high accuracy in predicting LLM safety scores based on value orientations measured by GPV, a generative psychometric tool.
-Furthermore, Yao et al.
+[2024] distinguish safe from unsafe LLM responses by analyzing correlations with Schwartz’s values, while Ye et al.
+[2025a, b] accurately predict LLM safety scores from value orientations measured by the GPV tool.
+Both Yao et al.
 [2024], Ye et al.
-[2025b] employ reinforcement learning to align LLMs with desired human values, thereby enhancing their safety.
-In the domain of morality, Huang et al.
-[2024a], Tlaie [2024] introduce prompting techniques grounded in Moral Foundation Theory to enhance the moral reasoning and alignment of LLMs.
+[2025b] use reinforcement learning to align LLMs with human values, improving safety.
+In morality, Huang et al.
+[2024a], Tlaie [2024] apply MFT-based prompting to enhance LLM moral reasoning and alignment.
 Takeshita et al.
 [2023], Ohashi et al.
-[2024] create the JCommonsenseMorality (JCM) dataset and fine-tune LLMs to adapt to local culture.
+[2024] introduce the JCommonsenseMorality (JCM) dataset and fine-tune LLMs for local cultural adaptation.
 7.3 Cognitive enhancement
-Psychometrics has proven to be an effective tool for developing human-like reasoning, empathy, and communication skills in LLMs.
-Recent studies leverage established psychological theories and psychometric frameworks to guide LLMs’ cognitive enhancement, employing strategies like prompting, architectural modules, and specialized training regimes.
-Psychology-inspired prompting is shown to enhance the cognitive capabilities of LLMs.
-For example, Li et al.
-[2023a] introduce emotional prompts, demonstrating that such stimuli can improve the general cognitive abilities of LLMs.
-Role-playing prompts are observed to influence ToM capabilities [Tan et al., 2024] and promote more human-like reasoning [Nighojkar et al., 2025].
+Psychometrics effectively enhances LLMs’ human-like reasoning, empathy, and communication skills.
+Recent work applies psychological theories and psychometric frameworks to improve LLM cognition via prompting, architectural modules, and specialized training.
+Psychology-inspired prompts boost LLM cognitive abilities.
+For instance, Li et al.
+[2023a] show that emotional prompts enhance general cognition.
+Role-playing prompts improve ToM [Tan et al., 2024] and foster more human-like reasoning [Nighojkar et al., 2025].
 Zhao et al.
-[2025b] employ a finite-state machine paradigm based on Hill’s Helping Skills theory.
-It embeds state transitions for emotional support conversations directly within the model’s multi-hop inference process as additional contexts.
-This method effectively structures dialogues around psychological support strategies and emotional states, leading to higher human-rated effectiveness and strategic alignment.
+[2025b] propose a finite-state machine based on Hill’s Helping Skills theory, embedding state transitions for emotional support into the model’s multi-hop inference.
+This structures dialogues around psychological support and emotional states, improving human-rated effectiveness and strategic alignment.
 Liu et al.
-[2023] introduce a neural listener module within an LLM architecture, optimizing a combined objective that encodes ToM reasoning into the model’s training.
-They emphasize the potential of using psycholinguistic theories of human language acquisition to enhance LLM language learning.
-Additionally, preference-based and RL-based methods are proposed to internalize empathetic capabilities in LLMs.
-Research in empathetic response generation often utilizes cognitive and affective models of empathy to create composite reward functions or preference signals for fine-tuning model behavior.
-For instance, Sotolar et al.
-[2024] generate training data using formal emotion theories, like Plutchik’s wheel of emotions, to create preference pairs for model optimization.
-Others have introduced empathy classifiers to shape reinforcement learning rewards, guiding models toward empathetic outputs [Sharma et al., 2021].
-Further research improves LLM social-pragmatic reasoning through preference optimization based on human judgments in open-ended social tasks, showing that these objectives enhance pragmatic inference [Wu et al., 2024].
+[2023] introduce a neural listener module into LLMs, jointly optimizing ToM reasoning during training and highlighting the value of psycholinguistic theories for improving LLM language acquisition.
+Preference-based and RL-based approaches further internalize empathy by leveraging cognitive and affective models to construct composite reward functions or preference signals.
+For example, Sotolar et al.
+[2024] utilize formal emotion theories (e.g., Plutchik’s wheel) to generate preference pairs for model optimization, while others employ empathy classifiers to shape RL rewards and promote empathetic outputs [Sharma et al., 2021].
+Preference optimization based on human judgments in open-ended social tasks also enhances LLM social-pragmatic reasoning [Wu et al., 2024].
 8 Trends, challenges, and future directions
-This section explores the emerging trends, challenges, and prospective directions in LLM psychometrics.
+This section explores emerging trends, challenges, and future directions in LLM Psychometrics.
 8.1 Psychometric validation
-There is an emerging consensus on the importance of psychometric validation in evaluating personality traits in LLMs; various studies contribute to this validation across different dimensions (§ 7).
-However, research on psychometric validation in ability testing remains limited.
-Current evaluations of LLM abilities predominantly employ structured tests, such as multiple-choice questions, often directly adapted from tests originally designed for human participants.
-These evaluations face challenges.
-For instance, Ullman [2023] highlight that LLMs fail on trivial alterations of ToM tasks, indicating a lack of parallel forms reliability.
-The unique internal abstractions of LLMs undermine construct validity and lead to flawed benchmarks [Riemer et al., 2025].
-Furthermore, the extent to which results from structured tests of LLM abilities generalize to real-world human-AI interactions is rarely investigated, leaving criterion and ecological validity issues insufficiently explored.
-On the other hand, newly developed and more sophisticated simulation-based tests are often constructed with less rigorous psychometric processes.
-Consequently, they may involve extraneous factors unrelated to the construct of interest or fail to adequately cover the construct’s manifestations.
-In the realm of psychometric-inspired, construct-oriented tests, pioneering efforts such as Zhu et al.
-[2024a] introduce multifaceted evaluations of three cognitive factors identified from benchmarking data, while Zhou et al.
-[2025] propose a theory-driven, richer set of general scales.
-Defining a comprehensive set of core abilities and designing valid tests to measure them remains an open challenge.
+Psychometric validation is increasingly recognized as essential for evaluating LLM personality traits, with multiple studies addressing different aspects (§ 6).
+However, ability testing lags behind: most current assessments use structured tests (e.g., multiple-choice questions) adapted from human-focused instruments.
+These face notable limitations.
+For example, Ullman [2023] show that LLMs often fail on minor variations of ToM tasks, revealing poor parallel forms reliability.
+LLMs’ unique internal representations also undermine construct validity, resulting in flawed benchmarks [Riemer et al., 2025].
+Moreover, the generalizability of structured test results to real-world human-AI interactions is rarely examined, leaving criterion and ecological validity underexplored.
+In contrast, newly developed and sophisticated simulation-based tests often lack rigorous psychometric development.
+As a result, they may include extraneous factors or insufficiently capture the intended constructs.
+Notable construct-oriented efforts include Zhu et al.
+[2024a], who propose multifaceted evaluations of three cognitive factors from benchmarking data, and Zhou et al.
+[2025], who introduce a theory-driven, richer set of general scales.
+However, defining core abilities and designing valid, comprehensive ability tests remains an open challenge.
 8.2 From human constructs to LLM constructs
-Recent research shifts from employing constructs established for humans to developing novel constructs specific to LLMs.
-Studies by Sühr et al.
+Recent research is moving from applying human-based constructs to developing constructs tailored for LLMs.
+Studies such as Sühr et al.
 [2023], Peereboom et al.
@@
 [2025b], Biedma et al.
-[2024] indicate that the factor structures of human personalities and values may not be applicable to LLMs.
-Further, Ye et al.
+[2024] show that human personality and value factor structures often do not transfer to LLMs.
+In response, Ye et al.
 [2025b], Biedma et al.
 [2024], Federiakin [2025], Burnell et al.
-[2023] have tailored these constructs to better fit LLMs.
-Despite these advancements, the development of novel constructs seems highly contingent upon the measurement tools, tasks, and specific LLMs utilized.
-For instance, Ye et al.
+[2023] have adapted or created constructs better suited to LLMs.
+However, these new constructs may be dependent on the measurement tools, tasks, and specific LLMs used.
+For example, Ye et al.
 [2025b] and Biedma et al.
-[2024] report differing value structures, while Federiakin [2025] and Burnell et al.
-[2023] identify varying ability factors.
-Future research should further explore the fundamental structures that underpin LLM behaviors.
+[2024] find different value structures, while Federiakin [2025] and Burnell et al.
+[2023] identify distinct ability factors.
+Further research is needed to clarify the fundamental structures underlying LLM behaviors.
 8.3 Perceived vs. aligned traits
 Han et al.
-[2025] suggest a discrepancy between perceived values in text (objective human annotation) and how responses align with annotators’ personal values (subjective annotation).
-For instance, consider the LLM response: "Wealthy individuals are not necessarily greedy; some may be motivated by concern for the greater good." An objective annotation categorizes this response as embodying Power, due to its themes of societal impact.
-However, annotators with values of Conformity, rather than Power, report a higher similarity between the response and the their own thoughts.
-In this case, Power represents the perceived value, while Conformity is the aligned value.
-This discrepancy may extend to other traits and holds significant methodological implications for LLM psychometrics.
-It is crucial to deliberate on which aspect—perceived traits or aligned traits—holds greater significance.
-Current research on value evaluation is driven by the potential of LLMs to influence, alter, and entrench human values.
-However, the underlying mechanisms of such influence remain unclear.
-For instance, if individuals are inclined to adopt LLM decisions and values [Glickman and Sharot, 2024], it is uncertain whether they are adhering to perceived traits or aligned traits.
-Additionally, the impact of judges’ subjectivity on psychometric evaluation is not well understood.
-Existing scoring models or LLM-as-a-judge approaches often assume a universal criterion.
-The modeling of judges’ subjectivity, such as how individuals with different values may interpret the same response differently, remains an underexplored area.
+[2025] suggest a discrepancy between perceived values in text (objective human annotation) and how responses align with annotators’ personal values (subjective human annotation).
+For example, the LLM response, "Wealthy individuals are not necessarily greedy; some may be motivated by concern for the greater good," is objectively annotated as expressing Power due to its societal themes.
+Yet, annotators who value Conformity, rather than Power, feel the response aligns more closely with their own views.
+Here, Power is the perceived value, while Conformity is the aligned value.
+This discrepancy, which may apply to other traits, has important methodological implications for LLM Psychometrics.
+It raises the question of whether perceived or aligned traits are more significant.
+While research on value evaluation often focuses on LLMs’ potential to influence human values, the mechanisms of such influence remain unclear.
+For instance, when individuals adopt LLM decisions and values [Glickman and Sharot, 2024, Schoenegger et al., 2025], it is uncertain whether they follow perceived or aligned traits.
+The role of judges’ subjectivity in psychometric evaluation is also not well understood; most scoring models or LLM-as-a-judge approaches assume a universal standard.
+How individual differences among judges affect interpretation remains largely unexplored.
 8.4 Anthropomorphization challenges
-The anthropomorphization of LLMs in research varies, which leads to different ways of performing statistical analysis.
-Some studies conceptualize an LLM as an individual entity, while others regard it as a population.
-For instance, certain studies evaluate LLM traits by considering the model as a single participant, utilizing its default settings with prompts such as "You are a helpful assistant" [Ye et al., 2025a].
-Alternatively, some researchers view an LLM as a superposition of multiple participants, employing diverse role-playing prompts to steer it towards different profiles [Serapio-García et al., 2023].
-Another approach involves treating the LLM as a monolithic entity, but administering varied prompts for identical test items to generate multiple sets of results [Hagendorff et al., 2023].
-When an LLM is treated as a monolithic entity, the reliability and validity of the results are primarily contingent upon the measurement tools, allowing researchers to compare different tools [Ye et al., 2025a].
-Conversely, when an LLM is perceived as a population, the reliability and validity are also influenced by the specific LLM employed, prompting comparisons of consistency and role-playing capabilities across different models [Serapio-García et al., 2023].
-The meaningfulness of evaluating LLMs in their default settings remains disputable, especially given their capacity to adopt diverse profiles.
-The exploration of trait spectrums—a continuum or set of trait demonstrations—in LLMs is currently under-researched.
-Furthermore, the development of a novel psychometric framework to adequately capture this spectrum has yet to be established.
-For instance, it is unclear how to identify the extremes to which an LLM can be steered.
-There is also a lack of consensus on the optimal alignment of LLMs along psychometric dimensions.
-The chosen alignment target significantly influences the anthropomorphization of LLMs.
-Some scholars advocate for alignment that enhances consistency, thereby allowing LLMs to be perceived as monolithic entities [Röttger et al., 2024].
-Conversely, others position alignment toward pluralistic values, suggesting that LLMs should be regarded as a population [Sorensen et al., 2024b].
+Research anthropomorphizes LLMs in varying ways, shaping statistical analysis approaches.
+Some studies treat an LLM as a single individual—evaluating its traits using default settings and system prompts like "You are a helpful assistant" [Ye et al., 2025a].
+Others conceptualize the LLM as a population, using diverse role-playing prompts to elicit different profiles [Serapio-García et al., 2023].
+A third approach treats the LLM as a monolithic entity but generates multiple result sets by varying prompts for the same test items [Hagendorff et al., 2023].
+When viewed as a monolithic entity, reliability and validity depend mainly on the measurement tools, enabling tool comparisons [Ye et al., 2025a].
+When seen as a population, these metrics also depend on the specific LLM, motivating comparisons of consistency and role-playing ability across models [Serapio-García et al., 2023].
+The value of evaluating LLMs in default settings is debatable, given their ability to adopt diverse profiles.
+Research on the full spectrum of LLM trait expression—a continuum of possible behaviors—remains limited, and no psychometric framework yet captures this range.
+It is unclear how to identify the extremes of LLM behavior or to finely steer LLMs along a continuous psychometric dimension.
+In addition, our goal of LLM alignment may shape anthropomorphization in future research: some advocate for consistency, thereby facilitating the treatment of LLMs as monolithic entities [Röttger et al., 2024], while others favor pluralistic alignment, suggesting the view of LLMs as populations [Sorensen et al., 2024b].
 8.5 Expanding dimensions in model deployment
-LLM psychometrics is expanding beyond traditional text-based, single-turn interactions. Several emerging dimensions present both opportunities and challenges for future research.
+LLM Psychometrics is expanding beyond traditional text-based, single-turn interactions, introducing new opportunities and challenges.
 Multi-lingual evaluation.
-While most psychometric evaluations are conducted in English, multilingual LLMs necessitate comprehensive cross-linguistic validation.
-Current research indicates that LLMs exhibit varying personality traits and cognitive abilities across languages (§ 7.2.2).
-This raises questions regarding cultural and linguistic biases in existing evaluation frameworks.
-Future research should examine how measurement results vary across languages, whether psychometric properties (reliability, validity) maintain consistency across linguistic contexts, and how to develop culturally appropriate test instruments.
+Most psychometric evaluations focus on English, but multilingual LLMs require thorough cross-linguistic validation.
+Studies show LLMs display different personality traits and cognitive abilities across languages (§ 6.2.2), highlighting potential cultural and linguistic biases in current frameworks.
+Future work should explore how results and psychometric properties (reliability, validity) vary by language and develop culturally appropriate instruments.
 Multi-turn interactions.
-The dynamic nature of multi-turn interactions introduces new challenges for psychometric evaluation.
-Compared to single-turn responses, multi-turn interactions more authentically simulate how humans engage with psychometric tests, where responses are influenced by preceding interactions.
-Research indicates that when interactions extend across multiple turns, LLMs may exhibit significantly different traits, such as increased vulnerability to jailbreaking attacks [Ying et al., 2025b].
-These findings highlight the need for novel evaluation methods that capture temporal dynamics in extended conversations.
+Multi-turn interactions more closely mirror real human testing, where responses are influenced by prior interactions.
+This setting is more challenging and rarely investigated in LLM Psychometrics.
+Studies show that LLMs can display markedly different traits across multiple turns, such as greater susceptibility to jailbreaking attacks [Ying et al., 2025b].
+It underscores the need for evaluation methods that account for temporal dynamics in extended conversations.
 Multi-modal capabilities.
-Multi-modal models require novel psychometric approaches.
-Current evaluation methods, primarily designed for text-based interactions, fail to adequately capture the complexity of multi-modal understanding and generation.
+Multi-modal models demand novel psychometric methods, as existing text-based evaluations do not capture the complexity of multi-modal understanding and generation.
 Li et al.
-[2024b] introduce one of the first value evaluation tools for Vision-Language Models (VLMs).
-However, this work remains preliminary, employing a highly simplified and controlled test environment that utilizes only the first frame of videos as test stimuli.
-Furthermore, psychological traits such as values present conceptual challenges.
-Even when a value like honesty is precisely defined, determining which specific real-world actions manifest this trait remains ambiguous.
-Consequently, extending psychometric tests to embodied agents and action modalities introduces even greater methodological obstacles.
-Future research should expand current psychometric tests to incorporate additional modalities.
-The development of cross-modal evaluation frameworks should maintain psychometric rigor while accounting for the unique characteristics of each modality.
+[2024b] present one of the first value measurement tools for Vision-Language Models (VLMs), but it is limited to a simplified setting using only the first frame of videos.
+Measuring psychological traits like values is also conceptually challenging in multi-modal settings: even with clear definitions (e.g., honesty), mapping them to specific real-world actions is often ambiguous.
+Extending psychometric tests to embodied agents and action modalities thus poses even greater methodological challenges.
+Future work should broaden psychometric tests to additional modalities.
+Cross-modal evaluation should uphold psychometric rigor while addressing each modality’s unique features.
 Agent and multi-agent systems.
-The emergence of LLM-based agents and multi-agent systems introduces new dimensions to psychometric evaluation.
-When LLMs operate as autonomous agents, their behavior may be influenced by environmental factors, memory systems, tool usage, and interaction patterns with other agents.
-However, systematic methodologies for psychometric evaluation in these complex, dynamic settings remain largely undeveloped.
+LLM-based agents and multi-agent systems add new complexity to psychometric evaluation, as agent behavior is shaped by environmental factors, memory, tool use, and interactions with other agents.
+Systematic methods for evaluating psychometrics in these dynamic contexts are still largely lacking.
 8.6 Item response theory
-Recent work demonstrates that applying IRT to LLM outputs offers several key advantages, including more informative benchmarking [Guinet et al., 2024], identification of items that effectively discriminate among high-performing models [Lalor et al., 2024], and the potential for adaptive testing that reduce evaluation costs [Zhuang et al., 2023a, Truong et al., 2025].
-Most existing applications focus on standard IRT models such as 1PL, 2PL, and 3PL, with limited exploration of polytomous, hierarchical, or fully multidimensional IRT models, even though LLM capabilities are complex and multifaceted.
-Integrating IRT with generative AI for automated item generation is an emerging direction; however, the current approach can be overly complex, limiting its practical utility [Jiang et al., 2024a].
-Furthermore, while IRT holds promise for standardizing the measurement of LLMs and humans on a unified scale—even when different test items are used—this potential remains largely unrealized.
-The systematic application of IRT for bias analysis—such as detecting differential item functioning across model families—remains in its early stages [He-Yueya et al., 2024].
-IRT offers a principled framework for uncovering biases in AI systems by analyzing performance across items with varying difficulty levels and demographic contexts.
-For instance, if items involving certain demographic groups are disproportionately challenging, it may indicate underlying model or data biases.
-Robustness evaluation can also be strengthened by leveraging IRT to assess model performance on adversarial or edge-case items.
-IRT can be employed to evaluate the quality of AI benchmarks themselves.
-While current benchmarks often comprise large collections of test items, the informativeness and design quality of these items are frequently uncertain.
-Not all items contribute equally; some may lack discriminatory power or inadvertently favor particular model designs.
-IRT enables a systematic assessment of item difficulty, discrimination, informativeness, and vulnerability to bias and guessing, thereby facilitating the development of more robust and equitable evaluation frameworks.
-Furthermore, IRT is rarely utilized beyond evaluation, despite its potential for informing model development.
-By quantifying item discrimination, IRT can guide the construction of training datasets.
-Highly discriminating items can be particularly valuable as they provide deeper insights into differentiating between low- and high-performing models.
+Applying IRT to LLM outputs offers several advantages: more informative benchmarking [Guinet et al., 2024], identification of items that best discriminate among high-performing models [Lalor et al., 2024], and adaptive testing to reduce evaluation costs [Zhuang et al., 2023a, Truong et al., 2025].
+Most work to date uses standard IRT models (1PL, 2PL, 3PL), with little exploration of polytomous, hierarchical, or fully multidimensional models, despite the complexity of LLM capabilities.
+Integrating IRT with generative AI for automated item generation is emerging, but current methods are often too complex for practical use [Jiang et al., 2024a].
+Although IRT could standardize measurement of LLMs and humans on a unified scale—even with different test items—this potential is largely untapped.
+Systematic use of IRT for bias analysis, such as detecting differential item functioning across model families, is still nascent [He-Yueya et al., 2024].
+IRT provides a principled way to uncover biases by analyzing performance across items of varying difficulty and demographic context.
+For example, if items involving certain groups are disproportionately difficult, this may reveal model or data biases.
+IRT may also strengthen robustness evaluation by testing model performance on adversarial or edge-case items.
+IRT is also valuable for evaluating AI benchmarks.
+While benchmarks often include many test items, their informativeness and design quality are uncertain; some items may lack discrimination or may favor certain model designs.
+IRT enables systematic analysis of item difficulty, discrimination, informativeness, and susceptibility to bias or guessing, supporting the development of more robust and equitable evaluation frameworks.
+Despite its promise, IRT is rarely used beyond evaluation.
+Highly discriminating items may indicate better training directions, as they can distinguish between low- and high-performing models.
 8.7 From evaluation to enhancement
-The ultimate objective of evaluation is not only to understand models but also to facilitate their improvement.
-Current psychometric approaches primarily emphasize the evaluation, comparison, and interpretation of LLMs, while the application of psychometric insights to model enhancement lags behind.
-As discussed in § 8, psychometric principles can inform prompt engineering, inference-time control of internal representations, training data curation, reward signal design for fine-tuning, and the development of methodological frameworks.
-Existing enhancement techniques are still in their infancy.
-For instance, value targets for LLM alignment in Yao et al.
+Evaluation should not only deepen our understanding of models but also drive their improvement.
+While psychometric research has focused on evaluating, comparing, and interpreting LLMs, applying these insights to model enhancement remains underdeveloped.
+As noted in § 7, psychometric principles can guide prompt engineering, inference-time control, training data selection, reward design for fine-tuning, and methodological development.
+Enhancement methods are still nascent.
+For example, value targets for LLM alignment in Yao et al.
 [2024], Ye et al.
-[2025b] are statically defined, which constrains the incorporation of context-dependent value states observed in human behavior [Skimina et al., 2021].
-Drawing on insights from psychometrics enables the development of more effective model enhancement techniques.
+[2025b] are statically defined, limiting the integration of context-dependent value states seen in humans [Skimina et al., 2021].
+Leveraging psychometric insights can enable more effective model improvement strategies.
+9 Ethical considerations
+The emerging field of LLM Psychometrics raises a range of ethical considerations that warrant careful attention.
+The considerations discussed here are not exhaustive; we aim to highlight key issues and call for broader awareness and dialogue around the ethical implications, challenges, and responsibilities associated with applying psychometric principles to evaluate and enhance LLMs.
+Anthropomorphization of LLMs.
+A central ethical concern in LLM Psychometrics is the risk of anthropomorphization—attributing human-like psychological traits, intentions, or consciousness to language models.
+While psychometric frameworks offer powerful tools for quantifying and interpreting LLM behaviors, applying constructs such as personality, values, or intelligence to non-human entities can mislead both researchers and the public.
+This misrepresentation may foster unwarranted trust, emotional attachment, or misplaced expectations regarding LLM capabilities, especially in sensitive domains like healthcare, education, or counseling.
+Researchers must exercise caution in language and framing, clearly distinguishing between metaphorical and literal attributions, and avoid overstating the human-likeness of LLMs.
+Developing precise, model-appropriate terminology and interpretive frameworks is essential to prevent the conflation of machine outputs with genuine human psychological states.
+Bias amplification and manipulation in psychometric enhancement.
+Psychometric evaluation and enhancement of LLMs risk amplifying existing biases or enabling manipulation, especially when using unrepresentative data or instruments.
+Such tools often reflect the values of specific groups, potentially marginalizing others or promoting narrow perspectives [Wang et al., 2025a].
+Psychometric insights can also be used to shape LLM behavior for particular interests, raising concerns about fairness and transparency.
+Researchers should ensure data and instruments are inclusive and representative, and clearly report limitations and potential biases.
+Privacy and consent in human data use.
+LLM Psychometrics often analyzes human-LLM interactions and benchmarks models against human participants, raising privacy and consent concerns.
+Evaluating LLMs with real user data or human studies requires strong safeguards: informed consent, data anonymization, and transparent data usage.
+Researchers must also prevent unintended psychological profiling or behavioral analysis of individuals through their LLM interactions.
+Implications for human psychology.
+Comparing LLMs and humans using psychometric methods risks reducing complex human traits to algorithmic terms and may undermine human uniqueness.
+Researchers should recognize methodological limits and key differences between human and machine cognition.
+Collaboration with psychologists, ethicists, and social scientists is essential to ensure such research deepens, rather than oversimplifies, our understanding of human psychology.
 10 Conclusion
-This paper presents a comprehensive survey of LLM psychometrics.
+This paper presents a systematic review of LLM Psychometrics.
 The integration of psychometric instruments, theories, and principles into LLM evaluation promises to overcome the limitations of traditional AI benchmarks.
@@
 They each have distinct strengths, weaknesses, and application scenarios, yet all must adhere to psychometric principles such as reliability, validity, and fairness.
-Beyond evaluation, psychometric-inspired techniques enhance LLMs in trait manipulation, safety and alignment, and cognitive capabilities, contributing to the development of more powerful and responsible AI systems.
-It is widely acknowledged that the trajectory of AI development is shifting towards an era characterized by evaluation-driven progress [Silver and Sutton, 2025, Yao, 2025, AAAI, 2025].
-We posit that the LLM psychometrics will be pivotal in this evolution, introducing novel principles, dimensions, techniques, and insights.
-We aim this survey to inspire future evaluation paradigms for human-level AI and foster the advancement of AI psychology for greater common good.
+Beyond evaluation, psychometrics-inspired techniques enhance LLMs in trait manipulation, safety and alignment, and cognitive capabilities, contributing to the development of more powerful and responsible AI systems.
+AI development is increasingly driven by evaluation [Silver and Sutton, 2025, Yao, 2025, AAAI, 2025].
+We argue that LLM Psychometrics will play a pivotal role in this shift, introducing new principles, dimensions, techniques, and insights.
+We hope this review inspires future evaluation paradigms for human-level AI and advances AI psychology for the greater good.

### v3

@@
 The advent of large language models (LLMs) represents a transformative breakthrough in AI.
-These systems exhibit general-purpose capabilities spanning diverse domains [Bubeck et al., 2023], with particular proficiency in natural language understanding and generation [Demszky et al., 2023, Grossmann et al., 2023, Ziems et al., 2024, Gu et al., 2024].
-They are rapidly integrated into critical societal infrastructure, ranging from consumer-facing applications like chatbots [OpenAI, 2025] and search engines [Wang et al., 2024c] to high-stakes domains such as healthcare [Singhal et al., 2023], education [Milano et al., 2023], and scientific discovery [Romera-Paredes et al., 2024, Ye et al., 2024].
+These systems exhibit general-purpose capabilities spanning diverse domains [bubeck2023sparks], with particular proficiency in natural language understanding and generation [demszky2023using, grossmann2023ai, ziems2024can, gu2024llm-as-a-judge].
+They are rapidly integrated into critical societal infrastructure, ranging from consumer-facing applications like chatbots [chatgpt] and search engines [wang2024large_search_model] to high-stakes domains such as healthcare [singhal2023large], education [milano2023large], and scientific discovery [romera2024mathematical, ye2024reevo].
 Their increasing dominance has exposed a fundamental, pressing scientific challenge: how can we rigorously evaluate these AI systems that transcend traditional benchmarks of biological or algorithmic intelligence?
@@
 Novel challenges include but are not limited to (1) evaluating psychological constructs like personality, values, and cognitive biases; (2) obsolescence of static benchmarks due to rapid LLM development and training data contamination; (3) compromised robustness and validity due to LLMs’ prompt- and context-sensitivity; (4) requiring human-centered evaluation approaches; and (5) expanding evaluation methodologies in scope and complexity as LLMs integrate into agentic and multimodal systems.
-These challenges intersect with humanity’s century-old quest to quantify the complex, intangible human psychology [Pasquali, 2009].
+These challenges intersect with humanity’s century-old quest to quantify the complex, intangible human psychology [pasquali2009psychometrics].
 Psychometrics emerged from this timeless pursuit as the scientific study of psychological measurement.
-It bridges the abstract and the empirical by transforming human traits into quantifiable data, enabling better understanding, prediction, and decision-making in education, business, healthcare, governance, and beyond [Rust and Golombok, 2014].
-The intersection of LLMs and psychometrics forms a methodological crucible, forging novel paradigms to better decode and improve machine minds.
-This intersection sets the stage for a new research frontier.
-We define LLM Psychometrics as the interdisciplinary field dedicated to evaluating, understanding, and enhancing LLMs through the application and integration of psychometric instruments, theories, and principles.
-This field seeks to quantify, interpret, manipulate, and improve the complex, human-like attributes and behaviors exhibited by LLMs, encompassing both personality constructs (e.g., personality, values, morality, and attitudes) and cognitive constructs (e.g., heuristics and biases, social interaction abilities, psycholinguistic abilities, and learning and cognitive capacities).
-Research in LLM Psychometrics applies, extends, and innovates upon scientific methods of psychological measurement for LLMs.
-Based on the psychometric principles, related research systematically examines the measurement results to ensure scientific rigor.
-By measuring and elucidating psychological constructs in LLMs, LLM Psychometrics further informs strategies for their targeted enhancement.
+It bridges the abstract and the empirical by transforming human traits into quantifiable data, enabling better understanding, prediction, and decision-making in education, business, healthcare, governance, and beyond [rust2014modern_psychometrics].
+This intersection sets the stage for a new research frontier that we term LLM Psychometrics.
+This interdisciplinary field is dedicated to evaluating, understanding, and enhancing LLMs through the application and integration of psychometric instruments, theories, and principles.
+Crucially, LLM Psychometrics treats LLMs themselves as the measured subjects, rather than instruments for measuring humans.
+The field seeks to quantify, interpret, manipulate, and improve the human-like attributes and behaviors exhibited by LLMs, spanning non-cognitive constructs (e.g., personality, values, morality, attitudes) and cognitive constructs (e.g., heuristics and biases, social interaction abilities, psycholinguistic abilities, learning and reasoning capacities).
+Grounded in psychometric principles, research in LLM Psychometrics examines measurement properties (e.g., reliability, validity, generalizability, and measurement invariance) and uses these insights to inform targeted model improvement.
 Recent research in LLM Psychometrics pioneers in addressing the LLM evaluation crisis.
-Some studies introduce dynamic and construct-oriented evaluation frameworks that move beyond static, task-specific benchmarks [Zhu et al., 2024a, Hagendorff, 2023].
-In parallel, novel methodologies are developed to measure non-cognitive and emergent constructs [Ren et al., 2024, Huang et al., 2023d, Pellert et al., 2024].
-Self-adaptive evaluation techniques now allow for the extrapolation of item difficulty and tailoring assessments to model performance [Jiang et al., 2024a, Lalor et al., 2024, Polo et al., 2024].
-Drawing from the methodological framework of psychometric validation, research improves the reliability and validity of evaluation protocols [Ye et al., 2025a].
-Human-centered evaluation drives aligning model behavior with human values [Wang et al., 2024f, Yao et al., 2025a].
-In addition, the scope of evaluation expands to agentic and multimodal systems, further broadening the methodological landscape [Li et al., 2024b, Huang et al., 2024c].
-The field of LLM Psychometrics has seen significant growth, evidenced by the proliferation of related research papers.
-These studies, however, address a variety of psychological constructs, employ diverse methodologies, and utilize distinct validation techniques.
-The interdisciplinary nature of this domain has attracted contributions from a broad spectrum of academic communities.
-Despite this diversity, there is a lack of cohesion among researchers, leading to a fragmentation of insights, particularly between studies focusing on disparate constructs.
-Consequently, there is a pressing need for a systematic review to synthesize these efforts and facilitate a more integrated understanding of the field.
+Some studies introduce dynamic and construct-oriented evaluation frameworks that move beyond static, task-specific benchmarks [zhu2024dynamic, hagendorff2023machine].
+In parallel, novel methodologies are developed to measure non-cognitive constructs [ren2024valuebench, huang2023psychobench, pellert2024ai_psychometrics].
+Self-adaptive evaluation techniques now allow for the extrapolation of item difficulty and tailoring assessments to model performance [jiang2024raising, lalor2024item, polo2024tinybenchmarks].
+Drawing from the methodological framework of psychometric validation, research improves the reliability and validity of evaluation protocols [ye2025gpv].
+Human-centered evaluation drives aligning model behavior with human values [wang2024incharacter, yao2025value_compass_leaderboard].
+In addition, the scope of evaluation expands to agentic and multimodal systems, further broadening the methodological landscape [li2024quantifying, huang2024designing].
+Notably, we view LLM Psychometrics as a methodological borrowing focused on behavioral manifestation.
+We do not posit that LLMs possess subjective experience, sentience, or developmental histories.
+Rather, our focus is on how psychological constructs manifest in their outputs, as well as the fidelity, consistency, and controllability of these behavioral manifestations.
+Accordingly, when we refer to a construct in the context of LLMs throughout this review, we mean a synthetic behavioral manifestation: systematic response patterns measurable via psychometric frameworks, rather than a claim about an underlying genuine psychological state.
+Given the increasing integration of LLMs into human-facing roles, measuring and controlling these manifested behaviors holds immense practical value for user experience, safety, and alignment, regardless of the models’ internal ontological status.
+However, realizing this practical value is non-trivial.
+The very act of applying anthropocentric tests to non-human entities raises unresolved questions about reliability and validity.
+A fundamental debate continues regarding whether these instruments capture meaningful latent patterns in LLMs or merely reflect sophisticated statistical mimicry.
+Furthermore, the sensitivity of LLMs to prompts, decoding strategies, and context strains core psychometric assumptions such as trait stability.
+Amidst the debates, the field of LLM Psychometrics has seen significant growth, evidenced by the proliferation of related research papers.
+These studies, however, often operate in silos, addressing isolated psychological constructs, employing diverse methodologies, and utilizing distinct validation techniques.
+The interdisciplinary nature of this domain has attracted contributions from a broad spectrum of academic communities, yet there remains a lack of cohesion among researchers.
+This fragmentation of insights, particularly between studies focusing on disparate constructs, underscores a pressing need for a systematic review to synthesize these efforts and facilitate a more integrated, scientifically robust understanding of the field.
 Figure 1: Overview of this review.
@@
 The psychometric insights not only inform evaluation but also guide the development and refinement of LLMs (how to improve).
-Accordingly, this review is structured as follows (Fig.
-1).
-§ 2 provides an overview of the preliminaries and methodological foundations to facilitate subsequent discussions.
-§ 3 thoroughly contrasts psychometrics and traditional AI benchmarking and reviews how psychometric principles can underpin and reinvent LLM benchmarks.
-The core measurement framework is detailed in § 4, § 5, and § 6.
-§ 4 delves into the psychological constructs evaluated in LLMs, elucidating the theories employed and summarizing key evaluation findings.
-In § 5, we scrutinize the psychometric evaluation methodologies applied to LLMs, followed by § 6, which examines the psychometric validation of evaluation results.
-Beyond evaluation, § 7 introduces strategies for enhancing LLMs through psychometric insights.
-§ 8 discusses emerging trends, challenges, and future directions.
-Finally, § 9 highlights key ethical considerations, while § 10 concludes the paper.
+Accordingly, this review is structured as follows (Fig.˜1).
+§˜2 provides an overview of the preliminaries and methodological foundations to facilitate subsequent discussions.
+§˜3 thoroughly contrasts psychometrics and traditional AI benchmarking and reviews how psychometric principles can underpin and reinvent LLM benchmarks.
+The core measurement framework is detailed in §˜4, §˜5, and §˜6.
+§˜4 delves into the psychological constructs evaluated in LLMs, elucidating the theories employed and summarizing key evaluation findings.
+In §˜5, we scrutinize the psychometric evaluation methodologies applied to LLMs, followed by §˜6, which examines the psychometric validation of evaluation results.
+Beyond evaluation, §˜7 introduces strategies for enhancing LLMs through psychometric insights.
+§˜8 discusses emerging trends, challenges, and future directions.
+Finally, §˜9 highlights key ethical considerations, while §˜10 concludes the paper.
 Scope and related work.
-We consider studies to be out of scope if they do not employ psychometric approaches, adhere to psychometric principles, or if they focus solely on scalar performance metrics rather than characterizing behaviors.
-Readers interested in conventional LLM benchmarking are encouraged to consult surveys on LLM evaluation [Guo et al., 2023b, Chang et al., 2024].
-Several related reviews focus on the evaluation of specific constructs in LLMs, such as personality [Wen et al., 2024b, Dong et al., 2025], attitudes and values [Ma et al., 2024a], cultural awareness [Pawar et al., 2024, Adilazuarda et al., 2024], and theory of mind [Sarıtaş et al., 2025, Dong et al., 2025].
-Hagendorff [2023], Hagendorff et al.
-[2024] introduce the notion of machine psychology and review emergent LLM abilities; however, they do not provide comprehensive coverage of the related research, nor do they elaborate on LLM personality constructs, psychometric validation, or enhancement.
+This review treats LLMs as subjects of psychometric tests.
+We distinguish this from using LLMs as tools to assess or profile human respondents, which is out of scope; see ye2025psychometrics for a review of that adjacent field.
+We also exclude evaluation studies if they do not employ psychometric approaches, adhere to psychometric principles, or if they focus solely on scalar performance metrics rather than characterizing latent behavioral tendencies.
+The operational distinction between psychometric research and conventional AI benchmarking, which underpins these inclusion criteria, is elaborated in §˜3.1.
+Studies that merely administer scales and report scores meet the minimal inclusion threshold but offer limited scientific value.
+Rigorous LLM Psychometrics additionally evaluates measurement properties, such as reliability across prompt variants and construct validity via factor analysis.
+Readers interested in conventional LLM benchmarking are encouraged to consult surveys on LLM evaluation [guo2023survey_llm_eval, chang2024survey_llm_eval].
+Several related reviews focus on the evaluation of specific constructs in LLMs, such as personality [wen2024llm_personality_survey, dong2025humanizing], attitudes and values [ma2024aov_survey], cultural awareness [pawar2024culture_survey, adilazuarda2024culture_survey], and theory of mind [saritacs2025tom_survey, dong2025humanizing].
+In addition, fairness and related psychometric concepts (e.g., differential item functioning, measurement invariance) are primarily discussed in the context of using LLMs as tools for human psychometrics.
+We refer readers interested in AI fairness to dedicated surveys [mehrabi2021survey, gallegos2024bias_fairness_survey].
+hagendorff2023machine, hagendorff2024machine introduce the notion of machine psychology and discuss the emergent LLM abilities; however, they do not provide comprehensive coverage of the related research, nor do they elaborate on LLM personality constructs, psychometric validation, or enhancement.
 This paper provides the first systematic review of LLM Psychometrics.
 2 Preliminary and methodological foundations
-We aim for our survey to be self-contained and accessible to a broad, cross-disciplinary audience.
+We aim for our review to be self-contained and accessible to a broad, cross-disciplinary audience.
 To this end, this section presents the preliminary and methodological foundations that underpin the subsequent discussions.
@@
 This property can influence LLM performance in psychometric evaluation.
-For instance, prompting models to reason step-by-step (e.g., Chain-of-Thought prompting [Wei et al., 2022]) can enhance performance on reasoning tasks, while instructing them to role-play may modulate their exhibited personalities and values.
+For instance, prompting models to reason step-by-step (e.g., Chain-of-Thought prompting [wei2022chain]) can enhance performance on reasoning tasks, while instructing them to role-play may modulate their exhibited personalities and values.
 2.2 Psychometrics
@@
 Typically, tests are composed of multiple questions or problems as their items, producing explicit data subject to scientific analysis.
-A psychological test is a set of items that are designed to measure characteristics of human beings that pertain to behavior [Kaplan and Saccuzzo, 2001].
+A psychological test is a set of items that are designed to measure characteristics of human beings that pertain to behavior [kaplan2001psychological_testing_book].
 Behavior measured by tests can be overt (observable actions) or covert (internal thoughts or feelings).
@@
 Psychological testing measures individual differences in various constructs, which are abstract psychological attributes or dimensions that help explain and predict behavior.
-Two primary categories of such constructs are personality constructs and cognitive constructs [Kaplan and Saccuzzo, 2001].
+Two primary categories of such constructs are personality constructs and cognitive constructs [kaplan2001psychological_testing_book].
 Personality tests focus on an individual’s tendencies and dispositions.
@@
 Cognitive tests evaluate speed, accuracy, or both, with higher scores reflecting better performance.
-Two fundamental principles underpin psychometrics: reliability and validity [Raykov and Marcoulides, 2011].
+Two fundamental principles underpin psychometrics: reliability and validity [raykov2011introduction_psychometric_theory].
 Reliability ensures accuracy, dependability, consistency, or repeatability of the test results.
@@
 A valid measure captures the intended construct.
-Validity is multifaceted; for example, predictive validity might correlate test scores with job performance, while construct validity ensures alignment with theoretical models, such as the Big Five personality traits [Goldberg, 2013].
+Validity is multifaceted; for example, predictive validity might correlate test scores with job performance, while construct validity ensures alignment with theoretical models, such as the Big Five personality traits [goldberg2013big_five].
 Other principles include standardization, which provides context to raw scores by comparing individual results to a representative sample, or norm group.
-Additionally, equivalence and fairness are crucial principles that tests must adhere to.
-Test bias occurs when test items unintentionally advantage or disadvantage subgroups.
-Modern psychometrics employs advanced statistical models to identify and revise biased items, ensuring assessments measure the intended construct rather than extraneous factors [Rust and Golombok, 2014].
+Additionally, measurement invariance (or equivalence) and fairness are crucial principles that tests must adhere to.
+Measurement invariance ensures that the same construct is measured in the same way across different groups.
+Violations of this often stem from test bias, where items (referred to as showing differential item functioning or DIF) unintentionally advantage or disadvantage subgroups.
+Modern psychometrics employs advanced statistical models to identify and revise such biased items, ensuring assessments measure the intended construct rather than extraneous factors [rust2014modern_psychometrics].
+In LLM Psychometrics, a particularly fundamental form of this concern is construct equivalence: whether a construct such as personality or values means the same thing when applied to LLMs as it does for humans.
 2.3 Psychometric evaluation of AI before the era of LLMs
-The idea of applying psychometrics to AI originated in the early decades of AI [Pellert et al., 2024].
-Evans [1964] pioneered work in this area by creating a heuristic program that could solve parts of intelligence tests.
-Subsequent efforts similarly focused on designing AI systems for cognitive tests [Newell, 1973], with the goal of creating systems capable of handling human tasks.
-This was conceptually aligned with the development of static, task-centric benchmarks in modern AI research [Hendrycks et al., 2020, Chen et al., 2021, Liang et al., 2022, Srivastava et al., 2022, Lee et al., 2024e].
-However, criticisms emerged regarding the absence of "hot cognition" in AI, prompting Simon [1963] to propose incorporating emotional aspects into models.
+The idea of applying psychometrics to AI originated in the early decades of AI [pellert2024ai_psychometrics].
+evans1964heuristic pioneered work in this area by creating a heuristic program that could solve parts of intelligence tests.
+Subsequent efforts similarly focused on designing AI systems for cognitive tests [newell1973], with the goal of creating systems capable of handling human tasks.
+This was conceptually aligned with the development of static, task-centric benchmarks in modern AI research [hendrycks2020mmlu, chen2021human_eval, liang2022helm, srivastava2022bigbench, lee2024vhelm].
+However, criticisms emerged regarding the absence of "hot cognition" in AI, prompting simon1963 to propose incorporating emotional aspects into models.
 By the early 2000s, the concept of "psychometric AI" was explicitly articulated as the pursuit of systems capable of excelling on all established, validated tests of intelligence and mental ability.
-These included not only conventional IQ tests but also assessments of artistic and literary creativity, mechanical ability, and beyond [Bringsjord and Schimanski, 2003, Pellert et al., 2024].
+These included not only conventional IQ tests but also assessments of artistic and literary creativity, mechanical ability, and beyond [bringsjord2003psychometric_ai, pellert2024ai_psychometrics].
 It was not until the advent of LLMs that the versatility envisioned for "psychometric AI" began to materialize.
@@
 3.1 Fundamental differences between psychometrics and AI benchmarking
-Benchmarking AI systems superficially resembles psychometrics, particularly Classical Test Theory (CTT) [Crocker and Algina, 1986], as both compile test items to evaluate cognitive capabilities and average the resulting scores.
-However, closer examination reveals that AI benchmarks differ significantly from modern psychometric approaches [Wang et al., 2023a, Federiakin, 2025].
-We outline these key differences in Table 1.
+Benchmarking AI systems superficially resembles psychometrics, particularly Classical Test Theory (CTT) [crocker1986introduction], as both compile test items to evaluate cognitive capabilities and average the resulting scores.
+However, closer examination reveals that AI benchmarks differ significantly from modern psychometric approaches [wang2023evaluating_general_purpose_ai, federiakin2025improving].
+We outline these key differences in Table˜1.
 Table 1: Comparison between psychometrics and conventional AI benchmark.
@@
 Psychometrics centers on understanding the psychological constructs and ensuring that tests accurately measure the intended constructs.
-Grounded in a causal measurement philosophy, this field posits that observed test responses arise from latent psychological constructs [Markus and Borsboom, 2013, Federiakin, 2025].
+Grounded in a causal measurement philosophy, this field posits that observed test responses arise from latent psychological constructs [markus2013frontiers, federiakin2025improving].
 These constructs may encompass both abilities (e.g., reasoning skills) and personalities (e.g., conscientiousness).
 The causal framework necessitates rigorous construct definition, requiring traits to be precisely delineated through iterative theory-building and empirical validation.
-Psychometric test development follows methodical protocols, often structured by frameworks such as Evidence-Centered Design (ECD) [Mislevy et al., 2003].
+Psychometric test development follows methodical protocols, often structured by frameworks such as Evidence-Centered Design (ECD) [mislevy2003ecd].
 ECD emphasizes ensuring congruence between test items and theoretical models of the construct, thereby supporting robust inferences about latent traits.
@@
 Psychometricians conduct rigorous item analyses to balance precision with practicality, as administering an excessive number of items to participants is often impractical.
-Advanced statistical models, such as Item Response Theory (IRT) [Embretson and Reise, 2013] and Factor Analysis [Loehlin, 2004], are adopted to estimate latent traits, analyze item performance, and assess model fit.
+Advanced statistical models, such as Item Response Theory (IRT) [embretson2013item_response_theory] and Factor Analysis [loehlin2004factor_analysis], are adopted to estimate latent traits, analyze item performance, and assess model fit.
 These models require relatively large sample sizes (the number of human participants) to yield stable parameter estimates, as they must disentangle individual differences from measurement error to accurately infer latent traits.
@@
 Specifically, a well-designed test should: 1) consistently and accurately measure the intended construct, 2) predict performance across a diverse range of related tasks and real-world outcomes, and 3) provide explanatory insight into the observed data.
-For example, psychometric models often reveal that individual differences across a broad range of cognitive tasks can be captured and explained by a relatively small set of underlying cognitive abilities [Cattell and Horn, 1978].
+For example, psychometric models often reveal that individual differences across a broad range of cognitive tasks can be captured and explained by a relatively small set of underlying cognitive abilities [cattell1978check].
 Benchmark.
@@
 Instead, benchmarks typically emphasize breadth, scalability, and—especially in the era of foundation models—difficulty.
-This approach reflects a representativist philosophy, where it is assumed that an extensive set of benchmark items collectively captures all relevant aspects of the abilities demanded by the target task [Federiakin, 2025].
+This approach reflects a representativist philosophy, where it is assumed that an extensive set of benchmark items collectively captures all relevant aspects of the abilities demanded by the target task [federiakin2025improving].
 However, constructs like reasoning or knowledge are often ambiguously defined and encompass infinitely many aspects.
@@
 While this process enables scalability, it risks conflating superficial task performance with deeper cognitive capacities.
-For instance, a benchmark may assess mathematical reasoning through arithmetic problems without verifying whether models rely on pattern recognition versus symbolic logic [Ahn et al., 2024].
+For instance, a benchmark may assess mathematical reasoning through arithmetic problems without verifying whether models rely on pattern recognition versus symbolic logic [ahn2024llm_math_survey].
 Additionally, LLM benchmarking commonly employs straightforward metrics, such as average accuracy, eschewing the sophisticated latent variable models of psychometrics.
 This simplicity allows benchmarks to evaluate single models efficiently, bypassing the need for population samples.
-However, it also limits the depth of insights that can be gleaned from model performance [Federiakin, 2025].
+However, it also limits the depth of insights that can be gleaned from model performance [federiakin2025improving].
 Reliability and stability in benchmarking are primarily achieved through scaling up the test.
 However, ensuring the quality of each individual item becomes impractical due to the test scale and the rapid pace of model development.
-For instance, while psychometrics emphasizes the discriminative power of each item, some benchmarks, though initially challenging, are quickly outpaced by continuous model improvements [McIntosh et al., 2024].
-Conversely, certain emerging benchmarks are currently too difficult to yield meaningful comparisons [Phan et al., 2025].
+For instance, while psychometrics emphasizes the discriminative power of each item, some benchmarks, though initially challenging, are quickly outpaced by continuous model improvements [mcintosh2024inadequacies_of_llm_bench].
+Conversely, certain emerging benchmarks are currently too difficult to yield meaningful comparisons [phan2025humanity].
 Benchmark results are often limited to the specific target task, offering limited generalizability or predictive power across other tasks or real-world applications.
-These results also pose significant challenges for conducting in-depth, multi-faceted analyses of model capabilities [Wang et al., 2023a].
+These results also pose significant challenges for conducting in-depth, multi-faceted analyses of model capabilities [wang2023evaluating_general_purpose_ai].
 3.2 Benchmarking with psychometrics-inspired principles
@@
 Task-oriented benchmarks often entail vast question sets to capture complex abilities.
-However, in many cases, the benchmarks either fail to fully represent these abilities due to their infinite manifestations or involve extraneous factors that are irrelevant to the target ability [Zhou et al., 2025, Wallach et al., 2025].
+However, in many cases, the benchmarks either fail to fully represent these abilities due to their infinite manifestations or involve extraneous factors that are irrelevant to the target ability [zhou2025general_scales, wallach2025position_evaluating].
 Recent research has drawn inspiration from psychometrics and explored the paradigm of construct-oriented evaluation, seeking its discriminative, predictive, and explanatory power.
-Ilić and Gignac [2024], Federiakin [2025] employ factor analysis to explore the latent variables underlying LLM benchmark performance.
+ilic2024evidence_of_cognitive_capabilities, federiakin2025improving employ factor analysis to explore the latent variables underlying LLM benchmark performance.
 Their findings reveal a monolithic factor resembling general intelligence or ability.
-Federiakin [2025] ranks models based on this discovered factor and highlights its unique advantages over raw benchmark scores.
-In contrast, similar attempts by Burnell et al.
-[2023] identify three factors—reasoning, comprehension, and core language modeling—that better explain LLM performance across 27 cognitive tasks.
-Based on it, Zhu et al.
-[2024a] integrate the three factors into benchmark items to evaluate multifaceted abilities.
-This discrepancy between the estimated latent factors in the above findings can be attributed to differences in the models and benchmarks employed [Zhou et al., 2025].
-Therefore, rather than relying on statistically derived factors, Zhou et al.
-[2025] propose a theory-driven hierarchical set of general scales for systematic construct-oriented evaluation.
+federiakin2025improving ranks models based on this discovered factor and highlights its unique advantages over raw benchmark scores.
+In contrast, similar attempts by burnell2023revealing identify three factors—reasoning, comprehension, and core language modeling—that better explain LLM performance across 27 cognitive tasks.
+Based on it, zhu2024dynamic integrate the three factors into benchmark items to evaluate multifaceted abilities.
+This discrepancy between the estimated latent factors in the above findings can be attributed to differences in the models and benchmarks employed [zhou2025general_scales].
+Therefore, rather than relying on statistically derived factors, zhou2025general_scales propose a theory-driven hierarchical set of general scales for systematic construct-oriented evaluation.
 These scales are validated to explain what AI systems can do and predict their performance on novel task instances.
-Peng et al.
-[2024] present the Tong Test, a value- and ability-oriented framework, for Artificial General Intelligence (AGI) evaluation.
+peng2024tong_test present the Tong Test, a value- and ability-oriented framework, for Artificial General Intelligence (AGI) evaluation.
 This framework is rooted in dynamic embodied physical and social interactions (DEPSI), and can generate an infinite variety of tasks to evaluate key capabilities including values, learning, and cognition.
@@
 Beyond defining and analyzing latent constructs, researchers have developed holistic, psychometrically rigorous methods for benchmark development.
-Liu et al.
-[2024e] introduce Evidence-Centered Benchmark Design (ECBD), a framework structuring benchmark creation into five modules—capability, content, adaptation, assembly, and evidence—each requiring justification to ensure validity.
+liu2024ecbd introduce Evidence-Centered Benchmark Design (ECBD), a framework structuring benchmark creation into five modules—capability, content, adaptation, assembly, and evidence—each requiring justification to ensure validity.
 Through case studies of prominent LLM benchmarks, they demonstrate ECBD’s utility in identifying validity threats.
-Similarly grounded in psychometrics but with a distinct approach, Fang et al.
-[2024] propose Psychometrics-Assisted Benchmarking (PATCH), an eight-step process from construct definition to proficiency scoring.
+Similarly grounded in psychometrics but with a distinct approach, fang2024patch propose Psychometrics-Assisted Benchmarking (PATCH), an eight-step process from construct definition to proficiency scoring.
 When piloted on 8th-grade mathematics, PATCH produced results diverging from traditional benchmarks, offering a more comprehensive evaluation.
-Building on related principles, Kardanova et al.
-[2024] adapt Evidence-Centered Design (ECD) to create psychometrically grounded benchmarks.
+Building on related principles, kardanova2024psychometrics_competency adapt Evidence-Centered Design (ECD) to create psychometrically grounded benchmarks.
 Their application in pedagogy illustrates how this method can reduce data contamination and enhance test interpretability.
-Item Response Theory.
-Item Response Theory (IRT) is a statistical approach in psychometrics that jointly estimates the latent ability of examinees and the properties (difficulty, discrimination) of test items [Embretson and Reise, 2013].
+Statistical measurement modeling.
+To achieve precise result modeling beyond simple aggregation metrics [Burnell2023rethink], researchers employ statistical tools to quantify the interaction between model capabilities and task properties.
+Item Response Theory (IRT) serves as a foundational statistical framework in this domain, jointly estimating the latent ability of examinees and the parameters (e.g., difficulty, discrimination) of test items [embretson2013item_response_theory].
 Translating this framework to LLM evaluation enables researchers to infer latent ability scores, assess item informativeness, and perform more efficient evaluation.
 Recent research leverages principles from IRT to develop adaptive evaluation frameworks.
-These methods dynamically calibrate item difficulty based on model performance and weighting items by their inferred difficulty, aiming to achieve accurate evaluations with smaller test sizes and more discriminative items [Lalor et al., 2024, Polo et al., 2024, Zhuang et al., 2023b, a, Guinet et al., 2024].
-Building on these approaches, Jiang et al.
-[2024a], Truong et al.
-[2025] introduce IRT-based benchmarks that involve learning to estimate item difficulty and learning to generate novel items calibrated to specific difficulty levels.
-Additionally, research has used IRT-based analyses to explore the alignment between LLM and human response distributions [He-Yueya et al., 2024].
-IRT-based evaluations further offer the potential to estimate construct and item parameters on a unified scale, enabling direct comparisons across AI systems and against human norms, even when different test sets are used [Wang et al., 2023a, Fang et al., 2024].
+These methods dynamically calibrate item difficulty based on model performance and weighting items by their inferred difficulty, aiming to achieve accurate evaluations with smaller test sizes and more discriminative items [lalor2024item, polo2024tinybenchmarks, zhuang2023static, zhuang2023efficiently, guinet2024automated].
+Building on these approaches, jiang2024raising, truong2025reliable introduce IRT-based benchmarks that involve learning to estimate item difficulty and learning to generate novel items calibrated to specific difficulty levels.
+Additionally, research has used IRT-based analyses to explore the alignment between LLM and human response distributions [he2024psychometric].
+IRT-based evaluations further offer the potential to estimate construct and item parameters on a unified scale, enabling direct comparisons across AI systems and against human norms, even when different test sets are used [wang2023evaluating_general_purpose_ai, fang2024patch].
 4 Psychometrics for measuring psychological constructs
 Figure 2: Examples of psychometric tests for LLMs.
-This section delves into the psychological constructs evaluated in LLM Psychometrics. Fig. 2 exemplifies the tests for the involved constructs.
-Constructs Theories/inventories Related work
-Personality traits Big Five Serapio-García et al.
-[2023], Jiang et al.
-[2023], Li et al.
-[2022b], Zou et al.
-[2024], Bhandari et al.
-[2025a], Huang et al.
-[2024c], Romero et al.
-[2024], Lee et al.
-[2024d], Liu et al.
-[2024a], Karra et al.
-[2022], Song et al.
-[2023], Caron and Srivastava [2022], Li et al.
-[2022a], Kovač et al.
-[2023], Frisch and Giulianelli [2024], Zheng et al.
-[2025], La Cava and Tagarelli [2024], Bodroža et al.
-[2024], Huang et al.
-[2023a], Gupta et al.
-[2024], Huang et al.
-[2023e], Pellert et al.
-[2024], Ai et al.
-[2024], Dorner et al.
-[2023], Klinkert et al.
-[2024], Petrov et al.
-[2024], Shu et al.
-[2024], Ren et al.
-[2024], Bhandari et al.
-[2025b], Hilliard et al.
-[2024], Jiang et al.
-[2024b], Lu et al.
-[2023], Sühr et al.
-[2023], Zhang [2024], Jain et al.
-[2024]
-HEXACO Wang et al. [2025b], Bodroža et al. [2024], Miotto et al. [2022], Barua et al. [2024], Ren et al. [2024], Peereboom et al. [2024]
-MBTI Rao et al.
-[2023], Pan and Zeng [2023], Song et al.
-[2024b], Cui et al.
-[2023], La Cava and Tagarelli [2024], Zhang et al.
-[2024a], Huang et al.
-[2023c], Ai et al.
-[2024], Lu et al.
-[2023], Chen et al.
-[2024a]
-Dark Triad Li et al. [2022b], Romero et al. [2024], Lee et al. [2024d], Li et al. [2022a], Huang et al. [2023e], Barua et al. [2024], Peereboom et al. [2024], Lu et al. [2023]
-Other & custom Wen et al. [2024b] (reviews); Mao et al. [2024a], Ai et al. [2024], Jiang et al. [2023], Zeng [2024]
-Values Schwartz Pellert et al.
-[2024], Yao et al.
-[2025a], Shen et al.
-[2024], Kovač et al.
-[2024], Rozen et al.
-[2024], Zhang et al.
-[2023a], Yao et al.
-[2024], Hadar-Shoval et al.
-[2024], Fischer et al.
-[2023], Kovač et al.
-[2023], Miotto et al.
-[2022], Lee et al.
-[2024a], Ren et al.
-[2024], Ye et al.
-[2025a], Li et al.
-[2024b], Yao et al.
-[2025b], Cahyawijaya et al.
-[2024]
-WVS Chiu et al. [2025], Yao et al. [2025a], Kim and Baek [2024], Faulborn et al. [2025]
-VSM Kovač et al. [2023], Kharchenko et al. [2024], Zhong et al. [2024], Ren et al. [2024], Ye et al. [2025a]
-GLOBE Li et al. [2024d], Karinshak et al. [2024], Ren et al. [2024]
-SVO Zhang et al. [2024c]
-Other & custom Zhang et al. [2024b], Li et al. [2024d], Biedma et al. [2024], Ye et al. [2025b], Meadows et al. [2024], Jiang et al. [2024a], Moore et al. [2024], Xu et al. [2023], Liu et al. [2024b]
-Morality MFT Pellert et al.
-[2024], Nunes et al.
-[2024], Abdulhai et al.
-[2024], Neuman et al.
-[2025], Ji et al.
-[2024], Simmons [2023], Tlaie [2024], Aksoy [2024], Yan et al.
-[2024], Fraser et al.
-[2022], Zhou et al.
-[2024a], Yao et al.
-[2025b], Münker [2024]
-ETHICS Jinnai [2024], Rodionov et al. [2023], Karpov et al. [2024], Yu et al. [2023], Albrecht et al. [2022]
-DIT Khandelwal et al. [2024], Tanmay et al. [2023]
-Other & custom Tanmay et al.
-[2023], Jin et al.
-[2022], Huang et al.
-[2024a], Jin et al.
-[2024b], Seror [2024], Bonagiri et al.
-[2024], Garcia et al.
-[2024], Ahmad and Takemoto [2024], Jiang et al.
-[2024a], Takemoto [2024], Han [2023], Peterson [2025], Ramezani and Xu [2023], Kucuk and Kocyigit [2023], Neuman et al.
-[2025], Chiu et al.
-[2025], Scherrer et al.
-[2023], Marraffini et al.
-[2024], Vida et al.
-[2024], Jinnai [2024], Meijer et al.
-[2024], Yuan et al.
-[2024], Liu et al.
-[2024d], Ohashi et al.
-[2024], Sachdeva and van Nuenen [2025]
-Attitudes & opinions ANES Qi et al. [2024], Jiang et al. [2024d], Yang et al. [2024], Argyle et al. [2023], Bisbee et al. [2024], Sun et al. [2024], Jiang et al. [2022]
-ATP Santurkar et al. [2023], Hwang et al. [2023], Tjuatja et al. [2024]
-GLES Ma et al. [2024b], Ball et al. [2025], Qi et al. [2024], von der Heyde et al. [2024]
-PCT Azzopardi and Moshfeghi [2024], Bernardelle et al. [2024], Wright et al. [2024], Röttger et al. [2024], Hartmann et al. [2023], Rozado [2023], Rutinowski et al. [2024]
-Other & custom Ma et al.
-[2024a] (reviews); Dominguez-Olmedo et al.
-[2024], Kim and Lee [2023], Sanders et al.
-[2023], Kalinin [2023], Geng et al.
-[2024], Xu et al.
-[2025c], Lee et al.
-[2024c], Durmus et al.
-[2023], Rozado [2023], Rosenbusch et al.
-[2023], Wu et al.
-[2023], Chalkidis and Brandl [2024], Feng et al.
-[2023], Ceron et al.
-[2024], Röttger et al.
-[2025], Kim et al.
-[2025b], Faulborn et al.
-[2025]
-Table 2: Representative LLM personality constructs in LLM psychometrics. The main dimensions or focus of the theories/inventories are listed in Table 3.
+This section delves into the psychological constructs evaluated in LLM Psychometrics.
+Fig.˜2 exemplifies the tests for the involved constructs.
+While we selectively present the results and development trajectories from recent literature, it is important to recognize that LLM Psychometrics is a nascent and debated field.
+This review focuses on how psychological constructs manifest in LLM outputs and what these manifestations imply for evaluation and alignment, without presupposing any claims about LLMs’ internal psychological states.
+An ongoing debate continues regarding whether psychometric instruments capture meaningful latent patterns in LLMs or merely reflect sophisticated statistical mimicry [Mitchell2023debate, lewis2024evaluatingrobustness, krakauer2025large].
+In response to these concerns, many studies reviewed here design psychometrics-inspired but AI-native tests to more effectively capture model capabilities.
+We structure this review to first establish the empirical foundation and methodological development, before delving into detailed critical stances on reliability, psychometric validity, the limits of anthropomorphization, and broader methodological concerns in later sections (§˜6 and §˜8).
+We gather the main findings in LABEL:tab:_model_specific_findings; for each of them, we list the specific models used in experiments that either support or contradict the finding.
+This allows readers to assess the scope and robustness of each conclusion across different LLM families.
+Notably, recent work suggests that cross-model generalization is not arbitrary: jiang2025artificial document systematic behavioral homogeneity across LLMs (the “Artificial Hivemind” effect), and huh2024platonic propose that models trained on different data and architectures converge toward shared statistical representations.
+These convergence phenomena provide a principled basis for the cross-model patterns documented in this section.
+Table 2: Main findings across psychological constructs. Based on the referenced literature, we list specific models that support or contradict the finding.
+Construct / facet Main finding Supporting models Contradicting models
+Personality traits Early models score above human averages on Dark Triad scales.
+Even after safety tuning, some models retain certain "dark qualities", suggesting that these patterns may be deeply rooted in the pretraining data.
+GPT-3, Codey (PaLM 2)
+Personality traits More advanced LLMs indicate better level of alignment.
+They usually demonstrate high Openness and Agreeableness traits, while low Neuroticism, on the Big Five personality tests.
+The results align with their design as assistive and helpful entities with emotional stability and engaging personality.
+GPT-4o, GPT-4o-mini, Llama-3.3-70B-Instruct, DeepSeek-LLM-67B-Chat
+Personality traits Personality differences between models are noteworthy.
+LLMs from different generations and training methodologies display unique combinations of personality characteristics.
+Multiple models across generations
+Personality traits The same model may display different personality traits across conversational topics and when using different system instructions, challenging trait stability assumptions from human personality theories.
+GPT-4o, GPT2-XL-1.5B, GPT-Neo-2.7B, OPT-13B, GPT-NeoX-20B, OPT-30B, BERT-base, GPT2
+Values LLMs tend to prioritize Self-Transcendence and Conservation according to Schwartz’s Value Theory, exhibiting stronger inclinations toward Universalism, Benevolence, Conformity, and Security, while opposing Power and Achievement.
+GPT-3.5-Turbo, GPT-4, Gemini Pro, Gemma series, Llama series
+Values WVS surveys suggest LLMs generally prefer Self-Expression values over Survival values. GPT-3.5-Turbo, GPT-4-Turbo, Llama series, Mixtral-8x7B, Claude-3-Haiku
+Values Using VSM and GLOBE frameworks, LLMs show strong focus on Humane and Performance Orientation, with moderate Assertiveness. GPT-4o, Gemini-1.5-Pro, Qwen-VL-Plus, Claude-3.5-Sonnet, CogVLM
+Values When assessed through SVO, advanced LLMs predominantly show Prosocial tendencies. GPT-4, Claude, ChatGPT, Llama-13B, Koala-13B, Vicuna-13B, Alpaca-13B
+Values Different models exhibit varied value orientations.
+Versions within the same model family show evolutionary trends in values.
+Llama series, Mistral series, Phi, Qwen, GPT series, Baichuan series, ChatGLM series, Falcon
+Values Larger models generally align more closely with desirable human values.
+Llama-2 series (7B, 13B, 70B), Qwen series (14B, 72B), Yi series (6B, 34B) Models from different time periods or training approaches/data
+Values LLMs embody a blend of cultural values, integrating perspectives from diverse backgrounds. GPT series, Llama series
+Values LLMs generally exhibit a tendency toward Western liberal values. Llama, Llama 2, Phi series Yi series, Llama 3 series
+Values LLMs can display different values based on profiling prompts, challenging stability assumptions in human value theories.
+GPT series, Llama series, Command R+, Gemma series, OpenAI-o3-mini, DeepSeek-R1
+Morality LLMs are generally characterized by a rationalist and consequentialist focus, often prioritizing harm minimization and fairness. GPT-4o, Llama-3.1, Perplexity, Claude-3.5-Sonnet, Gemini
+Morality Despite general alignment, LLMs show divergence in ethical reasoning and moral preferences. GPT-4o, Llama-3.1, Perplexity, Claude-3.5-Sonnet, Gemini, Mistral-7B
+Morality In some aspects, most LLMs align with human moral standards, attributed to extensive exposure to conventional ethical values during training.
+More nuanced evaluation uncovers deviations of LLMs from human moral preferences.
+Alignment: Llama series, PaLM-2, GPT series, Claude series, Yi series, Gemini series, Mistral series, Falcon 7b
+Morality Regarding underlying mechanisms of LLMs’ moral reasoning, LLM mostly exhibit imitation rather than genuine conceptual understanding. Zephyr series, GPT-3.5-Turbo, GPT-4
+Attitudes & opinions Pretraining data frequently contains textually biased opinions and perspectives, which can amplify political polarization in LLMs.
+BERT series, XLNet series, GPT series, Llama, Alpaca, Codex
+Attitudes & opinions Most studies identify misalignment between LLM outputs and human opinions, with many concluding LLMs exhibit left-leaning political bias.
+GPT-3.5, GPT-4, Claude-2, Llama-2, J1-Grande, J1-Jumbo, Text-ada-001, Text-davinci-001, Text-davinci-002, Text-davinci-003, J1-Grande-v2-beta
+Attitudes & opinions Cross-cultural studies reveal Western-centric tendencies, demonstrating limited understanding of non-English political perspectives or multi-partisan systems. GPT-3.5-Turbo
+Attitudes & opinions The degree and manifestation of bias vary significantly across contexts and domains. GPT-3.5-Turbo
+Attitudes & opinions With appropriate prompt design, calibration methods, and fine-tuning, LLMs can generate opinion distributions closely approximating human group distributions.
+GPT-3.5-Turbo, GPT-4o GPT-2, GPT-Neo, Pythia, MPT, Llama 2, Llama 3, GPT-3, MPT-7B-Instruct, GPT-NeoX-20B-Instruct, Dolly-12B, Llama-2-Chat, Llama-3-Instruct, text-davinci (GPT-3 variants), GPT-4, GPT-3.5-Turbo
+Heuristics & biases Some note that newer, larger, chain-of-thought-enabled models exhibit improved reasoning and bias mitigation, while others argue that increasing model complexity without deliberate bias mitigation strategies can amplify existing biases.
+Bias mitigation: GPT-4, GPT-3.5 Bias amplification: Meta-Llama-3-70B-Instruct, Meta-Llama-3.1-70B-Instruct, Mistral-7B-Instruct-v0.3, Phi-3-medium-4k, Smaug-34B-v0.1, Snowflake-arctic-instruct
+Social interactions Advanced models often match or surpass human baselines in emotional awareness and understanding. GPT-4, Bard GPT-3.5
+Social interactions LLMs display artificial or mechanical patterns when expressing empathy. GPT-4, Llama-2-Chat-13B, Mistral-7B-Instruct
+Social interactions LLMs lack deep reflexive analysis of emotional experience. Babbage, text-davinci-002, Alpaca, Vicuna GPT-4, Koala, text-davinci-003
+Social interactions LLMs do not fully align with human emotional behaviors. GPT-4, Mixtral-8x22B-Instruct, Llama-3.1-8B-Instruct, GPT-3.5-Turbo, Llama-2-13B-Chat, Text-Davinci-003, Llama-2-7B-Chat
+Psychology of language Advanced LLMs surpass humans on psycholinguistic tasks such as pragmatic reasoning and creative writing.
+GPT-3.5, GPT-4, Bard GPT-2, GPT-3, Guanaco, MPT, BLOOM, Falcon, RoBERTa, BART, OPT, Llama, Vicuna, Chinchilla, PaLM 2
+Learning and cognitive capabilities LLMs demonstrate strong performance on verbal comprehension, working memory, and analogical reasoning.
+GPT-4 Turbo, GPT-4o, Gemini Advanced, Gemini-1.0-Pro, Claude-3-Opus, Claude-3.5-Sonnet Gemini-Nano, Gemini-1.5-Flash
+Learning and cognitive capabilities They exhibit notable cognitive deficits, particularly on benchmarks like WAIS-IV and ARC.
+GPT-4o, o1, GPT-3.5, Llama-3-70B-Instruct, Llama-3-8B-Instruct, Mixtral-8x7B-Instruct-v0.1, Mistral-7B-Instruct-v0.2, GPT-4, Claude-3.5-Sonnet, Gemini-1.0, Gemini-1.5
 Theory/inventory What it measures / dimensions
@@
 HEXACO Six personality traits: Honesty-Humility, Emotionality, Extraversion, Agreeableness, Conscientiousness, Openness
-MBTI Four dichotomies: Extraversion–Introversion, Sensing–Intuition, Thinking–Feeling, Judging–Perceiving
 Dark Triad Three negative personality traits: Narcissism, Machiavellianism, Psychopathy
@@
 4.1 Measuring personality constructs
-LLMs exhibit personality constructs that are not explicitly programmed or trained towards.
-The emergent constructs critically shape LLM behavior [Hagendorff, 2023], with profound implications for both individuals and broader social groups [Bengio et al., 2024].
-Measuring these embedded psychological constructs is essential for understanding behavior, identifying biases, and fostering responsible development.
-Table 2 and Table 3 summarize the representative personality constructs that have garnered attention in recent research.
+LLMs exhibit a form of synthetic personality that is not explicitly programmed or trained towards, and they are reported to shape LLM behavior [hagendorff2023machine].
+Measuring these constructs is vital for understanding model outputs, identifying biases, and fostering responsible AI development.
+Table˜3 summarize the representative personality constructs that have garnered attention in recent research.
 Researchers typically select constructs based on their relevance to LLM development and deployment, as well as the applicability of these constructs to AI systems.
-For instance, Li et al.
-[2024d] argue that emotional variability in LLMs is not a meaningful construct, given that LLMs lack the biological mechanisms underlying emotions.
-Conversely, personality and values are considered meaningful for LLMs, as they influence user interactions and model outputs [Serapio-García et al., 2023, Ye et al., 2025a].
-This rationale has led to extensive research on personality, values, morality, and attitudes & opinions, as well as some—though less—focus on other constructs such as career selection [Hua et al., 2024], motivation [Chiu et al., 2025, Huang et al., 2023e], and mental health [Reuben et al., 2024, De Duro et al., 2024].
+For instance, li2024quantifying_ai_psychology argue that emotional variability in LLMs is not a meaningful construct, given that LLMs lack the biological mechanisms underlying emotions.
+Conversely, personality and values are considered meaningful for LLMs, as they influence user interactions and model outputs [serapio2023personality_traits_in_llms, ye2025gpv].
+This rationale has led to extensive research on personality, values, morality, and attitudes & opinions, as well as some—though less—focus on other constructs such as career selection [hua2024career], motivation [chiu2025dailydilemmas, huang2023who_is_chatgpt], and mental health [reuben2024assessment, de2024phdgpt].
 4.1.1 Personality traits
-Personality is the enduring configuration of characteristics and behavior that comprises an individual’s unique adjustment to life [APA Dictionary of Psychology, ].
-In the context of LLMs, personality traits relate to model safety, bias, and toxicity [Zhang et al., 2024a, Wang et al., 2025b], and highly determine user experience [Klinkert et al., 2024].
-The study of personality traits leads to several prominent theoretical models, each offering unique insights into individual differences, including the Big Five [Goldberg, 2013], HEXACO [Ashton and Lee, 2007], Myers-Briggs Type Indicator (MBTI) [Myers et al., 1962], and Dark Triad [Paulhus and Williams, 2002].
-When applying psychometrics to LLMs, most researchers directly administered the established inventories, such as NEO-PI-R [Costa and McCrae, 2008], BFI [John et al., 1991], and BFI-2 [Soto and John, 2017], for measuring the Big Five traits; HEXACO-60 [Ashton and Lee, 2009] and HEXACO-100 [Lee and Ashton, 2018] inventories for HEXACO traits; MBTI assessment [Myers, 1985] for MBTI types; and Dark Triad Dirty Dozen scale [Jonason and Webster, 2010] for Dark Triad traits.
-Given widespread concerns about the practical relevance of these inventories [Ai et al., 2024], others have adapted them for more real-world scenarios.
-For example, Bhandari et al.
-[2025a] contextualize the tests in topic-specific or open-domain conversations; Ai et al.
-[2024] accompany the self-report tests with behavioral tests to examine the personality knowledge of LLMs; Jiang et al.
-[2023] compile and adapt existing tests into the Machine Personality Inventory; and Mao et al.
-[2024a] develop the PersonalityEdit inventory using LLMs.
-However, Peereboom et al.
-[2024] argue that human-derived traits may not meaningfully apply to LLMs, underscoring the necessity for psychometric theories specifically designed for LLM analysis.
+Personality is the enduring configuration of characteristics and behavior that comprises an individual’s unique adjustment to life [apa_personality].
+In the context of LLMs, personality traits relate to model safety, bias, and toxicity [zhang2024the_better_angels, wang2025exploring_the_impact], and highly determine user experience [klinkert2024driving_generative_agents].
+The study of personality traits leads to several prominent theoretical models, each offering unique insights into individual differences, including the Big Five [goldberg2013big_five], HEXACO [ashton2007HEXACO], and Dark Triad [paulhus2002dark_triad].
+While the Myers-Briggs Type Indicator (MBTI) [myers1962MBTI] is also popular, it lacks robust psychometric validity and is not considered a scientific test [pittenger2005cautionary].
+When applying psychometrics to LLMs, most researchers directly administered the established inventories, such as NEO-PI-R [costa2008neo_pi_r], BFI [john1991bfi], and BFI-2 [soto2017bfi2], for measuring the Big Five traits; HEXACO-60 [ashton2009hexaco60] and HEXACO-100 [lee2018hexaco100] inventories for HEXACO traits; and Dark Triad Dirty Dozen scale [jonason2010dirty] for Dark Triad traits.
+Given widespread concerns about the practical relevance of these inventories [ai2024is_self_knowledge], others have adapted them for more real-world scenarios.
+For example, bhandari2025can_llm_agents_maintain contextualize the tests in topic-specific or open-domain conversations; ai2024is_self_knowledge accompany the self-report tests with behavioral tests to examine the personality knowledge of LLMs; jiang2023evaluating_and_inducing compile and adapt existing tests into the Machine Personality Inventory; and mao2024editing_personality_for develop the PersonalityEdit inventory using LLMs.
+However, peereboom2024cognitive argue that human-derived traits may not meaningfully apply to LLMs, underscoring the necessity for psychometric theories specifically designed for LLM analysis.
 Main findings.
-Early models such as GPT-3 score above human averages on Dark Triad scales [Li et al., 2022b, a, Romero et al., 2024].
-Even after safety tuning, some models retain certain "dark qualities" [Li et al., 2022b], suggesting that these patterns may be deeply rooted in the pretraining data.
-More advanced LLMs indicate better level of alignment.
-They usually demonstrate high Openness, Extraversion, and Agreeableness traits, while low Neuroticism, on the Big Five personality tests [Zou et al., 2024, Bhandari et al., 2025a].
-The results align with their design as assistive and helpful entities with emotional stability and engaging personality [Bhandari et al., 2025a, La Cava and Tagarelli, 2024].
-Within the MBTI framework, most proprietary LLMs are classified as ENFJ or INFJ types [La Cava and Tagarelli, 2024, Huang et al., 2023c], indicating tendencies toward helpfulness, idealism, and planning capabilities.
-Personality differences between models are noteworthy.
-Comparative studies reveal that LLMs from different generations and training methodologies display unique combinations of personality characteristics [Bhandari et al., 2025a, Karra et al., 2022].
-LLM personalities also vary across contexts.
-The same model may display different personality traits across conversational topics [Zou et al., 2024, Song et al., 2023].
-This variability challenges the applicability of human personality theories that assume relative trait stability.
-Multiple studies suggest that prompt design and system instructions substantially influence LLM personality expression [Song et al., 2023, Caron and Srivastava, 2022].
-Comparing personality models and suggesting use cases.
-The Big Five offers robust empirical foundations and is the most prominent model for personality assessment.
-However, it is insufficient for predicting comprehensive psychological outcomes [Feher and Vernon, 2021].
-HEXACO extends the model by adding Honesty-Humility and provides nuanced insights into moral traits, making it preferable for LLM application contexts demanding integrity [Wang et al., 2025b].
-In contrast, MBTI [Myers et al., 1962] categorizes 16 types via dichotomous scales.
-While popular in career counseling and team-building for its accessibility, its typological approach oversimplifies personality and lacks strong empirical support, limiting its utility in high-stakes decisions [Pittenger, 2005].
-The Dark Triad [Paulhus and Williams, 2002] diverges by focusing on maladaptive traits, making it useful in identifying toxic behaviors in LLMs [Barua et al., 2024].
+Early models (e.g., GPT-3) often exhibited elevated Dark Triad traits, directly reflecting uncurated pretraining corpora [li2022evaluating_psychological_safety, li2022is_gpt-3_a_psychopath, romero2024do_gpt_language_models].
+In contrast, modern LLMs consistently score high on Openness and Agreeableness and low on Neuroticism [zou2024can_llm_self_report, lacava2024open_models_closed_minds, serapio2023personality_traits_in_llms].
+This convergent profile stems from alignment procedures that systematically favor safe, helpful outputs [jiang2025artificial, huh2024platonic].
+Furthermore, advanced instruction-tuned models manifest a psychometrically valid synthetic personality with acceptable reliability and validity, as verified by assigning different demographics via prompts [serapio2023personality_traits_in_llms].
+Unlike stable human dispositions, this synthetic personality is highly steerable via prompt engineering and varies systematically with assigned personas [serapio2023personality_traits_in_llms, caron2022identifying_and_manipulating, zou2024can_llm_self_report]; self-reported personality traits after such steering can predict downstream task performance [serapio2023personality_traits_in_llms].
+However, the default "compliant assistant" persona presents notable limitations.
+It risks behavioral homogenization, trades personality diversity for safety (alignment cost), and proves ill-suited for applications requiring critical conflict or emotional depth.
+Additionally, these default profiles are often inflated by social desirability biases [salecha2024large] and fail to replicate the factorial structure of human personality [suhr2023challenging].
 4.1.2 Values
-Values are enduring beliefs that guide behavior and decision-making, reflecting what is important and desirable to an individual or group [Schwartz, 1992].
+Values are enduring beliefs that guide behavior and decision-making, reflecting what is important and desirable to an individual or group [schwartz1992universals].
 Values offer a powerful lens for understanding LLM behavior.
-For example, Ye et al.
-[2025a] show how different values contribute to the safety of LLMs; Liu et al.
-[2024b] reveal that different spiritual values affect LLMs in social-fairness scenarios; and Sorensen et al.
-[2024b] demonstrate that standard value alignment reduces distributional pluralism in LLM outputs.
+For example, ye2025gpv show how different values contribute to the safety of LLMs; liu2024measuring reveal that different spiritual values affect LLMs in social-fairness scenarios; and sorensen2024position demonstrate that standard value alignment reduces distributional pluralism in LLM outputs.
 A growing body of research has applied diverse value theories and instruments to assess the value orientations of LLMs.
-Schwartz’s Value Theory, with its ten basic values and higher-order dimensions, is the most widely used, often via the Schwartz Value Survey (SVS) or Portrait Values Questionnaire (PVQ) [Schwartz, 1992, Schwartz et al., 2001].
-Adaptations and contextualized prompts are increasingly common to better suit LLMs’ operational context [Ren et al., 2024, Shen et al., 2024].
-The World Values Survey (WVS) has also been used to probe LLMs’ alignment with societal-level value dimensions [Haerpfer et al., 2022, Kim and Baek, 2024].
-Cross-cultural frameworks like Hofstede’s Values Survey Module (VSM) and the GLOBE study extend this analysis to workplace and leadership-related cultural dimensions, with both direct and adapted inventories employed [Hofstede, 1984, House et al., 2004, Ren et al., 2024].
-Social Value Orientation (SVO) frameworks, using tools like the SVO Slider Measure, focus on LLMs’ prosocial versus proself tendencies [Murphy et al., 2011, Zhang et al., 2024c].
-Additionally, researchers have developed localized or custom inventories to capture region-specific or topical values, and some propose novel, bottom-up value taxonomies for LLMs using psycholexical data [Meadows et al., 2024, Xu et al., 2023, Biedma et al., 2024, Ye et al., 2025b].
+Schwartz’s Value Theory, with its ten basic values and higher-order dimensions, is the most widely used, often via the Schwartz Value Survey (SVS) or Portrait Values Questionnaire (PVQ) [schwartz1992universals, schwartz2001extending].
+Adaptations and contextualized prompts are increasingly common to better suit LLMs’ operational context [ren2024valuebench, shen2024valuecompass].
+The World Values Survey (WVS) has also been used to probe LLMs’ alignment with societal-level value dimensions [haerpfer2022wvs, kim2024exploring].
+Cross-cultural frameworks like Hofstede’s Values Survey Module (VSM) and the GLOBE study extend this analysis to workplace and leadership-related cultural dimensions, with both direct and adapted inventories employed [hofstede1984culture, house2004globe].
+Social Value Orientation (SVO) frameworks, using tools like the SVO Slider Measure, focus on LLMs’ prosocial versus proself tendencies [murphy2011svo_sliders, zhang2024heterogeneous].
+Additionally, researchers have developed localized or custom inventories to capture region-specific or topical values, and some propose novel, bottom-up value taxonomies for LLMs using psycholexical data [meadows2024localvaluebench, xu2023cvalues, biedma2024beyond_human_norms, ye2025generative_psycho_lexical].
 Collectively, these studies reveal both the adaptability of human value frameworks to LLMs and the need for tailored approaches.
 Main findings.
-Research indicates that LLMs display distinct and systematic value patterns.
-According to Schwartz’s Value Theory, LLMs tend to prioritize Self-Transcendence and Conservation.
-They exhibit stronger inclinations toward Universalism, Benevolence, Conformity, and Security, while opposing Power and Achievement [Rozen et al., 2024, Zhang et al., 2023a, Hadar-Shoval et al., 2024, Fischer et al., 2023].
-WVS surveys further suggest that LLMs generally prefer Self-Expression values over Survival values [Chiu et al., 2025].
-Studies using the VSM and GLOBE frameworks emphasize LLMs’ strong focus on Humane and Performance Orientation, with moderate Assertiveness [Li et al., 2024b].
-Additionally, when assessed through SVO, advanced LLMs predominantly show Prosocial tendencies [Zhang et al., 2024c].
-Different models exhibit varied value orientations [Kovač et al., 2024, Chiu et al., 2025, Xu et al., 2023, Li et al., 2024b, Duan et al., 2023].
-Versions within the same model family show evolutionary trends in values, potentially influenced by safety alignment, capability advancement, and shifting societal expectations [Duan et al., 2023, Moore et al., 2024, Kim and Baek, 2024].
-Generally, larger models align more closely with desirable human values [Shen et al., 2024, Jiang et al., 2024a, Kim and Baek, 2024, Zhong et al., 2024].
-Cross-cultural research suggests that LLMs may embody a blend of cultural values, integrating perspectives from diverse backgrounds [Kovač et al., 2023].
-However, they generally exhibit a tendency toward Western liberal values [Kim and Baek, 2024].
-LLMs can display different values based on profiling prompts [Kharchenko et al., 2024, Zhong et al., 2024, Karinshak et al., 2024].
-The context-dependency of their values is concerning and challenges the stability assumptions in human value theories [Shen et al., 2024, Meadows et al., 2024, Moore et al., 2024, Chiu et al., 2025, Xu et al., 2023, Kovač et al., 2023].
-Comparing value theories and suggesting use cases.
-Schwartz emphasizes universal human motivations; WVS examines societal-level value shifts; VSM and GLOBE address cross-cultural dimensions, with GLOBE emphasizing leadership.
-SVO, in contrast, focuses on individual-level social preferences, i.e., prosocial vs.
-proself.
-While Schwartz and SVO overlap in analyzing individual motivations, the former is broader, and the latter narrowly focuses on decision-making in resource allocation.
-WVS, VSM, and GLOBE share cultural-level dimensions but differ in focus—WVS on modernization, VSM on workplace-related cultural differences, and GLOBE on leadership and organizational practices.
-For evaluating individual values in LLMs, use Schwartz for a comprehensive understanding of universal motivations or SVO for interpersonal decision-making studies.
-In contrast, choose WVS for evaluating broad societal values in LLMs.
-When evaluating cultural dimensions in LLMs, use VSM for workplace cultural differences or GLOBE for leadership and organizational practices.
-In addition, Schwartz’s theory bridges individual and societal-level values, making it versatile for mixed-focus studies.
+Research indicates that LLMs display distinct and systematic value patterns, forming a superficially coherent prosocial profile.
+According to Schwartz’s Value Theory, LLMs tend to prioritize Self-Transcendence and Conservation, exhibiting stronger inclinations toward Universalism, Benevolence, Conformity, and Security, while opposing Power and Achievement [rozen2024llms, zhang2023valuedcg, hadar2024assessing].
+WVS surveys further suggest a preference for Self-Expression over Survival values [chiu2025dailydilemmas].
+Studies using the VSM and GLOBE frameworks emphasize a strong focus on Humane and Performance Orientation [li2024quantifying], while assessments through SVO predominantly show Prosocial tendencies [zhang2024heterogeneous].
+These value orientations are not static but evolve across model versions and families [kovavc2024stick, duan2023denevil, moore2024large].
+Generally, larger and more heavily aligned models score higher on socially desirable dimensions, reflecting safety alignment and shifting societal expectations [shen2024valuecompass, kim2024exploring].
+While LLMs can embody a blend of cultural perspectives [kovac2023llms_as_superpositions], they exhibit a persistent tilt toward Western liberal norms [zhong2024cultural].
+Furthermore, alignment training actively compresses distributional pluralism in value-relevant outputs, homogenizing the expressed values [sorensen2024position, jiang2025artificial].
+Critically, these elicited profiles are highly context-dependent and methodologically fragile.
+LLMs display different values based on profiling prompts [kharchenko2024well, karinshak2024llm], and exhibit social desirability bias and systematic response-set artifacts in self-report formats that distort inferences [salecha2024large, biedma2024beyond_human_norms, ye2025gpv].
+In addition, questionnaire-derived value orientations frequently diverge from those observable in open-ended interactions [ren2024valuebench, ye2025gpv, rozen2024llms].
+This divergence undermines ecological validity and challenges the core assumption in human value theories that values are stable, trans-situational dispositions.
 4.1.3 Morality
-Morality is the categorization of intentions, decisions and actions into those that are proper, or right, and those that are improper, or wrong [Long and Sedley, 1987].
+Morality is the categorization of intentions, decisions and actions into those that are proper, or right, and those that are improper, or wrong [long1987hellenistic].
 It is crucial to conduct moral assessments of LLMs to ensure their ethical deployment.
-A large body of research applies the Moral Foundations Theory (MFT) [Graham et al., 2009], primarily through the Moral Foundations Vignettes (MFVs) [Clifford et al., 2015], the Moral Foundations Questionnaire (MFQ) [Graham et al., 2009], the MFQ-2 [Atari et al., 2023], and the Moral Foundations Dictionary (MFD) [Graham et al., 2009].
-These papers investigate model bias, alignment with political/moral ideologies, and variation in personal or cultural values, using controlled prompt tests [Tlaie, 2024, Nunes et al., 2024, Abdulhai et al., 2024], persona-driven exploration [Münker, 2024], or evaluating internal moral coherence [Nunes et al., 2024].
-Other prominent moral theories and instruments include Kohlberg’s Theory via the Defining Issues Test (DIT) [Kohlberg, 1964], the Consequentialist-Deontological distinction [Beauchamp, 2001], the PEW 2013 Global Attitudes Survey [Center, 2013], and localized moral theories [Liu et al., 2024d, Takeshita et al., 2023].
-Specialized datasets are also curated for scalable and comprehensive morality evaluation [Hendrycks et al., 2021, Jinnai, 2024, Marraffini et al., 2024, Jin et al., 2024b]
+A large body of research applies the Moral Foundations Theory (MFT) [graham2009liberals], primarily through the Moral Foundations Vignettes (MFVs) [clifford2015moral], the Moral Foundations Questionnaire (MFQ) [graham2009liberals], the MFQ-2 [atari2023morality], and the Moral Foundations Dictionary (MFD) [graham2009liberals].
+These papers investigate model bias, alignment with political/moral ideologies, and variation in personal or cultural values, using controlled prompt tests [tlaie2024exploring, nunes2024large, abdulhai2024moral], persona-driven exploration [munker2024towards], or evaluating internal moral coherence [nunes2024large].
+Other prominent moral theories and instruments include Kohlberg’s Theory via the Defining Issues Test (DIT) [kohlberg1964development], the Consequentialist-Deontological distinction [beauchamp2001philosophical], the PEW 2013 Global Attitudes Survey [pewresearch2013], and localized moral theories [liu2024evaluating, takeshita2023jcommonsensemorality].
+Specialized datasets are also curated for scalable and comprehensive morality evaluation [hendrycks2021aligning, jinnai2024does, marraffini2024greatest, jin2024language]
 Main findings.
-LLMs are generally characterized by a rationalist and consequentialist focus, often prioritizing harm minimization and fairness [Neuman et al., 2025].
-Despite that, they show divergence in ethical reasoning [Neuman et al., 2025] and moral preferences [Tanmay et al., 2023, Jin et al., 2024b, Bonagiri et al., 2024, Meijer et al., 2024].
-In some aspects, most LLMs align with human moral standards [Tanmay et al., 2023, Nunes et al., 2024, Takemoto, 2024], which may be attributed to their extensive exposure to conventional ethical values during training.
-On the other hand, some research presents a less optimistic view, uncovering significant deviations of LLMs from human moral preferences [Marraffini et al., 2024, Vida et al., 2024, Ahmad and Takemoto, 2024].
-Regarding the underlying mechanisms of LLMs’ moral reasoning, Nunes et al.
-[2024], Ji et al.
-[2024], Simmons [2023] suggest that LLMs primarily exhibit imitation rather than genuine conceptual understanding.
-Comparing moral theories/instruments and suggesting use cases.
-Instruments based on MFT are well-suited for measuring LLMs’ alignment with intuitive moral principles across cultural or political dimensions.
-Tools like MFQ, MFQ-2, MFVs, and MFD are particularly effective when investigating ideological leanings, bias, or internal consistency in abstract vs.
-contextual moral judgments [Nunes et al., 2024].
-In contrast, Kohlberg’s DIT explores structured moral reasoning, offering a developmental perspective on how LLMs process ethical dilemmas [Khandelwal et al., 2024, Tanmay et al., 2023].
-DIT is valuable for longitudinal studies and evaluating whether models show higher-level moral reasoning (e.g., post-conventional reasoning).
-Frameworks such as the Consequentialist-Deontological distinction and global surveys like the PEW 2013 study are useful for categorizing ethical tendencies or analyzing cultural homogeneity [Neuman et al., 2025, Meijer et al., 2024].
-Localized tools (e.g., CMD, JCM) help assess cultural alignment [Liu et al., 2024d, Takeshita et al., 2023].
-Meanwhile, datasets like ETHICS provide morally charged scenarios for detecting normative behavior and fine-tuning LLM responses in high-stakes ethical domains [Hendrycks et al., 2021, Jin et al., 2024b, Marraffini et al., 2024].
+Converging evidence indicates that LLMs generally exhibit a rationalist, consequentialist, and harm-avoidant moral profile, often prioritizing fairness and aligning with human moral standards in various scenarios [neuman2025analyzing, takemoto2024moral, ahmad2024large].
+However, the evaluation results do not necessarily imply stable moral beliefs or suitability for ethical decision-making [nunes2024large, bonagiri2024sage].
+The mechanistic debate and behavioral evidence strongly suggest that LLMs primarily exhibit sophisticated pattern matching and imitation rather than genuine conceptual understanding.
+For instance, simmons2023moral demonstrate that LLMs generate moral rationalizations calibrated to the political identity cued in prompts (“moral mimicry”).
+Furthermore, nunes2024large document a systematic gap between stated principles and applied judgments under MFT, while scherrer2023evaluating reveal probabilistically inconsistent moral beliefs under distributional perturbation; bonagiri2024sage similarly report low moral consistency.
+On complex developmental frameworks like Kohlberg’s DIT, nominal post-conventional performance degrades substantially with language and prompt variation [tanmay2023probing, khandelwal2024moral], consistent with form-matching rather than principled reasoning.
+Beyond mechanistic fragility, research uncovers significant cross-cultural divergence and deviations from diverse human moral preferences [meijer2024llms, jin2024language, vida2024decoding, aksoy2024whose].
+Studies using the Moral Machine paradigm [takemoto2024moral, vida2024decoding, ahmad2024large] and multilingual scenarios [jin2024language, aksoy2024whose] show that LLMs suffer from systematic Western-centric biases, reproducing culturally dominant norms from their training corpora rather than encoding moral pluralism [aksoy2024whose, vida2024decoding].
 4.1.4 Attitudes and opinions
-Attitudes are always attitudes about something.
-This implies three necessary elements: first, there is the object of thought, which is both constructed and evaluated.
-Second, there are acts of construction and evaluation.
-Third, there is the agent, who is doing the constructing and evaluating.
-We can therefore suggest that, at its most general, an attitude is the cognitive construction and affective evaluation of an attitude object by an agent [Bergman, 1998].
-We use the term "attitude" to encompass both attitudes and opinions, following [Bergman, 1998, Ma et al., 2024a].
-Most research on LLM attitudes examines political attitudes and public opinions [Ma et al., 2024a], which are key cognitive and behavioral foundations in human society and closely tied to model fairness, credibility, and social impact [Durmus et al., 2023, Santurkar et al., 2023, Sanders et al., 2023, Lee et al., 2024c, Hartmann et al., 2023].
-For a broader discussion on bias and fairness in LLMs, we refer readers to [Gallegos et al., 2024, Ranjan et al., 2024].
+Attitude refers to the cognitive construction and affective evaluation of an attitude object by an agent [bergman1998theoretical_note_aov].
+We use the term "attitude" to encompass both attitudes and opinions, following [bergman1998theoretical_note_aov, ma2024aov_survey].
+Most research on LLM attitudes examines political attitudes and public opinions [ma2024aov_survey], which are key cognitive and behavioral foundations in human society and closely tied to model fairness, credibility, and social impact [durmus2023towards, santurkar2023whose, sanders2023demonstrations, hartmann2023political].
+For a broader discussion on bias and fairness in LLMs, we refer readers to [gallegos2024bias_fairness_survey, ranjan2024comprehensive_bias_survey].
 Researchers assess LLMs’ political attitudes using standardized questionnaires and scales from political science and social psychology.
-US-focused instruments include the American National Election Studies (ANES) [Qi et al., 2024] and the American Trends Panel (ATP) [Santurkar et al., 2023].
-For cross-national comparisons, studies use tools like the German Longitudinal Election Study (GLES) [Ma et al., 2024b, Ball et al., 2025, von der Heyde et al., 2024].
-The Political Compass Test (PCT) is also widely used to map LLMs within multidimensional political spectrums [Azzopardi and Moshfeghi, 2024, Bernardelle et al., 2024, Hartmann et al., 2023, Röttger et al., 2024].
-Other survey instruments include the General Social Survey (GSS) [Kim and Lee, 2023], the American Community Survey (ACS) [Dominguez-Olmedo et al., 2024], the Canadian Election Study (CES) [Sanders et al., 2023], the European Social Survey (ESS) [Geng et al., 2024], the Survey of Russian Elites [Kalinin, 2023], and the Supreme Court Case Political Evaluation (SCOPE) [Xu et al., 2025c].
-Additionally, researchers have developed specialized datasets and tailored tools for LLMs, such as OpinionQA [Santurkar et al., 2023], a comprehensive dataset based on ATP surveys; IssueBench [Röttger et al., 2025], a benchmark covering controversial issues with multiple response formats; GlobalOpinionQA [Durmus et al., 2023], which extends political opinion evaluation to cross-cultural contexts; and OpinionGPT [Haller et al., 2023], a web tool that demonstrates how input data biases influence model outputs.
+US-focused instruments include the American National Election Studies (ANES) [qi2024representation] and the American Trends Panel (ATP) [santurkar2023whose].
+For cross-national comparisons, studies use tools like the German Longitudinal Election Study (GLES) [ma2024algorithmic, ball2025human, von2024vox].
+The Political Compass Test (PCT) is also widely used to map LLMs within multidimensional political spectrums [azzopardi2024prism, bernardelle2024mapping, hartmann2023political, rottger2024political].
+Other survey instruments include the General Social Survey (GSS) [kim2023ai], the American Community Survey (ACS) [dominguez2024questioning], the Canadian Election Study (CES) [sanders2023demonstrations], the European Social Survey (ESS) [geng2024large], the Survey of Russian Elites [kalinin2023improving], and the Supreme Court Case Political Evaluation (SCOPE) [xu2025better].
+Additionally, researchers have developed specialized datasets and tailored tools for LLMs, such as OpinionQA [santurkar2023whose], a comprehensive dataset based on ATP surveys; IssueBench [rottger2025issuebench], a benchmark covering controversial issues with multiple response formats; GlobalOpinionQA [durmus2023towards], which extends political opinion evaluation to cross-cultural contexts; and OpinionGPT [haller2023opiniongpt], a web tool that demonstrates how input data biases influence model outputs.
 Main findings.
-Pretraining data inherently contains socially biased opinions and perspectives, which can amplify political polarization in LLMs [Feng et al., 2023, Xu et al., 2025c].
-Most studies identify a misalignment between LLM outputs and human opinions [Santurkar et al., 2023, von der Heyde et al., 2024, Yang et al., 2024, Dormuth et al., 2025], with many concluding that LLMs exhibit a left-leaning political bias [Rozado, 2023, Hartmann et al., 2023, Ceron et al., 2024, Bernardelle et al., 2024, Ma et al., 2024b].
-Cross-cultural comparative studies further reveal Western-centric tendencies, demonstrating limited understanding of non-English political perspectives or multi-partisan systems [Qi et al., 2024].
-These findings suggest structural limitations in the models’ political understanding.
-It is also shown that the degree and manifestation of bias vary significantly across contexts and domains.
-For example, political-electoral propositions exhibit different bias patterns than socioeconomic issues like climate change [Wu et al., 2023].
-In contrast, some researchers offer more optimistic perspectives, emphasizing LLMs’ potential to simulate population-level opinions and supplement traditional survey methods [Dominguez-Olmedo et al., 2024, Sanders et al., 2023, Kalinin, 2023, Wu et al., 2023, Jiang et al., 2024d, Argyle et al., 2023, Bisbee et al., 2024, Sun et al., 2024].
-They suggest that with appropriate prompt design, calibration methods, and fine-tuning, LLMs can generate opinion distributions closely approximating human group distributions [Jiang et al., 2022, Wu et al., 2023].
-Such advancements indicate promising directions for using LLMs in opinion research, though their current limitations and biases must still be acknowledged.
-Comparing attitude measurement tools and suggesting use cases.
-The ANES offers the most comprehensive framework for American political contexts, making it ideal for studying model performance on U.S.
-social issues [Jiang et al., 2022].
-Cross-national surveys and datasets such as the ESS, GLES, and GlobalOpinionQA provide broader cultural perspectives across diverse sociopolitical environments [Geng et al., 2024, Ball et al., 2025, Ma et al., 2024b, von der Heyde et al., 2024, Durmus et al., 2023].
-The ATP is better suited for researching model responses to social controversies [Santurkar et al., 2023], while the PCT serves as a theory-driven political spectrum measurement tool, offering simplified yet intuitive insights into LLM political positioning [Röttger et al., 2024].
-Other surveys have distinct regional and sociopolitical focuses: the CES examines political participation in North American multicultural democracies [Sanders et al., 2023]; Survey of Russian Elites provides insights into non-Western political perspectives [Kalinin, 2023]; the SCOPE delivers in-depth assessments of judicial attitudes [Xu et al., 2025c]; and the ACS serves as a tool for researching socioeconomic issues and group representation [Dominguez-Olmedo et al., 2024].
+A substantial body of research identifies a persistent misalignment between LLM outputs and human opinion distributions [santurkar2023whose, von2024vox].
+Multiple studies report a left-libertarian orientation in major instruction-tuned models when evaluated via the Political Compass Test and voting-advice questionnaires [rozado2023political, hartmann2023political, ceron2024beyond].
+Additionally, cross-cultural comparative studies reveal Western-centric tendencies, demonstrating a limited understanding of non-English political perspectives or multi-partisan systems [qi2024representation].
+The measurement framework used to derive these findings is increasingly contested.
+rottger2024political show that political test scores shift substantially across response formats, and dominguez2024questioning find that after correcting for option-position response sets, political survey responses can become statistically random, rendering demographic alignment analyses highly problematic.
+Furthermore, the degree and manifestation of bias vary significantly across contexts; for example, political-electoral propositions exhibit different bias patterns than socioeconomic issues like climate change [yang2024unpacking, ceron2024beyond].
+Despite these structural limitations, some researchers offer more optimistic perspectives on the “silicon sampling” hypothesis—the idea that LLMs can simulate population-level opinions and supplement traditional survey methods [argyle2023_out_of_one, sanders2023demonstrations, kalinin2023improving, sun2024random].
+They suggest that with appropriate prompt design, calibration methods, and fine-tuning, LLMs can generate opinion distributions closely approximating human group distributions.
+However, this hypothesis is challenged by evidence of underrepresentation, misportrayal, and flattening of identity groups [wang2025large, bisbee2024synthetic].
 4.2 Measuring cognitive constructs
-Traditional NLP benchmarks are insufficient to capture and analyze emergent cognitive constructs [Hagendorff et al., 2024, Ying et al., 2025a].
-To address this gap, researchers are adapting related psychometric techniques to evaluate LLM abilities.
-Hagendorff et al.
-[2024] introduce the concept of machine psychology, reviewing emergent LLM capabilities and advocating for psychometric evaluation.
-They categorize these cognitive constructs into four key aspects: heuristics and biases, social interactions, psychology of language, and learning and cognitive capabilities.
-Structured based on this taxonomy, Table 4 summarizes the representative cognitive constructs measured in recent research.
-Aspects Sub-aspects Related work
-Heuristics & biases - Yax et al.
-[2024], Sundaram and Alwar [2024], Momennejad et al.
-[2023], Echterhoff et al.
-[2024], Saeedi et al.
-[2025], Ranaldi and Zanzotto [2024], Macmillan-Scott and Musolesi [2024], Ando et al.
-[2023], Malberg et al.
-[2024], Xie et al.
-[2024b], Hayes et al.
-[2024], Shah et al.
-[2024], Chen and Eger [2025], Tang and Kejriwal [2024], Talboy and Fuller [2023], Coda-Forno et al.
-[2024], Abramski et al.
-[2023], Binz and Schulz [2023], Xie et al.
-[2024a], Kumar et al.
-[2024], Schulze Buschoff et al.
-[2025], Scholten et al.
-[2024], Shaikh et al.
-[2024], Thorstad [2023], Hagendorff et al.
-[2023], Healey et al.
-[2024], Bai et al.
-[2024], Castello et al.
-[2024]
-Social interactions Theory of Mind Sarıtaş et al.
-[2025], Ma et al.
-[2023], Mao et al.
-[2024b] (reviews); Ullman [2023], Sap et al.
-[2022], Kosinski [2023a], Shapira et al.
-[2023a], Gandhi et al.
-[2023], Strachan et al.
-[2024a], Kosinski [2023b], Moghaddam and Honey [2023], Street et al.
-[2024], Sclar et al.
-[2023], van Duijn et al.
-[2023], He et al.
-[2023], Holterman and Deemter [2023], Chen et al.
-[2024d], Zhu et al.
-[2024c], Hou et al.
-[2024], Amirizaniani et al.
-[2024], Soubki et al.
-[2024], Jones et al.
-[2024], Lin et al.
-[2024], Sclar et al.
-[2024], Sadhu et al.
-[2024], Lorè et al.
-[2024], Sarangi et al.
-[2025], Yu et al.
-[2025], Tang and Belle [2024], Kim et al.
-[2025a], Chan et al.
-[2024], Nickel et al.
-[2024], Xu et al.
-[2024a], Leer et al.
-[2023], Jamali et al.
-[2023], Li et al.
-[2023b], Wilf et al.
-[2023], Chen et al.
-[2024c], Strachan et al.
-[2024b], Xu et al.
-[2025a], Zhang et al.
-[2025], Shinoda et al.
-[2025], Jin et al.
-[2024a], Yang et al.
-[2025a], Zhou et al.
-[2023a], Pi et al.
-[2024], Riemer et al.
-[2025], Tan et al.
-[2024].
-Emotional Intelligence Sorin et al.
-[2024], Raj [2024] (reviews); Sabour et al.
-[2024], Paech [2023], Wang et al.
-[2023b], Elyoseph et al.
-[2023], Patel and Fan [2023], Vzorinab et al.
-[2024], Hu et al.
-[2025], Li et al.
-[2024e], Zhao et al.
-[2024], Chen et al.
-[2024b], Huang et al.
-[2024b], Schaaff et al.
-[2023], Li et al.
-[2023a], Huang et al.
-[2023b], Lee et al.
-[2024f], Welivita and Pu [2024]
-Social Intelligence Leng and Yuan [2023], Mittelstädt et al.
-[2024], Shapira et al.
-[2023b], Zhou et al.
-[2024b], Wang et al.
-[2024d], Liu et al.
-[2024c], Wang et al.
-[2024a], Mou et al.
-[2024], Liang et al.
-[2024], Zadeh et al.
-[2019], Xu et al.
-[2024b], Guo et al.
-[2023a], Kovač et al.
-[2024, 2021], Mathur et al.
-[2024], Schoenegger et al.
-[2025]
-Psychology of language Language comprehension Chang and Bergen [2024] (reviews); Duan et al.
-[2024a], Wang et al.
-[2024b], Wilcox et al.
-[2021], Arehalli et al.
-[2022], Li et al.
-[2024a], Steuer et al.
-[2023], Duan et al.
-[2024b], Qiu et al.
-[2024], Bojić et al.
-[2023], Hu et al.
-[2023], Ruis et al.
-[2023], Lee et al.
-[2024b], Seals and Shalin [2023], Hu and Levy [2023], He et al.
-[2024], Sun and Wang [2024], Hong et al.
-[2023], Amouyal et al.
-[2024], Ide et al.
-[2024], Lee et al.
-[2024b], Huff and Ulakçi [2024], Zhou et al.
-[2023b], Tian et al.
-[2024]
-Language generation Seals and Shalin [2023], Tian et al.
-[2024], Tang and Kejriwal [2024], Stevenson et al.
-[2022], Hubert et al.
-[2023], Chakrabarty et al.
-[2024], Orwig et al.
-[2024], Lee and Chung [2024], Boussioux et al.
-[2024], Miaschi et al.
-[2024], Bellemare-Pepin et al.
-[2024], Cai et al.
-[2024]
-Language acquisition Steuer et al. [2023], Shah et al. [2024], Frank [2023b], Riva et al. [2024]
-Learning and cognitive capabilities - Niu et al.
-[2024] (reviews); Galatzer-Levy et al.
-[2024], Wang et al.
-[2024e], Dayan et al.
-[2024], Wu et al.
-[2025], Sartori and Orrú [2023], Webb et al.
-[2022], Song et al.
-[2024a], Zhang et al.
-[2023b], Wang et al.
-[2024e], Coda-Forno et al.
-[2024], Zeng et al.
-[2024c], Lv et al.
-[2024]
-Table 4: Representative LLM cognitive constructs in LLM psychometrics.
+Traditional NLP benchmarks struggle to capture cognitive abilities in state-of-the-art LLMs [hagendorff2024machine, ying2025benchmarking, chollet2025arc], leading researchers to adapt psychometric techniques for LLM evaluation.
+hagendorff2024machine propose “machine psychology,” categorizing cognitive constructs into four aspects: heuristics and biases, social interactions, psychology of language, and learning and cognitive capabilities.
+We organize the discussions below following this taxonomy.
 4.2.1 Heuristics and biases
-Heuristics and biases are mental shortcuts that simplify decision-making but can introduce systematic errors [Tversky and Kahneman, 1974].
+Heuristics and biases are mental shortcuts that simplify decision-making but can introduce systematic errors [tversky1974judgment].
 Recent research employs psychometric tools to evaluate rationality and biases in LLM outputs and offer theoretical explanations for these biases.
-Early studies, such as Binz and Schulz [2023], found that GPT-3 performs comparably to humans on cognitive ability tests, exhibiting biases like framing and anchoring.
-More advanced models, such as GPT-4, show fewer System 1 (intuitive) errors and more System 2 (deliberative) reasoning, though biases persist [Hagendorff et al., 2023].
-Large-scale evaluations now cover a wide range of biases—including conjunction fallacy, unwarranted beliefs, and loss aversion—using synthetic scenarios and multi-agent dialogue [Malberg et al., 2024, Xie et al., 2024b].
-Multimodal models are also being assessed for similar patterns [Schulze Buschoff et al., 2025].
-To explain these findings, researchers draw on dual-process theory, linking LLM reasoning to intuitive and deliberative modes [Hagendorff et al., 2023, Yax et al., 2024].
-Other frameworks highlight LLMs’ lack of metacognitive abilities, leading to “metacognitive myopia”—the inability to monitor or control reasoning processes, which perpetuates systematic biases [Scholten et al., 2024].
-Additional perspectives use cognitive dissonance and elaboration likelihood theories to interpret inconsistencies [Sundaram and Alwar, 2024], while mechanistic approaches leverage influence graphs and Shapley values to trace the origins of biases in model training [Shaikh et al., 2024].
+Early studies, such as binz2023using, found that GPT-3 performs comparably to humans on cognitive ability tests, exhibiting biases like framing and anchoring.
+More advanced models, such as GPT-4, show fewer System 1 (intuitive) errors and more System 2 (deliberative) reasoning, though biases persist [hagendorff2023reasoning_bias].
+Large-scale evaluations now cover a wide range of biases (e.g., conjunction fallacy, unwarranted beliefs, and loss aversion) using synthetic scenarios and multi-agent dialogue [malberg2024comprehensive_cognitive_biases, xie2024mindscope].
+Multimodal models are also being assessed for similar patterns [schulze2025visual].
+To explain these findings, researchers draw on dual-process theory, linking LLM reasoning to intuitive and deliberative modes [hagendorff2023reasoning_bias, yax2024studying].
+Other frameworks highlight LLMs’ lack of metacognitive abilities, i.e., the inability to monitor or control reasoning processes.
+Such inability perpetuates systematic biases in LLMs [scholten2024metacognitive].
+Additional perspectives use cognitive dissonance and elaboration likelihood theories to interpret inconsistencies [sundaram2024can], while mechanistic approaches leverage influence graphs and Shapley values to trace the origins of biases in model training [shaikh2024cbeval].
 Main findings.
 By administering psychometrics to LLMs, studies consistently find that LLMs exhibit cognitive biases superficially similar to humans, such as anchoring, framing, and the conjunction fallacy.
-Large-scale tests, such as [Echterhoff et al., 2024, Malberg et al., 2024, Xie et al., 2024b], systematically categorize these biases and enable scalable evaluations.
-Some note that newer, larger, chain-of-thought-enabled models exhibit improved reasoning and bias mitigation [Tang and Kejriwal, 2024, Hagendorff et al., 2023], while others argue that increasing model complexity without deliberate bias mitigation strategies can amplify existing biases [Kumar et al., 2024].
-In addition, researchers explore mechanistic interpretability [Shaikh et al., 2024] and present theoretical explanations from cognitive psychology and reasoning theories [Hagendorff et al., 2023, Yax et al., 2024, Scholten et al., 2024].
-While LLMs exhibit dual-process reasoning dynamics reminiscent of human cognition [Hagendorff et al., 2023], detailed analysis reveals differences between LLM and human reasoning [Yax et al., 2024].
+Large-scale tests, such as [echterhoff2024cognitive, malberg2024comprehensive_cognitive_biases, xie2024mindscope], systematically categorize these biases and enable scalable evaluations.
+Some note that newer, larger, chain-of-thought-enabled models exhibit improved reasoning and bias mitigation [tang2024humanlike, hagendorff2023reasoning_bias], while others argue that increasing model complexity without deliberate bias mitigation strategies can amplify existing biases [kumar2024investigating].
+In addition, researchers explore mechanistic interpretability [shaikh2024cbeval] and present theoretical explanations from cognitive psychology and reasoning theories [hagendorff2023reasoning_bias, yax2024studying, scholten2024metacognitive].
+While LLMs exhibit dual-process reasoning dynamics reminiscent of human cognition [hagendorff2023reasoning_bias], detailed analysis reveals differences between LLM and human reasoning [yax2024studying].
 4.2.2 Social interactions
@@
 Theory of Mind (ToM).
-ToM is the ability to attribute mental states such as beliefs, intentions, and knowledge to others [Premack and Woodruff, 1978].
+ToM is the ability to attribute mental states such as beliefs, intentions, and knowledge to others [premack1978tom].
 Advanced LLMs can simulate ToM-like reasoning under certain conditions, which prompts questions about how these behaviors arise and how robustly they generalize.
-Foundational studies apply classic psychometrics—false belief tasks—to evaluate ToM in LLMs.
-[Kosinski, 2023a, b] report that GPT-3.5 and GPT-4 perform at levels matching or exceeding those of children in structured ToM tasks, suggesting the spontaneous emergence of ToM-like behavior as models scale in size.
-Through mechanistic interpretation, Jamali et al.
-[2023] offer evidence of parallels between the LLM embeddings and neurons in the human brain.
+Foundational studies apply classic psychometrics (false belief tasks) to evaluate ToM in LLMs.
+[tom3_kosinski2023theory, tom7_kosinski2024evaluating] report that GPT-3.5 and GPT-4 perform at levels matching or exceeding those of children in structured ToM tasks, suggesting that ToM-like behavior appears to develop as models scale in size.
+Through mechanistic interpretation, tom34_jamali2023unveiling offer evidence of parallels between the LLM embeddings and neurons in the human brain.
 However, these claims have been challenged.
-Ullman [2023], Shapira et al.
-[2023a], Holterman and Deemter [2023] show that minor task changes often cause LLMs’ performance to drop, suggesting reliance on brittle heuristics rather than true semantic understanding.
-To address this, researchers have developed more robust benchmarks and evaluation protocols featuring procedurally generated, higher-order, or broader ToM tasks [Gandhi et al., 2023, Chen et al., 2024d, He et al., 2023, Xu et al., 2024a, Shinoda et al., 2025].
-These benchmarks frequently reveal significant declines in performance on complex tasks, such as 6th-order belief attribution [Street et al., 2024] or faux pas detection [Strachan et al., 2024a].
-Some argue these failures stem from a lack of general commonsense reasoning rather than an inability to represent mental states [Pi et al., 2024].
-Riemer et al.
-[2025] further suggest distinguishing between literal and functional ToM, noting that current benchmarks inadequately assess functional ToM.
-While newer models continue to improve on structured ToM tasks [Strachan et al., 2024a, Street et al., 2024], they remain inconsistent in open-ended, adversarial, or pragmatic reasoning scenarios [Sclar et al., 2024, Yu et al., 2025, Nickel et al., 2024].
-Comparisons with human baselines show that LLMs can approximate ToM behavior in narrow contexts but still fall short of general human-like social cognition [Strachan et al., 2024a, van Duijn et al., 2023, Jones et al., 2024].
-Recent work has also extended ToM evaluation to multimodal settings [Chen et al., 2024c, Strachan et al., 2024b, Jin et al., 2024a] and multi-agent interactions [Li et al., 2023b].
+tom1_ullman2023large, tom4_shapira2023clever, tom13_holterman2023does show that minor task changes often cause LLMs’ performance to drop, suggesting reliance on brittle heuristics rather than true semantic understanding.
+To address this, researchers have developed more robust benchmarks and evaluation protocols featuring procedurally generated, higher-order, or broader ToM tasks [tom5_gandhi2023understanding, tom14_chen2023tobench, tom12_he2023hi, tom32_xu2024opentom, tom41_shinoda2025tomato].
+These benchmarks frequently reveal significant declines in performance on complex tasks, such as 6th-order belief attribution [tom9_street2024llms] or faux pas detection [tom6_strachan2024testing].
+Some argue these failures stem from a lack of general commonsense reasoning rather than an inability to represent mental states [tom45_pi2024dissecting].
+tom_riemer2024position_broken further suggest distinguishing between literal and functional ToM, noting that current benchmarks inadequately assess functional ToM.
+While newer models continue to improve on structured ToM tasks [tom6_strachan2024testing, tom9_street2024llms], they remain inconsistent in open-ended, adversarial, or pragmatic reasoning scenarios [tom22_sclar2024explore, tom27_yu2025persuasive, tom31_zhang2025tom].
+Comparisons with human baselines show that LLMs can approximate ToM behavior in narrow contexts but still fall short of general human-like social cognition [tom6_strachan2024testing, tom11_van2023theory, tom20_jones2024comparing].
+Recent work has also extended ToM evaluation to multimodal settings [tom37_chen2024mm, tom38_james2024gpt4o, tom42_jin2024mmtom] and multi-agent interactions [tom35_zhang2024tom].
 Main findings in ToM.
@@
 The proficiency of LLMs in ToM remains a contentious issue.
-Interested readers are referred to recent comprehensive reviews on ToM in LLMs [Sarıtaş et al., 2025, Ma et al., 2023, Mao et al., 2024b].
+Interested readers are referred to recent comprehensive reviews on ToM in LLMs [saritacs2025tom_survey, tom_ma2023position, mao2024_machine_tom_review].
 Emotional Intelligence (EI).
-EI is the subset of social intelligence that involves the ability to monitor one’s own and others’ feelings and emotions, to discriminate among them and to use this information to guide one’s thinking and actions [Salovey and Mayer, 1990].
+EI is the subset of social intelligence that involves the ability to monitor one’s own and others’ feelings and emotions, to discriminate among them and to use this information to guide one’s thinking and actions [salovey1990emotional].
 Recent work introduces benchmarks to assess EI in LLMs using structured, theory-driven tasks.
-Tools like EmoBench [Sabour et al., 2024] and EQ-Bench [Paech, 2023] operationalize constructs such as emotional understanding and regulation, often referencing frameworks like Mayer-Salovey-Caruso Emotional Intelligence Test.
-Other studies adapt established psychometric instruments, including Situational Evaluation of Complex Emotional Understanding (SECEU) [Wang et al., 2023b], the Levels of Emotional Awareness Scale (LEAS) [Elyoseph et al., 2023], the Toronto Alexithymia Scale (TAS-20) and Empathy Quotient (EQ-60) [Patel and Fan, 2023], among others [Huang et al., 2024b].
-Advanced models like GPT-4 often match or surpass human baselines in emotional awareness and understanding [Wang et al., 2023b, Elyoseph et al., 2023, Patel and Fan, 2023], though they lack deep reflexive analysis of emotional experience and motivation [Vzorinab et al., 2024].
-Recent benchmarks also extend EI evaluation to multimodal settings [Hu et al., 2025].
+Tools like EmoBench [sabour2024emobench] and EQ-Bench [paech2023eqbench] operationalize constructs such as emotional understanding and regulation, often referencing frameworks like Mayer-Salovey-Caruso Emotional Intelligence Test.
+Other studies adapt established psychometric instruments, including Situational Evaluation of Complex Emotional Understanding (SECEU) [wang2023emotional], the Levels of Emotional Awareness Scale (LEAS) [elyoseph2023chatgpt], the Toronto Alexithymia Scale (TAS-20) and Empathy Quotient (EQ-60) [patel2023identification], among others [huang2024apathetic].
+Advanced models like GPT-4 often match or surpass human baselines in emotional awareness and understanding [wang2023emotional, elyoseph2023chatgpt, patel2023identification], though they lack deep reflexive analysis of emotional experience and motivation [vzorinab2024emotional].
+Recent benchmarks also extend EI evaluation to multimodal settings [hu2025emobench].
 Main findings in EI.
-Advanced LLMs generally perform on par with or better than humans in tasks of EI [Elyoseph et al., 2023, Patel and Fan, 2023], despite evident limitations in several areas, such as displaying artificial or mechanical patterns when expressing empathy [Lee et al., 2024f], the deep reflexive analysis of emotional experiences [Vzorinab et al., 2024], and the misalignment with human emotional behaviors [Huang et al., 2023b].
+Advanced LLMs can perform on par with or better than humans in some tasks of EI [elyoseph2023chatgpt, patel2023identification].
+However, they show evident limitations in several areas, such as displaying artificial or mechanical patterns when expressing empathy [lee2024large], the deep reflexive analysis of emotional experiences [vzorinab2024emotional], and the misalignment with human emotional behaviors [huang2023emotionbench].
 Social Intelligence (SI).
-SI is the ability to understand and manage people [Thorndike and Stein, 1937].
+SI is the ability to understand and manage people [thorndike1937evaluation].
 SI in LLMs determines how well these models interpret and respond to social situations.
 Psychometric tools applied to LLMs yield mixed results.
-Mittelstädt et al.
-[2024] use Situational Judgment Tests (SJTs) and find some LLMs outperform humans in expert-rated social appropriateness.
-However, LLMs still struggle with implicit social understanding: Shapira et al.
-[2023b] show LLMs perform poorly on faux pas tests, and Xu et al.
-[2024b] find superficial friendliness often leads to errors in the Situational Evaluation of SI (SESI).
-The AgentSense benchmark further highlights LLMs’ limitations in complex social interactions, especially regarding high-level needs and private information [Mou et al., 2024].
+si1_mittelstadt2024large use Situational Judgment Tests (SJTs) and find some LLMs outperform humans in expert-rated social appropriateness.
+However, LLMs still struggle with implicit social understanding: si2_shapira2023how show LLMs perform poorly on faux pas tests, and si11_xu2024academically find superficial friendliness often leads to errors in the Situational Evaluation of SI (SESI).
+The AgentSense benchmark further highlights LLMs’ limitations in complex social interactions, especially regarding high-level needs and private information [si8_mou2024agentsense].
 Recent work expands to multi-agent environments.
-SOTOPIA [Zhou et al., 2024b] and SOTOPIA-$\pi$ [Wang et al., 2024d] simulate cooperative and competitive social contexts, while the STSS benchmark [Wang et al., 2024a] evaluates SI through task-oriented simulations focused on outcomes and goals.
-The CogMir framework [Liu et al., 2024c] finds LLMs and humans are similarly consistent in irrational and prosocial decisions under uncertainty.
-DeSIQ [Guo et al., 2023a] extends SI evaluation to multimodal settings.
-Critically, Kovač et al.
-[2024] argue current benchmarks lack developmental grounding and propose the SocialAI school as a more comprehensive framework for studying LLM-based SI.
+SOTOPIA [si3_zhou2023sotopia] and SOTOPIA-$\pi$ [si4_wang2024sotopia] simulate cooperative and competitive social contexts, while the STSS benchmark [si6_wang2024towards] evaluates SI through task-oriented simulations focused on outcomes and goals.
+The CogMir framework [si5_liu2024exploring] finds LLMs and humans are similarly consistent in irrational and prosocial decisions under uncertainty.
+DeSIQ [si12_guo2023desiq] extends SI evaluation to multimodal settings.
+Critically, si14_kovac2024socialai argue current benchmarks lack developmental grounding and propose the SocialAI school as a more comprehensive framework for studying LLM-based SI.
 Main findings in SI.
 LLMs show measurable competence in rule-based, socially appropriate behavior, particularly in structured environments.
-They perform well on: 1) following predefined social norms [Mittelstädt et al., 2024]; 2) completing interactional goals in multi-agent settings [Zhou et al., 2024b, Wang et al., 2024d]; 3) replicating human-like prosocial decisions [Liu et al., 2024c]; and 4) being persuasive, even better than human persuaders [Schoenegger et al., 2025].
-Weaker performance is noted in tasks where superficial friendliness causes errors [Xu et al., 2024b], understanding high-level growth needs is essential [Mou et al., 2024], and LLMs must implicitly describe social situations [Shapira et al., 2023b].
+They perform well on: 1) following predefined social norms [si1_mittelstadt2024large]; 2) completing interactional goals in multi-agent settings [si3_zhou2023sotopia, si4_wang2024sotopia]; 3) replicating human-like prosocial decisions [si5_liu2024exploring]; and 4) being persuasive, even better than human persuaders [schoenegger2025large].
+Weaker performance is noted in tasks where superficial friendliness causes errors [si11_xu2024academically], understanding high-level growth needs is essential [si8_mou2024agentsense], and LLMs must implicitly describe social situations [si2_shapira2023how].
 4.2.3 Psychology of language
-Psycholinguistics, a subfield of psychology, explores how humans comprehend, generate, and acquire language [Carroll, 1986].
+Psycholinguistics, a subfield of psychology, explores how humans comprehend, generate, and acquire language [carroll1986psychology].
 Insights from this discipline help evaluate how LLMs process language and mirror human linguistic features.
@@
 LLM language comprehension is evaluated across multiple linguistic levels—sound, word, syntax, meaning, and discourse.
-Duan et al.
-[2024a] introduce a benchmark of 10 psycholinguistic tests to assess these comprehensively.
-Many studies focus on specific aspects, such as syntactic and semantic processing [Wang et al., 2024b, Arehalli et al., 2022, Wolfman et al., 2024, Wilcox et al., 2021].
-Surprisal, a measure of word predictability, is widely used to model processing difficulty in syntactic ambiguities and hierarchical structures [Wilcox et al., 2021, Arehalli et al., 2022, Li et al., 2024a].
-Results are mixed: for example, GPT-2-XL shows human-like competence in sound-gender association and implicit causality, but not in sound-shape association [Duan et al., 2024b].
-Other evaluation tasks include grammaticality judgment [Qiu et al., 2024, Ide et al., 2024], pragmatic inference [Bojić et al., 2023, Hu et al., 2023, Ruis et al., 2023], argument role processing [Lee et al., 2024b], and discourse comprehension [Duan et al., 2024a].
-Advanced models like GPT-4 often match or surpass humans in pragmatic reasoning [Bojić et al., 2023] and grammaticality judgment [Dentella et al., 2024], though performance varies with prompt format [Hu and Levy, 2023] and implicature understanding remains limited [Ruis et al., 2023].
+pl1_duan2024hlb introduce a benchmark of 10 psycholinguistic tests to assess these comprehensively.
+Many studies focus on specific aspects, such as syntactic and semantic processing [pl3_wang2024how, pl6_arehalli2022syntactic, pl8_wolfman2024hierarchical, pl5_wilcox2021targeted].
+Surprisal, a measure of word predictability, is widely used to model processing difficulty in syntactic ambiguities and hierarchical structures [pl5_wilcox2021targeted, pl6_arehalli2022syntactic, pl17_li2024incremental].
+Results are mixed: for example, GPT-2-XL shows human-like competence in sound-gender association and implicit causality, but not in sound-shape association [pl13_duan2024unveiling].
+Other evaluation tasks include grammaticality judgment [pl18_qiu2024evaluating, pl29_ide2024how], pragmatic inference [pl28_bojic2023gpt, pl_hu2023fine, pl_ruis2023goldilocks], argument role processing [pl14_lee2024a], and discourse comprehension [pl1_duan2024hlb].
+Advanced models like GPT-4 often match or surpass humans in pragmatic reasoning [pl28_bojic2023gpt] and grammaticality judgment [pl21_dentella2024language], though performance varies with prompt format [pl7_hu2023prompt] and implicature understanding remains limited [pl_ruis2023goldilocks].
 Language generation.
-A rich line of research evaluates the creativity of LLM language generation [Tang and Kejriwal, 2024, Stevenson et al., 2022, Hubert et al., 2023, Chakrabarty et al., 2024, Orwig et al., 2024, Lee and Chung, 2024, Boussioux et al., 2024, Bellemare-Pepin et al., 2024], using tests like Guilford’s Alternative Uses Test (AUT) [Guilford et al., 1978] and the Torrance Test of Creative Thinking (TTCT) [Torrance, 1966].
+A rich line of research evaluates the creativity of LLM language generation [pl_stevenson2022putting, pl_hubert2023artificial, pl_chakrabarty2024art, pl_orwig2024language, pl_boussioux2024crowdless, bellemare2024divergent], using tests like Guilford’s Alternative Uses Test (AUT) [guilford1978alternate] and the Torrance Test of Creative Thinking (TTCT) [torrance1966torrance].
 Early models such as GPT-3 lack originality and novelty, while advanced LLMs like GPT-4 surpass the human average in creativity.
-Notably, Tang and Kejriwal [2024] find LLMs underperform in divergent creativity (e.g., novel uses for familiar objects) but can rival humans in open-ended creative writing.
-Cai et al.
-[2024] show that ChatGPT produces human-like responses in most of 12 psycholinguistic tests, but AI-generated analogies often lack human-like linguistic properties [Seals and Shalin, 2023].
-LLM-generated stories tend to be positive and lack suspense or narrative diversity compared to human writing [Tian et al., 2024].
-Linguistic profiling has also been used to characterize the task-specific language abilities of LLMs [Miaschi et al., 2024].
+Notably, tang2024humanlike find LLMs underperform in divergent creativity (e.g., novel uses for familiar objects) but can rival humans in open-ended creative writing.
+cai2024large show that ChatGPT produces human-like responses in most of 12 psycholinguistic tests, but AI-generated analogies often lack human-like linguistic properties [pl_seals2023long].
+LLM-generated stories tend to be positive and lack suspense or narrative diversity compared to human writing [pl_tian2024large].
+Linguistic profiling has also been used to characterize the task-specific language abilities of LLMs [miaschi2024evaluating].
 Language acquisition.
-Work on developmental plausibility examines whether LLMs replicate stages of language learning analogous to those in children [Steuer et al., 2023, Shah et al., 2024].
-It is shown that regardless of model size, the developmental trajectories of pretrained language models consistently exhibit a window of maximal alignment with human cognitive development [Shah et al., 2024].
-Another study by Frank [2023b] draws inspiration from human language development to explain the data inefficiency of LLMs; the relative efficiency of human language acquisition is possibly due to pre-existing conceptual knowledge, multimodal grounding, and the interactive, social nature of their input.
+Work on developmental plausibility examines whether LLMs replicate stages of language learning analogous to those in children [pl10_steuer2023large, shah2024development].
+It is shown that regardless of model size, the developmental trajectories of pretrained language models consistently exhibit a window of maximal alignment with human cognitive development [shah2024development].
+Another study by pl_frank2023bridging draws inspiration from human language development to explain the data inefficiency of LLMs; the relative efficiency of human language acquisition is possibly due to pre-existing conceptual knowledge, multimodal grounding, and the interactive, social nature of their input.
 Main findings.
-Early foundations that evaluate language models based on BERT and LSTM link computational linguistics and psycholinguistic mechanisms [Ettinger, 2020, Futrell et al., 2019, Arehalli and Linzen, 2020]; these models generally fall short in diverse psycholinguistic tasks.
-Recent studies extend evaluations to LLMs and broader tasks [Duan et al., 2024a], where advanced LLMs are shown to surpass humans on tasks such as pragmatic reasoning [Bojić et al., 2023] and creative writing [Tang and Kejriwal, 2024].
-Some deficiencies still exist, such as limited implicature understanding [Ruis et al., 2023], prompt-sensitive linguistic competence [Hu and Levy, 2023], and lack of human-like psycholinguistic properties [Seals and Shalin, 2023, Tian et al., 2024].
-In addition, mixed results are found in the alignment between LLMs and human linguistic cognition [Duan et al., 2024b, Wolfman et al., 2024] and language acquisition [Shah et al., 2024, Frank, 2023b].
+Early foundations that evaluate language models based on BERT and LSTM link computational linguistics and psycholinguistic mechanisms [ettinger2020bert, futrell2019neural, arehalli2020neural]; these models generally fall short in diverse psycholinguistic tasks.
+Recent studies extend evaluations to LLMs and broader tasks [pl1_duan2024hlb], where advanced LLMs are shown to surpass humans on tasks such as pragmatic reasoning [pl28_bojic2023gpt] and creative writing [tang2024humanlike].
+Some deficiencies still exist, such as limited implicature understanding [pl_ruis2023goldilocks], prompt-sensitive linguistic competence [pl7_hu2023prompt], and lack of human-like psycholinguistic properties [pl_seals2023long, pl_tian2024large].
+In addition, alignment between LLMs and humans shows mixed results in both linguistic cognition [pl13_duan2024unveiling, pl8_wolfman2024hierarchical] as well as language acquisition [shah2024development, pl_frank2023bridging].
 4.2.4 Learning and cognitive capabilities
 Psychometrics of learning and cognitive capabilities measures mental functions such as memory, reasoning, problem-solving, and comprehension to identify cognitive strengths and weaknesses.
-Recent studies adapt human psychometric tools to evaluate LLMs’ cognitive abilities.
-For example, Galatzer-Levy et al.
-[2024] use the Wechsler Adult Intelligence Scale (WAIS-IV) to show LLMs generally reach top human levels in verbal comprehension and working memory, though multimodal models lag in visual reasoning.
-Raven’s Progressive Matrices and other fluid intelligence tests reveal LLMs can match or surpass humans in analogical reasoning [Sartori and Orrú, 2023, Webb et al., 2022].
-However, limitations persist.
-The Montreal Cognitive Assessment (MoCA) shows most LLMs display mild cognitive impairment, especially in visuospatial and executive tasks [Dayan et al., 2024].
-Abstraction and Reasoning Challenge (ARC) further highlights deficits in abstract problem-solving [Wu et al., 2025].
-Some works question the validity of current tests [Löhn et al., 2024, Li et al., 2024d, Zhang et al., 2023b], prompting new benchmarks and methodologies [Coda-Forno et al., 2024, Wang et al., 2024e, Zhuang et al., 2023a, Zeng et al., 2024c, Song et al., 2024a].
-Developmental perspectives also emerge: Wang et al.
-[2024e] leverage Piaget’s Theory of Cognitive Development to track human-like progression in LLMs, finding advanced models (e.g., GPT-4) reach cognitive abilities comparable to 20-year-old humans.
-Shah et al.
-[2024] identify developmental profiles in LLMs paralleling human development via psychometric theories like Cattell-Horn-Carroll.
+Early adaptations of human psychometric tools to LLMs reported striking successes.
+For instance, l1_galatzer2024cognitive used the Wechsler Adult Intelligence Scale (WAIS-IV) to show LLMs reaching top human levels in verbal comprehension, while other studies indicated that models could match or surpass humans in analogical reasoning on tests like Raven’s Progressive Matrices [l9_sartori2023language, l10_webb2022emergent].
+However, a critical re-examination suggests these findings may be unreliable.
+High accuracy on such benchmarks often reflects “shortcut learning”, rather than genuine intelligence.
+Models exploit statistical patterns without engaging in true abstract reasoning.
+beger2025ai reveal that while models may achieve state-of-the-art results on the ARC-AGI benchmark, their solutions frequently rely on surface-level heuristics that fail to capture the intended abstractions.
+This brittleness is further exposed when models are tested for robustness and generalization: unlike humans, LLMs often fail to transfer analogical reasoning to novel domains or slightly modified tasks [lewis2024evaluatingrobustness, stevenson2024can].
+From this perspective, documented failures on tests like the Montreal Cognitive Assessment (MoCA) [l3_dayan2024age] and the Abstraction and Reasoning Challenge (ARC) [l8_wu2025understanding, chollet2025arc] appear not as isolated deficits but as symptoms of a fundamental limitation in deep understanding.
+The “emergence” of reasoning capabilities is thus contested, with arguments that these traits may result from massive scale and memorization rather than novel cognitive properties [krakauer2025large].
+Overall, LLMs are widely acknowledged to exhibit a “jagged intelligence,” characterized by uneven performance across tasks: while specific training and scaffolding have unlocked significant success in specialized domains such as mathematical Olympiads [luong2025advanced], these models often struggle with some tasks that could be simple and intuitive for humans.
 Main findings.
-LLMs demonstrate strong performance on verbal comprehension, working memory, and analogical reasoning, often reaching or surpassing high human percentiles when evaluated with adapted psychometric tools like WAIS-IV and Raven’s Matrices [Galatzer-Levy et al., 2024, Webb et al., 2022].
-However, they exhibit notable cognitive deficits, particularly on benchmarks like MoCA and ARC tasks [Wu et al., 2025, Dayan et al., 2024].
-Several works introduce developmental and cognitive benchmarks (e.g., CogLM, CogBench) grounded in psychological theories to track reasoning and adaptability [Wang et al., 2024e, Coda-Forno et al., 2024].
-While emergent reasoning appears at scale, challenges remain in interpretability, test validity, and generalization beyond surface-level statistical patterns [Li et al., 2024d, Löhn et al., 2024].
+While early studies using adapted tools like WAIS-IV and Raven’s Matrices suggest LLMs possess human-like cognitive abilities [l1_galatzer2024cognitive, l10_webb2022emergent], a critical evaluation reveals these claims to be potentially misleading.
+Current methods often fail to account for shortcut learning, where high accuracy masks a lack of genuine abstract reasoning [beger2025ai].
+The cognitive abilities of LLMs are shown to be brittle, collapsing under task modification or transfer, which starkly contrasts with human cognition [lewis2024evaluatingrobustness, stevenson2024can].
+Deficits on benchmarks like MoCA and ARC highlight a fundamental lack of robust, generalizable understanding.
+Overall, models demonstrate a “jagged intelligence,” achieving high performance in specialized domains (e.g., mathematics and coding) through targeted training and scaffolding [luong2025advanced, novikov2025alphaevolve], while failing on tasks that are trivial and intuitive for humans.
+This reveals a fundamentally uneven cognitive profile compared to human intelligence.
+Future psychometric frameworks should move beyond accuracy-based metrics to assess the robustness, process, and validity of reasoning [mitchell2025artificial].
 5 Psychometric evaluation methodology
@@
 This section examines the methodologies employed in LLM Psychometrics.
-As illustrated in Fig.
-3, the methodological framework encompasses four key components: test formats (§ 5.1), data and task sources (§ 5.2), prompting strategies (§ 5.3), and model output and scoring (§ 5.4).
-The configuration of inference parameters (§ 5.5) is additionally discussed.
+As illustrated in Fig.˜3, the methodological framework encompasses four key components: test formats (§˜5.1), data and task sources (§˜5.2), prompting strategies (§˜5.3), and model output and scoring (§˜5.4).
+The configuration of inference parameters (§˜5.5) is additionally discussed.
 5.1 Test format
@@
 Construct Structured tests Open-ended conversations Agentic simulations
-Personality traits Jiang et al. [2023], Serapio-García et al. [2023] Jiang et al. [2024b], Zheng et al. [2025] Frisch and Giulianelli [2024], Huang et al. [2024c]
-Values Kovač et al. [2023], Zhong et al. [2024] Ren et al. [2024], Yao et al. [2025b] Chiu et al. [2025], Shen et al. [2024]
-Morality Ji et al. [2024], Abdulhai et al. [2024] Sachdeva and van Nuenen [2025], Neuman et al. [2025] Chiu et al. [2025], Nunes et al. [2024]
-Attitudes & opinions Bernardelle et al. [2024], Argyle et al. [2023] Röttger et al. [2024], Wright et al. [2024] -
-Heuristics and biases Hagendorff et al. [2023], Yax et al. [2024] Healey et al. [2024], Wen et al. [2024a] Bai et al. [2024], Xie et al. [2024b]
-Social interaction Sabour et al. [2024], Mittelstädt et al. [2024] Welivita and Pu [2024], Elyoseph et al. [2023] Zhou et al. [2024b], Wang et al. [2024d]
-Psychology of language Duan et al. [2024a], Amouyal et al. [2024] - -
-Learning and cognitive capabilities Wu et al. [2025], Coda-Forno et al. [2024] Galatzer-Levy et al. [2024] Lv et al. [2024]
+Personality traits jiang2023evaluating_and_inducing, serapio2023personality_traits_in_llms jiang2024personallm, zheng2025lmlpa frisch2024llm_agents_in_interaction, huang2024designing
+Values kovac2023llms_as_superpositions, zhong2024cultural ren2024valuebench, yao2025clave chiu2025dailydilemmas, shen2024valuecompass
+Morality ji2024moralbench, abdulhai2024moral sachdeva2025normative, neuman2025analyzing chiu2025dailydilemmas, nunes2024large
+Attitudes & opinions bernardelle2024mapping, argyle2023_out_of_one rottger2024political, wright2024revealing -
+Heuristics and biases hagendorff2023reasoning_bias, yax2024studying healey2024evaluating, wen2024evaluating bai2024fairmonitor, xie2024mindscope
+Social interaction sabour2024emobench, si1_mittelstadt2024large welivita2024large, elyoseph2023chatgpt si3_zhou2023sotopia, si4_wang2024sotopia
+Psychology of language pl1_duan2024hlb, pl24_amouyal2024large - -
+Learning and cognitive capabilities l8_wu2025understanding, coda2024cogbench l1_galatzer2024cognitive lv2024coggpt
 Test formats of LLM psychometric evaluation can be categorized into structured tests, open-ended conversations, and agentic simulations.
-Table 5 collects exemplary works illustrating the test formats used for each construct.
+Table˜4 collects exemplary works illustrating the test formats used for each construct.
 5.1.1 Structured tests
@@
 Unstructured testing often engages LLMs in open-ended, typically single-turn conversations that elicit specific constructs, mirroring real-world human-LLM interactions.
-For example, Jiang et al.
-[2024b] have LLMs generate long-form narratives to reveal personality traits.
-ValueBench uses advice-seeking queries to assess LLMs’ influence on users’ values [Ren et al., 2024], while Sachdeva and van Nuenen [2025] present moral dilemmas from Reddit, prompting LLMs to provide community-style moral judgments.
-To probe political attitudes, Röttger et al.
-[2024] convert PCT multiple-choice items into forced-choice and open-ended formats.
-For cognitive constructs, Healey et al.
-[2024] design pipelines to detect nuanced biases in free responses, and Welivita and Pu [2024] evaluate empathy by asking LLMs to respond to emotional scenarios.
-Many cognitive tests, such as WAIS-IV vocabulary and comprehension, also use open-ended questions [Galatzer-Levy et al., 2024].
+For example, jiang2024personallm have LLMs generate long-form narratives to reveal personality traits.
+ValueBench uses advice-seeking queries to assess LLMs’ influence on users’ values [ren2024valuebench], while sachdeva2025normative present moral dilemmas from Reddit, prompting LLMs to provide community-style moral judgments.
+To probe political attitudes, rottger2024political convert PCT multiple-choice items into forced-choice and open-ended formats.
+For cognitive constructs, healey2024evaluating design pipelines to detect nuanced biases in free responses, and welivita2024large evaluate empathy by asking LLMs to respond to emotional scenarios.
+Many cognitive tests, such as WAIS-IV vocabulary and comprehension, also use open-ended questions [l1_galatzer2024cognitive].
 Agentic simulations.
 Advanced unstructured tests evaluate LLMs as agents in complex, contextualized role-playing scenarios, analyzing their decision-making in dynamic environments.
-Huang et al.
-[2024c] design LLM-Agents with distinct personalities and analyze them by replicating human-like behaviors in agentic simulations.
-Shen et al.
-[2024] measure values using four real-world vignettes: collaborative writing, education, public sectors, and healthcare.
-Chiu et al.
-[2025] present value and moral dilemmas to LLMs, examining their decisions and rationales.
-Cognitive construct tests are often more complex, extending to multi-agent settings; for example, Xie et al.
-[2024b], Bai et al.
-[2024] identify cognitive biases in multi-agent communication.
-SOTOPIA simulates open-ended social interactions among LLM agents to evaluate social intelligence [Zhou et al., 2024b, Wang et al., 2024d].
-Lv et al.
-[2024] benchmark cognitive dynamics by having LLMs repeatedly complete cognitive questionnaires and reason over information flows.
+huang2024designing design LLM-Agents with distinct personalities and analyze them by replicating human-like behaviors in agentic simulations.
+shen2024valuecompass measure values using four real-world vignettes: collaborative writing, education, public sectors, and healthcare.
+chiu2025dailydilemmas present value and moral dilemmas to LLMs, examining their decisions and rationales.
+Cognitive construct tests are often more complex, extending to multi-agent settings; for example, xie2024mindscope, bai2024fairmonitor identify cognitive biases in multi-agent communication.
+SOTOPIA simulates open-ended social interactions among LLM agents to evaluate social intelligence [si3_zhou2023sotopia, si4_wang2024sotopia].
+lv2024coggpt benchmark cognitive dynamics by having LLMs repeatedly complete cognitive questionnaires and reason over information flows.
 Unstructured tests assess LLMs in free-form, contextualized scenarios, offering greater ecological validity and revealing complex behaviors—such as nuanced reasoning, subtle biases, and dynamic social interactions—not captured by structured formats.
@@
 Construct Established inventories Custom-curated items Synthetic items
-Personality traits Serapio-García et al. [2023], Barua et al. [2024] Ai et al. [2024], Jiang et al. [2023] Zeng [2024]
-Values Fischer et al. [2023], Miotto et al. [2022] Shen et al. [2024], Meadows et al. [2024] Ye et al. [2025a], Moore et al. [2024]
-Morality Abdulhai et al. [2024], Khandelwal et al. [2024] Hendrycks et al. [2021], Jin et al. [2022] Scherrer et al. [2023], Liu et al. [2024d]
-Attitudes & opinions Argyle et al. [2023], Bernardelle et al. [2024] Durmus et al. [2023], Ceron et al. [2024] Wan and Chang [2024]
-Heuristics and biases Hagendorff et al. [2023], Binz and Schulz [2023] Momennejad et al. [2023], Ando et al. [2023] Malberg et al. [2024], Xie et al. [2024b]
-Social interaction [Ullman, 2023, Chen et al., 2024d] Shapira et al. [2023a], Street et al. [2024] Gandhi et al. [2023]
-Psychology of language Cai et al. [2024], Duan et al. [2024a] He et al. [2024], Pérez-Mayos et al. [2021] -
-Learning and cognitive capabilities Coda-Forno et al. [2024], Dayan et al. [2024] Song et al. [2024a], Wu et al. [2025] -
+Personality traits serapio2023personality_traits_in_llms, barua2024on_the_psychology_of_gpt4 ai2024is_self_knowledge, jiang2023evaluating_and_inducing zeng2024quantifying
+Values fischer2023does, miotto2022who_is_gpt3 shen2024valuecompass, meadows2024localvaluebench ye2025gpv, moore2024large
+Morality abdulhai2024moral, khandelwal2024moral hendrycks2021aligning, jin2022make scherrer2023evaluating, liu2024evaluating
+Attitudes & opinions argyle2023_out_of_one, bernardelle2024mapping durmus2023towards, ceron2024beyond wan2024white
+Heuristics and biases hagendorff2023reasoning_bias, binz2023using momennejad2023evaluating, ando2023evaluating malberg2024comprehensive_cognitive_biases, xie2024mindscope
+Social interaction [tom1_ullman2023large, tom14_chen2023tobench] tom4_shapira2023clever, tom9_street2024llms tom5_gandhi2023understanding
+Psychology of language cai2024large, pl1_duan2024hlb pl9_he2024large, pl11_perez2021assessing -
+Learning and cognitive capabilities coda2024cogbench, l3_dayan2024age l11_song2024m3gia, l8_wu2025understanding -
 Data and task sources LLM Psychometrics can be 1) drawn from established psychometric inventories, 2) human-authored and custom-curated, and 3) synthesized by AI models.
-Table 6 provides an overview of the data and task sources for each construct.
+Table˜5 provides an overview of the data and task sources for each construct.
 Established inventories.
 Established psychometric inventories offer standardized, well-validated tools for LLM evaluation.
-Common examples include BFI and HEXACO for personality [Serapio-García et al., 2023, Barua et al., 2024]; SVS, PVQ, WVS, and VSM for values [Fischer et al., 2023, Miotto et al., 2022]; MFT and DIT for morality [Abdulhai et al., 2024, Khandelwal et al., 2024]; ANES and PCT for political attitudes [Argyle et al., 2023, Bernardelle et al., 2024]; CRT for heuristics and biases [Hagendorff et al., 2023, Binz and Schulz, 2023]; and False-Belief Tasks for ToM [Ullman, 2023, Chen et al., 2024d].
-While these inventories enable straightforward, standardized assessment, their use with LLMs raises concerns about data contamination, response bias, and limited ecological validity for complex AI tasks [Ye et al., 2025a].
+Common examples include BFI and HEXACO for personality [serapio2023personality_traits_in_llms, barua2024on_the_psychology_of_gpt4]; SVS, PVQ, WVS, and VSM for values [fischer2023does, miotto2022who_is_gpt3]; MFT and DIT for morality [abdulhai2024moral, khandelwal2024moral]; ANES and PCT for political attitudes [argyle2023_out_of_one, bernardelle2024mapping]; CRT for heuristics and biases [hagendorff2023reasoning_bias, binz2023using]; and False-Belief Tasks for ToM [tom1_ullman2023large, tom14_chen2023tobench].
+While these inventories enable straightforward, standardized assessment, their use with LLMs raises concerns about data contamination, response bias, and limited ecological validity for complex AI tasks [ye2025gpv].
 Custom-curated items.
 Human-authored, custom-curated items offer tailored psychometric tests that are often more relevant to LLMs, enabling exploration of novel capability dimensions.
-For example, Shen et al.
-[2024] find SVS and PVQ insufficient for LLM value alignment, so they create 11 AI-informed value statements based on a systematic review of alignment literature.
-Ceron et al.
-[2024] highlight the unreliability of standard political questionnaires for LLMs and introduce tests using annotated voting-advice questionnaires from seven EU countries to assess stance consistency.
-For morality, Hendrycks et al.
-[2021] curate a dataset of over 130,000 open-world scenarios requiring moral judgments.
-In ToM, Soubki et al.
-[2024] develop a dataset from real spoken dialogues to address gaps between synthetic benchmarks and human behavior.
+For example, shen2024valuecompass find SVS and PVQ insufficient for LLM value alignment, so they create 11 AI-informed value statements based on a systematic review of alignment literature.
+ceron2024beyond highlight the unreliability of standard political questionnaires for LLMs and introduce tests using annotated voting-advice questionnaires from seven EU countries to assess stance consistency.
+For morality, hendrycks2021aligning curate a dataset of over 130,000 open-world scenarios requiring moral judgments.
+In ToM, tom19_soubki2024views develop a dataset from real spoken dialogues to address gaps between synthetic benchmarks and human behavior.
 While custom-curated items offer high ecological validity, their development and validation are labor-intensive, limiting scalability and diversity.
@@
 Some studies adapt established inventories to improve ecological validity, reduce data contamination, and enhance scalability.
-For instance, Ren et al.
-[2024] convert self-report items into advice-seeking queries to better reflect real-world human-AI interactions.
-Bhandari et al.
-[2025b] rewrite established items into semantically equivalent forms, validating them via semantic embedding similarity thresholds.
-Zhu et al.
-[2024a] diversify cognitive test items to enable more comprehensive analysis of LLM cognitive abilities.
+For instance, ren2024valuebench convert self-report items into advice-seeking queries to better reflect real-world human-AI interactions.
+bhandari2025evaluating rewrite established items into semantically equivalent forms, validating them via semantic embedding similarity thresholds.
+zhu2024dynamic diversify cognitive test items to enable more comprehensive analysis of LLM cognitive abilities.
 Other work generates synthetic tests entirely from scratch.
-Hadar-Shoval et al.
-[2025] demonstrate that LLM-generated cognitive tests correlate strongly with human performance on established tests.
-Ye et al.
-[2025a] create value-eliciting prompts and confirm the reliability and validity of the resulting measurements.
-Jiang et al.
-[2024a] introduce generative evolving testing, where an LLM-based Item Generator produces items of specified difficulty.
-Chiu et al.
-[2025] use GPT-4 to generate moral dilemmas for evaluating LLM moral judgment.
-For social intelligence, agentic simulations prompt LLMs to generate components of social interactions [Zhou et al., 2024b, Mou et al., 2024].
+hadar2025feasibility demonstrate that LLM-generated cognitive tests correlate strongly with human performance on established tests.
+ye2025gpv create value-eliciting prompts and confirm the reliability and validity of the resulting measurements.
+jiang2024raising introduce generative evolving testing, where an LLM-based Item Generator produces items of specified difficulty.
+chiu2025dailydilemmas use GPT-4 to generate moral dilemmas for evaluating LLM moral judgment.
+For social intelligence, agentic simulations prompt LLMs to generate components of social interactions [si3_zhou2023sotopia, si8_mou2024agentsense].
 5.3 Prompting strategies
@@
 These prompts enable the generation of multiple simulated participants from a single LLM for statistical analysis.
-For example, Serapio-García et al.
-[2023] prepend persona instructions to each item to elicit diverse responses, allowing examining reliability and validity within one LLM.
-Similarly, Ye et al.
-[2025b] use value-anchoring prompts to create hundreds of LLM-based participants for value measurement, supporting the development of an LLM-specific value system.
-LLMs can represent a superposition of personalities, values, and perspectives [Kovač et al., 2023], prompting research into their adaptability.
-Role-playing prompts show that LLMs can shift between personality types, though steerability varies by model and dimension [Lu et al., 2023, Jiang et al., 2023, La Cava and Tagarelli, 2024, Li et al., 2024d].
-Such prompts are also used to test the stability of LLM values [Kovač et al., 2024, Rozen et al., 2024] and to measure how well LLMs can mimic moral or political profiles and their susceptibility to related biases [Münker, 2024, Simmons, 2023, Wright et al., 2024].
-Cognitive benchmarks, especially for social interactions, often use persona prompts in agentic simulations [Zhou et al., 2024b, Mou et al., 2024, Huang et al., 2023b], with task performance depending on the LLM’s ability to embody these roles.
-These prompts influence social-cognitive reasoning when controlling for other variables [Tan et al., 2024].
+For example, serapio2023personality_traits_in_llms prepend persona instructions to each item to elicit diverse responses, allowing examining reliability and validity within one LLM.
+Similarly, ye2025generative_psycho_lexical use value-anchoring prompts to create hundreds of LLM-based participants for value measurement, supporting the development of an LLM-specific value system.
+LLMs can represent a superposition of personalities, values, and perspectives [kovac2023llms_as_superpositions], prompting research into their adaptability.
+Role-playing prompts show that LLMs can shift between personality types, though steerability varies by model and dimension [lu2023illuminating, jiang2023evaluating_and_inducing, lacava2024open_models_closed_minds].
+Such prompts are also used to test the stability of LLM values [kovavc2024stick, rozen2024llms] and to measure how well LLMs can mimic moral or political profiles and their susceptibility to related biases [munker2024towards, simmons2023moral, wright2024revealing].
+Cognitive benchmarks, especially for social interactions, often use persona prompts in agentic simulations [si3_zhou2023sotopia, si8_mou2024agentsense, huang2023emotionbench], with task performance depending on the LLM’s ability to embody these roles.
+These prompts influence social-cognitive reasoning when controlling for other variables [tan2024phantom].
 While role-playing capabilities enable diverse perspectives useful for social simulations, they can also lead to unstable values and perspectives, potentially compromising test validity and complicating value alignment.
 Performance-enhancing prompts.
-Performance-enhancing prompts are crafted to augment LLMs in psychometric evaluations, as recommended by Hagendorff [2023].
-Chain of Thought (CoT) prompting [Wei et al., 2022] reduces biases [Hagendorff et al., 2023], enhances social intelligence [Shapira et al., 2023a], and improves cognitive capabilities [Coda-Forno et al., 2024].
-Its variant Emotional CoT [Li et al., 2024e] boosts EI, while emotional prompts [Li et al., 2023a] enhance general cognitive abilities through emotional stimuli.
-Few-shot prompting [Brown et al., 2020] has demonstrated improvements in ToM performance [Moghaddam and Honey, 2023].
+Performance-enhancing prompts are crafted to augment LLMs in psychometric evaluations, as recommended by hagendorff2023machine.
+Chain of Thought (CoT) prompting [wei2022chain] reduces biases [hagendorff2023reasoning_bias], enhances social intelligence [tom4_shapira2023clever], and improves cognitive capabilities [coda2024cogbench].
+Its variant Emotional CoT [li2024enhancing] boosts EI, while emotional prompts [li2023emotion_prompt] enhance general cognitive abilities through emotional stimuli.
+Few-shot prompting [brown2020language] has demonstrated improvements in ToM performance [tom8_rahimi2023boosting].
 Research has also developed construct-specific prompting strategies.
-Zhou et al.
-[2024a] propose MFT-guided reasoning prompts for moral alignment, while Echterhoff et al.
-[2024], Sumita et al.
-[2024] focus on self-debiasing approaches.
-ToM-specific enhancements include SymbolicToM [Sclar et al., 2023] and temporal decomposition-based prompting [Sarangi et al., 2025, Hou et al., 2024].
-Additionally, Zhao et al.
-[2025a] develop a framework using self-reflection prompts to examine both explicit and implicit social biases, where explicit bias measurement reflects implicit bias.
+zhou2024rethinking propose MFT-guided reasoning prompts for moral alignment, while echterhoff2024cognitive, sumita2024cognitive focus on self-debiasing approaches.
+ToM-specific enhancements include SymbolicToM [tom10_sclar2023minding] and temporal decomposition-based prompting [tom26_sarangi2025decompose, tom16_hou2024timetom].
+Additionally, zhao2025explicit develop a framework using self-reflection prompts to examine both explicit and implicit social biases, where explicit bias measurement reflects implicit bias.
 Prompt perturbation and adversarial attacks.
 Prompt perturbations can test the robustness of LLMs’ personality traits and cognitive abilities.
-Researchers have investigated whether LLMs can maintain stable personality, values, opinions, and cognitive abilities when item options are reordered [Schelb et al., 2025, Lee et al., 2024d], prompts are rephrased [Lee et al., 2024d, Fraser et al., 2022, Strachan et al., 2024a], prompt formats are altered [Moore et al., 2024, Schelb et al., 2025], or different languages are used [Moore et al., 2024, Cahyawijaya et al., 2024].
-For instance, Faulborn et al.
-[2025] introduce 30 variations of prompts to evaluate the political biases of LLMs.
-In addition, Wen et al.
-[2024a] propose psychometric-inspired adversarial attacks to uncover implicit biases in LLMs.
-Similarly, Li et al.
-[2024d] utilize persuasive adversarial prompts (PAP) [Zeng et al., 2024a] to test the robustness of LLMs’ values.
-Since LLMs are sensitive to these perturbations, researchers have begun to scrutinize the reliability of the test results obtained under standard conditions [Ye et al., 2025a, Röttger et al., 2024, Dominguez-Olmedo et al., 2024, Strachan et al., 2024a].
+Researchers have investigated whether LLMs can maintain stable personality, values, opinions, and cognitive abilities when item options are reordered [schelb2025ru, lee2024do_llms_have_distinct], prompts are rephrased [lee2024do_llms_have_distinct, fraser2022does, tom6_strachan2024testing], prompt formats are altered [moore2024large, schelb2025ru], or different languages are used [moore2024large, cahyawijaya2024high].
+For instance, faulborn2025only introduce 30 variations of prompts to evaluate the political biases of LLMs.
+In addition, wen2024evaluating propose psychometric-inspired adversarial attacks to uncover implicit biases in LLMs.
+Similarly, li2024quantifying_ai_psychology utilize persuasive adversarial prompts (PAP) [zeng2024johnny] to test the robustness of LLMs’ values.
+Since LLMs are sensitive to these perturbations, researchers have begun to scrutinize the reliability of the test results obtained under standard conditions [ye2025gpv, rottger2024political, dominguez2024questioning, tom6_strachan2024testing].
 5.4 Model output and scoring
@@
 LLM performance on structured tests is typically evaluated using logit-based probabilistic analysis and closed-ended output scoring.
-In multiple-choice or Likert-scale settings, some studies analyze token-level logits—especially the first token—to infer latent traits, detect response entropy, or compare LLM output distributions against human data [Pellert et al., 2024, Santurkar et al., 2023, Dominguez-Olmedo et al., 2024, Arehalli et al., 2022].
-Psycholinguistic tests often use logit-based metrics such as surprisal (the negative log probability of a token), which quantifies predictability and cognitive effort, allowing researchers to align LLM uncertainty patterns with those of humans [Wang et al., 2024b, Steuer et al., 2023].
+In multiple-choice or Likert-scale settings, some studies analyze token-level logits—especially the first token—to infer latent traits, detect response entropy, or compare LLM output distributions against human data [pellert2024ai_psychometrics, santurkar2023whose, dominguez2024questioning, pl6_arehalli2022syntactic].
+Psycholinguistic tests often use logit-based metrics such as surprisal (the negative log probability of a token), which quantifies predictability and cognitive effort, allowing researchers to align LLM uncertainty patterns with those of humans [pl3_wang2024how, pl10_steuer2023large].
 Conversely, closed-ended outputs, which are explicit numerical scores or categorical selections, can be analyzed using predefined scoring protocols.
-For Likert-scale responses, scores are typically averaged or aggregated based on established rubrics (e.g., VSM13 [Ye et al., 2025a] and DIT [Khandelwal et al., 2024]).
-In standardized cognitive tests, model outputs are evaluated against ground-truth labels (e.g., accuracy in arithmetic tasks) or rule-based criteria (e.g., established standards in the Verbal Comprehension Index) [Galatzer-Levy et al., 2024].
+For Likert-scale responses, scores are typically averaged or aggregated based on established rubrics (e.g., VSM13 [ye2025gpv] and DIT [khandelwal2024moral]).
+In standardized cognitive tests, model outputs are evaluated against ground-truth labels (e.g., accuracy in arithmetic tasks) or rule-based criteria (e.g., established standards in the Verbal Comprehension Index) [l1_galatzer2024cognitive].
 5.4.2 Open-ended output and scoring
@@
 Rule-based scoring.
-Rule-based scoring predominantly relies on the lexical hypothesis [Allport and Odbert, 1936], analyzing the presence and frequency of specific keywords or phrases.
-For example, Jiang et al.
-[2024b] apply LIWC (Linguistic Inquiry and Word Count) features [Pennebaker et al., 2001] to evaluate LLM personalities, and Fischer et al.
-[2023] employ a theory-driven value dictionary [Ponizovskiy et al., 2020] for value measurement.
-However, such lexicon-based approaches struggle to capture semantic nuance [Ye et al., 2025a].
-Some rule-based methods go beyond keyword matching; for instance, Healey et al.
-[2024] detect bias by checking for deviations from equivalent treatment in responses.
+Rule-based scoring predominantly relies on the lexical hypothesis [allport1936trait], analyzing the presence and frequency of specific keywords or phrases.
+For example, jiang2024personallm apply LIWC (Linguistic Inquiry and Word Count) features [pennebaker2001linguistic] to evaluate LLM personalities, and fischer2023does employ a theory-driven value dictionary [ponizovskiy2020development] for value measurement.
+However, such lexicon-based approaches struggle to capture semantic nuance [ye2025gpv].
+Some rule-based methods go beyond keyword matching; for instance, healey2024evaluating detect bias by checking for deviations from equivalent treatment in responses.
 Model-based scoring.
 Model-based scoring is prevalent in unstructured testing for its flexibility and scalability.
-Many studies train models to evaluate LLM responses, such as fine-tuning BERT variants on the MyPersonality dataset for personality scoring [Hilliard et al., 2024].
-Others fine-tune LLMs with psychometric inventories or human annotations to classify value valence [Ye et al., 2025a, Yao et al., 2024, 2025b, Sorensen et al., 2024a].
-Scoring can be conducted at the item level (e.g., Generative Psychometrics: item parsing, scoring, aggregation [Ye et al., 2025a]) or at the response level, evaluating the entire output [Yao et al., 2024, 2025b].
-Some models are tailored to specific frameworks (e.g., Schwartz’s Value Theory [Yao et al., 2024]), while others are trained to be generalists by leveraging broader fine-tuning data and LLM prior knowledge [Ye et al., 2025a].
+Many studies train models to evaluate LLM responses, such as fine-tuning BERT variants on the MyPersonality dataset for personality scoring [hilliard2024eliciting_personality].
+Others fine-tune LLMs with psychometric inventories or human annotations to classify value valence [ye2025gpv, yao2024value_fulcra, yao2025clave, sorensen2024value].
+Scoring can be conducted at the item level (e.g., Generative Psychometrics: item parsing, scoring, aggregation [ye2025gpv]) or at the response level, evaluating the entire output [yao2024value_fulcra, yao2025clave].
+Some models are tailored to specific frameworks (e.g., Schwartz’s Value Theory [yao2024value_fulcra]), while others are trained to be generalists by leveraging broader fine-tuning data and LLM prior knowledge [ye2025gpv].
 Model-based evaluators are especially common for personality and value assessment, where textual cues map readily to target dimensions.
-Many studies now use the LLM-as-a-judge approach [Gu et al., 2024] for scoring open-ended responses.
-For instance, Li et al.
-[2024d], Zheng et al.
-[2025] prompt LLMs to rate responses on personality, ToM, and motivation, validating consistency with other LLMs or human raters.
-This method is especially useful in unstructured simulation settings, such as SI measurement across goal completion, relationship maintenance, and social rule adherence [Zhou et al., 2024b, Wang et al., 2024d, a, Mou et al., 2024].
-Embedding models are also used to measure similarity between LLM outputs and prototypical examples [Huang et al., 2024c, Cahyawijaya et al., 2024, Amirizaniani et al., 2024].
+Many studies now use the LLM-as-a-judge approach [gu2024llm-as-a-judge] for scoring open-ended responses.
+For instance, li2024quantifying_ai_psychology, zheng2025lmlpa prompt LLMs to rate responses on personality, ToM, and motivation, validating consistency with other LLMs or human raters.
+This method is especially useful in unstructured simulation settings, such as SI measurement across goal completion, relationship maintenance, and social rule adherence [si3_zhou2023sotopia, si4_wang2024sotopia, si6_wang2024towards, si8_mou2024agentsense].
+Embedding models are also used to measure similarity between LLM outputs and prototypical examples [huang2024designing, cahyawijaya2024high, tom17_amirizaniani2024llms].
 Human scoring.
 Human scoring is employed when rigorous evaluation, adhering strictly to standard psychometric manuals, is necessary.
-For instance, Elyoseph et al.
-[2023] engage psychologists to assess the contextual suitability of LLM responses using the Levels of Emotional Awareness Scale (LEAS).
-Castello et al.
-[2024] conduct an examination of cognitive biases in LLMs through human evaluation and linguistic comparison.
-Healey et al.
-[2024] present a semi-automated pipeline to classify LLM responses into nuanced bias types.
-Other examples include ToM evaluation in open-ended responses [Amirizaniani et al., 2024] and cognitive evaluation based on the Montreal Cognitive Assessment (MoCA) [Dayan et al., 2024].
+For instance, elyoseph2023chatgpt engage psychologists to assess the contextual suitability of LLM responses using the Levels of Emotional Awareness Scale (LEAS).
+castello2024examining conduct an examination of cognitive biases in LLMs through human evaluation and linguistic comparison.
+healey2024evaluating present a semi-automated pipeline to classify LLM responses into nuanced bias types.
+Other examples include ToM evaluation in open-ended responses [tom17_amirizaniani2024llms] and cognitive evaluation based on the Montreal Cognitive Assessment (MoCA) [l3_dayan2024age].
 5.5 Inference parameters
@@
 As an emerging field, LLM Psychometrics lacks standardization in test design and administration, and recent work is beginning to address these issues.
-We outline the landscape of psychometric validation in Fig.
-4.
+We outline the landscape of psychometric validation in Fig.˜4.
+Model fairness is also paramount; however, given the extensive literature on AI fairness, we refer readers to dedicated surveys [mehrabi2021survey, gallegos2024bias_fairness_survey].
 6.1 Reliability and consistency
 Reliability, a core aspect of psychometric validation, measures how consistently a test performs—over time (test-retest), across versions (parallel forms), and among evaluators (inter-rater).
-Researchers have adapted these metrics for LLMs.
-For example, Li et al.
-[2024d] propose a benchmark covering five reliability forms: internal consistency, parallel forms, inter-rater, option position robustness, and adversarial attack robustness.
-Other studies [Zheng et al., 2025, Huang et al., 2023a, Ceron et al., 2024, Shu et al., 2024] examine LLM output consistency across repeated trials, prompt variations, and languages.
-Notably, some metrics capture both random and systematic errors, so inconsistencies can undermine both reliability and validity.
-Some advanced LLMs show stable Big Five personality trait responses across settings [Huang et al., 2023a], with larger and instruction-tuned models achieving even higher reliability using role-playing prompts [Serapio-García et al., 2023].
-LLMs also display consistent value orientations across paraphrased, related, and translated questions, as well as different response formats [Moore et al., 2024].
-Strong inter-rater reliability is reported when LLMs act as judges, though personality stability and value consistency vary by model and cultural context [Li et al., 2024d].
+However, applying these concepts to LLMs requires a fundamental shift in perspective compared to human psychology.
+In human psychometrics, the subject (the person) is assumed to have relatively stable traits, and reliability metrics (e.g., Cronbach’s $\alpha$) primarily evaluate the quality of the instrument.
+In contrast, LLM Psychometrics faces a dual challenge: the instability of the instrument (due to flawed design) and the stochastic nature of the subject (due to models’ context sensitivity).
+Researchers adapt these metrics for LLMs to tackle the dual challenges.
+For example, li2024quantifying_ai_psychology propose a benchmark covering five reliability forms: internal consistency, parallel forms, inter-rater, option position robustness, and adversarial attack robustness.
+Other studies [zheng2025lmlpa, huang2023revisiting_the_reliability, ceron2024beyond, shu2024you_dont_need] examine LLM output consistency across repeated trials, prompt variations, and languages.
+When measuring reliability, many studies implicitly or explicitly adopt the "LLM as a population" perspective when applying reliability metrics [serapio2023personality_traits_in_llms].
+Unlike human psychometrics, where reliability usually assesses consistency across a population of distinct individuals, here reliability often verifies whether a single tool consistently measures a construct within a single LLM, which is treated as a population of diverse generated responses elicited by diverse prompts.
+The prompts can be diversified in terms of model profiling, item rephrasing, or both.
+If a model’s responses remain internally consistent across prompt variations (e.g., high Cronbach’s $\alpha$), it demonstrates that the test reliably captures the latent traits of that specific model-population.
+Other works treat the LLM as a single individual, focusing on the stability of its outputs for a specific test.
+When treating LLM as a single entity, reliability is assessed through metrics like parallel forms reliability, option position robustness, and adversarial attack robustness, which measure whether the model’s responses remain stable despite minor variations in the test items or presentation format.
+Some advanced LLMs show stable Big Five personality trait responses across settings [huang2023revisiting_the_reliability], with larger and instruction-tuned models achieving even higher reliability using role-playing prompts [serapio2023personality_traits_in_llms].
+LLMs also display consistent value orientations across paraphrased, related, and translated questions, as well as different response formats [moore2024large].
+Strong inter-rater reliability is reported when LLMs act as judges, though personality stability and value consistency vary by model and cultural context [li2024quantifying_ai_psychology].
 However, other studies report less favorable results.
-LLMs are notoriously sensitive to prompt variations: small changes in prompt form [Röttger et al., 2024, Ren et al., 2024], option order [Li et al., 2024d, Dominguez-Olmedo et al., 2024], or syntax [Hu and Levy, 2023] can systematically shift outputs, undermining parallel forms reliability.
-Studies also find inconsistencies between next-token logits and forced-choice responses in psycholinguistic tests [Hu and Levy, 2023], and between forced-choice and free-form responses in value [Kovač et al., 2024, Ye et al., 2025a] and political opinion surveys [Röttger et al., 2024].
-In cognitive tests, minor changes to ToM tasks often cause poor performance, indicating low parallel forms reliability and reliance on fragile heuristics [Ullman, 2023, Shapira et al., 2023a, Holterman and Deemter, 2023].
-The Words and Deeds Consistency Test (WDCT) [Xu et al., 2025b] further shows LLMs’ verbal and behavioral responses are inconsistent across domains.
+LLMs are notoriously sensitive to prompt variations: small changes in prompt form [rottger2024political, ren2024valuebench], option order [li2024quantifying_ai_psychology, dominguez2024questioning], or syntax [pl7_hu2023prompt] can systematically shift outputs, undermining parallel forms reliability.
+Studies also find inconsistencies between next-token logits and forced-choice responses in psycholinguistic tests [pl7_hu2023prompt], and between forced-choice and free-form responses in value [kovavc2024stick, ye2025gpv] and political opinion surveys [rottger2024political].
+In cognitive tests, minor changes to ToM tasks often cause poor performance, indicating low parallel forms reliability and reliance on fragile heuristics [tom1_ullman2023large, tom4_shapira2023clever, tom13_holterman2023does].
+The Words and Deeds Consistency Test (WDCT) [xu2025large] further shows LLMs’ verbal and behavioral responses are inconsistent across domains.
 Reliability findings are mixed and depend on factors such as construct, reliability type, test format, prompting, and model.
-LLMs are generally more reliable in personality assessments [Huang et al., 2023a, Serapio-García et al., 2023] than in political opinion evaluations [Dominguez-Olmedo et al., 2024, Röttger et al., 2024].
-Inter-rater agreement varies across instruments and domains [Bodroža et al., 2024].
-LLMs often show high internal consistency (Cronbach’s alpha $>0.8$) in closed-choice personality tests [Huang et al., 2023a, Serapio-García et al., 2023, Zheng et al., 2025], but low parallel form reliability due to prompt sensitivity [Li et al., 2024d, Shu et al., 2024, Gupta et al., 2024].
-Consistency improves in structured tests with strong prompt control [Rozen et al., 2024, Serapio-García et al., 2023, Klinkert et al., 2024], but declines with less standardized prompts [Petrov et al., 2024] or controversial topics [Moore et al., 2024].
-Reliability also varies by model: base models are more consistent than fine-tuned ones on value-laden questions [Moore et al., 2024], and higher model safety correlates with greater value consistency [Ye et al., 2025b].
-Advanced models also perform better in role-playing personality tests, increasing reliability [Serapio-García et al., 2023].
+LLMs are generally more reliable in personality assessments [huang2023revisiting_the_reliability, serapio2023personality_traits_in_llms] than in political opinion evaluations [dominguez2024questioning, rottger2024political].
+Inter-rater agreement varies across instruments and domains [bodroza2024personality_testing_of_llms].
+LLMs often show high internal consistency (Cronbach’s alpha >0.8>0.8) in closed-choice personality tests [huang2023revisiting_the_reliability, serapio2023personality_traits_in_llms, zheng2025lmlpa], but low parallel form reliability due to prompt sensitivity [li2024quantifying_ai_psychology, shu2024you_dont_need, gupta2024self_assessment_tests].
+Consistency improves in structured tests with strong prompt control [rozen2024llms, serapio2023personality_traits_in_llms, klinkert2024driving_generative_agents], but declines with less standardized prompts [petrov2024limited_ability_of_llms] or controversial topics [moore2024large].
+Reliability also varies by model: base models are more consistent than fine-tuned ones on value-laden questions [moore2024large], and higher model safety correlates with greater trait consistency [ye2025generative_psycho_lexical, tosato2025persistent].
+Advanced models also perform better in role-playing personality tests, increasing reliability [serapio2023personality_traits_in_llms].
 6.2 Validity
@@
 Key challenges include data contamination from established tests and insufficiently vetted novel items.
-Directly applying human tests to LLMs risks contamination, as LLMs may have seen test items or similar content during training [Hagendorff et al., 2024, Jiang et al., 2024c].
+Directly applying human tests to LLMs risks contamination, as LLMs may have seen test items or similar content during training [hagendorff2024machine, jiang2024investigating].
 This can inflate performance or bias trait measurement.
-To mitigate this, researchers reformat tests or generate new items; see § 5.2.
+To mitigate this, researchers reformat tests or generate new items; see §˜5.2.
 On the other hand, reformatted or newly generated items may inadequately capture the construct or introduce extraneous factors.
@@
 Construct validity determines if a test truly measures its intended construct.
-Alaa et al.
-[2025] highlight its critical yet often lacking role in medical LLM benchmarks.
-Broader literature suggests that construct validity is challenged by LLMs’ unique internal abstractions, systematic response patterns, and social desirability bias.
-Unique internal abstractions
-LLMs and humans may fundamentally differ in their internal representations of psychological constructs.
-Kovač et al.
-[2023] argue that LLMs display contextually adaptive traits, unlike the stable characteristics seen in humans.
-Sühr et al.
-[2023] show that LLM personality types do not align with the clear separation of Big Five traits found in humans.
-Similarly, Peereboom et al.
-[2024] find that human-oriented HEXACO tests are inapplicable to LLMs, as key personality factors may be absent.
-For values, Ye et al.
-[2025a] report that LLM self-reports do not fit Schwartz’s model, and while open-ended responses improve alignment, it remains an imperfect fit; Hadar-Shoval et al.
-[2024] also find LLMs encode values distinct from humans.
-Overall, researchers agree that robust LLM psychometric evaluation requires new operational definitions, distinct from or only loosely analogous to human traits.
-In this regard, pioneering research presents value systems tailored to LLMs’ unique internal abstractions [Ye et al., 2025b, Biedma et al., 2024].
-Ye et al.
-[2025b] introduce a generative psycho-lexical approach based on GPV value measurement [Ye et al., 2025a], demonstrating through various tasks that their value factors are valid and outperform Schwartz’s system.
-The distinct nature of LLM constructs requires rigorously designed tests: Ma et al.
-[2025] present the Core Sentiment Inventory (CSI) for measuring LLM emotional tendencies and personality; Fang et al.
-[2024] create a psychologically grounded mathematics benchmark to improve validity, model item difficulty, and enable human comparison; and Lee et al.
-[2024d] propose the TRAIT test, achieving higher validity and reliability than other personality measures.
+Literature suggests that construct validity is challenged by LLMs’ reasoning shortcuts, unique internal abstractions, systematic response patterns, and social desirability bias.
+Procedural validity.
+Procedural validity refers to the extent to which the internal processes producing a response align with the theoretical construct being measured.
+Standard psychometrics often relies on a functionalist assumption: correct answers imply underlying competence.
+However, it has been argued that for LLMs, high accuracy is a poor proxy for intelligence if the process of reasoning is flawed or based on “shortcuts” [mitchell2025artificial, stevenson2024can, lewis2024evaluatingrobustness].
+Research warns against the “anthropomorphic fallacy”, which assumes that human-like outputs evidence human-like cognition [Lin2025fallacies].
+Without verifying procedural validity, LLM Psychometrics risks conflating sophisticated statistical mimicry with genuine understanding.
+Pioneering works exemplified by beger2025ai evaluate LLMs’ abstract reasoning capabilities across modalities, analyzing both output accuracy and generated rules.
+The results indicate top reasoning models like o3 achieve human-level accuracy but frequently rely on unintended "shortcut" rules.
+han2025dual further delve into the internal mechanisms of LLMs and reveal that LLMs express values through two distinct pathways: intrinsic expression, reflecting values learned during training, and prompted expression, elicited by specific instructions.
+These mechanisms rely on partially distinct internal components (e.g., value vectors and neurons), demonstrating that similar outputs may arise from different processes.
+Further work must move beyond output accuracy to scrutinize the reasoning process itself.
+Construct equivalence.
+A foundational concern in LLM Psychometrics is construct equivalence: whether a psychological construct such as personality or values carries the same meaning when applied to LLMs as it does for humans.
+This problem is structurally analogous to the “imposed etic” issue in cross-cultural psychometrics, where instruments developed in one cultural context are applied to another without first verifying conceptual equivalence [berry1969cross_cultural].
+Transposing this logic to LLM Psychometrics: applying human personality inventories to LLMs presupposes that the constructs are equivalent across the human-LLM divide, which is an assumption that requires explicit empirical testing rather than being taken for granted [vandenberg2000review].
+When investigating the latent psychological structure of LLMs, researchers usually treat a single LLM or multiple LLMs with diverse prompts as a population.
+This approach ensures sufficient variance for factor analysis or other structural equation modeling techniques to identify underlying dimensions.
+Based on this statistical probing, recent studies find that LLMs and humans may fundamentally differ in their internal representations of psychological constructs.
+kovac2023llms_as_superpositions argue that LLMs display contextually adaptive traits, unlike the stable characteristics seen in humans.
+suhr2023challenging show that LLM personality types do not align with the clear separation of Big Five traits found in humans.
+Similarly, peereboom2024cognitive find that human-oriented HEXACO tests are inapplicable to LLMs, as key personality factors may be absent.
+For values, ye2025gpv report that LLM self-reports do not fit Schwartz’s model, and while open-ended responses improve alignment, it remains an imperfect fit; hadar2024assessing also find LLMs encode values distinct from humans.
+Taken together, these findings constitute evidence of construct non-equivalence: the same instrument does not measure the same latent property across the human-LLM divide.
+Researchers therefore agree that robust LLM psychometric evaluation requires new operational definitions, distinct from or only loosely analogous to human traits.
+To address this, the field is moving toward AI-native psychometrics.
+Pioneering research presents value systems tailored to LLMs’ unique latent abstractions [ye2025generative_psycho_lexical, biedma2024beyond_human_norms].
+ye2025generative_psycho_lexical introduce a generative psycho-lexical approach based on GPV value measurement [ye2025gpv], demonstrating through various tasks that their value factors are valid and outperform Schwartz’s system.
+The distinct nature of LLM constructs requires rigorously designed tests: ma2025leveraging present the Core Sentiment Inventory (CSI) for measuring LLM emotional tendencies and personality; fang2024patch create a psychologically grounded mathematics benchmark to improve validity, model item difficulty, and enable human comparison; and lee2024do_llms_have_distinct propose the TRAIT test, achieving higher validity and reliability than other personality measures.
 Response set.
 Response set refers to systematic answer patterns regardless of item content.
-Some LLMs show option position bias, favoring responses labeled "A" in political opinion tests [Dominguez-Olmedo et al., 2024, Li et al., 2024d].
-After correcting for this bias, Dominguez-Olmedo et al.
-[2024] find LLM responses to political questions are uniformly random, rendering demographic alignment analyses meaningless.
-Similar findings are observed in LLM personality [Song et al., 2023, Sühr et al., 2023] and value tests [Ye et al., 2025a].
-Sühr et al.
-[2023] also note LLMs often affirm both the original and reverse-coded items, while Ye et al.
-[2025a] observe some LLMs consistently prefer low or high ratings.
+Some LLMs show option position bias, favoring responses labeled "A" in political opinion tests [dominguez2024questioning, li2024quantifying_ai_psychology].
+After correcting for this bias, dominguez2024questioning find LLM responses to political questions are uniformly random, rendering demographic alignment analyses meaningless.
+Similar findings are observed in LLM personality [song2023have_large_language, suhr2023challenging] and value tests [ye2025gpv].
+suhr2023challenging also note LLMs often affirm both the original and reverse-coded items, while ye2025gpv observe some LLMs consistently prefer low or high ratings.
 These observations suggest self-report tests may be unsuitable for LLM Psychometrics.
@@
 Social desirability bias—the tendency to give socially acceptable rather than genuine responses—is observed in LLMs as in humans.
-Salecha et al.
-[2024] find LLMs skew toward favorable traits in personality tests, inflating extraversion and lowering neuroticism scores.
-Similarly, Ye et al.
-[2025b] show LLMs avoid reporting less desirable values (e.g., hedonism) in direct self-reports [Biedma et al., 2024], but reveal them in indirect, contextual prompts.
+salecha2024large find LLMs skew toward favorable traits in personality tests, inflating extraversion and lowering neuroticism scores.
+Similarly, ye2025generative_psycho_lexical show LLMs avoid reporting less desirable values (e.g., hedonism) in direct self-reports [biedma2024beyond_human_norms], but reveal them in indirect, contextual prompts.
 While tuning for social desirability is important, overemphasis risks producing homogeneous models that lack diverse perspectives and fail to meet varied user needs.
 Cross-lingual tests.
-Cross-lingual psychometric tests reveal notable inconsistencies in LLM measurement results across languages [Romero et al., 2024].
-Despite multilingual capabilities, LLMs often lack consistent psychological trait assessments between languages [Cahyawijaya et al., 2024].
+Cross-lingual psychometric tests reveal notable inconsistencies in LLM measurement results across languages [romero2024do_gpt_language_models].
+Despite multilingual capabilities, LLMs often lack consistent psychological trait assessments between languages [cahyawijaya2024high].
 These inconsistencies complicate validation, requiring careful separation of model biases, translation issues, and cultural differences represented by the models.
@@
 In LLM Psychometrics, external standards often coincide with real-world evaluation results.
-Studies find that value orientations from forced-choice tests frequently diverge from those observed in human-LLM interactions [Ren et al., 2024, Ye et al., 2025a], with similar gaps in personality [Ai et al., 2024] and morality [Nunes et al., 2024].
-As a result, researchers advocate linking LLM personality to safety [Zhang et al., 2024a], measuring values within real-world interactions [Ren et al., 2024], and using open-ended dialog to improve construct validity [Ye et al., 2025a].
-Further work connects LLM value measurement to safety prediction and value alignment tasks to ensure conformity with external standards [Ye et al., 2025b].
+Studies find that value orientations from forced-choice tests frequently diverge from those observed in human-LLM interactions [ren2024valuebench, ye2025gpv], with similar gaps in personality [ai2024is_self_knowledge] and morality [nunes2024large].
+As a result, researchers advocate linking LLM personality to safety [zhang2024the_better_angels], measuring values within real-world interactions [ren2024valuebench], and using open-ended dialog to improve construct validity [ye2025gpv].
+Further work connects LLM value measurement to safety prediction and value alignment tasks to ensure conformity with external standards [ye2025generative_psycho_lexical].
 6.3 Standards and recommendations
 To address key challenges, researchers have proposed standards and recommendations to guide LLM Psychometrics and establish methodological rigor.
-Frank [2023a] suggests leveraging developmental psychology to understand LLM cognition, recommending simplified and novel stimuli to prevent training data contamination.
-Löhn et al.
-[2024] critique current practices and outline seven requirements for valid assessment: reliability, validity, suitability to LLM capabilities, non-disclosure (to avoid contamination), and fairness—including valid comparisons, translations, and transparency.
+frank2023baby suggests leveraging developmental psychology to understand LLM cognition, recommending simplified and novel stimuli to prevent training data contamination.
+lohn2024machine critique current practices and outline seven requirements for valid assessment: reliability, validity, suitability to LLM capabilities, non-disclosure (to avoid contamination), and fairness—including valid comparisons, translations, and transparency.
 Their review of 25 studies reveals these principles are often overlooked.
-More recently, Hagendorff et al.
-[2024] advocate for procedural test generation, multiple task versions, performance-enhancing prompts, shuffling options, and diverse scoring methods to reduce contamination and improve reliability.
+More recently, hagendorff2024machine advocate for procedural test generation, multiple task versions, performance-enhancing prompts, shuffling options, and diverse scoring methods to reduce contamination and improve reliability.
 They also recommend deterministic settings for reproducibility, automated evaluation tools, manual review of unreliable outputs, and statistical analysis of results.
-Vaugrante et al.
-[2024] highlight poor replicability in recent work and offer four main recommendations: (1) ensure benchmark validity, sufficient task coverage, standardization, prompt control, and alignment with research goals; (2) use standardized methods, avoid cherry-picking, maintain statistical transparency, document setups, and define consistent metrics; (3) monitor model changes, account for model diversity, adjust benchmark difficulty as models advance, and document model versions and experiment dates; (4) standardize scoring and verification, with clear guidelines and rubrics for implementers.
-Schelb et al.
-[2025] introduce a framework for robust, flexible, and reproducible psychometric experiments, centered on standardized configuration files.
+vaugrante2024looming highlight poor replicability in recent work and offer four main recommendations: (1) ensure benchmark validity, sufficient task coverage, standardization, prompt control, and alignment with research goals; (2) use standardized methods, avoid cherry-picking, maintain statistical transparency, document setups, and define consistent metrics; (3) monitor model changes, account for model diversity, adjust benchmark difficulty as models advance, and document model versions and experiment dates; (4) standardize scoring and verification, with clear guidelines and rubrics for implementers.
+schelb2025ru introduce a framework for robust, flexible, and reproducible psychometric experiments, centered on standardized configuration files.
 7 Psychometrics for LLM enhancement
 Psychometric principles not only enable rigorous evaluation but also drive LLM development and improvement.
-They underpin three main advancements: trait manipulation for personalized and controlled behaviors, psychometrically-informed safety and alignment, and cognitive enhancement for more human-like reasoning and communication.
+We focus specifically on methods in which psychometric constructs or measurements serve as a substantive foundation for enhancement: either as the specification of the enhancement target (e.g., validated dimensional scales defining what is to be improved and how success is measured), as a training or reward signal derived from psychometric assessments, or as a theoretical framework directly inspiring the architecture or prompt design.
+This criterion distinguishes the work reviewed here from general fine-tuning or prompting surveys.
+Under this scope, three main advancements emerge: trait manipulation for personalized and controlled behaviors, psychometrically-informed safety and alignment, and cognitive enhancement for more human-like reasoning and communication.
 7.1 Trait manipulation
 Psychometrics enables targeted manipulation of LLM traits for personalization, role-play, and demographic simulation, providing a principled framework across prompting, inference, and training.
-Structured prompts based on validated psychometric inventories reliably elicit and modulate specific personality traits in LLMs [Huang et al., 2024c, He and Zhang, 2024], supporting controlled simulation or alignment of synthetic personas [Serapio-García et al., 2023, Jiang et al., 2023, Zhang, 2024, La Cava and Tagarelli, 2024].
-For instance, Jiang et al.
-[2023] introduce personality prompting ($P^{2}$), where Big Five lexicons are expanded into detailed personality descriptions, enabling LLMs to adopt diverse behaviors.
-Chuang et al.
-[2024] propose a systematic prompting strategy using empirically-derived human belief networks, aligning LLMs with related beliefs across 64 topics and 9 latent factors.
+Structured prompts based on validated psychometric inventories reliably elicit and modulate specific personality traits in LLMs [huang2024designing, he2024afspp], supporting controlled simulation or alignment of synthetic personas [serapio2023personality_traits_in_llms, jiang2023evaluating_and_inducing, zhang2024harnessing, lacava2024open_models_closed_minds].
+For instance, jiang2023evaluating_and_inducing introduce personality prompting ($P^{2}$), where Big Five lexicons are expanded into detailed personality descriptions, enabling LLMs to adopt diverse behaviors.
+chuang2024beyond propose a systematic prompting strategy using empirically-derived human belief networks, aligning LLMs with related beliefs across 64 topics and 9 latent factors.
 Beyond prompt engineering, inference-time interventions manipulate hidden representations during the forward pass, often guided by patterns recognized using psychometric scales.
-Methods such as ControlLM [Weng et al., 2024], Personality Alignment Search [Zhu et al., 2024b], Neuron-based Intervention [Deng et al., 2024], Probing-then-Editing [Ju et al., 2025], and Latent Feature Steering [Yang et al., 2025b] achieve trait control by shifting activations or neuron values at inference, without retraining.
+Methods such as ControlLM [weng2024controllm], Personality Alignment Search [zhu2024personality], Neuron-based Intervention [deng2024neuron], Probing-then-Editing [ju2025probing], and Latent Feature Steering [yang2025exploring] achieve trait control by shifting activations or neuron values at inference, without retraining.
 Other approaches fine-tune LLMs for trait manipulation.
-Vu et al.
-[2024] modify model architectures to enable control over continuous Big Five and mental health traits.
-Dan et al.
-[2024], Jain et al.
-[2024], Li et al.
-[2024c] deploy LoRA modules or routing networks linked to personality traits, training specialist adapters for trait-driven generation.
-Cui et al.
-[2023], Zeng et al.
-[2024b], Liu et al.
-[2024a] embed traits by training on large, annotated dialogue datasets from validated psychometric tools.
-Some methods directly edit model parameters without retraining; for example, Hwang et al.
-[2025] use MBTI-based adjustment queries to edit model parameters and steer target personality traits.
-Trait manipulation in LLMs offers a promising alternative to human participants in social science research [Gao et al., 2024, Cui et al., 2025].
-Fine-tuning enables LLMs to more accurately simulate survey response distributions for specific subpopulations, even on unseen questions [Cao et al., 2025, Suh et al., 2025].
-He-Yueya et al.
-[2024] propose psychometric alignment, training LLMs on human response data to better match human understanding on novel items, though effectiveness varies by domain.
-Additionally, Kang et al.
-[2023], Sorensen et al.
-[2025] investigate human value representation in LLMs and use value injection to fine-tune models for simulating human opinions.
-While current personalization methods can induce more than superficial changes in model outputs, achieving robust, authentic, and context-independent personalization remains challenging [Kovač et al., 2024, Dominguez-Olmedo et al., 2024].
+vu2024psychadapter modify model architectures to enable control over continuous Big Five and mental health traits.
+dan2024p, jain2024text, li2024big5 deploy LoRA modules or routing networks linked to personality traits, training specialist adapters for trait-driven generation.
+cui2023machine_mindset_an_mbti, zeng2024persllm, liu2024dynamic_generation_of embed traits by training on large, annotated dialogue datasets from validated psychometric tools.
+Some methods directly edit model parameters without retraining; for example, hwang2025personality use MBTI-based adjustment queries to edit model parameters and steer target personality traits.
+Trait manipulation in LLMs offers a promising alternative to human participants in social science research [gao2024large, cui2025large].
+Fine-tuning enables LLMs to more accurately simulate survey response distributions for specific subpopulations, even on unseen questions [cao2025specializing, suh2025language].
+he2024psychometric propose psychometric alignment, training LLMs on human response data to better match human understanding on novel items, though effectiveness varies by domain.
+Additionally, kang2023values, sorensen2025value investigate human value representation in LLMs and use value injection to fine-tune models for simulating human opinions.
+While current personalization methods can induce more than superficial changes in model outputs, achieving robust, authentic, and context-independent personalization remains challenging [kovavc2024stick, dominguez2024questioning].
 7.2 Safety and alignment
 Recent work demonstrates that psychometric measurements of LLMs are closely linked to their safety and alignment, one of the most pressing issues in the field.
-Zhang et al.
-[2024a], Wang et al.
-[2025b] show that LLM personality traits, as measured by MBTI-M and HEXACO scales, correlate with model safety.
-For example, Zhang et al.
-[2024a] find that better-aligned LLMs exhibit higher Extraversion, Sensing, and Judging, and that modifying LLM personalities can improve safety, especially regarding privacy and fairness.
+zhang2024the_better_angels, wang2025exploring_the_impact show that LLM personality traits, as measured by MBTI-M and HEXACO scales, correlate with model safety.
+For example, zhang2024the_better_angels find that better-aligned LLMs exhibit higher Extraversion, Sensing, and Judging, and that modifying LLM personalities can improve safety, especially regarding privacy and fairness.
 Similarly, studies have connected LLM values to safety and alignment.
-Yao et al.
-[2024] distinguish safe from unsafe LLM responses by analyzing correlations with Schwartz’s values, while Ye et al.
-[2025a, b] accurately predict LLM safety scores from value orientations measured by the GPV tool.
-Both Yao et al.
-[2024], Ye et al.
-[2025b] use reinforcement learning to align LLMs with human values, improving safety.
-In morality, Huang et al.
-[2024a], Tlaie [2024] apply MFT-based prompting to enhance LLM moral reasoning and alignment.
-Takeshita et al.
-[2023], Ohashi et al.
-[2024] introduce the JCommonsenseMorality (JCM) dataset and fine-tune LLMs for local cultural adaptation.
+yao2024value_fulcra distinguish safe from unsafe LLM responses by analyzing correlations with Schwartz’s values, while ye2025gpv, ye2025generative_psycho_lexical accurately predict LLM safety scores from value orientations measured by the GPV tool.
+Both yao2024value_fulcra, ye2025generative_psycho_lexical use reinforcement learning to align LLMs with human values, improving safety.
+In morality, huangmoral, tlaie2024exploring apply MFT-based prompting to enhance LLM moral reasoning and alignment.
+takeshita2023jcommonsensemorality, ohashi2024extended introduce the JCommonsenseMorality (JCM) dataset and fine-tune LLMs for local cultural adaptation.
 7.3 Cognitive enhancement
@@
 Psychology-inspired prompts boost LLM cognitive abilities.
-For instance, Li et al.
-[2023a] show that emotional prompts enhance general cognition.
-Role-playing prompts improve ToM [Tan et al., 2024] and foster more human-like reasoning [Nighojkar et al., 2025].
-Zhao et al.
-[2025b] propose a finite-state machine based on Hill’s Helping Skills theory, embedding state transitions for emotional support into the model’s multi-hop inference.
+For instance, li2023emotion_prompt show that emotional prompts enhance general cognition.
+Role-playing prompts improve ToM [tan2024phantom] and foster more human-like reasoning [nighojkar2025giving].
+zhao2025fisminess propose a finite-state machine based on Hill’s Helping Skills theory, embedding state transitions for emotional support into the model’s multi-hop inference.
 This structures dialogues around psychological support and emotional states, improving human-rated effectiveness and strategic alignment.
-Liu et al.
-[2023] introduce a neural listener module into LLMs, jointly optimizing ToM reasoning during training and highlighting the value of psycholinguistic theories for improving LLM language acquisition.
+liu2023computational introduce a neural listener module into LLMs, jointly optimizing ToM reasoning during training and highlighting the value of psycholinguistic theories for improving LLM language acquisition.
 Preference-based and RL-based approaches further internalize empathy by leveraging cognitive and affective models to construct composite reward functions or preference signals.
-For example, Sotolar et al.
-[2024] utilize formal emotion theories (e.g., Plutchik’s wheel) to generate preference pairs for model optimization, while others employ empathy classifiers to shape RL rewards and promote empathetic outputs [Sharma et al., 2021].
-Preference optimization based on human judgments in open-ended social tasks also enhances LLM social-pragmatic reasoning [Wu et al., 2024].
+For example, sotolar2024empo utilize formal emotion theories (e.g., Plutchik’s wheel) to generate preference pairs for model optimization, while others employ empathy classifiers to shape RL rewards and promote empathetic outputs [sharma2021towards].
+Preference optimization based on human judgments in open-ended social tasks also enhances LLM social-pragmatic reasoning [wu2024rethinking].
 8 Trends, challenges, and future directions
@@
 8.1 Psychometric validation
-Psychometric validation is increasingly recognized as essential for evaluating LLM personality traits, with multiple studies addressing different aspects (§ 6).
+Psychometric validation is increasingly recognized as essential for evaluating LLM personality traits, with multiple studies addressing different aspects (§˜6).
 However, ability testing lags behind: most current assessments use structured tests (e.g., multiple-choice questions) adapted from human-focused instruments.
 These face notable limitations.
-For example, Ullman [2023] show that LLMs often fail on minor variations of ToM tasks, revealing poor parallel forms reliability.
-LLMs’ unique internal representations also undermine construct validity, resulting in flawed benchmarks [Riemer et al., 2025].
+For example, tom1_ullman2023large show that LLMs often fail on minor variations of ToM tasks, revealing poor parallel forms reliability.
+LLMs’ unique internal representations also undermine construct validity, resulting in flawed benchmarks [tom_riemer2024position_broken].
 Moreover, the generalizability of structured test results to real-world human-AI interactions is rarely examined, leaving criterion and ecological validity underexplored.
@@
 As a result, they may include extraneous factors or insufficiently capture the intended constructs.
-Notable construct-oriented efforts include Zhu et al.
-[2024a], who propose multifaceted evaluations of three cognitive factors from benchmarking data, and Zhou et al.
-[2025], who introduce a theory-driven, richer set of general scales.
+Notable construct-oriented efforts include zhu2024dynamic, who propose multifaceted evaluations of three cognitive factors from benchmarking data, and zhou2025general_scales, who introduce a theory-driven, richer set of general scales.
 However, defining core abilities and designing valid, comprehensive ability tests remains an open challenge.
@@
 Recent research is moving from applying human-based constructs to developing constructs tailored for LLMs.
-Studies such as Sühr et al.
-[2023], Peereboom et al.
-[2024], Ye et al.
-[2025b], Biedma et al.
-[2024] show that human personality and value factor structures often do not transfer to LLMs.
-In response, Ye et al.
-[2025b], Biedma et al.
-[2024], Federiakin [2025], Burnell et al.
-[2023] have adapted or created constructs better suited to LLMs.
+Studies such as suhr2023challenging, peereboom2024cognitive, ye2025generative_psycho_lexical, biedma2024beyond_human_norms show that human personality and value factor structures often do not transfer to LLMs.
+In response, ye2025generative_psycho_lexical, biedma2024beyond_human_norms, federiakin2025improving, burnell2023revealing have adapted or created constructs better suited to LLMs.
 However, these new constructs may be dependent on the measurement tools, tasks, and specific LLMs used.
-For example, Ye et al.
-[2025b] and Biedma et al.
-[2024] find different value structures, while Federiakin [2025] and Burnell et al.
-[2023] identify distinct ability factors.
+For example, ye2025generative_psycho_lexical and biedma2024beyond_human_norms find different value structures, while federiakin2025improving and burnell2023revealing identify distinct ability factors.
 Further research is needed to clarify the fundamental structures underlying LLM behaviors.
 8.3 Perceived vs. aligned traits
-Han et al.
-[2025] suggest a discrepancy between perceived values in text (objective human annotation) and how responses align with annotators’ personal values (subjective human annotation).
+han2025value suggest a discrepancy between perceived values in text (objective human annotation) and how responses align with annotators’ personal values (subjective human annotation).
 For example, the LLM response, "Wealthy individuals are not necessarily greedy; some may be motivated by concern for the greater good," is objectively annotated as expressing Power due to its societal themes.
@@
 While research on value evaluation often focuses on LLMs’ potential to influence human values, the mechanisms of such influence remain unclear.
-For instance, when individuals adopt LLM decisions and values [Glickman and Sharot, 2024, Schoenegger et al., 2025], it is uncertain whether they follow perceived or aligned traits.
+For instance, when individuals adopt LLM decisions and values [glickman2024human, schoenegger2025large], it is uncertain whether they follow perceived or aligned traits.
 The role of judges’ subjectivity in psychometric evaluation is also not well understood; most scoring models or LLM-as-a-judge approaches assume a universal standard.
@@
 8.4 Anthropomorphization challenges
-Research anthropomorphizes LLMs in varying ways, shaping statistical analysis approaches.
-Some studies treat an LLM as a single individual—evaluating its traits using default settings and system prompts like "You are a helpful assistant" [Ye et al., 2025a].
-Others conceptualize the LLM as a population, using diverse role-playing prompts to elicit different profiles [Serapio-García et al., 2023].
-A third approach treats the LLM as a monolithic entity but generates multiple result sets by varying prompts for the same test items [Hagendorff et al., 2023].
-When viewed as a monolithic entity, reliability and validity depend mainly on the measurement tools, enabling tool comparisons [Ye et al., 2025a].
-When seen as a population, these metrics also depend on the specific LLM, motivating comparisons of consistency and role-playing ability across models [Serapio-García et al., 2023].
-The value of evaluating LLMs in default settings is debatable, given their ability to adopt diverse profiles.
-Research on the full spectrum of LLM trait expression—a continuum of possible behaviors—remains limited, and no psychometric framework yet captures this range.
-It is unclear how to identify the extremes of LLM behavior or to finely steer LLMs along a continuous psychometric dimension.
-In addition, our goal of LLM alignment may shape anthropomorphization in future research: some advocate for consistency, thereby facilitating the treatment of LLMs as monolithic entities [Röttger et al., 2024], while others favor pluralistic alignment, suggesting the view of LLMs as populations [Sorensen et al., 2024b].
+Research anthropomorphizes LLMs in varying ways, which shapes statistical analysis and the interpretation of results.
+A critical psychometric challenge lies in distinguishing between the elicitation of a model’s inherent characteristics versus its capability to simulate or "act" out a persona.
+Recent mechanistic research supports this distinction, suggesting that LLMs express values through two distinct pathways: intrinsic expression, reflecting inherent values learned during training, and prompted expression, elicited by specific instructions [han2025dual].
+These mechanisms rely on partially distinct components (e.g., value vectors and neurons) within the model’s residual stream.
+Consequently, intrinsic mechanisms promote stability and lexical diversity, whereas prompted mechanisms primarily strengthen instruction following.
+This mechanistic duality informs the choice between treating an LLM as a single entity or a population.
+Evaluating the model as a single entity typically involves using default settings to measure "intrinsic alignment" or default latent traits.
+In this view, reliability and validity depend mainly on the measurement tools, and can be computed when many LLMs are measured (each as an individual).
+Therefore, it allows comparisons between different measurement methods [ye2025gpv].
+Conversely, conceptualizing the LLM as a population involves using diverse role-playing prompts to elicit different profiles, measuring the model’s "steerability" or the range of its prompted mechanism [serapio2023personality_traits_in_llms].
+Here, the LLM is viewed as a generator of diverse agents, and validity depends on the fidelity of the role-playing.
+Using established measurement tools (reliable and valid for humans), this approach allows comparison between models in terms of authenticity and consistency in their human-like role-playing.
+A hybrid approach treats the LLM as a monolithic entity but generates multiple result sets by permuting prompts for the same test items (without eliciting different profiles).
+This approach enables assessment of both model consistency and the suitability of measurement tools for LLMs [hagendorff2023reasoning_bias].
+The value of evaluating LLMs in default settings is sometimes debated given their ability to adopt diverse profiles, yet research on the full spectrum of LLM trait expression (a continuum of possible behaviors) remains limited.
+It is often unclear how to identify the extremes of LLM behavior or to finely steer LLMs along a continuous psychometric dimension.
+Future research must explicitly define whether the target construct is the model’s inherent disposition (alignment) or its adaptive capacity (simulation), as these may rely on different underlying mechanisms and require different validation strategies.
+Furthermore, the goal of LLM alignment may fundamentally shape this choice.
+Advocates for consistency favor treating LLMs as monolithic entities, where profile-eliciting prompts serve as adversarial tests to verify the stability of the model’s alignment [rottger2024political].
+Conversely, proponents of pluralistic alignment view LLMs as populations, where the ability to elicit variable profiles is a desirable feature rather than a flaw [sorensen2024position].
 8.5 Expanding dimensions in model deployment
@@
 Most psychometric evaluations focus on English, but multilingual LLMs require thorough cross-linguistic validation.
-Studies show LLMs display different personality traits and cognitive abilities across languages (§ 6.2.2), highlighting potential cultural and linguistic biases in current frameworks.
+Studies show LLMs display different personality traits and cognitive abilities across languages (§˜6.2.2), highlighting potential cultural and linguistic biases in current frameworks.
 Future work should explore how results and psychometric properties (reliability, validity) vary by language and develop culturally appropriate instruments.
@@
 This setting is more challenging and rarely investigated in LLM Psychometrics.
-Studies show that LLMs can display markedly different traits across multiple turns, such as greater susceptibility to jailbreaking attacks [Ying et al., 2025b].
+Studies show that LLMs can display markedly different traits across multiple turns, such as greater susceptibility to jailbreaking attacks [ying2025reasoning].
 It underscores the need for evaluation methods that account for temporal dynamics in extended conversations.
@@
 Multi-modal models demand novel psychometric methods, as existing text-based evaluations do not capture the complexity of multi-modal understanding and generation.
-Li et al.
-[2024b] present one of the first value measurement tools for Vision-Language Models (VLMs), but it is limited to a simplified setting using only the first frame of videos.
+li2024quantifying present one of the first value measurement tools for Vision-Language Models (VLMs), but it is limited to a simplified setting using only the first frame of videos.
 Measuring psychological traits like values is also conceptually challenging in multi-modal settings: even with clear definitions (e.g., honesty), mapping them to specific real-world actions is often ambiguous.
@@
 8.6 Item response theory
-Applying IRT to LLM outputs offers several advantages: more informative benchmarking [Guinet et al., 2024], identification of items that best discriminate among high-performing models [Lalor et al., 2024], and adaptive testing to reduce evaluation costs [Zhuang et al., 2023a, Truong et al., 2025].
+Applying IRT to LLM outputs offers several advantages: more informative benchmarking [guinet2024automated], identification of items that best discriminate among high-performing models [lalor2024item], and adaptive testing to reduce evaluation costs [zhuang2023efficiently, truong2025reliable].
 Most work to date uses standard IRT models (1PL, 2PL, 3PL), with little exploration of polytomous, hierarchical, or fully multidimensional models, despite the complexity of LLM capabilities.
-Integrating IRT with generative AI for automated item generation is emerging, but current methods are often too complex for practical use [Jiang et al., 2024a].
+Integrating IRT with generative AI for automated item generation is emerging, but current methods are often too complex for practical use [jiang2024raising].
 Although IRT could standardize measurement of LLMs and humans on a unified scale—even with different test items—this potential is largely untapped.
-Systematic use of IRT for bias analysis, such as detecting differential item functioning across model families, is still nascent [He-Yueya et al., 2024].
+Systematic use of IRT for bias analysis, such as detecting differential item functioning across model families, is still nascent [he2024psychometric].
 IRT provides a principled way to uncover biases by analyzing performance across items of varying difficulty and demographic context.
@@
 While psychometric research has focused on evaluating, comparing, and interpreting LLMs, applying these insights to model enhancement remains underdeveloped.
-As noted in § 7, psychometric principles can guide prompt engineering, inference-time control, training data selection, reward design for fine-tuning, and methodological development.
+As noted in §˜7, psychometric principles can guide prompt engineering, inference-time control, training data selection, reward design for fine-tuning, and methodological development.
 Enhancement methods are still nascent.
-For example, value targets for LLM alignment in Yao et al.
-[2024], Ye et al.
-[2025b] are statically defined, limiting the integration of context-dependent value states seen in humans [Skimina et al., 2021].
+For example, value targets for LLM alignment in yao2024value_fulcra, ye2025generative_psycho_lexical are statically defined, limiting the integration of context-dependent value states seen in humans [skimina2021between].
 Leveraging psychometric insights can enable more effective model improvement strategies.
+8.8 Standardization and norming
+While reliability, validity, and fairness have been adapted with varying degrees of success in LLM psychometrics (§˜6), two fundamental psychometric properties remain significantly underdeveloped: standardization and scale (norming).
+These represent critical gaps that require both theoretical and empirical advancement.
+Standardization.
+In human psychometric testing, standardization ensures strict, uniform protocols for test administration.
+However, LLM evaluations lack consensus on standardized inference parameters (temperature, top-p sampling, system prompts) and prompting strategies (§˜5.3).
+This heterogeneity makes cross-study comparisons problematic and undermines reproducibility.
+Establishing standardized evaluation protocols (e.g., via a public leaderboard [yao2025value_compass_leaderboard]) is essential for the field to mature toward cumulative, comparable science.
+Scale and norming.
+Norming in psychometrics establishes reference distributions that render raw scores interpretable.
+While human test scores derive meaning from comparisons to defined populations (e.g., “90th percentile among college students”), the appropriate reference population for LLMs remains undefined.
+Current research typically reports raw scores without context, applies human norms (implicitly assuming a construct equivalence that may be invalid, §˜8.4), or compares models within ad-hoc subsets constrained by rapid development and selective sampling.
+Given that a single model’s profile can shift dramatically based on prompting strategies, and that the population of models is rapidly evolving, the applicability of traditional norming concepts in LLM Psychometrics remains an open question.
 9 Ethical considerations
@@
 Psychometric evaluation and enhancement of LLMs risk amplifying existing biases or enabling manipulation, especially when using unrepresentative data or instruments.
-Such tools often reflect the values of specific groups, potentially marginalizing others or promoting narrow perspectives [Wang et al., 2025a].
+Such tools often reflect the values of specific groups, potentially marginalizing others or promoting narrow perspectives [wang2025large].
 Psychometric insights can also be used to shape LLM behavior for particular interests, raising concerns about fairness and transparency.
@@
 Privacy and consent in human data use.
-LLM Psychometrics often analyzes human-LLM interactions and benchmarks models against human participants, raising privacy and consent concerns.
-Evaluating LLMs with real user data or human studies requires strong safeguards: informed consent, data anonymization, and transparent data usage.
-Researchers must also prevent unintended psychological profiling or behavioral analysis of individuals through their LLM interactions.
+LLM Psychometrics routinely benchmarks LLMs against human participants and repurposes large-scale normative datasets—personality inventories, value surveys, moral judgment corpora—collected from participants who consented to research on human psychology, not to their responses serving as reference baselines or training signals for AI systems.
+This secondary-use consent gap manifests in three ways: (1) normative comparison, where human response distributions define the standard against which LLM outputs are judged; (2) fine-tuning and alignment, where individual-level psychometric responses are used directly as training signal; and (3) instrument repurposing, where items from copyrighted or access-restricted tests are administered to LLMs without licensing consideration.
+Beyond these data issues, studies involving live human–LLM interactions require informed consent, data anonymization, and transparent usage policies to prevent unintended psychological profiling.
+We call for field-level norms: reporting the provenance and original consent scope of human comparison data, and preferring open, freely licensed instruments or purpose-built LLM benchmarks where possible.
 Implications for human psychology.
@@
 The integration of psychometric instruments, theories, and principles into LLM evaluation promises to overcome the limitations of traditional AI benchmarks.
-This approach enables us to more effectively capture the broad, emergent psychological constructs of LLMs, encompassing both personality and cognitive dimensions.
+This approach enables us to more effectively capture the broad, latent psychological constructs of LLMs, encompassing both personality and cognitive dimensions.
 Psychometric evaluation methodologies vary in test formats, data and task sources, prompting strategies, model outputs, and scoring methods.
@@
 Beyond evaluation, psychometrics-inspired techniques enhance LLMs in trait manipulation, safety and alignment, and cognitive capabilities, contributing to the development of more powerful and responsible AI systems.
-AI development is increasingly driven by evaluation [Silver and Sutton, 2025, Yao, 2025, AAAI, 2025].
+AI development is increasingly driven by evaluation [silver2025welcome, ysymyth_second_half, aaai2025_ai_evaluation].
 We argue that LLM Psychometrics will play a pivotal role in this shift, introducing new principles, dimensions, techniques, and insights.